A Bibliography of Publications in Information Processing Letters: 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 October 2023 Version 2.27

Title word cross-reference

#P [Jer94]. (0, 1) [SS92d, SS92c]. (Δ/2) [Pas92a]. (m, n) [AŽ91]. (n, k) [CC95b]. (n^{2} − 1) [Par95]. (s, t) [Cou97]. 0 [Lac96, MR95, Spi93, Zim96]. 0.5 [MT92b, SCGD92b]. 0/1 [MT92b, SCGD92b]. 1 [Den99, Esp92, HHS99, IIT92, Lac96, LL94a, Spi93, Vid91, WHS97, WHH98, Yen91, Zim96]. 1 + ϵ [IL97]. 13/12 [DKST98]. 2 [Acc94, ALLS96, Che92b, CC94a, CLLR97, Clo96, FMF95, FV90, Héb95, HTC92, IR91, KAL93, KP96, LL99b, Lu93, MNT92, MSW93, Mid93, PS95a, PS95b, RS91, Sri99, SR98, Ver99]. 2.5 [AR98]. 2m + 1 [Rhe98]. 2m^{2} [Rhe98]. 2 × 2 [Bsh95]. 2|E| [CT93]. 3 [ACS99, BK97, CP97, CS97a, DKK92, Dem95, Gro98, HR97, Juk95, Kan91, MVR96, Mid93, OS95, Raj92, RW93, Val95b, VS93]. 4 [CS97a, HHK91, SHC^{+}93]. ^{+} [XU99a]. _{5} [Khu90]. _{=} [YY99]. _{C} [YY99]. _{k} [XU99a]. A ∗ ×{b}∗ [GOPR92]. A^{∗} [HK96a]. A^{+} [Wan92a]. AC^{0} [IZ94, Zim97]. α [TU99]. Apx [KLP96]. B [AT95, BS94a, MSV96, DCD96]. B^{+} [MS90a]. β [Obe95]. c [Rac93, THI91]. C^{∞} [Car94]. C_{=}P [Gup93]. ∩ [Jur98]. CNF [Che92b]. d [Bar96]. Δ + 1 [Joh99]. ♢W [Chu98]. E [DF96a, DF96b, Rac93]. E^{n} [Mat91a]. ϵ [JM94a]. f [DGL98, GJLR96]. FewP [RRW94]. G [MN92]. H [BM91, Kan94b, Li94, DMM98]. hv [CD99]. i [RR99b]. k [ACS99, AR98, ASA97, AW93, CGL98, Cha90, CGR98, Che94a, DW96, DK94b, Dre93, Fuj98, Gär93, Gas91, GS98b, GP98b, HL95, IIO97, KFYA94, KKK97, KLP96, KLP98, Kor94, KL98, KL99, KH96, LCC99, Lim96, Mat95b, NM94, NM96, RW94, RWZW91, RW92, SK98b, SB90, ST90, She91b, SSC98, SGB95, WS96, Woe91, XL93]. k − 1 [SGB95]. K_{1,3} [Klo96]. K_{r,r} [TV93]. kμ [Fla94]. L [Che96, Tuz90, TB97]. L_{p} [Sho91a]. LALR(k) [SC93]. λ [Obe95]. λ_{β′} [Obe95]. log(n) [RR92]. log_{2} N [TCL93]. log_{2} n + O(1) [KS93b]. LR(k) [LC91a, LC91b, LK97, PLC93, Pit95]. m [Bar91, Bla94, GVFS99, Gro98, HK97a, LCW94b, LCW94a]. R^{2} [Gai97]. μ [KTS96]. N [BJ91, CST94, ETA92, Kal90, Oh93, AZ90, ASA97, Blu96, CK90, JS95, Mnu91, Par91b, RZ92, She91b, XL93]. n^{log n} [CST92]. n^{Ω(log n)} [RW93]. NC [GKL96, RL96]. NC^{1} [Sax94]. O((log n)^{2}) [TY94]. O(1) [CH93b, PKP93, Yao95]. O(log log n) [MN90]. O(log m) [SB93]. O(log N) [RL94, Kus97, WCL95, KK94]. O(m) [PS95a, PS95b]. O(n) [FPT97, LR96, MN90, WT97, Par91b, YY93a]. O(n^{1.5}\sqrt{m}/logn) [ABMP91]. O(N^{2}) [MC92, KS93b]. O(n^{3}) [SR98]. O(n^{3/14}) [BK97]. O(n log n) [AS97, AZ90, Rub90, Blu96]. O(N log N + cN) [CJKJ98]. O(NP) [WMMM90]. O(\sqrt{N}) [KH96]. O(|V | ∗ |E|) [ACJ96]. O(|V |^{2}) [Khu99, Khu00]. Ω [Chu98, CL91a, CDK93a, HLV92, JLP96, Lit91, MY96, Pin95]. Ω(n log n) [Wag94]. ⊕L [Dam90]. ⊕P [Gre91]. P [Käm90, Rao95, BTv93, CGR98, Kru95, Sho92, SW92]. P2//C_{max} [Zha97]. P_{4} [FPT97, Gia96]. P_{5} [Mos97]. π [VR99]. Π_{2}^{0} [CS90a]. P P^{P H} [Gre91]. R [Dam92, Blu92, DH98]. s [CM94, YT95]. SAT = SAT [Dem95]. Σ^{∗}\overline{L} [Bir96]. Σ_{n}^{0} [Kob93]. Σ_{1}^{1} [Ott95b]. SLR(k) [LC91a, LC91b]. \sqrt{N} [KC91b]. T [Wei91, AT99a, CM94]. t + 1 [AT99a]. τ [CS91a]. Θ(n log n) [MMS90]. ε [AM95, KA97a]. W_{4} [Klo96]. wlp [MM96]. wp [MM96]. x^{2} mod p [VMD99]. x^{n} [NL94]. z [Hil95]. Z_{m} [Sat98, Sat97].

-acyclic [GS98b]. -algorithm [GKL96].

-apex [TB97]. -approximation [ACS99, KA97a, Pas92a]. -ary [ASA97, Bar96, CK90, DCD96, Kor94, KL98, KL99, SB90, XL93]. -automata [CDK93a]. -biased [JM94a]. -Bisimulations [CS91a]. -bit [CT93]. -bounded [Esp92, MT92b, SCGD92b, Yen91]. -boxes [YT95]. -buffer [Hil95]. -calculus [KTS96, Obe95, VR99]. -Center [Kru95, CGR98, HK97a]. -codes [JLP96]. -colorable [BK97, CS97a]. -coloring [Raj92, BK97, Joh99]. -colourable [DMM98]. -colouring [BM91]. -complete [Sho92, SW92, CS90a, Jer94, KLP96, Kob93, Sax94]. -connected [Dre93]. -connectivity [IIO97]. -context-free [CL91a]. -convex [CD99]. -coterie [KFYA94]. -coteries [KH96, NM94, NM96]. -coupled-context-free [Pit95]. -cube [ASA97, JS95, She91b, XL93]. -Cut [KLP98]. -D [CC94a, FMF95, Lu93, MNT92]. -decision [Blu92]. -dense [SSC98]. -dependent [JM94a]. -dimensional [HL95, Kan91, CST94]. -discrepancy [AM95]. -DNF [Den99, Fla94]. -edge [HR97, WHS97]. -evader [KP96]. -factor [AR98]. -fault-tolerant [HHS99]. -free [Klo96, Mos97, TV93]. -generability [Dam92]. -generators [JLP96, Lit91]. -gons [RWZW91, RW92]. -grid [AŽ91]. -Hamiltonian [WHH98]. -immunity [Käm90]. -independent [HTC92]. -inkdot [IIT92]. -join [Wei91]. -laden [Gia96]. -languages [HLV92, MY96, Pin95]. -level [ACS99]. -limited [WS96]. -line [DW96]. -matching [DGL98, Kan94b]. -matchings [GJLR96]. -matrix [SS92d, SS92c]. -minor [Khu90]. -modular [FV90]. -MST [AR98]. -mutual [Fuj98, KFYA94]. -nearest [LCC99]. -neighbor [CGR98]. -nonattacking [Kal90]. -oriented [THI91]. -pairwise [GP98b]. -paths [Cou97].

-patterns [DF96a, DF96b]. -puzzle [Par95]. -QBF [ALLS96]. -queens [RZ92, ETA92]. -quorum [Fuj98]. -reduction [Obe95]. -REF [Clo96]. -regular [Mid93]. -resilient [AT99a, DK94b]. -SAT [Ver99]. -Satisfiability [Héb95]. -SCS [BJ91]. -selective [Rao95]. -selector [BTv93]. -sequences [Bla94, GVFS99]. -server [CLLR97, IR91]. -sets [KLP96]. -shaped [Che96]. -space [Li94]. -spanners [MVR96]. -spiral [LL94a]. -star [CC95b, SHC^{+}93]. -state [Blu96]. -Steiner [OS95]. -subsets [RWZW91]. -term [CGL98, Kus97]. -threshold [PS95a, PS95b]. -tidy [FPT97]. -tree [AT95, MS90a]. -tree-like [MSV96]. -trees [Cha90, BS94a]. -triangulated [RR99b]. -valued [Val95b]. -vertex [Par91b]. -violation [RW94]. -way [AW93]. -words [Car94]. -writer [Vid91].

0L [DPS93].

1 [DS95]. 1-Approximation [DS95]. 16 [SY93]. 180 [RRW00].

20 [NS96a]. 252 [NM96]. 264 [yCGlW99b]. 27 [Flo91]. 2DPDAs [Mog94].

3 [PT92]. 3-D [PT92]. 319 [CK96a]. 32 [Dun90, Gro91a]. 33 [BM90b]. 35 [Pet91]. 36 [AAA^{+}91]. 37 [SH93, Sny91a]. 38 [Mis92]. 3D [SI95].

40 [BCM92, HS94, KC92]. 41 [BDDS94, Pac92a]. 42 [SS92d]. 45 [Pro94]. 48 [Ahu96]. 49 [SS95a].

50 [BP95c, BP95b, BMP94a]. 51 [Har96, LCW94b]. 52 [Cha02, RRW00, Rit95]. 53 [BU95a, Sub95a]. 55 [NAG96]. 57 [DF96a]. 59 [HAM11].

60 [YT96].

70 [yCGlW99b]. 72 [Khu00].

802.4 [OCSF97]. 84m [SY93]. 89d [Flo91].

90k [Gro91a]. 91m [SH93]. 92m [HS94]. 93k [Pro94]. 94c [SS92d]. 95a [SS95a]. 95c [BMP94a]. 95f [Har96]. 95k [Sub95a]. 96d [NAG96]. 96m [DF96a].

= [YY99].

A. [Pro94]. Abadi [DW91]. Abelian [Rit95, Rit94]. abound [GLR99]. absolutely [MS99a]. absorption [UH95]. Abstract [BFSV99, Fv97, CCH93a, CLM96, Kam92, Nys96, THS97]. abstraction [CS91a, RST91]. accelerated [SL97]. access [BCDM97, Cai99, CY97, DF91, Gol99, HR98c, KCK93, Mar94, MZ93, Ogi96, OCSF97, Rob90, SW98]. accesses [FP91a]. accessibility [Mei90]. account [Har94, Har96]. Achieving [Shv91]. across [Bil93, SK91]. action [GW96, LC94a]. actions [CS91a, LvBDW92, Mas91]. active [Hor98]. ACU [Wal98]. ACU-compatible [Wal98]. acyclic [BS99, FGN97, GS98b, HR94, Mal98, QM90, SHS98]. Ádám [LM98]. adaptive [AN93, CC91, LP91a, LL94b]. add [GH94]. Addendum [BP95c, BP95b, Khu00, HAM11]. Adding [YL99, GJS97]. addition [CK93a, TM93]. additions [ITY96]. Additive [Tri92, Bsh95]. address [Lee90]. addressing [Wan90a]. adjacency [Kim96b]. adjoining [DBT90]. adjoint [The95]. adjusting [HM96, Mar91]. adjustment [DP97b, KLR97]. ADT [CYY^{+}92]. advantage [MS99a]. adversary [CL91b, Kuč97]. advice [Ram95]. AESA [Vil95]. affine [SV99]. after [DP97a]. against [Ada92, KOT95, YL97b]. Agent [DCD96, Mas91]. aggregate [CL99, KKK99]. Agreement [BOR96, Das98, FPT95, SW93b, Zhe95, GP90b]. aided [Hor98]. aids [Lau90b]. al. [DW91]. Algebra [Ust96, BP98, Dij96, FZ98, JH93a, LSS92, Lys94]. Algebraic [CMM95, BGM90, Bol93, Che90, CLW98, Gol95b, HP90, Kar90, KAL93, Lan91, Mül92, Rif98, Zip97, von91a]. algebraic-code [CLW98]. algebras [Bac98, Jac91, Tau91b]. Algorithm [CHM92, Sas91, ACS99, AS97, AJG95, AZ90, ABK98, AV94, Alb92, Alb94, AvSW95, All94, ACJ96, AD97, AM90, AR96a, AR90, AM99, AR98, BFR98, BNMN92, BP95a, BP94, BP95c, BP95b, Bax93, BP93, BW90, BK97, BU98, BKMT97, BCM91, BCM92, BDF^{+}95, Bsh96, BF96b, BR98, BC95, CM98a, Car91, CC91, CDF96, Cha98a, CHT90, CWK93, CY94, CY98, CHL97, Che91a, CYH91, Che92b, Che92c, CT93, CST94, CC94a, Che94b, Che95d, CW97a, hC98, CC99, CWL99, Che91b, CC94b, CM99b, CY92, Chi98, CJKJ98, yCGlW99a, yCGlW99b, CP95, Chu97b, CSS92, CK90, CS91b, COS95, Cou97, CL97c, DKK92, DIK97, Dam95, DHS95, DS95, DHO92, DKST98, Den90, DJM94, DKS97, Der95, DK94b, DS93a, DJM98, DIR99, ERC94, Ess92, FF90, Flo91, FM97b, FPT97, FMF95, GP90a]. algorithm [GJT98, GKL96, GJLR96, GL98, GIZ96, GM90, GM92a, Gol91, GBY90, Gon95b, GP98a, GP98b, Gua98, GS91, GLS92, HM98, Han93, Har98, HR97, HR98b, He91, HMP95, Her93, HP96, HW90, Hir96, HS93a, HWJ96, HAM11, Hoo90, HTC92, HH92, HCH94, HC92, HHW94, HMPS96, IR91, IIO97, Jan92, KFYA94, KR92a, Kan94a, Kei91, Kel94, Kha94a, Khu99, Khu00, Kim90b, KC91a, KC92, KAL93, Kim94b, KPC^{+}95, KLK97, KC98, KS93b, KTS96, KPRS97a, KS90, Kle94a, KP92, KG99, KIS90, KC91b, KS97c, Kun93, Kus97, LMT99, LL98, Lam97, LWS99, Lar90, LR94, LS99a, LS95a, LLM97, LY98, LP91a, LC94b, LL94a, LS95b, LO90, Lin90, LK92, LCL93, LS90, Low98a, Lu93, LCW94b, LCW94a, LT97, LP91b, MP91, MSW93, MMS90, MBY91, MW94, Man96, Mat91b, Meh88, MT93, MAUY98]. algorithm [MS94, MS95, NLW96, iNRN97, iNM97, NR99, Nuu94, OS95, Ols97, PM94, Par95, PKP93, PMC97, PH92, Pas92a, Pas92b, Pau90, PC92, Per94, PPK98, Pre93, Pri95, Pri97, Pro95, PN94, RK96, RS99, Rag91, RI91, Raj98, Raj92, Rav94, Ray89, RL94, RL96, Rif98, RR99b, Saa97, SCGD92b, SM92, SL97, Sha93, She92, ST90, She91b, SLC93, SY96, Shv92, Ski96, Sky91, Sla97, SR92, ST93, SR98, SB93, STN90, Tan91, TT92b, TY94, Tip90, TU99, Tsa92, TTC94, UG92, VT90, VKM92, VA97b, Ver87, Vil95, Vis92, Wan90a, Wan92a, Wan94, WCL95, WJ95, WMMM90, WT97, WCT99, YH94, Yan93a, YDL94, Yan98, Yen91, Yen96, YL99, YY92, YY93b, YY93a, Zel93, Zha97, Ziv91, dB93, van91a, VT94, HSC93, TKI99]. algorithmic [CL92b, SS95b]. Algorithms [Bax94, Her96, How95, KA97b, MT98a, MS96a, SKKC98, AM97, ADY91, Alo90, Ama93, AE98, ASA97, BYF96, BHMS90, BC94, BS90, BM90a, BJ91, CEG95, CR98a, CD93a, Cha95a, Cha95b, CW92b, CL93a, CvW94, CJ90, CS95, CR91, CW97b, DV98a, DG95, DD90, EYK95, Epp94, Fv97, FL97, Gar92, GM94b, Gol92, GR97, Gon98, Gre95, GJS99, HK96c, HL95, HCF91, HT91, HM96, HP94, IY98, IK93, JW95, Kar93, Khu90, Kim90c, KK94, KT97, Kle94b, Lat91, Lia95a, LMS96, Lin91a, LV92c, LY99, LSZ93, Mak90a, MH96, MM92, MR96a, MT92b, Mar98, Meg93, Mes91, ME97, Mos97, MN93b, MW97, Pan96b, PS93, Pru94, RR90, RR94, RPG98, RW96, RLDL94, SS94a, SS95a, SR95, Sch92a, Sch93c, SL96, She91a, SM95, SK92b, Tei93, TC99, TM93]. algorithms [TT93, VHV93, VCW95, VSS96, WB93, Wei91, Wen94, Wha92, YTC93, van92b]. aligned [FH94]. alignment [PAD99]. alignments [All94]. all-pairs [LP93a]. all-pairs-shortest-length [HC99a]. allocating [HH91]. Allocation [Rou97, VI98, WS94, FS98, SC90, VR96, WJ95]. allocations [BFL^{+}95]. Almost [Cla92, Luc98b, Amb99, Bur92, Kal90]. almost-constant [Bur92]. ALOHA [GS90]. Alpern [Gum93]. Alpern-Schneider [Gum93]. Alphabet [AFM94, BV92, Jed90]. alphabetic [HT98]. Alternating [Mos90, MP95, Slo90]. Alternation [GGK91]. Alternative [Wha92, BB93, FP94]. Always [Pet91, VT90]. Always-true [Pet91, VT90]. ambiguity [AL90b, DBT90]. ambiguous [Leu92]. among [BP95a, CK96a, CK96b, CM94, Grä90]. amortized [IL97]. Amplification [MS99a]. analog [RP99, TV96, Weg96]. analog-to-residue [RP99]. Analysis [AP93, Hac93, Jag97, JW95, KLL93, LPP91, Par91a, Ped94, PS94, Pom92, Rag91, CD93a, Cha98a, CWK93, CW98a, CM99b, CY92, DIK97, FJJR95, GM94a, GBY90, GS96b, GP98a, GS91, HL95, How95, KK92a, KM96, Lau90b, LS99a, LC99b, LBL98, Mat90, NB96, PS92, PG95b, PG95a, Pau90, PS93, Pro94, RR93, Sch93a, She93, TT90, VM97, ZD90]. Analyti [Pro94]. analytic [WC92a]. analytical [Oh93]. analyzer [Naw91]. Analyzing [HM96, SW90, SW98]. ancestors [Lin91a, NU94]. AND-parallel [LKC94]. AND/OR [KC91a, KC92]. ANDs [Zwi96]. angle [WCW90]. animate [BYF96]. animating [Tsa90]. annotated [LH92]. anonymous [CS95, MR96b]. ANSV [XU98]. answer [Pin95]. answering [CL99]. antipodal [JT94]. Antwerpen [Min91]. any [SHC^{+}93, Sho91a]. aperiodic [DB95]. apex [TB97]. Apostolico [CL97c]. Apostolico-Giancarlo [CL97c]. Appel [Ess92]. Applicability [SL96]. Application [KC98, WWT90, AM95, BP93, Che92c, Che94b, DB91, HP94, Ken91, Kle94a, LS95a, LBL98, MW94, MW91, RL96, SDG92, SW90, SW99]. applications [Dan91b, DCD92, DTW97, Dom96, FLSS93a, FLSS93b, JS95, MT98b, Raj92, Rub92, SCH95b, SSCK94, TB97]. applicative [Goe90]. Applying [Chu99]. approach [BLM90, CH97, CDA91, CYY^{+}92, Chu91, CDK93a, Fuj96, Li99, LRDL91, Miz99, RP99, RP91, Tho95, TKI99, WCC91, WT92]. approaching [LC99b]. Approximability [KLP96, AP95, GMM94, PS98]. approximable [AT99b]. Approximate [Aku95, Ber93, Hac93, KLP98, Sch92a, Tol93, VR96, Ver97, AL95, BYP96, BJ92, CFGG97, DG97, GL98, Hal93c, Hef91, HR98c, JW95, Lia95a, MW94, Meg92, NBY99, Sen90, Yao92, Wag94]. Approximately [Mye95, NO96, Kar93]. Approximating [AHH93, Chl96, FGHP93, GMSS99, GLY98, Hal93b, KKK97, NL94, Ser91, SW92, ABF94, BMST97, CKK97, Irv91, Rou97, SHS98, TB97]. Approximation [DGP94, LV92c, ACS99, Ali96, AE98, AR98, BNHK99, CR98a, CT97, CFS91, DGS97, DKST98, DKS97, Flo91, Fuj96, Gon95b, Gon98, GR98, Gus94, HPS93, HR97, HR98b, Jan92, KA97a, Kle94a, KG99, MR96a, Meh88, MMN91, Ols97, Pas92a, Rav94, VA97a, Vis92, Zel93, dlVK99, DS95]. Approximations [HR94, Luc98a]. April [BP95c, BP95b, Mis92]. arbitrarily [GR98, NU94]. arbitrary [GI91, Mil92a, RK96, SHL94]. Arbitrated [HW98]. arborescences [Woe92a]. Arboricity [Epp94, GKL96]. Arc [Che99, BM90a, ELS93, HT91, LR93]. architecture [KM96, PDBF91]. architectures [BS90, Gei97, KHJ98, LHW99, Man94, Pom92]. arcs [Kim90b, SG92]. Area [KLTT97, BCDM94, Kim97, LCC93, Lim96, PP90, Tre96]. area-time [PP90]. area-universal [BCDM94]. arising [PS93]. arithmetic [AM98, Bar91, Dam99, Hea90, HV96, Mei91, Weg93]. Arrangement [DT92, CC98, Ram98]. arrangements [Alf99, DS94a, HS91a, HS94, IT94, Sha96]. array [AO92, KC90, Kri93, Lin90, Lin91a, Lin96, MG90, WCL90, WC90]. arrays [GJT98, HI90, PP94, TM93]. ary [ASA97, Bar96, CK90, DCD96, Kor94, KL98, KL99, SB90, XL93]. ASCEND [VHV93]. ascent [Lee92]. aspect [HLV97]. aspects [ED91]. assay [SJKL95]. assembly [KLW93, KK95a]. Assertional [Pet91, VT94]. Assertions [Ahu93, Ahu96]. Assessment [MZM99]. assignment [Che92b, DB95, GK98a, WH94]. assignments [Gra96b]. Associative [Bac92, CH92b, CK90, RS97, WWT90]. Associative-commutative [Bac92]. associativity [Jac91]. assumption [CS90a, Col94]. assumption-commitment [Col94]. asteroidal [KMW98]. asymmetric [Vis92]. Asymptotic [FGW98, GP98a, HR98a, Pro94, Tsa92, Woe98]. asynchronous [Ahu90, FMT93, HL91a, MN98, Nas95, Rag91, Sin92a, SM95]. AT-free [COS95]. ATM [VI98]. atomic [KST96, LV92b, Vid91]. atomicity [Ser96]. attack [Hor98, KOT95, Low95, RS98, SGSD99]. Attacks [GGOQ98, Hwa92a, TJ99, GSSD97, HLL^{+}95, Tze99, TH99, YL97b]. attractor [ESW97]. attribute [SO97, Sid90, TW94]. attributed [Pav92, TW95]. attributes [HS93b]. augmentation [Kao95]. August [Dun90]. Authenticated [HCY96, LC97a]. authentication [Hor95, HC95, HLL^{+}95, HW98, LC96, LL95, Low95, RS98, TH99, YL97a, YL97b]. Author [Ano96a]. autoepistemic [Yua94]. Automata [CMR98, GM92b, Blu96, CDK93a, CI95, CR90, GS96a, Her98, JR91, Kai95, Kut90, LP94a, Leu92, MR95, Mor92, Mor94, Mos90, Pet94a, Pet94b, RV96, RS95a, Sal96, Sat93, Sat97, Sat98, SDB97, Slo90, Tze96, Wan92b, WIIO99]. automata-models [SDB97]. Automatic [McC98]. Automorphisms [IOU94]. availability [AW98, CW99, KC91b]. Avalanche [Cus96, CS96, O’C94, SZZ94, SLCS98, SCP99, YT96]. Average [BMI97, LV92a, Pru94, TTC94, BS94a, Ber96, Bop97, CEG95, CWK93, Dan93, GGH94, Man94, NCK99, THS97, Ven98]. Average-case [Pru94, Ber96, CWK93]. AVL [GM98, Kim97]. avoid [AKM95, LvBDW92]. Axiom [Qia90, RF93]. Axiomatic [OH91]. axiomatization [Fok94, Mal98, SR99]. Axiomatizing [Sza92]. axioms [BIN92]. axis [FH94, SK98b]. axis-aligned [FH94].

Bachmair [Lys94]. back [Bar96, Lar98]. backtrack [Tri92]. backward [KC91a, KC92]. Backwards [Cha98a]. bag [KR92b]. bags [LW95b]. balance [BY91]. Balanced [BFL^{+}95, Chl96, Ber94, Ber99, DMM98, GZ91, Mah98, Pug90, SLCS98, SCP99, YT95]. balancing [Coh94, Kar90, Kle94b, LS90, MP97, XL93]. banded [ADY91]. Bandwidth [KKM98, Mac95, MSW93, SGB95, VI98]. bandwidth- [MSW93, SGB95]. BANG [THS97]. Banker [HSC93]. Bardzin [ESW96]. barrier [KHJ98, KA97c, UH95]. Barrington [Cau96]. based [AB92, BCLM92, CCK93, CST94, hC98, CM96, Cus95, DK94b, EV97, FLMSS98, Fuj98, GS96b, HI97, HHL95, HWT99, Hwa92a, Hwa92b, Hwa93, JCH94, KKK99, KB90, LCH93, LHW99, Li99, MZM99, Mah91, ML95, PCK95, Pom92, Ros93, SGD98, SM92, SM90, WLF94, Zhe95]. bases [Acc94, ABF94, Dan96]. basic [BIN92]. basically [BL96]. basis [CGH95]. Batch [KS97a, Rub92]. BCH [LCW94b, LCW94a]. BCNF [MC92, VS93]. BCNF-decomposition [MC92]. be [All92, CY96, IL97, LSW98, Mog94, Par91b, van97]. behavior [LCH93, She93]. behaviour [AN93, Mes91]. behavioural [BIN92, But95]. behaviours [Kuč99a]. bends [Bie98]. benefits [BL91]. benevolent [CL91b]. best [GP90b, TU99]. best- [TU99]. Better [AE98, BKR94, CY96, CLLR97, Hal93c, Yan95]. between [BCMS97, CXZ99, CRC98, CDK93a, Fok97, GJS99, Hal93a, IIT92, JM94a, JM94b, Kem99, LC97a, LRH^{+}95, Liś93, NPW98, Rot91, RS95b, SRW95, TT91, XU99a, ZSS92]. beyond [Han92]. BFS [Jia91b]. bi [DS98]. bi-directional [DS98]. biased [Dij90, JM94a, Ueh95]. bicategories [Kle94a]. bichromatic [vB97]. biconnected [AR96a, ERC94, LR96, MSW93, Man96, PN94, STN90]. Bidirectional [GM91]. Biham [Ada92]. Billiard [AEE97]. bin [Cha92, CWK93, DKST98, IL96, IL97, van92b]. bin-packing [CWK93]. Binary [Gho91, Ser98, AS92a, BR94a, Bar91, BP94, BP95c, BP95b, BV92, Bra90, CC94a, CN94, CC95a, CK90, DL92, DD93, DP95, Dev98, FMF95, GS92a, GL98, GZ91, Gol95b, HW93a, HT98, KR92a, Kem99, KK92b, Kor93, Kru99, LR91, LCW94b, LCW94a, Mah98, NCC98, Pal90, RG93, SKKC98, Vaj98, Wen94, Yan93b, YLW91, dF92, van91a]. Binding [Han92, MABD90]. binomial [Fag96, PS93]. bins [DKST98]. Bipartite [DLR90, ABMP91, Che95b, CDP96, CKU98, DMK90, Epp94, HC99a, JW95, Kao95, KK95b, LW93, LL98, LB95, MVR96, MN93b, Ros99, SMN^{+}95]. bipartite-completeness [Kao95]. bipartition [STN90]. biprefix [Pir94]. biquaternions [DCD92]. bisecting [She92]. bisection [YH94]. Bisections [Sto91b]. bisimilarity [Bas96]. Bisimulation [BŞ94b, AI97, FGR91, Gro92, HSZF93]. Bisimulations [CS91a, Mow92]. bispecial [dM94]. Bit [BT90, Lin91b, ACH97, CT93, KC96, TKI99, Wis93, AvSW95]. bit-level [TKI99]. Bit-optimal [BT90]. bit-sliced [KC96]. bitonic [Pau94]. Bits [DMS94, yC90, KS93b, New91]. bivalency [AT99a]. Black [Enc97, Amb99, CW92a]. Black-box [Enc97, Amb99]. block [Ber96, Cha02, Chr96, Chu91, FP91a, HS93b, Ven98, YC94, Chi93]. block-interchanges [Chr96]. blocked [KT91]. Blocking [Bar90]. blocks [LC97a]. bodies [Ker92]. body [San96]. bolts [ABF96]. Boolean [NS96a, Amb99, Baa98, BÖS91, Ber99, Bij93b, BB94b, Cla92, Dam99, Dan96, GM94a, IZ94, MS99b, Mnu91, NS96b, O’C95, PLSK97, Ros97, SLCS98, SCP99, TV96, Vei97, Zwi96]. both [GP90b, Tse96]. bottleneck [DS95, Man96, PPK98, PN94]. bottom [CCH93a, RW93, SB95]. bottom-up [CCH93a, SB95]. Bound [Liu90, AL90a, Al 96, AGM92, BNHK99, BKR94, BMP94a, BMP94b, BCDM94, BR94c, Bsh92, BMST97, BdW99, CvW94, CL93c, CLLR97, CP90, Den96, DZ98, GS96a, GH92, GHY93, Gro98, KS91, Kir94, Kle94b, LPK90, MP97, O’C94, O’C95, Oya95, PS98, PP90, Rab98, SCGD91, Sch97, Sga97, SW90, Sub95a, Sub95b, SYUK98, Tak92, Tei93, TT91, TU99, YW96, van92b]. boundary [FP98, JM94b, KAL93]. Bounded [HK97a, IZ94, MN90, BOR96, Bop97, Esp92, GH98, Gup92b, Gup93, HS93a, How91, Kan91, Kan94b, LX99, MS98, MT92b, Mei90, MP95, SS94a, SS95a, SCGD92b, Sim91, Sny91a, Sny91b, Yen91]. bounded-degree [MS98]. bounded-depth [Bop97]. bounded-probability [Gup93]. bounded-space [HS93a]. Bounded-width [IZ94]. Bounding [ESW97]. Bounds [CS93a, Cus96, CS96, YT96, AERT92, Alt90, Ama93, Bas91, BD97, BCT93, BCDM97, yC90, CEG95, Che92a, CM98b, CL97c, Fra99, Gei97, GH94, HR90, HI90, HR98a, IP95, Joh98, KO96, LW95a, LK96, LO94, LP97b, Man94, Mei91, PLSK97, PK93, Rad92, Raj94, RV93, RW93, RW97, Sal92a, Sny91a, Sny91b, TTV91, TV93, TT94, Tom95, Vol96, Weg93, WY95]. box [Amb99, Enc97]. boxes [YT95]. BPA [Gro92]. BPA-processes [Gro92]. BPP [Can96, Ueh97]. branch [SCGD91]. Branching [Bas96, AN93, BW98, BTW98, Cha95a, Dam92, Gál97, GW96, Juk99]. Breadth [CD92, CHL97, HC92, Man90, RLDL94, Sim95]. Breadth-first [CD92, CHL97, HC92, Man90, RLDL94]. Breaking [BBR99, HR98a]. BRGC [Wan94]. bridge [CXZ99]. bridges [SK92b]. brittle [Gia96]. Broadcast [WC97, Ahu93, Ahu96, Fra98, GP98a, Mar94, Mid93, SY96, SW99, Yan91a]. Broadcasting [DV98b, DDK^{+}98, FP91b, Pel91, Stö91a, Bar96, BGOS95, Chi98, DV99, Gur93, Gur94, IOU94, LHW99, SLC93]. broadcasts [ST93]. brother [GM98]. Bruijn [HP97a, HS95, Oht93, Sch92b, SSO94]. BSP [JW96, KHJ98]. BSR [XU99b, XU99a]. BSRs [XU98]. BST [Mes91]. bubblesort [Vys90]. Büchi [Kai95]. bucket [Mat90]. Buckets [Hac93]. budget [SK98a]. budget-constrained [SK98a]. budgeted [KMN99]. buffer [Hil95, LPK90, Zha97]. Building [RZ91, DW94, LP93b, NMS97b, NR99]. built [CS99]. built-in [CS99]. bus [Pom92, TM93, WCL90, WC90]. bus-based [Pom92]. bused [CL93c]. buses [Lin91a, Vai92]. Bushiness [SYUK98]. butterflies [HS97, Luc98b]. butterfly [BR94b, CDK93b, Sch93c, Stö91a]. Byzantine [GP90b, TIK99].

C [SH93, YY99]. cache [SS98b]. cached [Pug90]. cactus [LWS99]. Calculating [BCMS97, Bij93a, CC99]. calculation [Küh98]. calculational [Dij95a, Hoo90]. calculi [BFSV99, LW94]. calculus [CHR91, IP95, Kai95, KTS96, Mat95a, Mas91, Mül92, Obe95, VR99, vH95, von91b, Rab98]. call [Tom95]. calls [Bij93a, FP91b, Pel91].

Campbell [SH93]. can [All92, Asp95, HR98c, IL97, Mog94, Par91b, Var90]. Cancellation [Sen97]. cannot [LSW98]. canonical [MMS90]. capacity [KLK97, Li90]. cardinality [ABMP91, Ken91]. cards [DKVW95]. care [Aku95]. cares [MBY91]. Carlo [HK96b]. Cartesian [Wei94, Wei95]. case [Ber96, BR94c, BDF^{+}95, CWK93, CY92, Chu99, Cri91, DvR94, FGW98, GHY93, KS97a, KT90, LS99a, LV92a, Par91b, Pau90, PS94, Pru94, Sak91b, San90, SYUK98, Vos93]. cases [ALLL98, But95, Chu99]. CASOP [CK90, Hoo90]. causal [AB92, AV94, RST91]. causality [GW96]. cause [Ust96, Ust98]. cause-effect [Ust96]. Caution [Mul91]. cautionary [RS98]. Cayley [BV98, CP97, THHC99, VS95a, VS96, WM96]. CCS [Mow92]. cell [dDDS97]. cells [LCC93, Lim96]. cellular [CI95, Mor92, Sat93, Sat97, Sat98]. Census [Hem94]. Center [Kru95, CGR98, Her93, HK97a, HHW94, LWS99, Ver97]. centering [AR96a, ERC94]. centres [MS96b]. certain [CS99, Cim90, Gup92b, MM92]. certificate [AM99, YS96]. CFL [RR92]. chain [Bil93, HNC92, Kim96a, KKM98, PK93, SC90]. chain-like [SC90]. chained [HNC92]. chains [Alt90, BM92, DKD97, HS91a, HS94]. changes [BLW96]. channel [FS98, GJK92, GM92c, Li90, TS95a]. channels [Jac90b, YL97a]. character [Per94]. Characteristic [DP95, Acc94]. characteristics [SZZ94, SLCS98, SCP99]. Characterization [Ito91, LK97, AI97, BS96a, BNN91, BS97, FP94, Gum93, Kum95, von91a]. characterizations [AQS97, Maz95]. characterize [SL95]. characterized [HPS93]. Characterizing [eC91]. characters [Aku95]. chart [Ros93]. checkable [RS95b]. Checker [Wie97]. Checking [LNP^{+}90, Cad92, Jac91, MW94, Rub92, Vol96]. checkpointing [TWK99]. checkpoints [WLF94]. Chernoff [HR90]. Chervonenkis [Sim91]. Chinese [hC98]. choice [GL98, MW91]. choices [Hd91]. Choosing [BD96]. chordal [CL93a, CJ97, DMK90, HC99a, IKOI99, KM91, KK95b, NAG95, NAG96, Pan96b, PC92, Sim95, VKM92, dFMdM99]. chords [CL92a]. chosen [Tze99]. chosen-message [Tze99]. Christmas [HHS99]. chromatic [Rhe98, dFMdM99]. Chvatal [Ola90a]. cipher [Gol99]. circle [CM98a, CL92a, Kim90b, Mat95b, Sky91]. circles [DG95]. Circuit [Mil92b, BL96, FGW98, Joh98, MMS90, Mnu91, MRR96, TNB96, TT94]. Circuits [Lan91, Bop97, yC90, Dam99, Dun90, Gro98, HHK91, IT94, Juk95, Koc89, LB95, LL99b, RW91, RW93, Ros97, Sin98, SC94, VCS96, Weg90, Weg93, Weg96, Zwi96]. circulant [CY98, KC98, Mic97]. Circular [LB95, BM90a, HT91, Kim90b, LR93]. circular-arc [BM90a, HT91, LR93]. circularly [BS98]. circularly-hidden [BS98]. circumscription [Yua94]. circumscriptive [Cad92]. Clark [Gra97]. Clark-Jacob [Gra97]. class [CW98a, DT92, Dem95, Gia96, Kem99, LR96, SGB95, Tri92, Xue96]. Classes [Mul91, BL96, BS97, Cai99, CS99, Cre98, Gup92b, GPV90, Hem94, Kos99, Kum95, MV95, Mäk92a, Mos97, Ram95, RRW94, RRW00, Sze98]. classification [BS96a, Cai99, GVFS99, MABD90]. classifier [WWT90]. classifiers [Kun93]. clausal [WC92a]. clauses [BCLM92]. clipping [FNTV90]. Clique [BNR96, CW92b, Che92a, CJ97, GH92, Kim90b, Lip96, RL96, CFS91]. clock [HL98]. clocks [CB91, HG95, MK94, SK92a]. closed [BHW91, CS90a, SSO94]. closeness [Rao95]. closest [Ber93, GMSS99, SS92d, SS92c, Woe92b]. closest-point [Ber93]. closure [ADY91, DP97a, Fv97, Gup92b, KP98a, MP96, Nuu94, Ola90b, Tho95]. closures [Mon96, Zha91]. CLUMPS [Cam98]. cluster [GP95b]. clustering [Chu91, DCGP99, DKS97, KP94, YH94]. Cn [Che90]. co [AB98, Jur98, Kim94a, YY99]. co-C [YY99]. co-minima [Kim94a]. co-prime [AB98]. co-UP [Jur98]. coalescing [LCL93]. coarsest [CH92c]. cocomparability [RL96]. code [CLW98, EST99, Kim96a, Rif98]. coded [BC95, Sun96]. Codes [Sta96, Bar95, Bra90, Bur93, JS95, Juk99, JLP96, JP97, KL99, Lam95, LS95a, LCW94b, LCW94a, Măn98, Pir94, Yan91b, YL97a]. coding [DML98, HV96]. coefficients [KA97b]. cographs [YY93a]. coin [Dij90, HCH94, Ueh95]. Collapse [Ueh97, Hem94]. collapsing [Boo94]. collection [Dev98, Tho95, YW96]. collections [BU95a, BU95b]. Collisionful [Gon95a]. collisions [Gon95a]. color [CR98a, Ola91, RR99b]. colorable [BK97, CS97a]. Colored [Ust98]. Coloring [Khu90, KS98b, Alb92, BNMN92, BNHK99, BK97, DGP94, GZ97, Hal93c, HH94, JW95, Joh99, Lia95a, iNM97, Raj92, SS94a, SS95a, Sax94, VR96, YY93b]. colorings [RS94b]. colors [HH94]. colour [CL92b]. colourability [DZ98]. colourable [DMM98]. colouring [BM91]. combination [OiMI98]. Combinatorial [DCH96, KO96, Ber94, TIK99]. combinators [Lar95]. combined [AvSW95]. combining [KKK97, KW95]. Comment [Dun90, TM98, YT96, Pro94]. Comments [Pet91, VT90, WHL99, AAA^{+}91, DW91, Dub90, Gro91a, LP93b]. commitment [Col94, Raz94]. commodity [KRY94]. Common [BM90c, BM90b, Tze99, But95, HM98, LLM97, Lin91a, New91, NU94, PP90]. common-multiplicand [HM98]. communicating [Goe90, Jac90a, Jac92]. Communication [ASA97, JW96, VR99, Bar90, BBGT96, CK93a, CMM95, CFGG97, Den96, Gia94, GH92, GH98, GP98a, Lin94, New91, PDBF91, Res96, SY96, TWK99, Yua90]. communication-induced [TWK99]. communication-privacy [CK93a]. communications [Hwa93, Yan91a]. commutative [Bac92, BL90, CDÉ99, MP96, Sat93]. commute [Hec91]. Compact [CC95a, DHO92, Kim96b, Que98, Ziv91]. compacted [LCC93, Lim96]. compaction [Hag92]. comparability [BNR96]. Comparison [DML98, BS96b, BCT93, BCDM97, HK96b, SS98a, WMMM90]. comparisons [DK94c, HK96c]. compatible [Wal98]. competitive [BMI97, CLLR97, HCH94, IR91, PMC97, RR94]. competitiveness [TT93]. compilation [Dan91a]. compilers [LCH93]. complement [IY98, Omo91]. complementation [PK99]. Complete [BIN92, Ste91, BBGT96, CS90a, Dam90, Den99, DK91, Fok94, HR98a, Hea90, HR98c, IT94, Jer94, Jon98, Kan91, Kan94b, KLP96, KK95a, Kin92, Kob93, Kui95, LW93, LR91, LCW94b, LCW94a, Mal98, MS96a, Mid93, Pel91, Rad92, RRS98, RG93, Sax94, Sei94, Ser91, Sho92, SW92, Tre96, VI98, YL91, YP98, SR99]. Completeness [MR95, Dre93, Kao95, LH92, Pac92a, Pac92b, Ram90, TCC91, WCL98]. completion [Ray95, NN91]. completions [Lam95]. complexities [HP90]. Complexity [CCK93, Cre98, CKM97, Dan96, HLW93, HH91, SS92d, SS92c, Acc94, ACH97, Agg92, AGM92, ÀG91b, Amb99, Ano94, AD96, Baa98, Bac97, Ber99, Bir92, Bir96, BR94c, Bop94, BU91, Bsh95, Cad92, CY94, CL92b, CL93b, CH93b, CW98b, CM94, CD98b, Chi90, CH92c, CJKJ98, Clo96, CMM95, CDF90, CS90b, CS93b, CL97c, DGL98, DS94a, DB91, DW94, Fei92, Fis95, FMS97, GM94a, GH98, Hag95, HKW90, Her97, How91, How95, IS98, JR91, KKS91, KLW93, KV94, Kir92, KST96, Lar90, LV92a, LC97b, Lin91b, Man94, Mei90, MP98b, Mon96, MK92, New91, PS94, PK99, PP90, RS97, Res96, RS91, San90, SR99, Sho90, SS92b, Sny93, Sze98, Tak92, TNB96, Tod90, TS95b, Tys97, VMD99, Wan91, Wan97a, Wan99b, Weg96, Woe92a, Yan93b, Yao95, Yua90, Zim96]. complexity [Zim97]. complicated [Grä90]. component [Che94b]. components [Fer94, HL91a, Jia93, LR96, MM99, NSS94]. componentwise [Kao95]. composite [BBR99]. composition [BMC96, Col94, CK90]. Compositionality [FGR91]. Compressibility [Her97]. compression [DS96a, HV96, YL99]. Computability [SV99]. computable [Maz95, Ogi96, Sch95a]. Computation [Mor92, BGOS95, BCM91, BCM92, Cam98, CH92b, CM94, DS96a, DBM97, Fer94, Gur93, Gur94, Han95, Hor98, HWT99, JCH94, KHJ98, KK92b, PR91, SW98, TV96]. Computation-universality [Mor92]. Computational [Höf93, CL93b, DG97, Gol90, HP90, LC97b, LP97b, MM99, Mon96, Nas98, van93]. computationally [Pău96]. computations [Fid98, FHT93, FJJR95, HL92, KB90, Nys96, Tho98]. compute [COS95, IZ94, KAL93, MS94, MS95]. computer [BLP98, Dan91b, NF92]. computers [CST94, FY93, GZ91, JW96, MP91, SC90]. Computing [AS90, ABMP91, CXZ99, CRC98, CT91, CD94a, Dan93, FH94, Juk95, KK95b, Kuč97, LP97a, Mat91a, MR92, NMS97a, Rot91, Sho91a, Sri99, SW93b, Vel94, Wis92, Zhu92, dDDS97, vB97, Acc94, Agg92, BR97, BU98, BM92, BC95, CHL97, Che92c, CC94b, CK90, DKK92, GM90, GM92a, Hir96, Lan91, MN98, Pri95, Sen97, Sky91, Wan92a]. concave [GP90a, HS91a, HS94]. concept [BL96]. concepts [SS98a, Wan99a].

Concerning [CB91, DT93b, GHS93, MR95, WS96, ZJ94]. concise [Yod91]. Concurrency [KLR97, ÀG91b, Bar90, Che91b, KK98, SZ92a, Yen91]. concurrent [BFSV99, CCK93, Gum93, HMPS96, Miz99]. condition [CY96, KT90, Sta96, TCL93]. conditional [SV99]. conditions [AF96, CL99]. conference [HI97]. configuration [HI97]. configurations [RL93]. Conflict [Naw91, Esp92, Yan91a, Yen91]. conflict-free [Esp92, Yen91]. conflicts [HK97b, LC94a]. Confluence [Mül92, KMTd94]. confluent [Ohl95, Ott95a, STKK97]. conformance [CLBW90]. congruence [ETA92, Hef91, IN90, Sch92a, Val95a]. Congruent [HSZF93]. conjecture [Alf99, GHS93, LM98, Lub90]. conjunctions [BCBF95, GM97, MP98a]. Conjunctive [DS96b, War98]. Connected [AR92, Chl96, CH98, CS91b, Dre93, Fer94, FHL97, GZ91, HL91a, HI90, Jia93, KC90, KK95a, Kri93, Lia95b, MM99, iNRN97, NSS94, Pas92b, iT96, WLK92, dBDL96]. connectedness [Khu99, Khu00]. connection [Bir96, Lip96, NF95]. connections [Bac98]. Connectivity [Kul97, Lau90a, ASS97, BC93, Epp97, IY98, IIO97, KKS91, SR95, Sin92b]. conquer [PAD99, WF96]. Consecutive [DG99, CW99]. Consensus [YWC99, AT99a, CY94, FMT93, Tau91a, Nei94]. conservative [How91, VCW95]. Conservativity [LW94]. consistency [AF96, Che99, DK94c, LNP^{+}90, Lys94, Vel96]. Consistent [WLF94, Den99, SHS98]. Constant [Nik95b, Sch93c, WCL90, BS94a, Bur92, yC90, DR94, DP97b, FS97, LL99b, RW91, Sen90]. constant-depth [yC90, RW91]. Constant-slowdown [Sch93c]. Constant-time [Nik95b]. constants [Bac97]. Constrained [DTT92, MMN91, Bir90, NDM99, SK98a, SK91, TT92a, WCT99, Rou97]. Constraint [How95, Nik95a, BP93, BCS97, How98, HM95, KLK97, Li99, Nie93]. constraint-based [Li99]. constraints [Amo97, BK98, BU98, BJ97, Cha98b, CW99, DS96b, FS98, HL92, HM95, LCL93, WY92a]. construct [JS95, KS97c, KH96, SI95, ZI98]. constructed [HS97, Pet95]. Constructing [Cha94, The95, Hir97, Lei95, TIK99, CHT90, CYH91, CS90b, HC92, Sha93, WL93, WCL95]. Construction [AC90b, Man99, VB96a, CR90, Hes98, Kho93, RU93, TR95, Vid91, Wei94, Wei95]. constructions [Hec91, OM97]. constructive [CP90, MG92]. Constructor [DS93b, DS94b]. contact [CKU98, RS94a]. Containment [HS97, DS96b, GG96, Her96, ZÖ94]. content [BS96a]. contention [JH93b]. contention-free [JH93b]. context [AGM92, AL90b, CL91a, DB91, Gál94, Hùy91, JCH94, Mai94b, Mäk92b, Mäk92a, Mye95, Pit95, RK96, RS99, Ros98, Slo90, TW94, Zha92]. context-free [AGM92, AL90b, Gál94, Hùy91, Mai94b, Mäk92b, Mäk92a, Mye95, RK96, RS99, Ros98, TW94, Zha92]. context-sensitive [DB91, Slo90]. continuations [Wan90b]. contour [Bea91]. contraction [PR91, PRS96]. contribution [LS99b]. Control [KLR97, Che91b, FLMSS98, KK98, OCSF97, Tom95]. Control-Time [KLR97]. Convergence [HK96c, CD93a, Coh94, PK93]. conversion [BP94, BP95c, BP95b, KA97b, RP99]. convex [BGOS95, CXZ99, CD99, CDP96, DMK90, DvR94, DW96, DG95, DS94a, FH94, GJS99, He91, Ker92, Lar90, Lê97, LB95, MRSW95, Nie96, RWZW91, RW92, Saa97, Sto91b]. convex-hull-and- [DW96]. convex-hull-and-line [DvR94]. cooperating [BH99]. cooperative [Dan91b, LL94a]. coprocessors [Tsa93]. copy [Wis93]. copying [BV92, Tho95]. core [Alb94, OSK96]. cores [DB91]. correct [THHC99]. correcting [LS95a, LCW94b, LCW94a, YL97a]. correction [Mae90, MS95]. Corrections [KIS90]. correctly [GH94]. correctness [AM99, EV97, JH93b, Sis91]. correctors [GS92b]. correlation [GSSD97, MS99b, PLSK97, SGSD99]. correspondence [Kem99]. corridor [Che96, SSC98]. Corrigenda [BM90b, KC92, LC91a, Mis92, Myo92, SS92d]. Corrigendum [BMP94a, BDDS94, BU95a, BCM92, Cha02, FLSS93a, HS94, LCW94b, Pac92a, RRW00, Rit95, SS95a, Sny91a]. cost [BU98, CS90b, Cri91, CDK93b, JCT99, LCC99, MS99a, PR91, Ser98, SW92]. costs [DCD96]. coterie [KFYA94]. coteries [DKK^{+}94, KH96, NM94, NM96]. cotriangulated [SMN^{+}95]. count [Bax94, BF96a]. Counter [Woj99, HS91b, KS97c, AM97]. counter-example [KS97c]. Counterexample [Lub90, She91a, SD91]. counterexamples [McC98, Sak91a]. counterfeit [HCH94]. counters [RS90]. Counting [BGM90, Jer94, Kor93, Loz92, MRSW95, PP94, RWZW91, RW92, SSO94, Bus92, yCH91, Chi90, GJS99, Kar93, KM90, Lin92, Lin93b, MWL90, PG95a, SS92a, Wis93]. countries [Riv93]. coupled [MS90a, Pit95]. cover [AR90, BFR98, Che91a, Fuj96, Gon95b, LS95b, Low98b, RL96, Rub90, Sla97]. coverage [KMN99]. Covering [Gon91, Hea90, BK95, Esp92, KCLA99, LC91a, LC91b, Ov98, Rad92, GHY93]. coverings [LC94a]. covers [AHH93, MS94, MS95, Saa95, Yen95]. cpos [Hec91]. CPU [Jia93]. CPU-optimal [Jia93]. CRCW [Kuč97, Shv91, WC90]. CREW [dB93]. criterion [SCP99, YT96, Cus96, CS96, O’C94]. critical [BC94, NPW98, Ray89, Sim90, SR92]. cross [Gup92a]. cross-product [Gup92a]. crossbar [Den96]. crossed [Kul97]. Cryptanalysis [Cus95, Ada92, YT95]. cryptographic [SZZ94]. cryptography [Hwa93]. cryptosystem [SM90]. CSP [KS97c]. CTL [Lar95]. cube [ASA97, Che95a, FHL97, JS95, Kul97, She91b, XL93]. cube-connected [FHL97]. cubes [CL95, HC99b]. curve [Jag97, KOT95]. curved [KAL93, Sri99]. curves [BCMS97, IT94, KB90, ZI98]. Cusick [YT96]. cut [AM99, DIK97, SK92b, KLP98]. cuts [HW96, Sto91b]. cutset [Ola90a]. Cutting [Joh98]. cycle [CGH95, DN93, Gon98, KW92, WCL98]. cycles [Alb92, BR94b, Bax94, BF96a, CC97, Cim90, FW92, FHL97, HHL99, KS90, Mic97, QMA91, Sub95a, Sub95b, WM96, dBDL96]. cycletrees [VB96a, VB96b]. Cyclic [Alf99, Lin92, MWL90, HC99b, Lin93b, Mae90, NCC98, Nuu94, Wan94, Wha92, WC93]. cyclic-cubes [HC99b]. cyclic-permutation [Wha92]. cyclotomic [DH98].

D [SH93, CP97, CS97a, CC94a, FMF95, KAL93, Lu93, MNT92, PT92, RS91, Sri99]. D-image [RS91]. DAG [SB96]. Data [SS98b, SW98, AR91, AR99, CD93b, hC98, CMS90, CD96, Her91b, HK97b, HWT99, KS98a, Kle94a, KC91b, Lee96, LCC99, Lut90, MZM99, Mäk97, MSK96, Mat90, Mot91, Nau94, Nep99, Slo95, Tys97, WS94, Wie97, WJ95, XSR98, Yan98, YL99, de 92]. data-parallel [HWT99, Yan98]. Database [Qia90, Che91b, GS98b, Hof94, TL93]. databases [AE90, BCS97, CY97, Das96, FM97b, GPV90, HK97b, LNP^{+}90, LH93, LW95b, MZM99, OV96]. DATALOG [WY92a, LRH^{+}95, MN93a]. datamodel [DT93b]. date [GK98a]. deadline [Mid93]. Deadlines [Fre83, DB95, KS97a, Ple90, SY93].

Deadlock [SB90, UH95, Hof94, LKCNW91, PDBF91, TC99, Val95a]. Deadlock-free [SB90, UH95, PDBF91]. deadlock-preserving [Val95a]. deadlocks [LK95]. debate [RS95b]. debit [SG92]. debugging [SO97]. decentralized [Hof94, Yua90]. Decidability [GOPR92, Sat93, WS96, Gro92, LRH^{+}95, Sin94a]. decide [Yen91]. Deciding [Jur98, MN92, Her96]. Decimal [DP97b]. decision [Blu92, BV98, BLW96, Bop94, GM94a, Her93, JR91, Sch92a, SHS98, Sim91, Vas97, Wan91]. decision-tree [Bop94]. decisions [KKK97]. decode [Rif98]. decoders [IC98, TM98]. Decoding [Bar95, CWL99, Chu97a, LCW94b, LCW94a, TM98]. decomposable [GJT98]. decomposed [Gar92]. Decomposing [QMA91]. Decomposition [CYY^{+}92, Fra99, MC92]. decompositions [CDF96, dv94]. Decreasing [Mac95]. decremental [ASS97]. dedicated [BBDD93, BDDS92, BDDS94]. deduction [NN91]. deductive [LNP^{+}90, LH93]. deepening [HK96a, SCGD92a]. defaults [LS92]. defective [LC94b]. defectives [Dam98]. Deferred [AR91, CMS90]. definability [Dam92]. definite [Nep99, WY92a]. definitely [Sac96]. definition [DK94c, Yod91, ZD90, RT99]. definitions [Tys97]. deflection [Sym99]. degradation [JCT99]. Degree [TT92a, BOR96, MZM99, MS98, Mei90, SS94a, SS95a]. degrees [JS93]. Delaunay [AHR99, Dil90, Lam97, Sug94, SI95]. delay [BS94a, JT94]. delays [BBGT96]. deleting [Car91]. deletion [HM95, Mes91]. deletions [DP97a, HM96]. delivery [AB92, BP95a]. demand [ML95, PG95b]. demand-based [ML95]. Denotational [BMC96]. Dense [RG93, HSZF93, SSC98, Spi93, Ver97, dlVK99]. Density [Sas91, SV96]. dependence [AFM94]. dependencies [GW96, LL99a, MC92, Mal98, QB99]. dependency [KKK97, ME97, WLF94]. dependent [CD98b, Gaw96, HLW93, JM94a, Woe95]. deplanarized [FHT93]. Depth [Dam99, MAUY98, BB94b, Bop97, yC90, CDA91, CW97a, Chi90, CD94b, DN93, Gro98, HHK91, Joh98, Juk95, KIS90, MH96, Mil92b, Mnu91, RW91, RW93, RLDL94, Tho95, TT94, TV96, Weg93, FGN97]. depth- [Gro98, Juk95, RW93]. Depth-efficient [Dam99]. Depth-first [MAUY98, CDA91, CW97a, CD94b, DN93, KIS90, MH96, RLDL94]. Depth-First-Search [FGN97]. deque [CD93b]. deques [CD93b]. Derivation [Bij91, BF96b, Cha95b, Fv97, Hoo90, KS91, SI99]. derived [Mas91]. Deriving [Bol93, VCW95]. DESCEND [VHV93]. descendants [SB96]. describable [Sho92]. description [Kim96a, LCH93, SO97]. descriptive [CS93b]. descriptor [YL91]. Design [KM96, BS96b, HWT99, Hd91, HHS99, KCK93, Mur90, Pra96]. Designing [GM94b, KRY94]. designs [Ber94, TIK99]. Detecting [CH92a, FW92, GM97, RRS99]. detection [Gia94, Hof94, Jac90b, MK94, Naw91, RS90]. detectors [RT99]. determinacy [Pet94a]. determinant [Sen97]. determination [IY98]. determined [LC99a]. Determining [ARS93, BC93, Lin94, Kei91, She91a]. deterministic [AD97, CY95, FG93, LP94a, LC94a, Liś93, Mor97, Nys96, Sho90, Slo90, Tuz90]. detour [NPW98]. detour-critical [NPW98]. deunions [GI91]. developing [Miz99]. development [Bri93b]. developments [Wha92]. devices [Hwa92b]. DEXPTIME [Sei94]. DEXPTIME-complete [Sei94]. DFA [SHS98]. DFS [Jia91a, Jia91b]. DFS-traversing [Jia91a]. diagnosis [KCLA99]. diagonal [AS90, Zhu92]. diagram [SHS98, Tip90]. diagrams [BLW96, CHT90, HP90, HKK92, Küh98]. diameter [CC93, FHL97, HT95, JM94b, Lat93, QM90, RS93, THHC99]. diameters [Che92c]. diametral [DKS97]. Diamond [CD93b]. dictionaries [GM98, MSK96, MN90, Owo93]. dictionary [AF92]. difference [BF96a, Ski96, Var90, Wag98]. differences [GPS92]. different [AD97]. differential [Ada92, YT95]. difficult [SI95]. difficulty [GJK92]. Diffie [BBR99]. diffusion [XL93]. digital [Epp97, Hwa93]. digraph [HS97, KV94, Pas92b, VCS96]. digraphs [HP97b, HS95, KW92, Nuu94, TS95c]. dilation [PSSV93]. dimension [CKM97, DKK92, Lit95, NS96a, NS96b, Pan96a, Sim91, SR98]. Dimensional [WC90, AF92, BYR93, BCS97, CST94, CL93c, CL97b, CGR93, ESW96, GP90a, HI90, HL95, Jag97, Kan91, KP98b, MT98a, Mor92, NMS97b, Pet95, Rob90, VT93, Woe98]. dimensions [Ber93, Cha94, DK91, KB95]. dining [HP92]. Diophantine [Cus95]. direct [DS94b, RP99, TT94, WLF94]. Directed [RS94a, Ama93, BC93, CT93, Dan91a, FGN97, JM94b, Mei90, MS96a, NSS94, NU94, QM90, SC94, Vos93, vB97]. direction [BV98, FMS97, Flo97]. directional [DS98, Ham95]. directory [AP93, Pro94, THS97]. disabling [Mas91]. Disassembly [Maz93]. discrepancy [AM95]. discrete [CNV92, DNPS98, HI97, KKK97]. discrete-state [CNV92]. Disjoint [Kha94b, Mic97, Agg92, Alb92, BLHDW99, BR94b, CM94, Cou97, DS95, GP97, GP98b, HS91a, HS94, LKCNW91, PM94, QMA91, RG93]. disjunctions [Cal94]. disjunctive [Ger97, Käm90]. disk [CNV92, Che95c, MK96, Oya95, SR96a]. disks [MV91, VM97]. dispatching [Que98]. dissimilar [AS97]. Distance [SSCK94, Al 96, BCMS97, BCM91, BCM92, BC95, CC94a, Dan93, FMF95, Hir96, JS95, JT94, KK92b, Lê97, Leu92, MN93b, OS95, Pal90, Rot91, Sho91a, ZSS92]. distance-hereditary [MN93b]. distance-three [JS95]. distance-transitive [JT94]. Distances [GPS92, DS93a, MV91, MK96, Vis92, WT97]. distinction [DKL99]. distinctions [Hal93a]. distinctness [Bop94, LO94, Sze96]. Distributed [BC94, DF91, Erc91, Gia94, Lat91, MH96, MK94, Maz93, Maz97, RS90, Wan97b, AB92, Ahu90, AV94, AR96a, AF96, BH99, BCLM92, BF96b, CR98b, CG91, CB91, CT93, DK94b, Fid98, FJJR95, Gho91, Gre95, HW90, HP92, Joh99, KFYA94, KM91, KIS90, LK95, LC97b, Man94, PK90, RI91, RV93, Ray89, RW97, San98, SM92, SM95, SR92, Tri92, VT90, Ver87, WS94, WJ95, YW96, Nei94]. Distributing [MS90a]. distribution [BY91, Bsh97, Her91a, HI97, Hwa92a, Hwa92b, KKK97, Ker92, KP91, KLK97, LC96, LV92a, Sch95a, TH99]. distributions [Fla94, Lon94, MP98a]. distributive [Han91]. div [Mnu91]. divide [PAD99, WF96]. divide-and-conquer [WF96]. Dividing [CL96]. divisible [VA97b]. division [CS92, Gup92b, HP94, Mar90, PSL92, PH92, Tuz90, de 95]. divisor [PP90]. divisors [HP94]. DMC [YL99]. DNF [Bsh96, CGL98, Den99, Fla94, Kha94b, Kus97, PR93]. DNFs [HK92]. do [Zha91]. document [Chu99]. Does [Cha92, Hüh93]. domain [Hec91]. domains [FV90]. domatic [KS94, PC92, YY92]. dominance [GK98b, MT98a, SR96a]. dominances [Mat91a]. dominant [KP94]. dominating [COS95, DKS97, Irv91, KT97, Par91a]. Domination [DMK90, AR92, CL93a, GSSH93, KS97b, LRDL91, Lia95b, SMN^{+}95, YL90]. Dominations [Lia94a]. dominators [ACJ96, Gup95]. domino [TWK99]. domino-free [TWK99]. don’t [Aku95, MBY91]. double [AR99, yCGlW99a, yCGlW99b, Gua98, LM98, LCW94b, LCW94a, MG90, SLCH98, Eng92]. double-ended [AR99]. double-error-correcting [LCW94b, LCW94a]. double-loop [yCGlW99a, yCGlW99b, Gua98]. Doubly [Spi93]. down [FG93, KS91, SB95, Sch97]. dragging [Kim90c]. Drawing [BG92, LL96, BH98, CS97a, CP95, Kim97, SKKC98, Tre96, WCY99]. drawings [Bie98, DLR90, ESW96, TTV91]. driven [Lin93a]. dual [BG92, Rav94, Tan91]. dually [dFMdM99]. due [GK98a]. duo [DGS97]. duo-processor [DGS97]. Duration [Rab98, MP97]. durations [CHR91]. Dynamic [CMS90, CL97b, Epp97, FS97, Gup92a, Mes91, Myo91, Myo92, All92, BGP^{+}97, Blu95, BG94, BJ97, CJ90, yCGlW99a, yCGlW99b, CLM96, DS96a, DV98b, DV99, Fer94, FRS94, GP90a, GIZ96, HS91b, IL96, IL97, Jan91a, KS98a, KLR97, LR94, LCH93, LY98, Que98, RR94, RZ92, SS94b, Tam91, WJ95, XSR98]. dynamic-programming [All92]. dynamical [SDB97]. dynamically [LMS96]. dynamics [Tsa90]. dynamization [Cri91].

Eades [BR98]. eager [All92]. Ear [Fra99]. Earley [Lee92]. easier [BBR99]. East [Riv93]. Easy [But95]. eavesdropper [Al 96]. ECFSM [HHL95]. ECFSM-based [HHL95]. economics [Tho98]. ED0L [WS96]. Edge [NF92, SMN^{+}95, Alb92, BNMN92, BLHDW99, BR94b, BJ92, CR98a, CM94, Cou97, Dam95, DP97a, GSSH93, HR97, HJL^{+}91, IK93, JW95, KPC^{+}95, KK95b, Lia95a, LX99, LKCNW91, MSNR96, NPW98, NCK99, PM94, RG93, SD91, WHS97, LW93, VSS96]. edge-coloring [Alb92, JW95, Lia95a]. edge-disjoint [Alb92, BLHDW99, BR94b, CM94, Cou97, LKCNW91, RG93]. edge-length [LX99]. edge-set [DP97a]. edges [HR98a, LC93, MS92, Rhe98, VB96b, Zuc91]. edit [BC95]. editing [ZSS92]. Editor [Pet91, DF96a]. Effect [Sas91, Bar90, BLW96, Ust96, Ust98]. Effective [SCGD92a, ELS93, Hùy91]. effectiveness [LBL98]. efficiency [CT97, DN93]. Efficient [AF95, ALLL98, BCBF95, CW92b, CCH93a, CY97, CJ90, Chu97a, CR91, Dun90, EKR96, FLMSS98, FRS94, GSSH93, Gue90, GJS99, Hwa92b, IK93, KS98a, KLK97, Koc89, Lee90, LY99, NN91, OM97, Owo93, RPG98, RLDL94, SC90, ST90, TM98, Ueh95, Weg90, AJG95, AR96a, BA95, BW90, BU98, Che92b, CST94, Che95d, CWL99, CD96, DKK92, Dam99, DS99, DJS96, FMT93, GJT98, Gol91, GP98b, GLS92, Han91, HK90, He91, HP96, HTC92, HMPS96, IC98, JCH94, KA97a, KC91a, KC92, LK92, LP91b, MSK96, MI94, NDM99, Nuu94, PCK95, Pav92, Shv92, SK92a, ST93, Ten90, TT92b, TC99, Tho95, TW95, Tsa93, Von92, WCT99, YDL94, ZI98]. efficiently [Che95b, DMS94, Gav93b, LSW98, NL94, SvK99]. egds [LMS94]. El-Yaniv-Fiat-Karp-Turpin [DS98]. Elastic [Hac93]. elastically [ND95]. election [BT90, GG96, KM91, Sin92b, Sin94b]. Element [CH95, AD97, Bop94, LO94, Sze96]. elementary [Eng92, Rab98, SI99]. elements [JP97, Pro95, WS94, Wie97]. Eliminating [Cal94, KS90]. elimination [Asp95, DN93, KT96, KK95b, MNU90, Pan96b]. ellipse [San96]. ellipses [GS98a]. elliptic [KOT95, ZI98]. embeddable [ND95]. embedded [CH92a, OV96]. Embedding [CC97, LR91, Sch92b, Tse96, CC95a, Gon98, HLV97, HHL99, MNT92, NB96, ONI96, RG93, Sen98, SS92b, SLCH98, TT92a, VCS96]. Embeddings [Gre98]. emptiness [FV90]. emptive [DB95]. empty [Che96, CH99]. emulations [MS98]. enabling [Mas91]. Enclosing [SK98b, GS98a, Mat95b, Sky91]. enclosure [BKMT97, LMT99]. encoding [Weg96]. encryption [CLW98, Gol99, HK90, HP94, HCY96, JT96, LC97a, SW99]. ended [AR99]. Endmarkers [Var90]. enough [LSW98]. Ensuring [MC92]. entailment [Che90]. entries [BC94, Ray89, SR92]. entropies [Hùy91]. entropy [CRC98, Gai97]. Enumerating [AW93, SC94, DS93a]. Enumeration [BR94a, Maz97, NDM99, Yan91b]. enumerative [yCH91, SS92a]. envelope [Kei91]. envelopes [DG95]. environment [GIZ96, Jia91a, Jia91b, MS90a, MI94, Pom90, Tho95]. environments [EV97, MN98, OCSF97, SV96]. equal [WY92b, AE98]. equal-subset-sum [WY92b]. equality [Bas94, NN91, ZÖ94]. equalization [PCK95]. equally [Ito91]. equals [Kin92, LV92a, Ven98]. Equational [GS95, Fok94]. equations [CDÉ99, Cus95, ETA92, Fok97, Mar98, Ott95a, RAGR94]. equity [Gre95]. equivalence [Bas96, BCBF95, Bij93b, BK95, Bsh96, CM94, CDK93a, DF96a, DF96b, Ist97, LMS94, LRH^{+}95, PR93, Rac93, Tze96]. equivalences [BIN92]. equivalent [Bas94, Bus92, DS93b, DS94b, LvBDW92, Zim96]. EREW [CHL97, KPRS97a, RR99a, Weg90]. ergodic [PK93]. Ergodicity [Sat97]. Erratum [Ahu96, BP95b, CK96a, yCGlW99b, HAM11, NS96a, NAG96, NM96, PS95a, Sub95a]. erroneous [Bag92]. error [CJKJ98, GS96b, JCH94, LS95a, LCW94b, LCW94a, YL97a, YCH93]. error-correcting [YL97a]. errors [Ser98, VR99]. escape [OH91]. estimate [CD93a, LC94b]. Estimating [Mou99, Al 96, CEG95, Chu91, RS99]. estimation [Mah91]. ET0L [Fer95, Ist97]. Ethernet [KK92a]. Euclidean [BU95a, BU95b, CC94a, FMF95, TTC94]. Euclidian [KK92b]. Euler [CT93].

European [Riv93]. evader [KP96].

Evaluating [Chu91, CS90b]. evaluation [Erc91, GS90, Gup92a, KC96, PSL92, Ron91, SL96, TW94, Wan90b, YL91, Yua90]. Evaluations [TWK99]. evaluator [TW95]. evaluators [Hal93a]. evasion [Daw92, Neu96]. even [Kor97, SL95]. event [MCR98]. every [Sch95a]. Everywhere [Dij96]. evidence [Oh93]. evolutionary [KB94, WT97]. Exact [DG97, GS98a, GLR99, NS96a, NS96b, RU93, BP93, Bsh96, KAL93, Luc98b, MRR96, VM97]. exactly [IZ94, Slo90]. example [KS97c]. examples [Hir97]. exceptions [Lei95]. Excess [Sha96]. exchange [LL95, Lit95, PT92]. exclusion [Bax93, BW90, DK94b, Fuj98, Gre95, KFYA94, MT93, NM94, NM96, SM92, Yen96]. executed [Gaw96]. execution [CCH93a, HLW93, KC91a, KC92, KMP90, LKC94, Ros93, Woe95]. executions [Gum93]. exercise [GG96, Van92a]. Existence [Bün90, JS93, MC92]. existential [LH93, Loe97]. EXP [Zim99]. expander [HW93b, Rub90]. expanders [KS93a, Kuz94]. expansion [Gue90]. expectations [GM94b]. Expected [GS91, Sin92b, DP93, DPP96, GH94, O’C95, VM97]. experiment [Dij95b]. explanatory [Col94]. Exploiting [AE90, QB99]. exponent [AH94, KOT95]. Exponential [IP95, BW98, BDF^{+}95, Gál97, LC94b, WWT90]. exponentially [Sen97]. exponentiation [DJM98, HM98, Von92]. express [Hal93a]. expressible [PW97]. expressions [Bai92, Bol93, Gue90, Jed90, Lar98, Mor97]. expressive [Lar95]. EXPSPACE [AKM95]. extendability [LL98]. extendable [DKST98]. Extended [CL99, Gál94, LK97, SAFS95]. Extendible [Chu92, Hac93]. Extending [BG94, Lys94, Wal98]. Extension [KTS96, BB93, GMPW98, Mog94, Zwi91]. Extensions [MW91, GK97, Hùy91, Măn98]. external [BL91, San98]. extremal [Pri97, YJ93]. extreme [Wan91, Woe92b].

fabrication [Bri93b, Riv93]. face [DS94a]. facilities [EGL96]. facility [ACS99]. factor [AR98, Chu91, CR90]. factored [Gue90]. factoring [BBR99, Sho90, Sho91b]. factorization [Mel96]. factorizations [JP97]. factorize [CR91]. factors [Car94, Yen95, dM94]. fail [FP91b, Shv92]. fail-stop [Shv92]. Failure [LvBDW92, RT99]. Failure-equivalent [LvBDW92]. failures [JCT99, SB90]. Fair [BGP^{+}97, LH95, Dij90, HSC93, PK90]. Fairness [Vog95, BK98, KW97, KvdA99, RV96]. Faithful [WHS97]. false [Fos93]. families [GS98b, KS94, KB92]. family [CR98a, GR98, Ito91]. fan [Dan96, HK97a, LL99b, Sin98]. fan-in [Dan96, Sin98]. fan-out [HK97a, LL99b]. Fast [BYR93, BYP96, CCG^{+}99, FY93, Gem92, GSSD97, HP94, KC90, Kar93, KK92b, Lia95a, Lin91a, Omo91, Que98, Ros98, Sei96, TB97, XSR98, AF96, vB99, Che92b, CC94a, Che95d, CC99, CWL99, Chu97b, CK90, Der95, DJS96, ELS93, GP90b, LO90, Low98a, MP91, Mak90a, MT93, NBY99, NR99, Pug90, PN94, SY96, SGSD99, Ziv91]. Faster [Gus94, ABF96, BB93, CC94b, FF90, Flo91, Her93, IIO97, LLM97, Meh88, Zel93]. Fastest [SK91]. fat [Ove92]. Fault [CC92, GVV93, JM94c, RS93, She91b, SLCH98, AO92, Cai99, Chi98, yCGlW99a, yCGlW99b, GG96, GP95b, HHS99, KCLA99, KB94, Lat93, MS92, PP94, Pru94, SV94, Shv91, VCS96, WLK92, Wan97b, WHS97, ZD90]. fault-containment [GG96]. fault-diameter [Lat93]. fault-tolerance [Shv91]. Fault-tolerant [CC92, She91b, AO92, Chi98, yCGlW99a, yCGlW99b, KB94, MS92, SV94, VCS96, WLK92, Wan97b]. faults [BOR96, DV98b, DV99, FKP93]. faulty [Das98, OM97, Pel91, SGB92, Sen98, Tse96]. FDMA [FS98]. FDs [LMS94]. feasibility [AZ90]. feasible [Low98b]. feasibly [CP90]. feature [Kun93]. February [BU95a]. feedback [ELS93, Fuj96, Lia94b, LT97, Luc98b]. few [Vel94]. fewer [KP99]. FewP [RRW00]. FFT [HS95]. Fiat [DS98]. Fibonacci [Dro95, PRS96, Tre96, de 95]. field [AH94, Lan91]. fields [Sho90, Sho91b, Von92]. FIFO [BP95a, Lin94]. file [DML98, GIZ96, KCK93, Mat90]. files [Gol99, KC96, MZ93]. fill [Asp95]. filter [GSSD97]. filtering [CCH93b]. filters [GVFS99]. final [Lys94]. find [BR94c, TT90, Pro95]. Finding [AL95, Als96, BJ92, CR98b, CL92a, Che94a, Cim90, DL92, Den99, FKP93, HJL^{+}91, Iba97, IT94, Kim90a, LR93, LR96, LL99b, MS90b, NPW98, NU94, SH94, Sen90, Spi91, Woe92b, Yao92, YJ93, AS97, Alb94, ACJ96, BTY97, CDF96, Che92b, Che94b, CW97a, CGH95, DHO92, DG95, He91, HP96, HWJ96, HAM11, HTC92, IK93, Kim90b, KPC^{+}95, Kim96b, MMS90, Mar94, NSS94, Par91a, Pri97, RLDL94, SS94a, SS95a, SAFS95, SY96, SK92b, TY94, Woe92a, WT97, Ziv91]. fine [Yen95]. finger [DR94]. Finitary [Mow92, AI97, BL90]. Finite [SDB97, BF96a, Blu96, BIN92, CI95, Fei92, Fok97, FGW98, GS96a, Gre98, GMPW98, Hir97, HLV92, IC98, Jed90, JR91, Jon97, Kuč99a, Kut90, Lam95, MK92, Ott95a, RS95a, Sat93, Sho90, Sho91b, Slo90, TM98, Tze96, Von92, WIIO99]. finite-difference [BF96a]. finite-state [IC98, TM98]. Finitely [BCW99, Leu92]. Finitistic [Lac96]. firm [DB95]. First [FGN97, OV96, CHL97, CDA91, CD92, CW97a, CD94b, DN93, DP97a, HC92, KIS90, MH96, Man90, MAUY98, NLW96, RLDL94, Sim95, Tho95]. First-order [OV96]. five [KB95]. Fixed [Cai96, DD90, Hoa90, Mat95a, Tur91a, AB98, BFR98, BNN91, Cha90, DB95, FW92, HP94, Kuč97, SV94, Tsa93, Woe92b]. Fixed-parameter [Cai96, BFR98]. Fixed-point [Mat95a]. Fixed-radius [DD90, Tur91a]. fixed-sized [Tsa93]. fixpoint [Nys96, Sza92]. fixpoints [Sei96]. flipping [JLD92]. floating [AM98, MRR96]. floating-point [MRR96]. flow [Alo90, AM98, CW98a, CM99b, DKD97, Gol91, Gol92, KRY94, LY90, SK98a, SW92, TV93, Tsa92, XSR98]. flowshop [CW99]. Flush [Ahu90, Ahu93, Ahu96]. flush-vector-time [Ahu93, Ahu96]. fly [FJJR95]. fold [Ger97]. forbidden [CMR98, CLP91]. forest [Rav94]. forests [AS94, She91a]. form [Lar98, War98, Eng92]. Formal [AG97, BF96b, BGM90, Don95, SI99]. formalism [LSS92]. formed [BR98, GP96]. former [Riv93]. forms [Mok97, O’C95, Pal93]. formula [BF97, Che92b, Joh98, YMC98]. formulae [BB94b, Cad92, Fla94]. formulas [Bün90, CL93b, Cre98, IZ94, Mos90, PR93, SAFS95, Vei97]. fossil [YW96]. Four [HL98, Slo95, iNRN97]. four-connected [iNRN97]. four-partitioning [iNRN97]. Four-state [HL98]. Fourier [Kha94b, Mah91]. fraction [LC94b]. fragment [Cha98b, NLW96]. fragments [Lac96]. frame [WF96]. framework [BYF96, BB94a, CLM96, Mur90, PCK95]. fraternal [UG92]. Frederickson [SY93]. free [AGM92, AL90b, CL91a, COS95, Esp92, Gál94, Hùy91, JH93b, Kha94a, Khu90, KK98, Klo96, KMW98, Kol94, LV92b, Lit91, LS92, Mai94b, Mäk92b, Mäk92a, Mos97, MP96, Mye95, Ola90b, Pit95, PDBF91, RK96, RS99, Ros98, SB90, Ste91, TV93, TW94, TWK99, UH95, Xue96, Yen91, Zha92, Gon95b]. freeness [CR91]. frequency [WWT90]. fringe [DPP96, Mah98]. fringe-balanced [Mah98]. full [CS91a, KP98a, Li90, Mal98, Wis92]. fully [BP94, BP95c, BP95b, IL96, LL94b, Nys96, Pet95]. function [BW98, Erc91, Gál97, Gup92b, Jon97, Kha94a, Kos99, LTT95, Loe97, Mou99]. functional [Bij91, Kor95, LNP^{+}90, LL99a, MC92, Pio90]. Functions [CHM92, Mul91, Ogi96, Amb99, Ber99, BCBF95, CYY^{+}92, Chi90, Cla92, Cus96, CS96, Dan96, Dun90, Fei92, GM94a, Gon95a, GHS93, Gup92b, HK92, Hoa90, IZ94, Jai99, Jon95, Juk95, Kob92, Koc89, Kum95, Lê97, LW94, LM93, MTV94, MS99b, Maz95, NI98, O’C94, O’C95, PLSK97, Pet95, RS97, RS94a, Rub92, SZZ94, SH94, SLCS98, SCP99, TNB96, TV96, Weg93, Yod91, YT96, Yua90, Zip97]. fundamental [IL96]. Further [Gro91a]. furthest [GM92a]. future [Ahu93, Ahu96]. fuzzy [ABK98, GAG91, KLL93, MZM99, YH94]. FX [KP91].

G [SY93]. Galois [Bac98, The95]. game [BB94a, Dan91b, DKVW95, GH92, LS95a]. Games [Bv95, Jur98, YS96]. gap [MR96b]. Garbage [Tho95]. gate [Gol97]. gates [Dan96, Gol95a, Gro98, Juk95, RW93]. GCD [Lan91, LS99a, SL97]. generability [Dam92]. general [BB94a, CDA91, CY96, GVFS99, Jan92, KAL93, Ott95a, Uhl91, WF96, Zha91]. generalised [MV95]. Generalization [DK94c, BW90, CJ97, Fag96, Kru95, Ola90a]. Generalized [AQS97, CS90a, EV97, Gup95, ZNN95, Ali96, BBR99, Che95a, CC95b, CL95, DT92, DCH96, FvW95, GJS97, KM96, KS97c, Lit95, MN93a, Obe95, Pau90, SSO94, TS95b, Wan92a]. generate [All94]. generated [BH99, BV92, SS92d, SS92c, Tuz90]. Generating [Alo90, AS92a, BS99, CHM92, KR92b, Mai94b, Mil92a, Ste94, SI99, AD97, Bai92, BS94a, CH97, GK97, Kor93, Lin90, Pan96b, van91a]. generation [ARS98, AS94, vB99, hC98, CK93b, Gol95b, Jon97, Kor94, KL98, KL99, Lee90, San90, SC93, Vaj98]. Generational [Lin93b]. generations [CD93a]. generator [DH98, Naw91]. generators [GSSD97, JLP96, Lit91, SGSD99]. generic [Fos93, LW94]. Genetic [Kun93, CD93a, Cha95a, Cha95b, SGD98, SL96]. genetic-programming [SGD98]. genus [CKK97]. geodetically [CH98]. geographic [CHT90]. Geometric [EYK95, CJ90, HHI97, LV92c, iT96]. geometry [DG97, HP90, van93]. Giancarlo [CL97c]. Girard [HM99]. given [YY92]. glance [Gum93]. glasses [PS98]. Global [Ahu96, SLCS98, SCP99, CW98b, GM97, Hof94, How98, Lin94, VCW95, WLF94, Ahu93]. globally [Slo90]. goal [Han92]. goals [DCD96, LC99a]. Golay [Rif98]. gons [RWZW91, RW92]. Good [RS94b, BTY97, BJ92, Hir97]. gossiping [BHMS90, FY93]. gp [LH92]. gp-resolution [LH92]. graceful [JCT99]. Graham [Liu90]. Grammar [LC94a, BH99, Chu99, SO97]. grammars [AL90b, DBT90, FP94, Gál94, Ist97, Jef91, LC91a, LC91b, LK97, Mak90b, Mäk92b, Mäk92a, Pav92, Pit95, RK96, RS99, SC93, Sid90, TW95, Zha92]. grammatical [Mak90b, Mäk92b, Mäk92a]. graph [ABMP91, AR96a, AHH93, Ban96, BS99, BR94b, Bie98, Bus92, Cai96, Cha91, CKK97, Che91b, CM94, CC95b, CC98, CP95, CSS92, COS95, Dan93, DCGP99, DGP94, ESW96, ERC94, FW92, Gav93b, Hal93c, HW96, HMP95, HWJ96, HAM11, HTC91, IY98, Jer94, JM94b, JM94c, JT94, KS94, Kim90a, Kin92, KK95b, Lat93, LR93, Lip96, Lit95, MM92, MN92, Mei90, Mon96, NB96, NSS94, OS95, PPK98, QMA91, RR90, Ray95, SDG92, SHC^{+}93, SLC93, Sub95a, Sub95b, SC94, VS96, WCY99, Wis92, YH94, YMC98, Yan91a, YLW91, YY92, Zuc91, Bas94]. graphical [Bol93]. Graphs [KKM99, Alb92, Ama93, AR90, AR92, BNR96, BG91, BM90a, BCDM94, BK97, BV98, BH98, BM98, BHWG99, BC93, CS97a, CH91, Cha94, CHT90, CW92b, CL93a, CC97, CH98, CNP98, Cha02, CHL97, Che95b, CW97a, Chl96, CJ97, Cim90, CKN^{+}95, Cou97, CDP96, CKU98, DMM98, DMK90, DS99, DT92, DT93a, DV98b, DKS97, DLR90, DTT92, DCH96, FBW91, Flo91, FPT97, FGN97, FHT93, GS98b, GM94a, Gia96, Gia94, GLY98, GSSH93, GP95b, GP97, GP98b, GZ97, HSY97, Han92, HR98a, HP97a, HP97b, HKW90, HC99a, HL96, HW93b, HTC92, HJL^{+}91, HT91, HLC92, HZS99, IWW93, IK97, IY98, JW95, Jia91a, Jia91b, JB99, Joh99, Kan94a, Kao95, Khu90, Kim96b, Kir92, KS90, Klo96, KKM98, KMW98, KS98b, LW93, LL98, LWS99, LRDL91, Lia94a, Lia94b, LB95, Lia95b]. graphs [LR96, LO90, LT97, LKCNW91, MVR96, Mai94a, MSW93, Meh88, Mei90, MS96a, Mic97, Mid93, Mos97, MS92, MN93b, iNRN97, iNM97, ND95, Nik95b, NR98, Ola90b, Ola91, PR91, PM94, Pan96b, Par91a, PC92, PS95a, PS95b, Raj94, RS96, RS94b, RLDL94, RL96, Rhe98, Ros99, RS93, RR99b, Rus95, SS94a, SS95a, SR95, Sch92b, SH93, SSO94, Sim95, SK92b, SMN^{+}95, SR98, STN90, SGB95, TTV91, TV93, TB97, TR95, iT96, THHC99, UG92, VS95a, VKM92, Vos93, WLK92, WM96, WCL95, WHS97, WHH98, Woe90, Xue96, YC94, YDL94, YP98, YY93b, YTC93, Zel93, dFMdM99, vB97]. Gray [KL99]. greatest [PP90]. greed [HR98c]. greedy [BNMN92, BNHK99, BTY97, Coh94, GHY93, KP92, MT92b, Ola91, Par91a, SCGD92b, SR96a, Sla97, TT90, TTC94]. Greibach [Eng92]. grid [AŽ91, AEE97, CS97a, CP95, KC91b, NB96, Neu96]. grids [CM98b, Daw92]. Gröbner [Acc94]. ground [KR95, PS98]. group [Dam98, DV98a, Hwa92c, WHL99]. groups [LYC99, Rit94, Rit95, Ste91]. guarantee [Hal93c, KLP98, Zis95]. guaranteeing [GP90b]. guards [LL94a, SD91]. guided [HR90].

Hajós [IP95]. Halfspace [Lon94]. Halin [PPK98]. halvers [Man99]. ham [Sto91b]. ham-sandwich [Sto91b]. Hamilton [WM96]. Hamiltonian [LW93, MS92, Als96, BR94b, Bax93, Cim90, FGW98, IT94, LB95, MMS90, MYL99, Mic97, MN93b, Ray95, VCS96, WHH98, WCL98]. Hamming [IK97, MS99b]. handle [BJ97]. handled [LSW98]. handshake [Hes98]. handshakes [PK90]. Hankel [Gem92]. Hanoi [WC92b, WC93]. hard [BTY97, BJ92, FP91a, Juk99, KS97a, MM99, OSK96, OCSF97, San90, SO93, Wan97c, Yao92, ZJ94, Zis95]. harder [GMSS99]. hardness [SHS98, dlVK99]. hardware [Zhe95]. Hash [CHM92, Mul91, Gon95a, PCK95, SH94]. hash-based [PCK95]. Hashing [AP93, Chu92, Hac93, LPP91, Pro94, TKI99]. Haskell [Sei94]. Hausdorff [BCMS97, Rot91, Sho91a]. having [Lam95]. hazard [Cha95b, Kha94a]. hazard-free [Kha94a]. Head [Kam92]. Head-strictness [Kam92]. headed [BCLM92, CNV92, MK96, SR96a]. heads [Che95c, MV91]. heap [AR99, Car91, CN94, DW94]. heap-ordered [CN94]. heaps [HMPS96, KS91, Kho93, LP93b, RZ91, Sch97]. heapsort [Pau90]. heaviest [WCT99]. heavy [CFGG97]. Height [BY91, Asp95, AS94, BDF^{+}95, THS97, TT92b]. Held [SW90]. Held-Karp [SW90]. Hellman [BBR99]. help [yC90, Cha92]. helping [Ogi95]. hereditary [Cai96, MN93b, OS95]. hetero [WWT90]. hetero-associative [WWT90]. heterogeneous [KLK97]. heuristic [Bai92, DCD95, ELS93, GM91, GHY93, Kal90, LW95a, Oh93, Ola91, Par91a, TT90]. Heuristics [Dij95a, LK96, Mur90, Vos93]. Hex [AQS97]. Hexagonal [TS95a, AO92]. Hidden [MNU90, BS98, Jac90b, KT96].

Hierarchical [Fra98, CY94, FBW91, KC91b, Pom92, San98]. hierarchies [eC91, FM97b, GGH94]. hierarchy [Boo94, BH99, Cai99, eCCM91, CW98a, DE92, GPV90, Jed90, Lar98, Wag98, Can96]. Higgins [SH93]. high [Ber93, KLL93, KC91b, Sen90, Zim96]. high-low [Zim96]. higher [Cha94]. highest [CM99b]. highest-level [CM99b]. Highly [GS92b, Yen96]. Highways [Ahu96, Ahu93]. Hilbert [Jag97]. hinge [HWJ96, HAM11]. hinges [Mil92a]. Hirschberg [All94]. histogram [PCK95]. Hoare [Pro95, Sun96]. hold [Hüh93]. holes [Spi91]. homogeneous [CEdK98]. homomorphism [Dom96]. homomorphisms [NO96]. homophonic [HYHW98]. Honest [WB93]. Honeycomb [MYL99]. Hopfield [WC97]. Horn [CH92a, CL93b, HK92, Héb95, NN91, Pre93, SAFS95]. Horn-completion [NN91]. Horton [DK94a, DK95, Kru99]. host [Bil93]. host- [Bil93]. Hough [KC90]. Huffman [TM98, Bur93, CWL99, Chu97a, HV96, IC98, Kir94]. hull [BGOS95, DvR94, DW96]. hulls [DG95]. human [BLP98]. human-computer [BLP98]. Hutchinson [BCM91, BCM92]. hyper [ONI96]. hyper-rings [ONI96]. Hypercube [Pan96a, AZ90, DCH96, Erc91, FP91b, Ham95, HCF91, KL90, KKM99, KG99, KB95, Lub90, Man91, NF92, Ple90, RG93, Sch93c, SB90, SR91, SHC^{+}93, SHL94, ST93, Tse96, WL93]. hypercubes [BLHDW99, CC93, CC95a, Chi98, DV98b, DDK^{+}98, DV99, Fuj95, HHL99, KCLA99, Kor95, Lat91, LR91, LL94b, ONI96, QM90, RL94, SV94, Sen98, Wan94]. hyperedge [Dre93]. hyperedge-replacement [Dre93]. hyperedges [Gon98]. hypergraph [DKK92, GKL96, Gra96a]. hypergraphs [IWW93, Rad92, Woe92a]. Hyperspherical [AR96b]. hypothesis [Fos93]. Hypothetical [BLM90].

I-serializability [EV97]. I/O [Jia93, RV96]. I/O- [Jia93]. ID [Hwa92a, Hwa92b, LHW99]. ID-based [Hwa92a, Hwa92b, LHW99]. identical [LYC99]. identification [DF91, JS93, Jai99, Lat91]. Identifying [FLSS93a, FLSS93b]. identities [McC98]. Identity [Ten90, WH94, LP90]. IEH [CC97]. IFS [ESW97]. Illumination [ECOUX95]. image [CY97, CC94a, CK93b, HV96, JT96, NB96, RS91, YLW91]. images [Epp97, FMF95, KK92b, RRS99]. immediate [ACJ96]. immortality [SY90]. immune [MS99b, PLSK97]. immunity [Ada92, Käm90]. impact [FMS97]. implement [RST91]. implementation [BA95, Blu96, BCLM92, Gol91, Gus94, IC98, LY92, Pav92, SGD98, SK92a, Tsa93, TM98]. Implementing [Gra96b, KW95, WCC91]. implication [ME97]. Implicit [YL91]. Impossibility [BM97]. impossible [OK96]. Improved [Ama93, AN93, CM98b, DV98a, Gon98, HS90, MG90, MW97, Ov98, Rad92, Sla97, TV93, ABK98, AD97, BFR98, BC95, CY95, DZ98, GHY93, MP97, Saa97, SC93, TT91, Wan92a, van92b]. improvement [DS98, Han95, Mar98, Saa95, SK98a]. Improvements [SL97]. Improving [CR98a, DN93, PLSK97, SZZ94, ME97]. In-place [HW99, KP99]. inadequacy [BJ91]. including [CYY^{+}92]. Inclusion [Bax93, CDK93a, LL99a, MZM99]. incomplete [HHL99, LW95b]. incompletely [Kha94a]. inconsistency [HK96c]. Incorporating [MM92]. increase [Asp95]. Increasing [Bla94, vB99, Hoa90]. increment [WS90]. Incremental [DG95, HM95, DS94a, FGN97, Tam91, Yan93a]. indeed [Bas96]. independence [BNR96, FvW95, Hal93b, von91a]. Independent [HL96, IKOI99, AZ90, BTY97, Boz95, CG91, CW92b, CL92a, CDP96, DKK92, GKL96, HR98c, HTC91, HTC92, Irv91, Jef91, KS92, Kim90a, MV91, Pas92a, Ple90, SS94a, SS95a, YTC93]. Index [Ano96a, Ano96b, Ano96c, Ano99a, Ano99b, Ano99c, Ano99d, Don95, Rhe98, dFMdM99]. Indexed [Chu92, TKI99]. indexes [LPK90]. indexing [BCS97]. indistinguishability [Gol90]. induced [TB97, TWK99]. induced-subgraph [TB97]. Inductive [Mot91, de 92, Lau90b, Nau94]. inductively [HLM98]. inequalities [DP95, War98, van97]. inequality [BG94]. inference [BP93, JS93, Mak90b, Mäk92b, Mäk92a, Mot91, PS92]. Inferring [Mäk97, EST99]. Infinite [Jed90, Pir94, SMDS94, Fok97, JP97, Kuč99a, Mel96, Nys96, OV96, Vas95]. infinite-state [Kuč99a]. infinitely [Ito91]. influence [Lit95]. Inform [AAA^{+}91, Ahu96, BP95c, BP95b, BMP94a, BDDS94, BU95a, BM90b, BCM92, DF96a, Dun90, Flo91, Gro91a, HS94, Har96, KC92, LCW94b, Mis92, NAG96, Pac92a, Pet91, Pro94, Rit95, SS95a, SH93, Sny91a, SY93, Sub95a, SS92d, YT96]. Information [Cha02, CK96a, yCGlW99b, HAM11, Khu00, NS96a, NM96, RRW00, SS98a, ABK98, BS96a, Bag92, BÖS91, CCK93, DBM97, FS97, KLL93, LW95b, LM93, Tsa90, XU99b, van97]. information-theoretic [CCK93]. inherently [DB91]. inheritance [CW98a, FM97b, Thi93]. inhibitors [DV98a]. initial [MW91]. initialization [FMT93]. initialized [Shv92]. initially [SR99]. injectivity [FG93]. inkdot [IIT92, IITY93]. input [Ger94, KT91, KS97c]. input-output [KS97c]. inputs [BFL^{+}95]. Inscribing [San96]. insertions [MSNR96]. insight [LP97b]. instance [Low98b]. Integer [Fre83, Kri93, CD92, Ger94, Gup92b, HS90, Hag95, Lit92, MMN91, SL97, SY93]. integers [GPS92]. integral [GS90]. integrated [RP91]. Integrating [LC96]. Integrity [ED91, WY92a]. Inter [TH99]. Inter-protocol [TH99]. interaction [Bir90, BLP98, LL99a]. interactive [Bur92, DCGP99, GH98]. interchanges [Chr96, KR92b]. interconnection [ASA97, Ber94, GVV93, JM94c, SLC93, VS95a, VCS96, Wan97b]. interesting [Bha99]. interface [Mar98]. interfaces [Das98]. interleaving [CTVC98, TH99]. Internal [CN94, BG91, LW94, LvBDW92]. Internet [HC95]. interpolation [AM95, GS91, PSL92, Vel96]. interpretation [BFSV99, CCH93a, CLM96, Fer95]. interpretations [Grä90]. interprocess [Bar90]. intersecting [Saa97, SSCK94]. Intersection [Bir92, GS98b, iT96, CJ90, MMN91, NF92, PM94, RI91, THI91]. intersections [BM90c, BM90b, GJS99, TT91]. interval [AT95, AR90, CNP98, CKN^{+}95, Dan93, DW94, HMP95, HTC92, KLR97, LT97, MVR96, Ola91, RR90, Ray95, SR95, SZ92b, SK92b, YY92, YY93b, KLR97]. intervals [BR94c, Che92a, MMS90]. intra [QB99]. intra-object [QB99]. intrees [BBGT96]. Introduction [Hd91, RS95a, VV97]. invariance [Pet91]. invariant [PW97, Pet91, VT90]. Invariants [Hes98, Göb99]. inverse [Mnu91, Wan92a]. Inversion [van91b, Gem92, von91b]. Inverting [Pre91]. investigation [SDB97]. involving [How91, Ste91]. IO [CT91]. IP [RS95b]. IPL [Ano94]. irreducible [Ito91]. irredundance [CNP98]. irredundant [O’C95]. irregular [Nas98]. irrelevant [GLR99]. Ising [PS98]. isomorphic [NAG95, NAG96]. Isomorphism [Cha90, CST92, DIR99, Dub90, Gro91a, Gro91b, Mäk89, VKM92, Bas94, GH96]. isooriented [MNU90]. isothetic [dF92]. issues [Tur91b]. item [VA97b]. iterable [Yen95]. iterated [HP97b, Lit92, Sho92]. iteration [Fok94, Nep99, WB93]. Iterative [HK96a, Saa95, Erc91, ITY96, Maz95, SCGD92a, Tip90].

J [SH93, HM99]. Jacob [Gra97]. January [BM90b, Pac92a, Sny91a]. Java [Miz99]. jobs [AZ90, GP95a, Gaw96, GK98a, LYC99, Ple90]. jog [LCC93]. jogs [TT90]. join [CST94, FP91a, HW90, LPK90, Mal98, PCK95, TL93, Wei91]. joining [LCC93, Lim96]. Jordan [FNTV90, IT94]. July [NS96a]. jump [Wan97a]. June [BMP94a, NM96]. justification [AG97].

K-M-P [RUG97]. Kaikoura [SW93b]. Kalé [Oh93]. Karger [Cha98a, DIK97]. Karger-Klein-Tarjan [Cha98a]. Karp [DS98, GBY90, SW90]. Karp-Rabin [GBY90]. Kautz [HP97a, HS95]. kernel [LC99b, PDBF91]. kernel-search [LC99b]. kernels [SC94, vB97]. key [Al 96, Cus95, HI97, Hwa92a, Hwa92b, Hwa93, LC96, LL95, Low95, SM90, Tze99, TH99, YL97b, Zhe95]. keyed [Gon95a]. keys [CH97, DT93b, Mok97]. Khrapchenko [Zwi91]. kinded [Vas95]. Kleene [BP98]. Klein [Cha98a]. Knapsack [Jen95, GL98, MT92b, SCGD92b, VA97b]. know [BTY97]. knowledge [Bur92, DCGP99, DP94, DF91, GAG91]. known [BKMT97, DPP96, Hes91, KCK93]. Ko [CS97b]. Kochol [Dun90]. Kolmogorov [Kir92, MP98b, Tys97, Zim96, Zim97]. Komolgorov [ESW96]. Kripke [Sin94a]. Kronecker [BM98].

L [MR95]. label [KW98]. labeled

[ZSS92, ZJ94]. Labeling

[SvK99, CL97a, PZC98, Wag94]. labelled [Sin94a]. laden [Gia96]. Lag [SY90]. LALR [Gál94]. lambda [Mül92]. Lamport [MT93]. language [BL96, BS97, CDK93a, Gol95b, JS93, KB92, Mai94b, NO96, Sho92]. Languages [Vei97, Bér95, BMP94a, BMP94b, Bir92, BH99, BV92, CRC98, CL91a, CLM96, CDÉ99, DPS93, DB91, Don95, Dre93, EST99, FP94, HLV92, Hùy91, Kim96a, KP98a, Lit91, Mak90b, Mäk97, MR95, MY96, MP96, Mye95, Nys96, Ogi95, PG95b, Pin95, Que98, Ros98, RS91, Sal96, SL95, Slo90, SI99, TW94, Wan98, Yen95]. Large [Zim97, Bai92, BH98, CD96, Kol94, LY98, Lip96, MSK96, Owo93, Spi91, Tho98]. large-memory [Tho98]. largest [AS92b, Gas91]. late [GK98a]. Latency [SW98]. Latency-of-Data-Access [SW98]. lateness [BBDD93]. later [Hen92]. Latin [JS95, KC98]. lattice [GMSS99, Hen97, Pal90, SS92d, SS92c]. lattice-theoretic [Pal90]. lattices [NR99]. law [BGKL98, Zim96]. laws [Lac96]. layer [GM92c, TS95a]. layers [Nie96]. layout [CM99a, GIZ96, MAUY98, SDG92, TT90, Yan98]. layouts [YP98]. Lazy [All92, BA95, Hil95, Lin92, PG95b, Wan90b]. LC [TW95]. LCA [Wen94]. LCF [VS95b]. leader [GG96, Sin92b, Sin94b]. leads [Cal94, Pac92a, Pac92b]. leads-to [Cal94, Pac92a, Pac92b]. leaf [BL96, BS97]. learn [Kha94b]. learnability [Fla94, KB92]. Learning [CGL98, Ker92, MP98a, AD96, BCBF95, Bsh96, Bsh97, BTW98, GR97, GLR99, Jai95, Kum95, Kus97, Lon94, Sak91a, Slo95]. learning-theoretic [Kum95]. least [BU98, Cri91, Ols97, WC90]. Leaves [GMM94, Mat90]. left [Mül92]. left-linear [Mül92]. leftist [Kem99]. Leftmost [Gur93, Gur94]. lemma [Oht93, Ola90a]. length [AH94, Alt90, vB99, BF96a, BC95, Bur93, CW92a, CN94, CM98b, DP93, DPP96, Fra99, HC99a, LX99, LYC99, MZ93, NCK99, RU93, WCT99]. length-constrained [WCT99]. lengths [YW96]. less [Loe97, PMC97, YJ93]. lETOL [Gär93]. Lett [AAA^{+}91, Ahu96, BP95c, BP95b, BMP94a, BDDS94, BU95a, BM90b, BCM92, DF96a, Dun90, Flo91, HS94, Har96, KC92, Mis92, NAG96, Pac92a, Pet91, Pro94, Rit95, SS95a, SH93, Sny91a, SY93, Sub95a, SS92d, YT96, LCW94b]. Lett. [Gro91a]. Letter [DF96a, Pet91, AKK97]. Letters [Cha02, CK96a, yCGlW99b, HAM11, Khu00, NS96a, NM96, RRW00]. level [ACS99, CM99b, TKI99, Ust96]. levels [Lar98]. Lexical [DBT90, Naw91, Spi93, Wan90a]. lexicographic [BS94a, Sim95]. lie [CLP91]. life [LC94b]. light [Sub95a, Sub95b]. lights [ECOUX95]. like [Bil93, BFL^{+}95, IP95, Joh98, MSV96, Ogi95, SC90]. limitation [Fid98]. Limitations [AF96, Höf93]. limited [Fer95, GM91, KPC^{+}95, Ogi96, TNB96, WS96]. line [ASS97, BMI97, BNMN92, BKR94, BGP^{+}97, Bre92, CS97a, Cha92, CWK93, CvW94, CW97b, CKU98, DS96a, DvR94, DW96, ES98, GJT98, GIZ96, HP97b, HS97, KT97, KT96, MNU90, MMN91, NLW96, Pru94, RR94, Ray95, RW96, Rot91, Sga97, TT93, TTC94, VCS96, WY95, Wis92, Zha97, van92b]. Linear [AR90, ETA92, HT91, Hüh93, IY98, Juk99, Kao95, RW91, VA97a, VMD99, Wan97a, Wei94, WY95, AŽ91, BMC96, CM98a, CDF96, CY98, CL93b, CW97a, CP95, CSS92, Cla92, COS95, CKN^{+}95, Cou97, DL92, Den90, DDK^{+}98, Dro96, ERC94, Esp92, FNTV90, GP90a, Gol95b, HMP95, Hir96, HWJ96, HS91b, HAM11, HM96, HM95, IK97, Kim97, KST96, KMP90, LWS99, LLM97, Lee96, Lin90, Lin96, Lon94, LT97, MSW93, Mak90b, MT98a, Man90, Man96, Meg92, Mog94, MK92, Mül92, Myo91, Myo92, MW97, iNRN97, Pan96b, PC92, PWL94, PPK98, PR93, Pre93, RL94, RW94, RR99b, RAGR94, Rub92, Sat93, Sat97, Sat98, SL95, Ser91, She92, Sho91a, SW93a, STN90, Tze99, VKM92, War98, Yan93b, YT95, Zhu92]. Linear-size [RW91]. Linear-time [Kao95, Wei94, CDF96, CL93b, CP95, FNTV90, GP90a, HMP95, Hir96, LWS99, LT97, MSW93, iNRN97, VKM92, War98]. linearly [Asp95, MR92]. lines [HS91a, HS94, Kei91, SS95b]. link [AL95, CM98b, EGL96, KLK97]. linkage [HCY96, LC97a]. linked [Sax94]. links [Sen98, Tse96]. LISP [Van92a]. List [WC91, AvSW95, AM97, AM90, BdW99, HM96, HW99, Ira91, Kim96b, LP94b, Ped94, RW96, Res92, SM94, Tei93]. list-ranking [AM90]. list-update [Ped94]. listing [Epp94]. lists [Blu92, HS91b, Sax94, Sen91]. Literal [MY96]. Liveness [KvdA99, AAA^{+}91, CL91a, DW90, Gum93]. LL [Yan93a]. load [CFGG97, Coh94, LS90, MP97, XL93]. Local [AP95, Küh98, Ali96, BLW96, Fis95, FS92, GM98, HCF91, MWL90, MW91]. Locality [Thi93, CL92c]. locally [KKM99]. location [ACS99, EGL96, Ove92, Tam91]. lock [JH93b]. locks [HSC93]. Löf [Kob93]. log [GKL96, Ser91]. log-space [Ser91]. logarithm [HI97]. Logarithmic [Kim97, Ram95, Ron91, Sax94]. Logic [Mei90, Yua94, BP98, Bij91, Bir96, BLM90, BCLM92, Cal94, CH92a, CCH93a, CL93b, Che90, CLM96, Dij96, Ger97, GS95, Kob92, Lac96, LKC94, Lee96, Loe97, NN91, Pac92a, Pac92b, RF93, SR99, Sun96]. logical [AE90, AQS97, BP93, CB91, Tys97, VM95, Vel96]. Logics [Sin94a, AP99, Dam92, Das96, Dem95, Jon95, LH92, Mow92, Pen92, Sza92]. LOGSNP [Den99]. LOGSNP-complete [Den99]. logspace [Ste91]. long [DP97b, ES98]. longest [AS90, DS93a, LLM97, NLW96, WCL98]. lookahead [SM94]. lookup [MSK96]. loop [CR98b, yCGlW99a, yCGlW99b, Gua98, PDBF91, SLCH98]. looping [LP94a]. Loopless [Kor94, KL99, KL98, Vaj98, van91a]. loops [LM98, Nik95a]. loosely [MS90a]. lossless [HV96]. LOTOS [BIN92]. Low [KOT95, SV96, Zim96, Zim97, dv94]. low-density [SV96]. Lower [BD97, yC90, CEG95, Hec91, Joh98, Man94, Mei91, Sny91a, Sny91b, TTV91, Tom95, Vol96, AL90a, AGM92, Alt90, BKR94, BMP94a, BMP94b, BCDM94, BR94c, BCDM97, Bsh92, BdW99, CvW94, CL93c, CLLR97, CM98b, CP90, Den96, DG95, GS96a, GH92, Gro98, HI90, IP95, Kle94b, KO96, LK96, LO94, MP97, O’C95, PS98, PP90, Rab98, Raj94, RV93, RW93, Sal92a, Sch97, Sga97, Tei93, TT94, TU99, Weg93, van92b]. lowest [Lin91a, NU94]. LP [Gon95b]. LP-free [Gon95b]. LPO [Nie93]. LR [AFG92, JCH94, LC99a, Pav92]. LR-attributed [Pav92]. LR-based [JCH94]. LRU [BA95, Jia91a]. Lucas [LTT95]. lumped [Oya95]. Lyndon [SM90, SMDS94].

M [RUG97]. M. [DW91]. machine [BCDM97, CW99, GK98a, IC98, LO94, MP98b, Pio90, TM98]. machines [Ger94, GGH94, Hüh93, IIT92, IITY93, KS97a, LYC99, MP95, Pet95, Rob90, Sze96, Woj99, Yod91]. magic [MN93a]. magic-sets [MN93a]. mail [BHMS90]. Main [AP93, Pro94]. Maintaining [CD98a, DP97a, MSNR96]. maintenance [FGN97, FS97, IL96, Jan91a]. majority [ARS93, FS97, Gol95a, OSK96]. make [Var90]. makespan [Tsa92]. Mäkinen [Gro91a]. Making [Dij90, Vys90, DS98]. male [KT90]. Mallat [Lu93]. malleable [ST99]. management [AF95, CL97b, CM96, FRS94, KB94]. Manhattan [GJK92, Kan94a]. manipulation [CY92, GM94a, RR93]. Many [DS94a, AD97, BF97, FPT95, Gav93b, GK98b, Ito91, Zim97]. many-element [AD97]. Many-face [DS94a]. Map [Wag94, Hef91, PZC98, SvK99]. Mapping [HNC92, Kor95, KL90]. maps [BH98, CC94a, FMF95, Hir96, KK92b]. March [CK96a, SS95a]. mark [Lin92, MWL90]. mark-scan [Lin92, MWL90]. marked [Raj94, TR95]. Markov [Nas95, PK93]. Martin [Kob93]. Martin-Löf [Kob93]. Master [Ano96b, Ano96c, HWT99]. master-slave [HWT99]. Mat [DP99]. match [KP91, KCK93]. matched [BNHK99]. Matching [ABF96, Aku95, ABMP91, AF92, AFM94, ALLL98, ABL99, BYR93, BYP96, BBL98, Bra90, BCT93, CM98a, Che95d, Chu97b, CGR93, CCG^{+}99, DS95, DG97, Der95, DGL98, FPT97, GBY90, Han95, He91, HCF91, HH92, Kan91, Kan94b, KT91, Kel94, KT90, LR93, LB95, LP91b, MBY91, Mye95, NBY99, Per94, RS96, RUG97, Tel94, TTC94, TCC91, Wei91, YY93a, dB93]. matchings [BG91, FGHP93, GJLR96, Iba97, Ros99]. materialized [CL99]. matrices [ADY91, BT97, Gem92, PKP93, SS98a, SR96b, Spi93]. matrix [Bsh95, BD98, CS93a, GMR90, KS93b, Lar90, Lin91b, Mac95, Man94, MG90, Pre91, SSS98, SS92d, SS92c]. matrix-vector [GMR90]. matroid [ABF94]. matroids [FBW91]. max [CM99b, TV93, CFS91, Kan91, ZJ94, AR99, Kan94b, KLP98]. max-flow [CM99b, TV93]. Maximal [Tel94, Agg92, Che91a, Che95d, CGH95, DKK92, GKL96, GJLR96, Hal93b, Han95, HH92, JS93, Jan91a, Kel94, Nie96, SS94a, SS95a, XSR98]. Maximally [Chl96]. maximization [CS90b, CS93b]. Maximizing [HH94]. Maximum [AM98, DMM98, Kan91, Kan94b, Mar94, WCY99, Ali96, Alo90, ABMP91, BG91, Bur93, CW92b, CL92a, Che92a, CW98b, CM94, CJ97, CJKJ98, CDP96, DGL98, FF90, FH94, FPT97, Gol91, HR94, HR97, HR98b, HR98c, HP96, HTC91, HTC92, KLP96, KMN99, Kim90a, Kim90b, LR93, LB95, Mos97, Pas92a, RS96, SY96, SW93b, SW92, TB97, TY94, YTC93, YY93a, GMM94]. maximum-weight [YTC93]. MAXSAT [RRS98]. Mazurkiewicz [GMPW98]. McCulloch [SL95]. McCulloch-Pitts [SL95]. McKay [BR98]. MCS [JH93b]. mean [DKD97, KK92a, LY90, LC94b]. means [PSL92]. measurable [Zim99]. measures [CCK93, Pal90, SSCK94, Sri99]. mechanism [Ros93]. mechanisms [MABD90]. media [SCH95b]. median [CD94a, LV92c, Ols97, Sen90]. Medians [MS96b, DNPS98]. medium [OCSF97]. Meeting [BHWG99]. Mehlhorn [Flo91]. membership [MW94]. memories [AF96, IS98, Rob90]. Memory [AP93, NDM99, BB93, CWL99, FMT93, HCF91, MP91, OM97, Pro94, Rou97, San98, SCGD92a, Shv92, SJKL95, Tho98, WWT90]. memory-Constrained [Rou97]. Memory-efficient [NDM99, CWL99]. merge [GS91]. mergeable [Bri93a]. mergesort [GS93]. merging [DJM94, GS91, Lin96, YL99]. Mesh [CL92c, CS91b, FY93, GZ91, HL91a, HI90, KC90, KK94, Kri93, PKP93, PSSV93, She93, SRW95]. mesh-connected [CS91b, GZ91, HI90, Kri93]. meshes [BGOS95, CCC96, CL93c, Gur93, Gur94, HLV97, Luc98b, Sch92b, Sym99, VT93]. Message [CY94, CS95, AB92, AV94, BP95a, yCGlW99a, yCGlW99b, FMS97, Gua98, HCY96, KHJ98, LC97a, LY99, Tze99]. message-passing [KHJ98, LY99]. messages [DDK^{+}98]. method [Bai92, vB99, Blu96, HM98, JT96, KP91, Lys94, Nep99, RAGR94, SO97, ST91, Tam91, XL93, YW96]. methodological [AAA^{+}91, DW90]. methods [Bij91, CY97, Nie93, Per94, VV97]. Metric [ND95, Kwi90, LBL98, Pal90, Pra96, Uhl91, Vil95]. metric-space [Vil95]. metrics [Bag97, BS96b, CCK93, GP96]. Metropolis [Sas91]. middle [MP95]. migrating [BP95a]. Miller [CST92]. Min [AR99, DP99, AM99, TV93]. min-cut [AM99]. min-multicut [TV93]. mincut [IWW93]. minima [GJT98, Kim94a]. Minimal [BV98, CH91, CHM92, Flo97, LL94b, SH93, vB99, Bas91, BBDD93, BM92, CR90, Den99, Jai95, Jai99, LS92, Ram98, SH94, She91a, TKI99]. minimax [BB94a]. minimise [BD96]. minimization [FL97, Kha94a, RP91, TT92b, Tsa92, ZÖ94]. minimize [DKD97, GK98a, LY90, ST99]. Minimizing [Asp95, Blu96, CJKJ98, NI98, Woe91]. Minimum [GM92c, Lim96, Mid93, AL95, Cha98a, CG91, CL97a, CW98b, DIK97, DGP94, Gai97, GM90, Gon95b, Hal93b, HT95, He91, HJL^{+}91, Irv91, IK93, KPRS97a, KW98, LL96, LK96, LL94a, LS95b, LCC93, LC93, LX99, Low98b, Luc98b, MR96a, MR92, MS92, NCK99, PR91, Pre91, Rot91, SHS98, SW92, SB93, TT90, Tre96, TTC94, Yan95, Yan98, YY93b]. minimum-area [Tre96]. minimum-cost [SW92]. minimum-link [AL95]. minimum-width [Yan98]. Minkowski [FP98]. minor [Khu90]. minterms [KA97b]. MIP [TS96]. mirror [KKM99]. MIS [YDL94]. mismatches [Der95, Kar93]. ML [VS95b]. mobile [Hwa93]. Mod [Bar91, Gro98, MV95]. Modal [Das96, Yod91, Dem95, Dij96, Kai95]. mode [LLP90]. Model [KC92, BNN91, BCDM97, BCLM92, Cad92, Cam98, CNV92, Cha95a, CW98a, CFGG97, HWT99, Jac90a, Jac92, KC91a, KHJ98, LCC99, MSV96, MN98, Nas98, Nas95, Pio90, SW98, VI98, WC90, YY92]. Modeling [IWW93, Gei97, LCH93]. Modelling [GW96, MP91, Tsa90, Lee96]. Models [RF93, DML98, LS92, Loz92, Sac96, SDB97, SRW95, ZD90]. modification [Cai96]. modified [CLW98, GHY93, Mah91]. modifying [RS95a, WIIO99]. Modular [KK90, CC99, CK93a, CLM96, DJM98, FV90, Gol95a, Mnu91, Ohl95, Ram90]. Modularity [KMTd94, Vel96]. Modularization [VM95]. Module [Rou97]. Modulo [Ram95, BCBF95, BBR99, Bsh92, CC99, Ott95a]. modulus [Tze99]. monadic [FV90, Ott95b]. money [DS98]. money-making [DS98]. Monge [BD98, GJT98]. monoid [AKK97]. monoids [HLV92]. monomials [NS96a, NS96b]. Monotone [Ros97, CGL98, Clo96, Dun90, Fla94, GH92, Joh98, Koc89, RS94a, TT94, Zhu92, Zwi96]. monotonic [Kam92, SO93, OCSF97]. monotonicity [CG93, GT96, SW90]. Monte [HK96b]. Montgomery [HM98]. morphisms [Kui95, Mit98]. Morse [BM92, TS95b, dM94]. mortal [BT97]. Mosaics [Maz93]. Most [VSS96, Cha94, HJL^{+}91, IK93, LC93]. mostly [GH94]. Motion [Mah91]. moves [KP99, WC92b, WC93]. moving [CD98a, HKK92, SK91]. MR [BMP94a, DF96a, Flo91, Gro91a, HS94, Har96, NAG96, Pro94, SS95a, SH93, SY93, Sub95a, SS92d]. MRU [Jia91a]. MST [AR98]. mu [Kai95]. mu-calculus [Kai95]. Muller [KA97b]. Multi [AP93, Kuz94, vB99, Boz95, CCG^{+}99, HI90, HW98, IITY93, KRY94, KM96, Li90, Mar91, Mat90, NR98, PK90, Pro94, Rob90, TL93, TCL93, YMC98]. multi- [TCL93]. multi-bucket [Mat90]. multi-channel [Li90]. multi-commodity [KRY94]. multi-dimensional [HI90, Rob90]. Multi-directory [AP93, Pro94]. multi-inkdot [IITY93]. multi-join [TL93]. multi-pattern [CCG^{+}99]. multi-priority [Boz95]. multi-process [PK90]. Multi-processor [Kuz94]. multi-ring [KM96]. multi-senders [HW98]. multi-stack [vB99]. multi-star [NR98, YMC98]. multi-way [Mar91]. multibus [Nas95]. multicast [Fuj95, KLK97]. multicommodity [Gol92]. multicomputer [Pom92]. multiconnection [TCL93]. multicut [TV93]. multidatabase [CM96]. Multidimensional [CWK93, CS91b, Gon91, LCC99, VT93]. Multidisk [KCK93]. multigraphs [CR98a]. multigrid [CM99a]. multihead [Kut90, Pet94b]. multimedia [KLK97, LY98, SCH95b]. multiobjective [HK96a]. multipaging [Lib98]. Multiple [CCC96, DCD96, Sac96, SCGD91, ACJ96, AF95, Bag97, BC94, Bil93, BGOS95, BCLM92, CLBW90, CY92, GL98, HK97b, KKK97, Mar94, NM94, NM96, Ray89, SR92, ST93, TL95, WL93, Wen91, Pro95]. multiple-chain-like [Bil93]. multiple-choice [GL98]. multiple-headed [BCLM92]. multiple-vertex [ACJ96]. multiplexer [Lin93a, SGSD99]. multiplicand [HM98]. multiplication [Bsh92, Bsh95, Bus92, CC99, DCD92, Ger94, HHK91, Man94, Omo91, TM93, WCC91]. multiplications [Vel94]. Multiply [KB95, ITY96]. multiply-additions [ITY96]. Multiply-twisted [KB95]. multiplying [PKP93]. multiport [LY99]. Multiprocessor [Ber94, DGL95, Amo97, BBDD93, BDDS92, BDDS94, LS90, Nas95, Pom90, Sga97]. multiprocessors [HCF91]. multireader [Vid91]. Multiresource [ST99]. multiselection [DJM94]. multiset [HL91b]. multisets [CDF90, Raj98]. multisignature [LHW99]. multistage [CC92, Wan97b]. Multisymbol [OSK96]. Multitolerant [KA97c]. multivalued [Vid91]. multivariate [Bsh97]. multiversion [KK98]. multiway [EKR96]. multiwriter [LV92b]. Mutex [KW97]. mutual [BW90, DK94b, Fuj98, Gre95, KFYA94, LPS99, MT93, NM94, NM96, SM92, Yen96]. mutually [Dam98]. MWT [AAH99].

N [CC99]. Nam [CLW98]. narrow [Li99]. natural [Gol92, GR98, VB96a]. navigation [MI94]. NC [Cha90, Che91a, Che95d, Gra96a, LMS96, MN92, VSS96]. NCE [Jef91]. Near [Jon98, DD90, Tur91a]. Near-optimal [Jon98]. nearest [AR91, GPS92, Hen97, LCC99, MS90b, Vil95, WCW90]. Nearly [Gra96a, CGL98, Dam95, Jai95, Jai99]. Nearly-perfect [Gra96a]. necessarily [DP99]. necessary [KT90]. needed [Sac96]. neededness [Luc98a]. Needham [Low95]. needs [KW97]. Negation [TNB96]. Negation-limited [TNB96]. negative [FGR91, MM92, Mot91, Pin95]. negative-weight [MM92]. neighbor [AR91, CGR98, LCC99, WCW90]. neighborhood [CHT90]. neighbors [DD90, GM92a, MS90b, Tur91a]. neighbour [Vil95]. neighbourhoods [AR96b]. nested [BCW99, GM90, GPV90, Lar98, LW94, RE94, Sei96]. nesting [BD90]. net [GAG91, SL95]. nets [Esp92, How91, KvdA99, Rac93, SG92, Yen91, ZD90]. network [Alb94, AM98, Ber94, CC92, CH93b, Che94a, CC94b, CDK93b, EGL96, GS90, JM94c, KC98, KB94, Mar94, Sch93c, SB90, SC90, Stö91a, VS96, VI98, Wan97b]. networks [ASA97, AR96b, BOR96, Bol93, Boz95, BM97, CP97, CM99a, CR98b, yCGlW99a, yCGlW99b, Dan91b, FS98, Flo97, FKP93, Fra98, GVV93, GM92b, Gua98, HS97, Höf93, Jan91b, JT92, KW95, KP98b, MS98, MW91, Pel91, RS94a, She91b, SLC93, Sin92b, SLCH98, Thi93, TCL93, VS95a, VCS96, WC97, XL93, Yao92, Yao95, Zis95]. Neuman [HLL^{+}95]. Neuman-Stubblebine [HLL^{+}95]. neural [AR96b, CH93b, Höf93, Jan91b, MW91, SL95, WC97, Yao92, Yao95, Zis95]. Newton [PH92]. next [PW97, SR99]. next-time [PW97]. NF [VS93]. NL [eCCM91, Jen95]. no [BMP94a, BBR99, BDDS94, Cla92, DF96a, Flo91, Gro91a, HS94, Har96, Lam95, Lip96, LCW94b, MS99a, NAG96, Nys96, Pro94, SS95a, SH93, SY93, Sub95a, SC94, SS92d, Woe98, YT96]. Node [GP95b, GP97, DP97a, EGL96, THS97]. node-set [DP97a]. Node-to-node [GP95b]. Node-to-set [GP97]. nodes [NPW98, Pel91, Sen98, Tse96]. noise [AD96, MP98a, Sak91a, Slo95]. noise-tolerant [AD96]. noisy [RRS99].

Non [CH92b, Mor97, Rab98, DZ98, FZ98, JS93, Lê97, Nys96, Reh96, RAGR94, SV99]. non-3-colourability [DZ98]. Non-associative [CH92b]. non-conditional [SV99]. Non-deterministic [Mor97, Nys96]. Non-elementary [Rab98]. non-existence [JS93]. non-linear [RAGR94]. non-smooth [Lê97]. non-SOS [FZ98]. non-structural [Reh96]. nonapproximability [Jon97, Jon98]. nonattacking [Kal90]. nonblocking [Bar90, TCL93]. nonclustered [LPK90]. Noncommutative [FL97]. noncommuting [BGM90]. nonconservative [HS90]. nonconvex [GM90]. nondeterministic [AGM92, CY95, GS96a, Gra96b, IIT92, IITY93, Ros98, Ros93, Sis91, Tze96]. Nondominated [NM94, NM96]. nonexistence [Tau91a]. nonlinear [Dan91a, GVFS99, GSSD97, LY92]. nonlinearity [SLCS98]. nonmalicious [Lon94]. nonpreemptive [KG99]. nonrecursive [LRH^{+}95]. nonregular [BMP94a, BMP94b]. nonuniform [Tys97]. nonuniformly [Tri92]. Normal [Eng92, Pal93, Dem95, LS92, Mok97, Sch93c, War98]. normalization [Reh96]. normed [Gro92, Kuč99b]. Not-All-Equal [AE98]. notation [Kor95, Oht93, Tau91b]. Note [Ano95a, Ano95b, Hen97, Liu90, AL90b, Amb99, Amo97, Bar90, Blu95, BM91, yCH91, CW92a, Cau96, CY95, CKK97, CL91b, Clo96, CFS91, CS93b, CDÉ99, DF96a, DF96b, DPP96, DGL98, DK94a, DK95, EST99, Fer95, Fis95, Flo91, Fuj95, Fuj96, GS92a, GMM94, GI91, Gál94, Gaw96, Goe90, Gol95a, Gol90, GHS93, Gro91b, Har96, Her91a, HS95, HC99b, IITY93, Jan99, JR91, JLD92, Kar90, KW95, Kim96b, Kos99, Kru99, KPRS97b, Lam95, LMS94, Leu92, LM98, LPS99, MTV94, MV95, MT92b, Mar90, Meg92, MP96, NI98, NAG95, NAG96, Ott95b, Pau94, PS93, QM90, Rao95, RV96, Sal92a, Sal92b, SM92, SH93, Sim95, SY93, Sym99, TS96, Tau91b, TR95, Tre96, Tys97, VS93, Wag98, Wan92b, WIIO99, Wei95, YL97a, Yao95, Yen95]. Notes [LP94a, HHK91]. notion [DBM97, FvW95, Jac92]. notional [Che90]. November [KC92]. NP [AT99b, AP95, BJ92, DK91, Dre93, FP91a, Hea90, HR98c, IT94, KK95a, Kin92, KS97a, LW93, Mid93, Ogi96, RRS98, San90, TCC91, Vav90, VI98, Wan97c, Yao92, Zis95]. NP-complete [DK91, Hea90, IT94, KK95a, Kin92, LW93, Mid93, RRS98, VI98]. NP-completeness [Dre93, TCC91]. NP-hard [BJ92, FP91a, KS97a, San90, Wan97c, Yao92, Zis95]. NP-optimization [AP95]. NPO [WCL98, Jon98]. NPO-completeness [WCL98]. NQP [YY99]. NTS [Zha92]. nuclei [Wan98]. NUMA [Gei97]. number [AŽ91, AH94, ABF94, Bac97, Bha99, BK95, BM92, CD93a, CM94, Cus96, CS96, DK94a, DK95, DDK^{+}98, DJS96, FP91a, Gaw96, GK98a, Hal93b, HH94, HW96, KS94, Kin92, Kru99, Lan91, LX99, MN92, MS92, NR98, O’C94, Ott95b, PP94, PLSK97, Ram98, Ray95, SB96, Sub95a, Sub95b, SYUK98, TT91, iT96, VB96b, YMC98, YS96, YT96, Zwi96, dv94, dFMdM99]. number-theoretic [Bac97]. numbers [DP97b, MRR96, PRS96]. numerical [KK94]. nuts [ABF96].

O [RV96]. O- [Jia93]. OBDDs [BJ97, LSW98, SW93a]. Object [Bir90, BS96b, Erc91, FM97b, KAL93, LCH93, LS99b, QB99, Que98, SL96, SB96]. Object-oriented [Bir90, BS96b, FM97b, LS99b, Que98, SL96]. objectives [CH95]. objects [SSCK94, Sri99, VA97a]. oblivious [Ger94, Juk99]. obscuring [Dam98]. observation [RS97]. observational [IN90, Mow92]. Observations [RR92, Sen91, UK93]. observers [Gou94]. obstacles [CK96a, CK96b, SK91]. Obtaining [Wan90b]. Occam [DTW97]. OCCTI [KLR97]. occurrences [FLSS93a, FLSS93b, FPS96]. occurring [GP98a]. October [AAA^{+}91, Ahu96, BCM92, Rit95]. Odd [Kor97, HL98, LCW94b, LCW94a, SC94]. Off [RW96, DS96a, TV96]. Off-line [RW96, DS96a]. offsets [AB98]. Okamoto [Hwa92a]. old [Pri97]. Omega [CC92]. omission [JCT99]. On-line [DS96a, KT97, NLW96, ASS97, BMI97, BNMN92, BKR94, BGP^{+}97, Bre92, Cha92, CWK93, CvW94, CW97b, GJT98, GIZ96, Pru94, RR94, Sga97, TT93, TTC94, WY95, Zha97, van92b]. On-the-fly [FJJR95]. once [BW98, Gál97, Juk99]. One [Loe97, Slo90, BGKL98, GP90a, Gur93, Gur94, Hwa92a, Kem99, Kut90, LM93, LO94, MN92, Mor92, RS99, RS97, STKK97, Sze96, Sze98, TT92b, Vis92, Wis93]. one-bit [Wis93]. one-dimensional [GP90a, Mor92]. one-rule [STKK97]. one-side [TT92b]. one-tape [LO94, Sze96]. one-to-one [Kem99]. One-way [Slo90, Hwa92a, Kut90, LM93, Mor92, RS97, Sze98]. one’s [Omo91, Dam99]. one’s-complement [Omo91]. Onion [AGHP^{+}96]. Online [FS98, KT91, LC99b, Mou99]. only [Vel94]. onto [Kor95]. opaque [She91a]. open [Sta96]. operation [BR94c, HS97, LPK90, Nas95]. operational [FZ98, JH93a]. operations [BÖS91, HL91b, LCL93, OiMI98, VV97]. operator [CK90, HM99, PW97]. operators [Gup93, KLL93, Mas91]. optical [KPRS97b, SCH95b, Yan91b]. Optimal [AERT92, ASS97, AW98, AFI91, Bar96, Che92a, CHM92, DB95, DKK^{+}94, DV99, EGL96, FM97a, HCF91, HS93b, HT98, JT92, JB99, KS93a, Kho93, KB94, LSZ93, LKCNW91, MNT92, PH92, PT92, Raj92, SS94a, SS95a, SR96b, SK92b, VT93, WLK92, WHH98, Weg93, XL93, AV94, AC90b, AR90, BNMN92, BCS97, BMP94a, BMP94b, BT90, BGOS95, BKMT97, CXZ99, CM99a, Car91, CC91, CHT90, CHL97, CT91, Che95c, Dam95, DHO92, Den90, DJM94, DP94, DGL95, FMF95, Gar92, GM90, GM92a, GK98b, Gua98, Har98, HL91b, HHL99, Jia93, Jon98, KM91, KR92a, Kao95, Kel94, Kim94b, KPC^{+}95, KPRS97a, Kir94, KT96, KT90, Lam97, Lar90, LR94, LL94a, LL99b, LBL98, Mat91b, MS94, MS95, ONI96, Ola91, PKP93, Pas92b, PAD99, RI91, Raj98, RU93, SL96, Shv91, WT97, YS96]. optimal [YY92, YY93b, dF92]. Optimality [LV92b, HS91b, SM95]. optimally [CG93, KKK97, Pug90]. Optimistic [KLR97]. Optimization [AFG92, AP95, CLBW90, CL96, GT96, GM92b, GR98, Jan91b, KA97a, SZ92b, Zis95]. optimize [GIZ96, HSY97]. optimizing [CYY^{+}92, CLM96, FP91a, PG95a]. Optimum [Măn98, CH95, CS90b, KMP90, MMS90]. Oracle [Gre91, HP90, Boo91, Cai99, Fos93, Mil92b, Zim99]. orbits [VMD99]. Order [Ess92, Jef91, Kob92, PG95b, AB92, AP99, BS94a, DCD95, DMS94, DH98, DP97a, Dre93, KKK97, KB90, KMP90, KLR97, Lac96, Loe97, MSNR96, OV96, Pri95, SB95, Vog95]. Order-of-demand [PG95b]. ordered [BLW96, BdW99, CN94, Dub90, Gar92, Gro91a, Gro91b, Mäk89, MN90, MR92, SH94, Sin94a]. ordering [AV94, BLW96, JB99, KK95b, RR94, RST91, Raz94, Spi93]. orderings [Bac92, CS92, Jan99, Mar90, Pan96a, Pan96b, Sny93, Ste94, Wal98]. orders [SZ92b]. organizations [Mat90]. organizing [BH98]. orientation [UG92]. orientations [BS99]. oriented [BS96b, Bir90, CHT90, FM97b, Hwa92c, LCH93, LS99b, Que98, RS94b, Ros93, SL96, THI91]. Orienting [HSY97, HR98a]. ORs [Zwi96]. orthants [Vas97]. orthogonal [Bie98, CD99, CKU98, ESW96, TTV91, Yan91b]. ortholattice [McC98]. other [CDF90, KLP96, Riv93, dlVK99]. outerplanar [GLY98, LL96]. Output [Nie96, Bla94, KS97c]. Output-sensitive [Nie96]. outstanding [AF95]. overhead [Vil95]. Overlap [Zha91, AS92b]. overlapped [Lin96]. overlay [dDDS97]. overloading [Hal93a, Sei94, Vol96].

P [YT96, RUG97, Ser91, YY99, Zim99]. P/poly [Zim99]. PA [Kuč99b]. PAC [Slo95]. packet [Boz95, Lin93a, WC97]. packet-switched [Lin93a]. packing [Cha92, CWK93, CW97b, DKST98, Gra96a, HR97, HL95, IL96, IL97, Woe98, van92b]. page [MN92]. pagenumber [SGB95]. paging [Jia91a, Jia91b, Lib98]. Pair [Bac98, BT97]. Pairs [GLS92, Erc91, HC99a, LP93a, LMS96, Tak92]. pairwise [Ber94, DK94c, GP98b, GJS99, HK96b, HK96c, SS98a, Yen91]. Palindromes [Dro95, DL92]. pancake [GVV93]. paper [Dun90, Pro94]. parabolas [DG95]. Parallel [Agg92, Alb94, BHMS90, BÖS91, BM90a, CR90, CDP96, GM98, HL91a, HV96, Ken91, Kim90c, Lin90, Lin93a, LPP91, PP91, PRS96, SGD98, SR95, TM93, Wen91, Alb92, AC90a, ÀG91b, Ama93, Amo97, AM90, AT99b, ALLS96, AS99, BP94, BP95c, BP95b, Bre94, BCDM97, Cam98, CC91, CY98, CHL97, eC91, CDA91, CD92, Che92b, CH92b, Che92a, Che92c, CG93, CST94, Che94b, CH92c, CR91, DKK92, Dam95, DHS95, DS99, Den90, DJM94, DGL98, Dro96, FM97a, Fer94, FGHP93, FMF95, GJLR96, GM92a, Gra96b, Gup92a, HS90, Hag95, Han95, He91, HR98c, HWT99, IIO97, Kel94, Kho93, KV94, Kim90b, KP91, KC98, KS90, Kor95, KS97a, KG99, Kun93, LR94, Lee90, LKC94, Lia95a, LS95b, LO90, Lin94, LY99, LL99b, LK92, LYC99, LP94b, LP93b, LBL98, MP91, MN98]. parallel [iNM97, Nik95b, PR91, PCK95, Pre91, QB99, RK96, RP99, Raj98, Raj92, RR90, RZ91, RR92, RS91, SS94a, SS95a, SK98b, Sen90, Sha93, SR96b, ST90, She91b, SW98, Ski96, SK92b, SB93, TL93, VHV93, Wag98, WC91, Wan92a, WCL95, WC92b, WC93, Yan98, YY92, YY93b, YTC93, van93]. parallel-random-access-machine [BCDM97]. parallelization [CST92]. Parallelizing [Lu93]. parameter [BFR98, Cai96, CH95]. parameterized [AFM94]. parameters [FGW98, OCSF97, XL93]. parametric [KA97a, Tol93]. Parametricity [HM99]. Parametrized [CK93b]. paranoid [GGOQ98]. parentheses [BR98, GP96]. Parity [LSW98, BHW91, Juk99, Jur98, Ogi95, Rus95]. parity-like [Ogi95]. parse [SB95]. parser [Gál94, LC94a, LC99a, Lee92, Ros93, SC93]. parsers [AFG92, PLC93]. parsimonious [MS96a]. parsing [NLW96, RK96, SB95, Yan93a]. part [SC93]. Partial [Jon95, KL90, KC96, AP99, DCD95, FV90, KP91, KCK93, PS92, Pal93, Pri95, Sla97, Vog95, DTW97]. Partially [IL97, Gar92, Sin94a]. partially-ordered [Sin94a]. particular [Mos97]. Partition [Chl96, Bha99, Cha02, CY96, CH92c, CSS92, DK91, Gav93b, LW95a, PC92, SDG92, YC94, YY92, dF92]. partitionable [LO90]. partitioned [Che91b]. Partitioning [Bil93, KK95a, KLW93, KB94, Meg93, iNRN97, PCK95, TB97]. Partitions [Gai97, AŽ91, BJ92]. parts [KK95a]. pass [RS99]. passing [KHJ98, LY99]. Password [Hor95, MW94]. past [Ahu93, Ahu96]. Path [Bax93, GS90, LW93, LKC94, ADY91, AL95, Als96, AR90, CW92a, CDF96, CC92, Cha02, CH93a, CN94, CK96a, CK96b, COS95, DS95, DP93, DPP96, DKS97, FHT93, GJT98, GP96, Her91b, Kan94a, Kin92, Kir94, LP93a, MMS90, NPW98, PSSV93, PWL94, SL96, Sny93, SK91, Tak92, Tze96, WCT99, YC94, dlVK99]. path- [CDF96]. path-decomposable [GJT98]. path-partition [Cha02, YC94]. path-width [Kin92]. pathmax [DCD95]. paths [Agg92, Bax94, BF96a, BU98, CR98b, Che94a, Che95b, CM94, CM98b, Cou97, Cri91, GP97, GP98b, HR97, LP97a, LMS96, PR91, PM94, RV93, Sim90, TT91, VSS96, Woe90]. pathwidth [GLY98]. Pattern [BBL98, ALLL98, AR96b, BYR93, CGR93, CCG^{+}99, CLP91, FPS96, Iba97, KCK93, Owo93, Per94]. pattern-matching [Per94]. patterns [Dan91a, DF96a, DF96b, HWT99, PP94, Ver92, YCH93]. PB [Jon98]. PB-complete [Jon98]. PCP [TS96]. pebble [GGK91]. peeling [Nie96]. penetration [Sri99]. Perceptrons [Fal90]. Perfect [BN96, CHM92, MK92, CL93a, DCGP99, DP94, Gra96a, Kal90, KS94, KK95b, Mai94a, Pan96b, Ros99, Rus95, SH94, TKI99, Wei91, Yan93b, YL90]. Perfectly [Lin96]. Performance [AP93, Gei97, Hac93, Mat90, Pro94, TT90, Hal93c, Nas98, Sla97, TTC94, Vos93, Zis95]. performing [FP91a]. period [Pas92b]. Periodic [Tuz90, FLSS93a, FLSS93b, RW97]. periods [AB98, BD96, NCC98]. permanent [BF97, Sal92b]. permutation [AR92, CW92b, CHL97, CST94, Che95b, CL92c, CL93c, HWJ96, HAM11, Kim90a, LRDL91, Lia94b, MVR96, RS96, RLDL94, SHL94, SMN^{+}95, WCL95, Wha92, YTC93]. permutation-based [CST94]. Permutations [Chi93, Alo90, vB99, BBL98, Chr96, HW99, Iba97, KR92b, Lin90, Pan96a, Res92, San98]. perturbation [CH95]. perturbations [SJKL95]. Peterson [BW90]. Petri [Esp92, GAG91, How91, KvdA99, Rac93, Yen91, ZD90]. Phase [Mis91, Mis92, BF96b, FGW98, HG95, HL98]. phase-clocks [HG95]. philosophers [HP92]. picture [Kim96a, TCC91]. pipeline [MG90, VHV93]. pipelines [Myo91, Myo92].

Pipelining [BS90]. Pitts [SL95]. Place [SG92, HW99, KP99, Res92]. Place/transition [SG92]. plactic [AKK97]. planar [BG92, CP95, Cim90, Cou97, CKU98, KW92, KK95a, MMN91, Mid93, MRSW95, iNRN97, RS94b, RW92, Tam91, TTV91, TV93, Tip90, Woe90]. planarity [DLR90]. plane [AFB95, AERT92, Den90, DS93a, GT96, Jan91a, Lau90a, ND95, RWZW91, SKKC98, TT92a, dF92]. Planes [Joh98]. planning [Li94]. Point [Ove92, AAA^{+}91, AM98, Ber93, BNN91, CD98a, CD94a, DW90, DG97, GS96b, Mat95a, MRSW95, MRR96, NF95, Rot91, RW92, Sho91a, Tam91, Ver97, Woe92b]. point-to-point [NF95]. points [BU95a, BU95b, DD90, Gon91, GK98b, He91, Hoa90, HKK92, Jan91a, KKK99, LX99, Mat95b, MR92, MW91, SK98b, TT92a, Ver97, Wan91]. polarity [KA97b]. policy [Ano94, Lin93a, Oya95, SR96a]. polling [Res96]. poly [GKL96, Zim99]. poly-log [GKL96]. Polygon [BD90, AS90, AL95, Als96, CG93, CD98a, CD94a, FH94, FNTV90, GM90, He91, Her91b, KPC^{+}95, She92, SYUK98, Zhu92]. polygonizations [AGHP^{+}96]. polygons [BM90c, BM90b, CXZ99, ECOUX95, GJS99, Her96, LL94a, MRSW95, Saa97, Sto91b, SD91, THI91, dDDS97]. polyhedra [Sha93, Sto91b]. polylog [Gav93a, RW91]. polylog-threshold [RW91]. polylogarithmic [IL97]. polymorphic [Har94, Har96]. Polynomial [CL93a, Mos97, MN93b, PSL92, AQS97, BHW91, Boo94, Bsh92, Cai99, CT97, CT91, DK91, Enc97, Gav93a, Göb99, IZ94, Jef91, KA97a, Kol94, LL98, Low98b, OS95, PZC98, RL93, Sch95a, Ste94, Sub95a, Sub95b, TY94, Vel94, VA97b, Weg93, Yen91, Can96]. polynomial-size [IZ94, Weg93]. polynomial-time [Boo94, Sch95a, VA97b, Can96]. polynomially [Gup92b]. polynomially-bounded [Gup92b]. polynomials [AM95, Bsh92, Bsh97, GPS92, GS92b, Ito91, Lan91, Sch91, Sho90, Sho91b]. polyominoes [Bea91, CD99, MS96b]. Polyominos [Mai94a]. Pop [AS99]. Pop-stacks [AS99]. portion [hC98]. poset [NMS97b, NMS97a]. position [WT97]. positive [Den99, Don95, Gav93a, KA97b, Mäk97, Mot91]. possibly [Dij90]. Postman [hC98]. power [BN96, BGM90, CS99, FS92, Gol95a, Gol97, Hec91, Lar95, SS92a, van93]. powerful [Ros97, Sen97, WC90]. powerlists [Kor97]. powersets [Jan99]. pp [AAA^{+}91, Ahu96, BP95c, BP95b, BU95a, BM90b, BCM92, Dun90, KC92, Mis92, Pac92a, Pet91, Rit95, Sny91a, Ueh97]. PR [PS94]. practical [BYP96, CCG^{+}99, LL95]. practice [AW98]. PRAM [Chi93, KPRS97a, Kuč97, LP97b, Rag91, RR99a, Shv91, Shv92, dB93]. Pramanik [Pro94]. PRAMs [VHV92, Vai92, Weg90, Sin98]. pre [DB95, LC99a, LS92, Sin94a, Zha92]. pre- [Sin94a]. pre-determined [LC99a]. pre-emptive [DB95]. pre-NTS [Zha92]. pre-requisite-free [LS92]. precedence [Amo97, DKL99, HK97b, LCL93, RPG98]. Precision [Bri93b, Riv93]. precondition [SZ92a]. preconditions [PR98]. predecessor [ACH97]. Predicate [YL91, Dij95b, Dij96, Kna90, Mor97, Nau94, de 92]. predicates [GM97, SL96]. predictable [Nas98]. preemptable [Kim94b]. Preemptive [BBDD93, LY90, Ple90, AZ90]. Prefix [FP94, GLS92, Lit91, BCT93, CH92b, Fok94, Hag95, LY99, LL99b, Măn98]. Prefix-free [Lit91]. prefix-matching [BCT93]. preflow [CM99b]. preflow-push [CM99b]. preprocessed [HKW90]. presence [Bag92, BOR96, yC90, Sak91a, SB90]. presentation [Col94]. Preservation [WY92a, Cho96, Nik95a]. preserves [MC92]. Preserving [AAA^{+}91, Ohl95, Val95a]. prevent [LL99a]. primal [Rav94]. primal-dual [Rav94]. prime [AB98, DJS96]. Primitive [Mit98, LCW94b, LCW94a, Maz95]. primitives [Ahu90, GS98a, JW96, Sin98]. principle [GMPW98]. priority [AR99, AB98, BC94, Boz95, Bri93a, CD93b, DB95, HMPS96, Li90, Lin93a, PP91, RR94, Tsa93]. priority-ordering [RR94]. privacy [CK93a, Dom96]. Private [New91]. Probabilistic [BHW91, Her90, HL95, Jac91, BK98, Cai99, CM94, DIK97, HK90, HHL95, KS93b, MS99a, Mes91, Wan92b]. probabilistically [RS95b]. Probability [BSG99, Al 96, GS96b, Gup93, KKK97, Sen90]. probably [ONI96]. Problem [Bax93, FP91a, ACS99, ACH97, AL90a, AR91, AvSW95, AR90, AR98, ABF94, Bas94, Bea91, BP93, BR94c, BU91, BJ91, BCT93, BD98, CM98a, CEdK98, Cha91, Cha02, CCC96, CGR98, Che91a, Che92a, CT93, CH93a, Che95b, CJ97, CL91b, CLLR97, CI95, Cou97, CFS91, Cri91, DMM98, DF96a, DF96b, DCGP99, DG97, DW96, DGS97, DGP94, Dub90, ELS93, ETA92, Esp92, FF90, Flo91, FV90, Gav93b, GL98, GHY93, Gon95b, Gro91a, GP97, GZ97, HPS93, HHI97, HR94, HH94, HT95, HR98b, Her91a, Her93, HI97, HK97a, HC99a, HP92, HTC91, HCH94, HH91, HHW94, Ira91, Jan92, Jan91b, Kal90, KS94, KKS91, KM90, KMN99, Kim90c, KL90, Kim94a, KT97, Kin92, KP96, KW98, LL98, LWS99, LP93a, LC99b, LW95a, LRDL91, Lia94b, Lia95a, LMS96, LL94a, LS95b, LC93, LC97b]. problem [LX99, LP90, Low98a, Low98b, LT97, LP94b, Mae90, Mäk89, Mak90b, Mäk92b, Mäk92a, Man96, MR96a, MT92b, MM99, Meh88, Mei90, Min91, Mos97, iNM97, NF95, Neu96, Nta90, Ov98, Oh93, Par91b, Pas92a, PC92, Pet94a, PPK98, PS98, Pra96, PN94, PP90, RI91, RR94, RRS98, RPG98, Rav94, RW96, RL96, RZ92, SY90, Sal92b, SCGD92b, SDG92, SGB92, SZ92b, SM94, SB93, Sze96, SS92d, SS92c, Tak92, THI91, Tan91, TTC94, VSS96, Vis92, Vos93, Wan91, WCW90, Wen94, Woe90, WY92b, WT92, WC92b, WC93, WCL98, WCT99, XU98, YC94, YL90, YY92, YY93b, YTC93, Yu95, Zel93, Chl96, DvR94, GMM94, GOPR92, GH96, GLS92, Kru95, LW93, Rou97, VI98]. Problems [Dam90, AERT92, Ali96, ÀG91b, AP95, BGM90, BV98, Cai96, CW92b, CL93a, CH95, CW98b, CJ90, CM94, CH92c, CSS92, CS90b, CS93b, DIK97, DE92, DS99, Daw92, DT93b, DM91, Fis95, Fuj96, GT96, Gav93b, GSSH93, Gro91b, GJS97, GR98, How91, Jen95, JR91, Jon98, KLP96, LR94, LMS96, LV92c, LP91b, Mei91, MN93b, Nas98, PK99, Sac96, San90, Sat93, SR95, SSC98, SS92b, Ste91, TB97, Tom95, TT93, Wei91, WY95, Yan91a, Yao92, Zis95, dlVK99]. procedure [Bij93a, SI99]. Procedures [Ser96]. Process [AAA^{+}91, Ahu96, BP95c, BP95b, BMP94a, BDDS94, BU95a, BM90b, BCM92, DF96a, Dun90, Flo91, Gro91a, HS94, Har96, KC92, LCW94b, Mis92, NAG96, Pac92a, Pet91, Pro94, Rit95, SS95a, SH93, Sny91a, SY93, Sub95a, SS92d, YT96, BP98, Bol93, Cha95a, FZ98, KC91a, Mes91, PK90, Tau91b]. processes [BK98, Goe90, Jac90a, Jac92, KTS96, Kuč99b, Gro92]. Processing [Cha02, CK96a, yCGlW99b, HAM11, Khu00, NS96a, NM96, RRW00, GS98b, GP95a, KP91, KKK99, LCC99, Mak90a, SCH95b, TL95, YLW91]. Processor [Gol91, Ten90, Von92, AA96, Ama93, BBDD93, CFGG97, DGS97, Gaw96, HI90, KC90, KP94, Kuz94, Lee90, Lin91a, LP90, RS96, SB90, TM93, WCL90, WC90, YDL94]. Processor-efficient [Gol91, Von92]. processors [Amo97, BDDS92, BDDS94, BBGT96, CS91b, GMR90, HNC92, Kri93, PT92, VHV93]. Producing [SB95]. product [BM98, CC99, CS93a, Fla94, GMR90, GMPW98, Gup92a, Lit92, MP98a]. products [KS93b, Lin91b, Yod91]. profile [MK92, Yan93b]. profiles [Wan97a]. Program [von91b, CS99, DHS95, HW93a, Lei95, Yua94]. programmed [Don95]. programming [All92, BG94, BLM90, BCLM92, Den90, Esp92, GP90a, KB94, LR94, Lon94, Meg92, Myo91, Myo92, RZ92, RW94, SGD98, Ser91, Van92a, Vav90]. programs [Bij91, BW98, BTW98, CCK93, CCH93a, CCH93b, CH99, Gál97, Ger97, Hir97, Jai95, Juk99, Kob93, Kor95, LKC94, LRH^{+}95, MN98, Miz99, PR98, Ros93, SJKL95, WY92a]. Progress [CH99, Kna90, Lin94]. progressions [Hea90]. projection [Sha93, van93]. projections [CD99]. Prolog [PR98]. proof [AT99a, Bur92, Cho96, CKU98, DCGP99, DKVW95, Dij95a, DS94b, Eng92, Gro92, JH93b, KMTd94, LRH^{+}95, Lys94, MG92, Oht93, Pac92a, Pac92b, PSSV93, Ram90, SM95, Sug94, VT90, Ver87, VS95b, Wan90a, Zhe95, VT94]. proofs [CP90, DP94, GH98, Lac96, Lau90b, Lut90, McC98]. propagation [Han92, SCP99]. proper [HMP95]. properties [BK98, Cai96, Cal94, CH99, DCH96, Ham95, HS95, IWW93, Kob93, LH93, Mei90, NO96, OiMI98, PW97, Sch95a, Tsa92, UK93, Yua91, dBDL96]. property [Cho96, DG99, Kam92, KP98a, Lin94, SW90, Zha92, de 95]. proposed [Gál94]. propositional [Bün90, CH92a, CL93b, Cre98, GS95, Loz92, Pen92, Rab98]. protein [FLSS93a, FLSS93b]. protocol [BU91, CLBW90, CT91, hC98, DKVW95, HHL95, KLR97, Li90, Low95, RS98, Ten90, TH99, Val95b]. Protocols [Hwa92c, CJ95, GP90b, Hor98, HLL^{+}95, LHW99, LL95, RT99, Tau91a, TWK99, TH99, VCW95, Zhe95]. Proving [Sis91, DF91, Lip96]. proximity [AERT92, Uhl91]. proximity/similarity [Uhl91]. Pruned [KP98b]. Pseudo [Ven98, Alo90, HS91a, HS94, SS95b]. pseudo-algorithmic [SS95b]. Pseudo-average [Ven98]. pseudo-lines [SS95b]. pseudo-random [Alo90]. PSPACE [RS95b]. PTAS [Woe98]. public [Cus95, LC96, Low95, SM90, Tze99]. public-key [Low95, Tze99]. Publisher [Ano95a, Ano95b]. pure [PR98]. purpose [Pom90]. pursuit [Daw92, Neu96]. pursuit-evasion [Daw92, Neu96]. push [CM99b]. pushdown [LP94a, Pet94a, Sal96]. pushout [Vel96]. puzzle [Par95]. pyramid [AJG95].

QBF [ALLS96]. quadrangle [BG94]. Quadratic [Vav90, Gas91, Pri95, Pri97, Tel94, YJ93]. quadtree [PS94]. qualitative [BK98]. quality [BS96b, KLL93, Wan99b]. quantifiers [LH93, Ott95b]. quantum [Amb99, BdW99]. Quasi [Bij93b, Rus95, ED91, Jan99]. Quasi-Boolean [Bij93b]. quasi-orderings [Jan99]. Quasi-parity [Rus95]. queens [ETA92, Kal90, Oh93, RZ92]. queries [AT99b, Ber93, BCBF95, Bsh96, CYY^{+}92, CK96a, CK96b, Gas91, GK98b, HKW90, Her91b, KP91, KC96, KMP90, OV96, Sak91a, SL96, Ser98, SS98b, TL93, Uhl91, Wag98, ZÖ94]. Query [SCH95b, CL99, DS96b, GS98b, LCC99, TL95, YL91]. querying [LW95b]. question [Jai95, Jai99, Pin95]. questions [BTv93]. Queue [GS93, AR99, HMPS96]. Queue-mergesort [GS93]. queues [Bri93a, Fag96, PP91, Ros98, Tsa93]. quickest [CH93a, Che94a, LP93a]. Quickselect [Pro95]. quorum [AW98, CY94, Fuj98, KH96, TIK99]. quorumcast [Low98a]. quorums [AEE97]. quotient [KP98a].

Rabin [GBY90]. radio [Boz95, WC97]. radius [DD90, Tur91a]. Radix [Man91, EKR96]. Ragde [Hag92]. ramp [SW99]. Ramsey [Lip96]. Rana [VT94]. Ranade [CDK93b]. Random [Gol95b, Res92, Rob90, San98, Ver99, All94, Alo90, AD97, AS92a, Bai92, Boo91, BCDM97, BDF^{+}95, BHWG99, CY96, DK94a, DK95, Dev98, Fei92, FS92, FKP93, FPS96, Gol99, HW99, JM94a, KS93b, Kir92, KS98b, Kru99, Mai94b, Mil92b, MZ93, New91, Sch93b, Ueh97, YDL94, Zim99]. randomization [Cha92, Gol92]. Randomized [Dam98, Mat91b, AM90, Che91a, CvW94, CLLR97, DIK97, FM97a, GJLR96, GP90b, Kuč97, LPS99, Sga97, Tei93, TY94, TT93, WT92]. randomizing [WB93]. randomly [Kor93]. randomness [Kob93, Wan99a]. Range [Bri93a, SV96, CST94, GJS97, KS98a, Kim94a, Mou99]. range-join [CST94]. Range-restricted [Bri93a]. Rank [Blu92, Sim91]. Rank- [Blu92]. ranking [AM90, KMW98]. rankings [ZNN95]. ranks [RR99a]. Rao [CLW98]. Raphson [PH92]. Rate [OCSF97, SO93, Bla94, PK93]. ratio [CLLR97, HLV97, PMC97, TV93]. rational [Bea91, HLV92, Wan98]. Rationality [CS92]. ratios [BMI97]. ray [KB90]. ray-tracing [KB90]. Raymond [SM92]. Razor [DTW97]. reach [Han92]. reachability [CI95, HSY97, HKW90, KV94]. reactive [AS95, Hen92]. read [BW98, Gál97, HSC93, Juk99, PR93]. read-once [BW98, Gál97, Juk99]. read-twice [PR93]. Real [ABL99, CJKJ98, Dro96, MS98, Sin94b, AH94, AS95, HL92, HSZF93, HK97b, Joh98, Kim94b, LCL93, MP98b, OCSF97, Par95, Ros97, SO93]. real-number [AH94]. Real-time [CJKJ98, Dro96, MS98, Sin94b, AS95, HL92, HSZF93, HK97b, Kim94b, LCL93, OCSF97, Par95, SO93]. Realizability [Dil90, Sug94]. Realization [SHL94]. realizer [NMS97a]. realizing [Lam97]. Rearrangeable [HZS99, TCL93]. Rearranging [XU99b]. reasoning [BLM90, GAG91, Pet91].

Recognition [PM94, AGM92, AR96b, Cai93, CH98, CKN^{+}95, HMP95, Jia93, MSW93, Nik95b, RR90, Ros98, RR92, RS91]. recognizability [Her98]. Recognizable [AKK97]. recognize [LO90, Slo90]. recognized [Jef91]. Recognizing [HR98c, IK97, Man90, PWL94, PS95a, PS95b, SR98]. reconfigurable [CCC96, Lin91a, PKP93, SRW95, TM93, Vai92, VT93, WCL90, WC90]. reconfiguration [Jac90a, Jac92]. Reconstructing [CD99, Fid98, KR92a]. reconstruction [GS92a, LH95, Tam91, Wen94]. Recording [AB92]. records [MZ93]. recovering [LLM97]. recovery [CG91, SM95]. rectangle [BKMT97, FH94, LMT99, SK98b, TT92b]. rectangles [KT96, MNU90, dF92]. Rectangular [DK91, CK96a, CK96b, HLV97, TT91]. Rectilinear [CK96a, CK96b, dv94, CD94a, CM98b, PZC98, Sal92a, Sny91a, Sny91b, SK91, SvK99]. recurrence [KvdA99, Ron91, Tze99]. recurrences [PS93, RL94, WF96]. recurrent [CK93b, SV99]. recursion [GP98a, Nau94, Tau91b]. recursive [HS95, Hir97, KC98, KTS96, Kum95, Lee92, LRH^{+}95, Mic97, RS98, Sny93, Vys90, van91b]. red [CW92a]. red-black [CW92a]. reduce [ST91]. Reduced [AFG92, DCD92]. reducers [Tho95]. reduces [Lib98]. reducibilities [eC91]. reducibility [AP95, Käm90, KR95, Wan95]. reducible [Gav93b, Gia94, LKCNW91]. Reducing [Chu98, Vil95, Yan91a]. Reduction [CDK93b, SW93a, Bac92, LC99a, NMS97b, Obe95, PLC93, RT99, ST91, Wal98, WS94]. reductions [AG91a, AKM95, BHW91, FS92, Sch93b]. redundancy [MZM99]. redundant [PP94]. Reed [KA97b]. REF [Clo96]. reference [Lin92, Lin93b, MWL90, PG95a, Wis93]. references [Har94, Har96]. referencing [Lar98]. Refined [Pet94b]. refinement [AS95, CS91a, Lut90, Ser96, Wri99, von91b]. regeneration [GK97]. regions [Gai97]. register [Hes98, VR96]. regression [Ols97, SGD98]. Regular [Lar98, Bir92, CRC98, CDÉ99, Flo97, FP94, KP98a, MVR96, Mäk97, Mid93, Rub90, SC93]. Regularity [Kuč99b, BV92, RRS99]. regulation [Gär93]. related [DBM97, Gol92, Gro91b, Jan91b, LMS96, NR98, RRW94, RRW00, Tur91b, YMC98]. Relating [Bie98, LS92, Käm90]. relation [DT93b, VS93, XU99a]. relational [Das96, DT93b, FP91a, GS98b, LSS92, LW94, MZM99]. relations [BCW99, CS99, DT93b, PCK95, Ron91, Tze99]. relationship [IIT92, JM94a, JM94b, Liś93, RS95b]. relationships [HK97b]. Relative [CS90b, Zim99, AS92b, CRC98, DMS94, Mil92b, WT97]. relativization [Gav93a]. relativized [RS95b]. Relaxed [KHJ98, BD98]. Release [Fre83, CD98b, GK98a, Ple90, SY93]. reliability [CD93a, LC97b]. reliable [KKS91, Lat91]. remainder [PSL92]. remark [Gär93, MP95, SSS98, TCC91]. Remarks [Kut90, Mäk92b, Rac93, BM98]. removal [BS98]. renaming [FM97a, Héb95]. repair [JCH94, YCH93]. repeated [Car94]. repeaters [HH91]. Repetitions [Hes91]. replacement [Dre93, Pru94, Wri99]. replay [YL97b]. replicated [AE90, KC91b]. replication [KS98a]. Reply [DW91]. reporting [BKMT97, GJS99, LMT99]. representable [BCW99]. representation [AH94, CP97, CN94, CKU98, GAG91, LW95b, YLW91]. Representations [Kui95, Bol93, BL96, DTT92, KLTT97, Kuč99a]. representatives [BR97]. represented [Kim96b, Vei97]. representing [Jag97]. requests [Har98]. requirement [KLTT97]. requirements [CG91]. requires [AT99a, BW98, Gál97]. requisite [LS92]. research [Bri93b]. residue [BP94, BP95c, Bar95, BP95b, RP99]. residues [DMS94]. resilient [AT99a, DK94b, GS92b, GHS93, Tau91a]. Resistance [YT95]. resistive [DM91]. resolution [CP90, LH92, Naw91, TC99, Yan91a]. resolutions [LKCNW91]. Resolving [HK97b, LK95]. resource [BGP^{+}97, Bir90, TL93]. resources [GM91, Pom90]. respect [DS96b, HJL^{+}91, IOU94, Sis91, VSS96]. response [ST99]. responses [CLP91]. Restricted [RT99, AS94, BM91, Bri93a, Che95c, Gas91, HCF91, WCW90]. restriction [IL96]. restrictions [GJS97]. result [CS97b, DGS97, FGR91, Jon97, Käm90, Pac92a, Pac92b]. resultants [Enc97]. results [AAH99, Ano94, BM97, DGP94, Ira91, Jon98, Man94, MR95, Wan95, WS96, Yan91b, ZJ94, ZÖ94]. retrieval [BÖS91, Chu99, DG99, KLL93]. retrieving [ABK98]. reuse [FS98]. reversal [AL90a, Liś93]. reversible [Mor92]. revisited [AG91a, BH91, Che99, DG99, Fal90, MCR98, Nei94, RUG97]. Revisiting [AM97, CS97b, How98]. rewriting [Ban96, DS93b, DS94b, Göb99, HLM98, IN90, KR95, KK90, Mül92, Ohl95, Ott95a, STKK97, Sun96, Zha91]. richness [Dev98]. right [JCH94, SC93]. rights [DF91]. rigidity [SSS98]. rigidly [HKK92]. ring [GMR90, KM96, MNT92, Sen98, SLCH98, Tse96]. rings [BT90, BL90, CS95, CFGG97, FLMSS98, HP96, HL98, HHS99, IKOI99, KM91, NAG95, NAG96, ONI96, Sat93, WH94]. river [TT90]. RNC [Wei91]. RNS [HP94]. robber [Nta90]. robot [MI94]. Robust [Pad98, BD97, Boz95]. robustness [BD90]. rollback [YW96]. Root [Luc98a, BMST97, Car91]. Root-neededness [Luc98a]. rooted [Raj92, SS94a, SS95a]. rooting [ITY96]. roots [GPS92, Sch91]. Rotation [KB90]. Rotation-based [KB90]. rotations [Mah98]. Roudneff [Alf99]. roughly [AC90a, MK94]. roulette [Dij90]. Round [DP94, Bur92]. Round-optimal [DP94]. rounds [AT99a, IOU94]. routable [Pan96a]. Route [Bag97, Nta90]. Routing [LLP90, Sim90, Sym96, BL91, CC92, CL92c, CL93c, yCGlW99a, yCGlW99b, CDK93b, GVV93, GJK92, GM92c, GP95b, Gua98, Jan92, JT92, KK94, KC98, KPRS97b, LL94b, Low98a, Lub90, RP91, SGB92, SB90, SV94, Sym99, TS95a, TT90, TT92b, VS96, Wan94, Wan97b, Wan99b]. routings [WLK92]. row [BL91, Gur93, Gur94]. RSA [CH97, GGOQ98, Hor98, HP94, KOT95]. Rudimentary [AG91a]. rule [CM99b, Han92, Obe95, STKK97]. rule/goal [Han92]. rules [Col94, GM98]. run [BC95]. run-length [BC95]. Running [VHV93, BDF^{+}95].

S. [Pro94]. S5 [AG97]. safe [VV97, Yen96]. safer [Hen92]. Safety [AAA^{+}91, CL91a, DW90, PS92, Gum93, Sis91]. Salesman [DvR94, BD98, DW96, HR98b, Par91b, PPK98, Vis92]. same [BL96, IWW93]. sample [AD96, CS93a, Den99]. samples [AD97, LC94b]. sampling [CEG95, CGR93, GS96b]. sandwich [CEdK98, Sto91b]. SAT [Ver99, AE98]. satellite [Bil93]. satisfaction [BP93, How95, How98]. Satisfiability [HW93a, Ali96, CL93b, DE92, KP92, Li99, MR96a, Pre93, Tan91, WT92, Héb95]. satisfied [NO96]. Satisfying [Uhl91, Che92b, Cus96, CS96, O’C94, YT96]. saturated [Reh96]. says [Loe97]. Scalability [GH96]. Scalable [Nas98, Pru94]. scale [LY98]. scaled [ABL99]. scan [Lin92, MWL90, Sin98, Oya95]. scattered [XU99b]. schedule [Boz95]. Scheduling [AA96, BDDS92, BDDS94, BBGT96, DKD97, Fre83, GP95a, HL92, LYC99, NCC98, SR91, SY93, Woe95, AZ90, Amo97, AB98, BKR94, BGP^{+}97, BBDD93, Bir90, CvW94, CD98b, CW99, CJKJ98, DGS97, DKL99, Dro96, Gaw96, GK98a, Har98, HLW93, Kim94b, KP94, KS97a, KG99, Kuz94, LY98, LY90, Lin93a, OCSF97, Oya95, Ple90, RS96, SO93, Sga97, ST99, TL93, TL95, WC91, WC97]. schema [Cha95b]. Scheme [Hwa93, CLW98, DS99, GK97, HK90, HS91b, HW98, JS95, KA97a, KH96, LY92, PK90, Wan90b]. schemes [Blu95, BD97, BDV98, BSG99, CT97, CS99, DT93b, DKK^{+}94, HCY96, IOU94, KO96, LC97a, MPSV99, OK96, Pad98, SW99, SS94b, TJ99, Tze99, VS93, WB93, ZI98]. Schneider [Gum93]. Schröder [ARS98]. Schroeder [Low95]. screening [SJKL95]. SCS [BJ91]. Search [TKI99, AFB95, Ali96, AP95, BY91, BB94a, BdW99, Cha91, CCC96, CHL97, CD94b, Dam95, DCD95, DCD96, DD93, Dev98, DD90, DR94, DN93, EKR96, Fis95, GL98, GM91, HT98, IY98, Kru99, KIS90, LC99b, Li99, LPS99, Low98a, MH96, MM92, Man90, Mar91, MW91, Pug90, RLDL94, SCGD92a, Ser98, ST91, Sim95, SYUK98, Tur91a, Wen91, FGN97]. searches [HM96]. Searching [AS92b, CLP91, CD96, CDF90, GJT98, GJS97, Han93, HTC91, KS98a, Lar90, MT98a, Owo93, Per94, SV96, THI91, Tol93, Vil95]. second [KB90, Lac96, LP97a, Loe97]. second-order [KB90, Lac96]. Secret [MT98b, Al 96, BD97, BDV98, BSG99, Hwa92c, KO96, LH95, MPSV99, OK96, Pad98, van97]. section [BC94, Ray89, SR92]. secure [Hwa93, HW98, KK98, YL97b]. security [CJ95, HC95, JT96, LTT95, LHW99, MW94]. Seek [MV91, MK96, Oya95]. seeks [VM97]. segment [Kim90c]. segmentation [CTVC98]. segments [AS92b, CKU98, DD90, ES98, Sha96, Wis92]. selectable [Gon95a]. Selecting [RR99a, Gas91, hC98]. Selection [Kol94, AJG95, Bag97, CM99b, KS93a, Kun93, Mat91b, SR96b, ST90, She91b, Wri99]. selective [Rao95]. selectivities [Chu91]. selectivity [Wan95]. selector [BTv93]. Self [BH98, CD94b, HC90, Mar91, RS95a, CYH91, DKVW95, Gho91, GG96, Goe90, Her90, HP92, HH92, HC92, HM96, Käm90, LC99b, MN98, Wan95, WIIO99, Yen96, TKI99]. Self-adjusting [Mar91, HM96]. self-approaching [LC99b]. Self-Indexed [TKI99]. Self-modifying [RS95a, WIIO99]. Self-organizing [BH98]. self-reducibility [Käm90, Wan95]. self-stabilization [Gho91, Her90]. Self-stabilizing [CD94b, CYH91, DKVW95, GG96, HP92, HH92, HC92, MN98, Yen96]. Selftype [Sak91b, HC90]. Semantic [Lee96, BLM90, FvW95, MZM99, TW95, YL91, von91a]. Semantics [CM96, Dan91a, LH93, BMC96, BFSV99, BNN91, Che90, FZ98, GW96, Hd91, JH93a, Nys96, OH91, SZ92a, Vog95, Yua94]. Semantics-based [CM96]. Semantics-directed [Dan91a]. Semelectivity [DV91]. Semi [Rit94, Rit95, Dam99, HP90, RS94b, Ueh97, Zha97]. semi-algebraic [HP90]. semi-random [Ueh97]. semi-strong [RS94b]. semi-unbounded [Dam99]. Semi-unification [Rit94, Rit95]. semiring [Kui95]. senders [HW98]. sense [BV98, FMS97, Flo97]. sensitive [DB91, Nie96, Slo90]. Sensitivity [Ber96, Bop97, Ven98]. Separability [Jac90b]. Separating [BU95a, BU95b, ES98, Gre91]. separation [GM92c, HL92, Kin92, RRW94, RRW00, SS95b, Wan99a]. separations [Boo91]. separator [Che94b]. separators [KW92]. September [Pet91]. Sequence [Bra90, hC98, KP94, MBY91, WMMM90]. sequences [AC90a, Bla94, CLBW90, CT91, GVFS99, HW93b, Kor94, MK92, RU93, UK93, Vaj98, WS90, Yan93b, van91a]. Sequential [YTC93, vH95, BMC96, DS94b, Ger94, HS90, HLM98, Jac90a, Jac92, LHW99, Pio90, RR90]. serial [MN98]. Serializability [Raz94, EV97, ED91, RE94]. serialization [KLR97]. series [BGM90, DS99]. Server [DM91, CL91b, CLLR97, EYK95, Hor98, IR91, Kle94b]. server-aided [Hor98]. servers [LY98]. service [Wan99b]. set [Agg92, AR92, BGP^{+}97, CEdK98, CW92b, CL92a, Cha98b, CH92c, CDP96, DKK92, DG97, DP97a, ELS93, Fuj96, GI91, GKL96, Gon91, GP97, Gup92a, Hea90, Hir97, HTC91, HTC92, HS93b, HL91b, Irv91, Kei91, Kim90a, Kim90b, KT97, LMS94, Lia94b, Lia95b, Luc98b, MMS90, MR92, Mos97, OiMI98, PMC97, Pas92a, RI91, Sho91a, Sta96, Wis92, YTC93, AE98, GHY93]. set-valued [HS93b]. sets [AM95, AKM95, AT99b, Bea91, BTY97, BIN92, DHO92, DNPS98, Fok97, Gar92, Gre98, HP90, HR98c, HL96, HKK92, KLP96, Kim96b, MS96a, MN93a, MRSW95, MT98b, Par91a, Pri97, Rao95, Reh96, Ros99, Rot91, RW92, SS94a, SS95a, Sch93b, Sch95a, Ver97, Wan97c, YJ93, Zim97, Ziv91, dF92]. Setting [OCSF97, MP98b, Mou99]. setup [SZ92b]. Several [LL95, Pro95]. shadowed [VM97]. Shamir [Ada92, GGOQ98, TT94]. shaped [Che96]. Shared [YL97b, AF96, FMT93, MP91, OM97, SJKL95, Tho95]. shared-environment [Tho95]. shared-memory [SJKL95]. shares [BSG99]. Sharing [Pom90, BD97, BDV98, BSG99, Hwa92c, KO96, MPSV99, OK96, Pad98, van97]. shattering [ES98]. Shelf [CW97b]. Shellsort [WS90]. shift [Per94]. shift-or [Per94]. Shifts [KL98, Wan94]. short [CK96a, CK96b, GMM94, Goe90, Gro92, PSSV93, Wan90a, YS96]. Shortest [ADY91, FHT93, VS96, Als96, CR98b, Che95b, CGH95, DP93, GP96, GMSS99, Gus94, Hen97, Her91b, HC99a, JLD92, Kan94a, LP97a, LMS96, NPW98, PWL94, RV93, Tak92, VSS96, Zhu92]. shortest-path [Kan94a, PWL94]. shortest-paths [Che95b]. showing [CDK93a]. shuffle [Jed90, MY96, Sho92]. side [TT92b]. sieve [BF96a, DJS96]. sigmoid [Höf93]. signal [Mak90a]. signature [DML98, Hor98, KC96, TJ99]. signatures [WHL99]. signcryption [ZI98]. signers [TJ99]. SIMD [BS90, Erc91, HCF91, MP91]. similarity [RL93, Uhl91]. Simon [Han93]. Simple [Cho96, CKN^{+}95, Grä90, Joh99, KST96, Nie93, PK93, AS90, AT99a, AL95, Als96, AM90, BW98, BNN91, Bün90, CDF96, CD93b, CH97, Che92c, CG93, Che94a, CD94a, Cou97, CKU98, DS93a, Gál97, GKL96, GJLR96, GH92, Gon95b, GS96b, GS91, Har98, Her91b, HP96, JH93b, Kei91, KH96, Kus97, LRH^{+}95, MSW93, Mas91, NBY99, NO96, Pac92a, Pac92b, PC92, Pri95, PN94, RST91, SCGD92b, Sch91, Sei96, Sky91, SYUK98, VA97a, Vid91, WCC91, Woe90, YS96, YY93b, Zha97, dDDS97, Che95c]. Simpler [Sug94, AR99, Cha95b]. simplex [BDF^{+}95]. Simplified [FNTV90, DKVW95, Har94, Har96, KMTd94, Pio90, RRS98]. simplifying [Sta96]. simply [Hir97]. simulated [Mog94]. Simulation [Sin98, BŞ94b, CY95, Cho96, Dam99, DGL95, Lin93a, Lin94, Pet94b, QB99, VT93, VCW95, Weg90]. Simulations [SRW95, Sch93c, Ueh95]. simulators [YW96]. simultaneous [ST93]. Single [Cha91, GK98a, LCC93, Per94, AA96, BGP^{+}97, BL91, BR94c, Gaw96, GM92c, HTC91, Khu99, Khu00, dDDS97, VSS96]. single-layer [GM92c]. single-operation [BR94c]. sixteen [BS96b]. Size [BB94b, Bie98, BW98, CB91, CS93a, CS95, FGW98, Fuj95, Gál97, GS96a, HS93b, HL98, IZ94, JM94b, Joh98, LPK90, Mes91, O’C95, RW91, RW93, ST91, TV96, VHV92, Weg93]. Size-depth [BB94b, TV96]. sized [Har98, Tsa93]. sizes [KH96, VA97b]. skew [KS91, Sch97]. skewing [LY92]. skip [Sen91]. skip-lists [Sen91]. slave [HWT99]. SLDNF [MT92a]. SLDNF-tree [MT92a]. sliced [KC96]. slicing [DHS95]. slight [MS99a]. slope [KS93a, Mat91b]. Slow [Pug90]. slowdown [Sch93c]. small [CDF96, CY95, GR98, HW96, IIT92, IITY93, KW92, Par91a, RR99a, VHV93, WCC91, Wan97c]. smaller [CC93, HLV97, Juk95]. smallest [GS98a, HW93a, SK98b, Sky91]. smooth [Lê97, Mnu91]. Smoothness [Sho91b]. Smoothsort [BH91]. snapshot [CW98b, IS98]. snapshots [AB92, AV94, KST96]. Snir [TT94]. SNP [Kan91, Kan94b, ZJ94]. SNP-complete [Kan91, Kan94b]. SNP-hard [ZJ94]. software [LS99b, ML95, Pra96]. solution [Esp92, HP92, KST96, PZC98, RZ92, Sac96, Woe90, Yu95]. solutions [DG97, ETA92, NB96, Ov98, RU93, SAFS95, Yao92]. Solvability [Ott95a]. solvable [DvR94, Gav93b, Low98b]. solve [DS99, HHW94, LL94a, RL94, RAGR94, Sac96]. solved [IL97, Par91b]. Solving [Che95b, HC99a, HTC91, WT92, LWS99, LMS96, Mar98, iNM97, Nie93, Yan91a]. Some [Bas91, BM98, Daw92, DT93b, Dub90, Gar92, HHK91, LP93b, Lut90, Sch95a, Sen91, Wan95, ZJ94, ZÖ94, dBDL96, AERT92, CJ90, FP91b, Hùy91, JR91, Jon98, JP97, KLP96, KS94, LR94, LP91b, Mäk92a, McC98, Mot91, Mow92, PK99, Sat93, Tur91b, TH99, UK93, Vos93, Weg93]. Sooner [Hen92]. sort [Han91, Kor97, Man91, Wie97]. sorted [AC90a, SR96b, YY92]. Sorting [AC90a, Chr96, Vai92, Bag92, CC91, CD92, CH93b, CS91b, FNTV90, HS90, HI90, KP99, Kri93, Kut90, LP91a, Lin96, LK92, Pau94, Raj98, Tod90, WCL90, Yao95]. SOS [FZ98]. sound [YL91]. soundness [VS95b]. source [FHT93, Ueh97]. source-deplanarized [FHT93]. Soviet [Riv93]. Space [FMT93, KT96, Ten90, Tho95, Alf99, AGM92, AC90b, AQS97, BKMT97, CY95, CS93a, CD96, DL92, DMS94, DJS96, Gav93a, Ger94, Gol95b, Gon91, GK98b, Han91, HKW90, Hem94, HS93a, HL91b, IK97, IIT92, IITY93, Jag97, JR91, LCC99, Li94, MS99a, MT98a, MN90, MP95, Pad98, PS94, Pet95, Sax94, Ser91, ST91, Sny91a, Sny91b, Sze98, TTC94, Vil95]. Space-efficient [FMT93, Tho95, DJS96]. Space-optimal [KT96, BKMT97]. spaced [Ito91]. spaces [BU95a, BU95b, DM91, ST91]. span [Coh94]. Spanners [HP97a, HP97b, Cha94, MVR96]. Spanning [GMM94, AR96a, BMI97, BLHDW99, Cha98a, CH95, CL97a, CYH91, CW97a, CM98b, ERC94, HT95, HJL^{+}91, IK93, IKOI99, KS92, KPRS97a, KW98, LC93, Man96, NCK99, NDM99, NR98, PN94, SB93, TT93, WL93, WCL95, YMC98]. SpanP [MTV94]. Sparse [AT99b, GK98b, Man94, AM95, Cha94, FKP93, Sch93b]. spatial [RRS99]. Special [HLC92, ALLL98, Cri91, KC98, Ott95a, Par91b, Pom90, Ram98, Sak91b, Tri92]. specification [LS99b]. specifications [Col94, Sis91, VM95, Vel96]. specified [Kha94a]. spectrum [Mah91]. speed [AA96, BG94, DGL95, Gaw96, HCF91, Hüh93]. speed-up [BG94, DGL95, HCF91, Hüh93]. Speeding [MT93, HM98, Kle94a]. speedup [Dro96, LSZ93]. spin [PS98]. spiral [LL94a]. splaying [Sch93a]. SPLICE [Gra97, HC95]. SPLICE/AS [Gra97, HC95]. Splicing [Pău96]. split [CL93a, Nik95b]. splitgraphs [BK95]. Splitsort [LP91a]. splitting [KLP96, AE98]. Square [ITY96, BMST97, CR91, FY93, Hef91, KC98, LLP90, Xue96]. square-free [Xue96]. square-freeness [CR91]. squares [JS95, Ols97]. SSS* [BB93]. stabbing [dv94]. stability [Cre98]. stabilization [Gho91, Her90]. stabilizes [Tel94]. Stabilizing [GH90, Gou94, HG95, CYH91, CD94b, DKVW95, GG96, HP92, HH92, HC92, HL98, MN98, Yen96]. Stable [HL91b, BNN91, KT90, Mos97, Sac96]. stack [vB99, SCGD91]. stacks [AS99, BA95, CY92, Ken91]. staircase [MK92, Yan93b]. stamps [Gup92a, KST96]. standard [Blu96]. Stanica [YT96]. star [CC95b, DT92, DT93a, DV98b, GVV93, GP95b, GP97, GP98b, HI97, JM94c, Lat93, MP96, NB96, NR98, Ola90a, QMA91, RS93, SHC^{+}93, SLC93, SD91, YMC98, YP98, dBDL96, GOPR92]. star-cutset [Ola90a]. star-free [MP96]. start [IOU94, Mos90]. starting [CD98b]. Starvation [KK98, Tur91b]. Starvation-free [KK98]. state [Bir92, Bir96, Blu96, CNV92, HL98, IC98, Kob92, Kuč99a, PS98, Ros93, TM98, Yod91, YL99]. state-oriented [Ros93]. Statecharts [Hd91]. statements [OH91]. States [Sas91, UK93]. Static [CCH93b, Fer94, PG95a, She93, DHS95]. statistical [SJKL95]. statistics [ML95, MW97]. Steiner [AR92, Flo91, Lia95b, LX99, Meh88, Ov98, OS95, Rav94, Sal92a, Sny91a, Sny91b, TT93, Vos93, WY95, Zel93]. step [Cha91, HTC91]. stereo [Chu97b, DG97]. still [Hal93c]. Stochastic [SR96a, RS99, Raj94, YS96]. Stoer [AM99]. Stop [Wis93, Shv92]. Stop-and-copy [Wis93]. stopping [YS96]. Storage [MSK96, KMP90, MSV96]. Storage-efficient [MSK96]. storages [Hüh93]. Strahler [DK94a, DK95, Kru99]. straight [CS97a, CKU98]. straight-line [CS97a]. straightforward [Tip90]. strategies [Bv95, CL96, DGL95, Pug90, YS96]. strategy [DS98, Gol92, IN90, LC99b, MI94, Ped94]. stratified [CCH93b, MN93a]. Straubing [Cau96]. Stream [Gol99, KLK97, PR91]. strength [CC94b]. strict [PS95a, PS95b, SZZ94, YT96, Cus96, CS96, O’C94]. strictly [LP97a]. strictly-second [LP97a]. Strictness [Lau90b, Kam92]. string [Aku95, BYF96, BYP96, Bre92, BCT93, Bre94, CD96, Der95, GBY90, Han93, Mae90, MBY91, MS94, MS95, NBY99, Ott95a, PAD99, RUG97, STKK97, Tuz90, dB93]. string-matching [dB93]. string-rewriting [Ott95a, STKK97]. string-to-string [Mae90]. Strings [Ver92, AS97, All94, AFI91, BR98, BC95, CR91, GP96, Mar90, Zim97]. strip [CW97b]. Strong [Reh96, DKL99, Ist97, RS94b, Sze98]. strong-weak [DKL99]. Strongly [HLM98, DS94b, Jia93, JT94, KA97a, NSS94, Pas92b, PC92]. Structural [AHR99, Mäk92b, Mäk92a, Reh96]. structurally [How91]. structure [AR91, AR99, BG91, CH92a, CD93b, CD96, FLSS93a, FLSS93b, Her91b, KKM99, Kle94a, Lit95, MSK96, MM99, OV96]. structured [BS90, Miz99]. structures [AE90, MCR98, MSV96, Nep99, Rac93, Sin94a, Ust96, Ust98, Wie97]. structuring [CMS90]. Stubblebine [HLL^{+}95]. study [Ber96, CTVC98, FGW98, GP98a, HK96b]. Stutter [PW97]. Stutter-invariant [PW97]. sub [Pri95, Pri97]. sub-quadratic [Pri95, Pri97]. subclass [Blu92, VKM92]. subcube [Lat91]. subdivisions [Ove92]. subexponential [Bsh96]. subgraph [DMM98, Epp94, HR94, KLP98, Man96, PN94, SHC^{+}93, TB97, Yan95]. subgraphs [AAH99, HLC92]. subhypergraphs [Mil92a]. Subject [Ano96c, BBDD93]. subliminal [YL97a]. sublinear [LR94]. submodular [NI98]. subnetwork [FKP93]. subprogram [Pra96]. subsequence [LLM97]. subsequences [vB99]. subset [Kol94, Pri95, WY92b]. subsets [AKK97, RWZW91]. substitution [Oht93, Ola90b, RF93]. substring [MW97]. subtree [DIR99, Dub90, Gro91a, Gro91b, Mäk89, SW93b]. subtrees [Dev98, GS98b]. subtyping [But95, Reh96]. Subword [TS95b]. Successful [BLP98]. successfully [Al 96]. Succinct [BL96, Maz95, SH94]. suffice [Zha91]. sufficiency [SS98b, SD91]. sufficient [CY96, DV91, KT90]. Suffix [GLS92, CL97b, CR90]. suite [CL96]. sum [BNHK99, KL90, Kol94, Woe91, WY92b]. sum-free [Kol94]. summation [Hag95, MRR96]. sums [FP98]. supercube [Yua91]. supercubes [SGB92]. superimposed [DML98]. superprimitivity [AFI91, Bre92, Bre94]. superstring [Gus94]. superstrings [JLD92]. superterse [Wan97c]. supplementary [MN93a]. Supporting [MZ93]. surface [BS98, MV91]. surfaces [KB90]. Surjective [Sat98]. swaps [ALLL98]. sweep [KAL93]. sweepline [HHW94]. switched [Lin93a]. switching [CL92b, Kha94a]. Symbolic [Nep99, SGD98]. Symmetric [GMR90, GHS93, Hwa93, San96, Ste91, TNB96, Wag98, AR99]. symmetric-difference [Wag98]. symmetries [Cla92]. Symmetry [LM93, HR98a]. synchronisations [UH95]. synchronization [BU91, BF96b, KHJ98, KA97c, Mis91, Mis92, PK90, RW97]. synchronized [MK94, Slo90]. synchronous [BT90, WH94]. Syntactic [YCH93, HLV92, Wan98]. synthesis [KTS96, Mak90a, Sid90]. synthetic [SJKL95]. System [Qia90, BBDD93, Bil93, BD96, Che95c, Cus95, HI97, Hwa92a, HC95, LC96, RW97, She93, WCL90, WC90]. Systematic [MW91, SO97, Sch93a]. systems [Ahu90, ABK98, AW98, AS95, Bar91, BH99, BV92, BD96, CG91, CY98, CB91, CM96, CK93b, DB95, DK94b, DS93b, DS94b, Fer95, Fuj98, Gär93, Gho91, Gre95, HLM98, Hen92, Hof94, HM95, Kar90, KG99, KR95, KK90, LK95, LC97b, LY99, LCL93, LS90, LvBDW92, MCR98, MK96, Nas95, NO96, Ohl95, Ott95a, Pău96, Pen92, Ram90, RS95b, SY90, SO93, SV99, SDB97, STKK97, Sin92a, SM95, SI99, TL93, TM93, Tsa90, TIK99, Tuz90, WS96, WJ95, Zha91]. Systolic [Meg93, AO92, LLM97, MG90, MR95, MW97, Tsa93]. Szilard [Mak90b]. Szymanski [Lub90].

T [YT96]. table [Hor95]. tableaux

[WC92a]. tables [FRS94, WCC91]. tailored [Tho95]. tale [RS98]. tally [Sch93b]. tamper [Zhe95]. tamper-proof [Zhe95]. tamperfree [Hwa92b]. Tanaka [Hwa92a]. tape [LO94, Sze96]. tardiness [FF90]. tardinesses [Woe91]. target [FKP93]. targets [Pău96]. Tarjan [Cha98a]. task [AB98, Bil93, BD96, DGS97, HNC92, KG99, ST99, SC90]. Tasks [Fre83, Amo97, BP95a, BGP^{+}97, BBDD93, BDDS92, BDDS94, CD98b, DB95, Dro96, HLW93, Kim94b, MP97, NCC98, SY93, WC91]. Technique [AP93, CLBW90, CST92, GZ91, GS96a, Göb99, GJS97, PAD99, Pro94, WC92a]. techniques [Ali96, Cho96, CMM95, ESW96, Hem94, WS94]. technology [Riv93]. template [FLSS93a, FLSS93b, HCF91]. temporal [Bir96, Cal94, CL93b, KKK99, PW97, Pen92, RPG98]. Term [Sun96, Ban96, Bas94, CGL98, DS93b, DS94b, HLM98, KK90, Kus97, Ohl95, Zha91]. terminal [DS99, Kar90]. terminating [CS95, EST99, Mäk97]. Termination [Ohl95, Gia94, GP90b, KK90, MK94, RS90, STKK97, Zha91]. terms [BF97, Fok97, GM94a, Rit94, Rit95]. ternary [Rif98]. tessellations [MS90b]. Test [hC98, Bre92, CL96, CR91, PR93, RU93, San90]. test&set [OM97]. Testing [Bre94, CG93, AFI91, ALLS96, BN96, BU91, Cha90, CLBW90, CY96, Dam98, DV98a, ML95, OS95, Zip97]. tests [TY94]. text [DS96a, FPS96]. texts [CL97b, CD96]. th [KKK97]. th-order [KKK97]. their [AM95, HK92, Juk95, SSCK94]. theorem [BG91, Cau96, Hag92, MG92, MR96b, Nau94, RE94, Sug94, XU99a, Zwi91, VM95]. theorems [SW93b]. theoretic [Bac97, CCK93, Kum95, Mei90, Pal90]. theoretical [van97]. theories [Grä90, Yua94]. theory [Cha98b, Das96, Goe90, Lip96, RS97, AW98]. There [Nys96, Woe98]. Thérien [Cau96]. thinning [MP91]. those [IZ94]. three [BP98, BDDS92, BDDS94, BNN91, CM94, DK91, ESW96, FW92, JS95, KP98b, MT98a, TS95a]. three-dimensional [ESW96, KP98b, MT98a]. three-layer [TS95a]. three-valued [BP98, BNN91]. threshold [Blu95, BM97, Dun90, DZ98, Gol97, HHK91, Juk95, Koc89, MPSV99, PS95a, PS95b, RS94a, RW91, RW93, SR98, SS94b, TJ99, Weg93]. thresholds [GR98, Juk95]. throughput [KK92a]. Thue [BM92, TS95b, Yu95, dM94]. tidy [FPT97]. Tight [Alt90, BCT93, CL97c, GH94, LC99b, LK96, AL90a, BMST97, Oya95, Sch97, SYUK98, TU99]. Tighter [LW95a, KS91]. Time [BGOS95, Fre83, Ger94, HI90, KLR97, RW97, TC99, AERT92, Ahu93, Ahu96, ABMP91, AC90b, AS95, BMC96, BHW91, BR94c, Boo94, BDF^{+}95, BD96, CM98a, Cai99, CDF96, CT97, CY95, CT91, Che92a, CL93b, CH93b, CW97a, CW98b, CD98b, CJKJ98, CP95, CSS92, CD96, COS95, CKN^{+}95, Dam92, DL92, DR94, DP97b, DKD97, Dro96, FPT97, FS97, FNTV90, GP90a, Gav93a, GH94, GGH94, GW96, GK98b, GS91, Gup92a, HL92, HKW90, HMP95, Hir96, HSZF93, HLW93, HWJ96, HK97b, HAM11, HT91, HL91b, IK97, IS98, IY98, IL97, Jef91, Kao95, Kim94b, KKK99, KS93b, KST96, Kol94, LWS99, Lar90, LR94, LY90, LR96, LL99b, Liś93, LCL93, Low98b, LT97, MS98, MC92, MSW93, Man90, Man96, MN90, Mid93, Mog94, MN93b, iNRN97, Nik95b, OCSF97, Oya95, Pan96b, Par95]. time [Par91b, PKP93, PW97, PC92, PWL94, PS95a, PS95b, PPK98, PR93, PZC98, Pre91, Pre93, PP90, RR92, Rub90, SO93, SZ92a, Sch95a, Sen90, ST99, Sho91a, SW93a, Sin94b, SY93, SR98, Tel94, Tsa92, VCW95, VKM92, VA97b, WCL90, War98, Wei94, Woe95, Yan91a, Yao95, YJ93, Yen91, YW96, YY93a, Zhu92, Zuc91, Can96]. Time- [BGOS95]. time-dependent [HLW93, Woe95]. Time-space [Ger94, HKW90]. time-stamps [Gup92a, KST96]. Timed [Her98, Bér95]. timeout [Hof94]. timeouts [AF95]. Times [Fre83, BHWG99, CD98b, GP95a, HLW93, LLP90, Ple90, SY93, Woe95]. timestamp [KLR97, AvSW95]. timestamps [Fid98]. Toeplitz [Gem92]. token [DK94b, FLMSS98, Her91a, HHS99, SLCH98, YL97b]. token-based [FLMSS98]. tolerance [AHR99, JM94c, Shv91]. tolerant [AD96, AO92, Cai99, CC92, Chi98, yCGlW99a, yCGlW99b, GVV93, GP95b, HHS99, KB94, MS92, SV94, She91b, SLCH98, VCS96, WLK92, Wan97b, WHS97]. tool [RR93, SGD98]. top [FG93, Gol97, KS91, SB95, Sch97]. top-down [FG93, KS91, SB95, Sch97]. Topological [Ham95, Yua91, MSNR96, Tod90, dBDL96]. topologies [MS92]. topology [Boz95]. topology-independent [Boz95]. tori [MYL99]. toroidal [JT92, KP98b, PSSV93]. torus [CST94, MNT92, MR96b, PT92]. Total [KS97b, dFMdM99, GT96, PT92, Ray95, Sac96, TT90]. Total-chromatic [dFMdM99]. total-jogs [TT90]. tour [AHH93, HR90]. tournament [BJ91]. Towers [Min91, WC92b, WC93]. trace [BFSV99, Gre98, Lin93a, Pen92]. trace-driven [Lin93a]. traceable [TJ99]. traces [GMPW98, Kwi90]. tracing [KB90]. tracking [WLF94]. tractability [Cai96]. tractable [DE92, Sch95a]. trade [TV96]. trade-off [TV96]. tradeoff [CK93a]. tradeoffs [BB94b, Ger94]. trading [DS98]. trail [CT93]. train [AL90a]. Training [TS95c, Höf93]. Trajectory [Li94]. transaction [CM96, EV97]. transaction-based [EV97]. Transactions [Qia90, Jac92, RE94]. transducers [FG93]. transform [KC90, Kha94b]. transformation [Che91b, KS97b, LvBDW92, MN93a, MN98, War98]. Transformational [AS95]. transformations [FV90, Ger97]. transformer [Kna90]. transformers [Dij95b, Mor97, Nau94, de 92]. Transforming [BCDM97]. transforms [Lu93]. transients [GS90]. Transition [MCR98, LvBDW92, Mac95, SG92]. transitions [FGW98, Yen91]. Transitive [Ban96, DP97a, Fv97, JT94, KB95, NMS97b, Nuu94]. translation [Hef91, Rot91]. transmission [Boz95, XSR98, YWC99]. transmitting [JT94]. transversal [BNR96]. trapezoid [CW97a, Lia94a, Lia95b, iNM97]. Traveling [FP98, BD98, HR98b, Par91b, PPK98, DvR94]. traveling-salesman [Par91b]. travelling [DW96, Vis92]. traversal [CD92, HS93a, HW93b, van91b]. traversals [AC90b, CDA91, GS92a]. traverse [Zuc91]. Traversing [Jia91b, Jia91a]. treaps [Wei94, Wei95]. Tree [DKL99, GMM94, MVR96, Alb94, ARS98, AC90b, AT95, AR96a, AHH93, Asp95, BS90, Bop94, BFL^{+}95, CRC98, Che91a, Che92c, CL97b, DL92, DBT90, DR94, ERC94, FGN97, Fuj95, FV90, FG93, GS92a, GM92a, GZ91, HT95, HS93a, HJL^{+}91, Hüh93, HHS99, IP95, IK93, Joh98, Kai95, KR92a, KPRS97a, Kor94, KW98, LC93, LX99, LP91b, MT92a, MS90a, MSV96, MR95, Mor94, PM94, PRS96, Ray95, Sal96, SM92, SS92b, Ski96, SW93b, SB93, TT92a, THS97, TT93, Vaj98, WCL95, Wen94, WT97, WCT99, Yan98, YLW91, Zel93, van91a, van91b]. tree-contraction [PRS96]. tree-like [BFL^{+}95, IP95, Joh98]. tree-structured [BS90]. trees [ASS97, AS92a, AW93, BY91, BMI97, BR94a, BLHDW99, Bar96, Bas91, BS94a, BB94a, Blu92, Cai93, CW92a, Cha98a, Cha90, CH95, CL97a, CHL97, CYH91, CDA91, CD92, CN94, CC95a, CW97a, DCD96, DD93, DP95, DPP96, DK94a, DK95, Dev98, Dub90, EKR96, FPT95, GM98, Gro91a, Gro91b, HT98, HC92, IKOI99, Jer94, KLTT97, Kem99, KS92, KRY94, Kim97, Kir94, Kor93, KL98, KL99, Kru99, KPRS97b, LR91, Li99, Mah98, Mäk89, Man90, Mar91, MAUY98, NCK99, NDM99, NR98, NU94, Pal90, PWL94, Raj92, RG93, RLDL94, SS94a, SS95a, Sal92a, Sch92b, SKKC98, Sim91, Sny91a, Sny91b, Sym96, Tre96, Tri92, Uhl91, Vas95, Vas97, Ver92, WL93, Wei94, Wei95, YMC98, ZSS92, ZJ94, ZNN95]. triangle [KM90, Ola90b]. triangle-free [Ola90b]. triangles [HL96, Ram98]. triangularisation [MG90]. triangulated [RR99b]. triangulation [AHR99, Lam97, LK96, MR92, SI95, WCY99, Yan95]. triangulations [Dil90, Lau90a, LL96, Sug94]. triclique [VR96]. triconnected [Che94b]. tridiagonal [CY98]. tridimensional [GJT98, PS98]. trie [Mur90]. tries [AN93]. triggered [BD96]. Trigonometric [Sch91]. triple [KMW98, Ov98]. triple-free [KMW98]. Trivalent [VS95a, CP97, THHC99, VS96, WM96]. true [Pet91, VT90]. truncation [WC92a]. truth [ALLS96, Che92b]. TSP [SW90, dlVK99]. Turing [GGH94, Hüh93, IIT92, IITY93, LO94, MP95, MP98b, Pet95, Sze96]. turn [BU98]. Turpin [DS98]. Twenty [BTv93]. twice [PR93]. twisted [Che95a, CL95, KB95]. Two [AF92, BR94b, CW99, CGR93, GP96, HWT99, HLL^{+}95, Ira91, KP94, Sax94, WC90, ABK98, ÀG91b, AKK97, BYR93, BŞ94b, BCS97, BBGT96, BTW98, CXZ99, CNV92, CH95, Che95c, CL93c, CL97b, DS99, DK91, EYK95, FS92, GT96, GS96b, Her93, HTC91, HHW94, Jag97, KK95a, Kle94b, LP94a, Lan91, LHW99, Lin96, LYC99, Loe97, MV91, Mes91, Mor94, MK96, NMS97b, NPW98, PMC97, Pet94a, Pet95, RS96, Rit94, Rit95, Ros98, Ros99, Rot91, Sal96, SR96a, SRW95, Sri99, TT91, iT96, Tom95, Ust96, VT93, Vis92, Wan92b, WHL99, Wan99a, Woe90, Woe98, dDDS97]. two-center [Her93, HHW94]. Two-coloring [Sax94]. Two-Dimensional [WC90, AF92, CGR93, BYR93, BCS97, CL93c, CL97b, Jag97, NMS97b, Pet95, VT93, Woe98]. two-headed [CNV92, MK96]. two-independent [HTC91]. two-level [Ust96]. two-local [FS92]. Two-machine [CW99]. two-parameter [CH95]. two-point [GS96b, Rot91]. two-processor [RS96]. two-server [EYK95, Kle94b]. two-terminal [DS99]. Two-way [KP94, BŞ94b, LP94a, Lin96, Mor94, Pet94a, Sal96, Wan92b]. type [PS92, VS95b, Vol96]. types [CDK93a, Ger94, Nau94, PS92, Pal93, Slo95, de 92].

UIO [CLBW90]. Ulam [LS95a]. Unambiguous [Kim96a, Kos99, RR92]. unary [Loe97, Maz95, Wei91]. unbounded [Dam99, Dan96, KST96, Sin92a, Sin98]. uncapacitated [ACS99]. unconditionally [HW98]. Undecidability [AP99, KR95, Pen92]. undecidable [Bea91, BLP98, Cha98b, FV90, Ott95a, VR99, Zha92]. underlying [HP97b]. underspecification [AG97]. understanding [Sin92a]. undirected [BS99, Cou97, KS90, Sub95a, Sub95b]. unequal [DCD96]. Unfold [Ger97]. Unfold/fold [Ger97]. uni [BC93]. uni-connectivity [BC93]. unicity [Al 96]. Unidirectional [DT93a, HL98]. Unification [BL90, Fok97, Vas95, Baa98, Rit94, Rit95]. unified [CDA91, CDK93a, CS91b, Fuj96, Hir96].

Uniform [ARS98, AS94, Lib98, LPP91, PCK95, BBGT96, Bsh97, DCD96, Her97, KB92, Ker92, KH96, MSV96, Rad92, WH94]. uniformly [Fer95, Mai94b, Pug90]. Unifying [MM96, Mur90]. union [Bir92, BR94c, Cai93, DPS93, GI91, Riv93]. union-find [BR94c]. unique [CC92, Fer95, Pre93, Sac96, Héb95]. unique-path [CC92]. uniquely [EST99, Mäk97]. unison [GH90]. Unit [Fre83, CKN^{+}95, LYC99, SY93]. Unit-Time [Fre83, SY93]. Unity [UK93, Dij95b, Pac92a, Pac92b, RF93, CH99]. Universal [HW93b, Mul91, BCDM94, Göb99, GR97, LV92a, Pău96, YS96]. universality [Mor92]. unknown [CS95, Hes91, Lar90, MP97, MI94, YL91]. unlabeled [DDK^{+}98, LC94b]. unless [Wan97c]. unordered [ZSS92, ZJ94]. unreliable [EGL96, Pel91, Sin92b, YWC99]. unrestricted [DG97]. unsatisfiability [Ver99]. unsuccessful [HM96]. Untiming [Bér95]. unused [HH94]. update [AvSW95, AM97, BA95, DR94, FS97, Ira91, LP94b, Ped94, RW96, SM94, Tei93]. Updating [KLR97, IS98, Mes91]. Upper [Raj94, Al 96, Bas91, DZ98, Hec91, LPK90, O’C94, Oya95, SYUK98, Tak92, TT91]. Upward [RRW94, RRW00, DLR90, Kim97, Tre96]. use [Dij95b, Mat90, SCGD92a]. used [YLW91, van97]. user [DF91]. Using [All94, BJ97, CD93a, GT96, Hal93a, SJKL95, AR96b, Ber94, Bsh96, CCH93a, CL99, CLBW90, DJM94, Esp92, GH92, Gup92a, Hor95, HV96, JS95, KFYA94, KS98a, Kha94b, KS93b, KB94, LPK90, Lee96, LC94b, MG90, Nas98, Pal90, PCK95, Pel91, Ros98, Tho95, Tsa93, VHV93, WT92, XL93, YCH93]. utilisation [BD96]. Utility [DCD95]. utilization [Li90, MSV96].

value [KK92a, PH92, Sal92b]. valued [BF97, BP98, BNN91, HS93b, Val95b]. values [Vel94, YL91]. Vandermonde [Pre91]. Vapnik [Sim91]. variable [AH94, AA96, BB93, BLW96, Har98, MZ93, Ohl95, Ser98, VHV92, Vid91]. variable-length [AH94]. variable-preserving [Ohl95]. variable-sized [Har98]. variables [BGM90, GLR99, JM94a, KR95, LV92b, Sin92a, VS95b]. variants [HM99]. variation [BGKL98, CC93]. varieties [Pin95]. various [CDK93a, Ger94]. varying [GP95a]. VAS [Yen95]. VC [NS96a, NS96b]. VC-dimension [NS96a, NS96b]. vector [Ahu93, Ahu96, Fid98, GMR90, Man94, Pad98, SK92a, SS92d, SS92c, Woe98]. vectors [GMSS99, Hen97]. Vegas [Wei91, LSZ93]. Verbs [Chu99]. verification [BK98, But95, CS93a, FM97b, HHL95, KPRS97a, Nep99]. verify [IN90]. verifying [KS93b]. versa [KA97b]. versatile [HHS99, PAD99]. version [BKMT97, CLW98, Gra97, TT94]. versions [HK97b]. versus [Bar90, CTVC98, Cre98, DS96a, Gav93a, PS92, Zwi96]. Vertex [KMW98, ACJ96, BFR98, BJ92, ECOUX95, Fuj96, Gon95b, HH94, JB99, Kin92, KK95b, KB95, Lia94b, LS95b, LT97, Luc98b, MM92, Par91b, Saa95, Woe92b, ZNN95]. vertex-rankings [ZNN95]. vertex-search [MM92]. vertex-transitive [KB95]. vertices [Che92c, FW92, HWJ96, HAM11, KCLA99, MM92, Rhe98, SK92b]. Very [NBY99, BW98, Vid91]. Veto [OK96]. via [Cho96, GW96, KS93a, RP91, Reh96]. vice [KA97b]. Viennot [Mel96]. view [AAA^{+}91, CD98a, DW90]. views [CL99, DS96b]. violation [RW94]. Virginity [LS99b]. virtual [VCW95]. visibility [AL95, CH91, CD98a, DTT92, HLC92, KLTT97, KPC^{+}95, SH93, Wis92]. visual [BLP98]. Vital [VSS96, HJL^{+}91, IK93, LC93]. Vizing [MG92]. VLSI [BGOS95, PP94, SDG92, YP98].

VLSI-optimal [BGOS95]. Volume [Ano96a]. Volumes [Ano96b, Ano96c]. Voronoı [HKK92, CHT90, HP90, Küh98, MS90b, Tip90]. vote [OSK96]. voting [DKK^{+}94]. vs [AS95, Ber96, Mei90, New91, Pug90, TS96].

W [YT96]. Wagner [AM99]. wait

[LV92b, LKCNW91]. wait-for [LKCNW91]. wait-free [LV92b]. wakeup [Val95b]. walks [BHWG99, SSO94]. warning [Jon95]. warp [YW96]. warped [TW95]. wavelet [Lu93]. way [AW93, BŞ94b, Bha99, Hwa92a, KP94, Kut90, LP94a, Lin96, LM93, Mar91, Mor92, Mor94, Pet94a, RS97, RST91, Sal96, Slo90, Sze98, Wan92b]. Weak [Sze98, BM97, DKL99, HSZF93, YL97b]. Weakest [PR98, SZ92a, Val95a]. weakly [LL94a]. Weber [LS99a]. weight [CW92b, Gon95b, He91, HTC92, LL96, LK96, MM92, MR92, WCY99, Yan95, YTC93]. Weighted [AFB95, CNP98, MPSV99, Ama93, AR92, CL93a, CJKJ98, FF90, Fuj98, GK98a, GZ97, HJL^{+}91, HLC92, LWS99, LY90, LS95b, LT97, Wei91, YL90]. weights [Kir94, MS99b, Tri92]. well [BR98, GP96, HPS93, Jan99]. well-characterized [HPS93]. well-formed [BR98, GP96]. where [Mas91]. whether [MN92, OS95]. Which [HLV92, Gav93b, KKM99, VS93, Wie97, YL91]. WIDE [HC95]. Widest [Che96, SSC98]. width [BTW98, CDF96, IZ94, Kim97, Kin92, Yan98]. Wilke [Pin95]. winner [Jur98]. winning [Bv95]. wires [BL91]. wiring [CDK93b]. within [IL97]. without [FMT93, Hor95, KK95b, PW97, Shv92, Sid90]. word [Alt90, BM92, Dro95, KR95, Ott95a, TS95b, VHV92, dM94, de 95]. word-size [VHV92]. words [Bea91, Car94, CMR98, Gol95b, JP97, Mai94b, Mel96, Pir94, SM90, SMDS94, Sta96]. working [PMC97]. Workload [TL95]. world [CS90a]. worlds [GP90b]. worm [Sym99, Mog94]. WORM-2DPDAs [Mog94]. wormhole [KK94, LL94b]. Worst [LS99a, Pau90, Vos93, BR94c, BDF^{+}95, CY92, GHY93, KT90, LV92a, PS94, SYUK98]. Worst-case [LS99a, Pau90, Vos93, BR94c, BDF^{+}95, KT90, LV92a, SYUK98]. WPRAM [Nas98]. wreath [GMPW98]. write [HSC93, Shv92]. Write-All [Shv92]. writer [Vid91].

X [CDF90].

Yaniv [DS98]. Yield [Sal96]. Yield-languages [Sal96].

Zemlyachenko [DIR99]. Zero [Zip97, BF97, BGKL98, Bur92, DCGP99, DP94, DF91]. zero-knowledge [Bur92, DCGP99, DP94, DF91]. zero-one [BGKL98]. zero-valued [BF97]. zigzags [HHI97].

References

Alidaee:1996:SSP

[AA96] Bahram Alidaee and Ahmad Ahmadian. Scheduling on a single processor with variable speed. Information Processing Letters, 60(4):189–193, November 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Abadi:1991:PLC

[AAA^{+}91] M. Abadi, B. Alpern, K. R. Apt, N. Francez, S. Katz, L. Lamport, and F. B. Schneider. Preserving liveness: comments on “Safety and liveness from a methodological point of view” [Inform. Process. Lett. 36(1), 1 October 1990, pp. 25–30]. Information Processing Letters, 40(3):

141–142, November 8, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [DW90, DW91].

Aichholzer:1999:NRM

[AAH99] Oswin Aichholzer, Franz Aurenhammer, and Reinhard Hainz. New results on MWT subgraphs. Information Processing Letters, 69(5):215–219, March 12, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/21/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/21/17/article.
pdf.

Acharya:1992:RDS

[AB92] Arup Acharya and B. R. Badrinath. Recording distributed snapshots based on causal order of message delivery. Information Processing Letters, 44 (6):317–321, December 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Audsley:1998:FPS

[AB98] N. C. Audsley and A. Burns. On fixed priority scheduling, offsets and co-prime task periods. Information Processing Letters, 67 (2):65–69, July 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Azar:1994:PAN

[ABF94] Y. Azar, A. Z. Broder, and A. M. Frieze. On the problem of approximating the number of bases of a matroid. Information Processing Letters, 50(1):9– 11, April 8, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alon:1996:MNB

[ABF96] Noga Alon, Phillip G. Bradford, and Rudolf Fleischer. Matching nuts and bolts faster. Information Processing Letters, 59(3): 123–127, August 12, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ait-Bouziad:1998:IAR

[ABK98] Ahmed Ait-Bouziad and Helen Kassel. An improved algorithm for retrieving fuzzy information from two systems. Information Processing Letters, 65(2):63–66, January 29, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amir:1999:RSM

[ABL99] Amihood Amir, Ayelet Butman, and Moshe Lewenstein. Real scaled matching. Information Processing Letters, 70(4): 185–190, May 28, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/20/
20/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/20/20/article.
pdf.

Alt:1991:CMC

[ABMP91] H. Alt, N. Blum, K. Mehlhorn, and M. Paul. Computing a maximum cardinality matching in a bipartite graph in time O(n^{1.5}\sqrt{m}/logn). Information Processing Letters, 37(4):237– 240, February 28, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Altman:1990:SRS

[AC90a] Tom Altman and Bogdan S. Chlebus. Sorting roughly sorted sequences in parallel. Information Processing Letters, 33 (6):297–300, February 10, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Andersson:1990:CTT

[AC90b] Arne Andersson and Svante Carlsson. Construction of a tree from its traversals in optimal time and space. Information Processing Letters, 34(1):21–25, February 22, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Acciaro:1994:CCG

[Acc94] Vincenzo Acciaro. On the complexity of computing Gröbner bases in characteristic 2. Information Processing Letters, 51(6):321–323, September 26, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Afek:1997:BCP

[ACH97] Y. Afek, M. Cohen, and E. Haalman. The bit complexity of the predecessor problem. Information Processing Letters, 63(2): 109–112, August 13, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alstrup:1996:AFI

[ACJ96] Stephen Alstrup, Jens Clausen, and Kristian Jørgensen. An O(|V |∗|E|) algorithm for finding immediate multiple-vertex dominators. Information Processing Letters, 59(1):9–11, July 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aardal:1999:AAK

[ACS99] Karen Aardal, Fabián A. Chudak, and David B. Shmoys. A 3-approximation algorithm for the k-level uncapacitated facility location problem. Information Processing Letters, 72(5–6):161– 167, December 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/60/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/60/25/27/article.
pdf.

Aslam:1996:SCN

[AD96] Javed A. Aslam and Scott E. Decatur. On the sample complexity of noise-tolerant learning. Information Processing Letters, 57 (4):189–195, February 26, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amir:1997:IDA

[AD97] Amihood Amir and Emanuel Dar. An improved deterministic algorithm for generating different many-element random samples. Information Processing Letters, 62(2):95–101, May 21, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Adams:1992:IAB

[Ada92] Carlisle M. Adams. On immunity against Biham and Shamir’s “differential cryptanalysis”. Information Processing Letters, 41 (2):77–80, February 14, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Allison:1991:SPC

[ADY91] L. Allison, T. I. Dix, and C. N. Yee. Shortest path and closure algorithms for banded matrices. Information Processing Letters, 40(6):317–322, December 30, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Agrawal:1990:ELS

[AE90] Divyakant Agrawal and Amr El Abbadi. Exploiting logical structures in replicated databases. Information Processing Letters, 33 (5):255–260, January 10, 1990.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Andersson:1998:BAA

[AE98] Gunnar Andersson and Lars Engebretsen. Better approximation algorithms for Set Splitting and Not-All-Equal Sat. Information Processing Letters, 65(6): 305–311, March 27, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Agrawal:1997:BQG

[AEE97] Divyakant Agrawal, Ömer Eğecioğlu, and Amr El Abbadi. Billiard quorums on the grid. Information Processing Letters, 64 (1):9–16, October 22, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aggarwal:1992:OTB

[AERT92] Alok Aggarwal, Herbert Edelsbrunner, Prahakar Raghavan, and Prasoon Tiwari. Optimal time bounds for some proximity problems in the plane. Information Processing Letters, 42 (1):55–60, April 27, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amir:1992:TDD

[AF92] Amihood Amir and Martin Farach. Two-dimensional dictionary matching. Information Processing Letters, 44(5):233– 239, December 21, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alvarez:1995:EMM

[AF95] Guillermo A. Alvarez and Marcelo O. Fernández. Efficient management of multiple outstanding timeouts. Information Processing Letters, 54(3): 139–145, May 12, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Attiya:1996:LFC

[AF96] Hagit Attiya and Roy Friedman. Limitations of fast consistency conditions for distributed shared memories. Information Processing Letters, 57(5):243–248, March 11, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Agarwala:1995:WSP

[AFB95] Richa Agarwala and David Fernández-Baca. Weighted search in the plane. Information Processing Letters, 54(2): 97–100, April 28, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ancona:1992:OLK

[AFG92] Massimo Ancona, Claudia Fassino, and Vittoria Gianuzzi. Optimization of LR(k) “reduced parsers”. Information Processing Letters, 41(1):13–20, January 21, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Apostolico:1991:OST

[AFI91] Alberto Apostolico, Martin Farach, and Costas S. Iliopoulos. Optimal superprimitivity testing for strings. Information Processing Letters, 39(1):17–20, July 12, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amir:1994:ADP

[AFM94] Amihood Amir, Martin Farach, and S. Muthukrishnan. Alphabet dependence in parameterized matching. Information Processing Letters, 49(3):111–115, February 11, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Allender:1991:RRR

[AG91a] Eric Allender and Vivek Gore.

Rudimentary reductions revisited. Information Processing Letters, 40(2):89–95, October 25, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alvarez:1991:PCT

[ÀG91b] C. Àlvarez and J. Gabarró. The parallel complexity of two problems on concurrency. Information Processing Letters, 38(2): 61–70, April 26, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aaron:1997:FJU

[AG97] Eric Aaron and David Gries.

Formal justification of underspecification for S5. Information Processing Letters, 64(3): 115–121, November 29, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aggarwal:1992:PCC

[Agg92] Alok Aggarwal. Parallel complexity of computing a maximal set of disjoint paths. Information Processing Letters, 41(3): 149–151, March 6, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Abellanas:1996:OP

[AGHP^{+}96] Manuel Abellanas, Jesús García, Gregorio Hernández-Peñalver, Ferrán Hurtado, Oriol Serra, and Jorge Urrutia. Onion polygonizations. Information Processing Letters, 57(3):165–173, February 12, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alt:1992:LBN

[AGM92] Helmut Alt, Viliam Geffert, and Kurt Mehlhorn. A lower bound for the nondeterministic space complexity of context-free recognition. Information Processing Letters, 42(1):25–27, April 27, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Akinari:1994:RNR

[AH94] Nakagawa Akinari and Hiroshi Hagiwara. On the real-number representation with variablelength exponent field. Information Processing Letters, 52(1): 1–6, October 14, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arkin:1993:ATT

[AHH93] Esther M. Arkin, Magnús M.

Halldórsson, and Refael Hassin. Approximating the tree and tour covers of a graph. Information Processing Letters, 47(6): 275–282, October 18, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Abellanas:1999:STD

[AHR99] Manuel Abellanas, Ferran Hurtado, and Pedro A. Ramos. Structural tolerance and Delaunay triangulation. Information Processing Letters, 71(5–6):221– 227, September 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/30/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/30/33/article.
pdf.

Ahuja:1990:FPA

[Ahu90] Mohan Ahuja. Flush primitives for asynchronous distributed systems. Information Processing Letters, 34(1):5–12, February 22, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ahuja:1993:AAP

[Ahu93] Mohan Ahuja. Assertions about past and future in Highways: Global flush broadcast and flush-vector-time. Information Processing Letters, 48(1):21–28, October 29, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [Ahu96].

Ahuja:1996:AAP

[Ahu96] Mohan Ahuja. Erratum: “Assertions about past and future in Highways: Global flush broadcast and flush-vector-time” [Inform. Process. Lett. 48(1), 29 October 1993, pp. 21–28]. Information Processing Letters, 58 (6):325, June 24, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Ahu93].

Aceto:1997:CFB

[AI97] Luca Aceto and Anna Ingólfsdóttir. A characterization of finitary bisimulation. Information Processing Letters, 64(3):127–134, November 29, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aggarwal:1995:ESA

[AJG95] C. C. Aggarwal, N. Jain, and P. Gupta. An efficient selection algorithm on the pyramid. Information Processing Letters, 53(1):37–47, January 13, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arnold:1997:RST

[AKK97] A. Arnold, M. Kanta, and D. Krob. Recognizable subsets of the two letter plactic monoid. Information Processing Letters, 64(2):53–59, November 5, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arvind:1995:RSA

[AKM95] V. Arvind, J. Köbler, and M. Mundhenk. On reductions to sets that avoid EXPSPACE. Information Processing Letters, 56 (2):109–114, October 27, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Akutsu:1995:ASM

[Aku95] Tatsuya Akutsu. Approximate string matching with don’t care characters. Information Processing Letters, 55(5):235–239, September 15, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

AlJabri:1996:UDU

[Al 96] A. Kh. Al Jabri. The unicity distance: An upper bound on the probability of an eavesdropper successfully estimating the secret key. Information Processing Letters, 60(1):43–47, October 14, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aggarwal:1990:TLB

[AL90a] Alok Aggarwal and Tom Leighton.

A tight lower bound for the train reversal problem. Information Processing Letters, 35(6):301– 304, September 15, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Altman:1990:NAC

[AL90b] Tom Altman and George Logothetis. A note on ambiguity in context-free grammars. Information Processing Letters, 35 (3):111–114, July 20, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alsuwaiyel:1995:FAM

[AL95] Muhammad H. Alsuwaiyel and D. T. Lee. Finding an approximate minimum-link visibility path inside a simple polygon. Information Processing Letters, 55(2):75–79, July 21, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Albacea:1992:PAE

[Alb92] Eliezer A. Albacea. A parallel algorithm for edge-coloring of graphs with edge-disjoint cycles. Information Processing Letters, 43(6):309–314, October 19, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Albacea:1994:PAF

[Alb94] Eliezer A. Albacea. Parallel algorithm for finding a core of a tree network. Information Processing Letters, 51(5):223– 226, September 12, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alfonsin:1999:CAR

[Alf99] J. L. Ramírez Alfonsín. Cyclic arrangements and Roudneff’s conjecture in the space. Information Processing Letters, 71(5–6):179–182, September 30, 1999. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/30/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/30/26/article.
pdf.

Alimonti:1996:NLS

[Ali96] Paola Alimonti. New local search approximation techniques for maximum generalized satisfiability problems. Information Processing Letters, 57(3):151– 158, February 12, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Allison:1992:LDC

[All92]  L. Allison.  Lazy dynamic-

programming can be eager. Information Processing Letters, 43(4):207–212, September 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Allison:1994:UHA

[All94] Lloyd Allison. Using Hirschberg’s algorithm to generate random alignments of strings. Information Processing Letters, 51(5): 251–255, September 12, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amir:1998:ESC

[ALLL98] Amihood Amir, Gad M. Landau, Moshe Lewenstein, and Noa Lewenstein. Efficient special cases of pattern matching with swaps. Information Processing Letters, 68(3):125– 132, November 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/ipl/cas_sub/browse/
                                        browse.cgi?year=1998&volume=
                                        68&issue=3&aid=2077.

Aspvall:1996:QTT

[ALLS96] Bengt Aspvall, Christos Levcopoulos, Andrzej Lingas, and Robert Storlind. On 2-QBF truth testing in parallel. Information Processing Letters, 57 (2):89–93, January 29, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alon:1990:GPP

[Alo90] Noga Alon. Generating pseudorandom permutations and maximum flow algorithms. Information Processing Letters, 35(4): 201–204, August 7, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alsuwaiyel:1996:FSH

[Als96] Muhammad H. Alsuwaiyel. Finding a shortest Hamiltonian path inside a simple polygon. Information Processing Letters, 59 (4):207–210, August 26, 1996.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Althofer:1990:TLB

[Alt90] Ingo Althöfer. Tight lower bounds on the length of word chains. Information Processing Letters, 34(5):275–276, May 7, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anderson:1990:SRP

[AM90] Richard J. Anderson and Gary L. Miller. A simple randomized parallel algorithm for list-ranking. Information Processing Letters, 33(5):269–273, January 10, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alon:1995:DST

[AM95] Noga Alon and Yishay Mansour. ε-discrepancy sets and their application for interpolation of sparse polynomials. Information Processing Letters, 54 (6):337–342, June 23, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Albers:1997:RCA

[AM97] Susanne Albers and Michael Mitzenmacher. Revisiting the COUNTER algorithms for list update. Information Processing Letters, 64(3):155–160, November 29, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Althaus:1998:MNF

[AM98] Ernst Althaus and Kurt Mehlhorn. Maximum network flow with floating point arithmetic. Information Processing Letters, 66(3):109–113, May 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arikati:1999:CCS

[AM99] Srinivasa R. Arikati and Kurt Mehlhorn. A correctness certificate for the Stoer–Wagner min-cut algorithm. Information Processing Letters, 70(5): 251–254, June 21, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/21/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/21/25/article.
pdf.

Amato:1993:IPB

[Ama93] Nancy Amato. Improved processor bounds for parallel algorithms for weighted directed graphs. Information Processing Letters, 45(3):147–152, March 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ambainis:1999:NQB

[Amb99] Andris Ambainis. A note on quantum black-box complexity of almost all Boolean functions. Information Processing Letters, 71(1):5–7, July 16, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/17/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/17/18/article.
pdf.

Amoura:1997:NSM

[Amo97] A. K. Amoura. A note on scheduling multiprocessor tasks with precedence constraints on parallel processors. Information Processing Letters, 63(3): 119–122, August 27, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Andersson:1993:IBT

[AN93] Arne Andersson and Stefan Nilsson. Improved behaviour of tries by adaptive branching. Information Processing Letters, 46(6): 295–300, July 26, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:1994:IPC

[Ano94] Anonymous. IPL’s policy on complexity results. Information Processing Letters, 50(6): 287, June 27, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:1995:NPPa

[Ano95a] Anonymous. Note from the publisher. Information Processing Letters, 56(5):i, December 8, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:1995:NPPb

[Ano95b] Anonymous. Note from the Publisher. Information Processing Letters, 56(6):i, December 22, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:1996:AIV

[Ano96a] Anonymous. Author index: Volume 60 (1996). Information Processing Letters, 60(6):322–324, December 23, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:1996:MIV

[Ano96b] Anonymous. Master index volumes 51–60. Information Processing Letters, 60(6):337–351, December 23, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:1996:MSI

[Ano96c] Anonymous. Master subject index: Volumes 51–60. Information Processing Letters, 60 (6):325–336, December 23, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:1999:Ia

[Ano99a] Anonymous. Index. Information Processing Letters, 71(5–6):243– 244, September 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/49/30/
                                        37/abstract.html;   http:

//www.elsevier.nl/gej-ng/
29/28/24/49/30/37/article.
pdf.

Anonymous:1999:Ib

[Ano99b] Anonymous. Index. Information Processing Letters, 71(5–6):245– 246, September 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/30/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/30/38/article.
pdf.

Anonymous:1999:Ic

[Ano99c] Anonymous. Index. Information Processing Letters, 72(5–6):219– 220, December 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/60/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/60/25/35/article.
pdf.

Anonymous:1999:Id

[Ano99d] Anonymous. Index. Information Processing Letters, 72(5–6):221– 222, December 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/60/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/60/25/36/article.
pdf.

Aykanat:1992:FHS

[AO92] C. Aykanat and F. Ozguner. A fault-tolerant hexagonal systolic array. Information Processing Letters, 42(4):187–196, June 19, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Analyti:1993:PAM

[AP93] Anastasia Analyti and Sakti Pramanik. Performance analysis of a main memory multidirectory hashing technique. Information Processing Letters, 45 (4):191–197, March 22, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also comment [Pro94].

Ausiello:1995:LSR

[AP95] Giorgio Ausiello and Marco Protasi. Local search, reducibility and approximability of NPoptimization problems. Information Processing Letters, 54 (2):73–79, April 28, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alur:1999:UPO

[AP99] Rajeev Alur and Doron Peled. Undecidability of partial order logics. Information Processing Letters, 69(3):137–143, February 12, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/47/19/
                                        24/abstract.html;   http:

//www.elsevier.nl/gej-ng/
10/23/20/47/19/24/article.
pdf.

Arratia-Quesada:1997:GHL

[AQS97] A. A. Arratia-Quesada and I. A.

Stewart. Generalized Hex and logical characterizations of polynomial space. Information Processing Letters, 63(3):147–152, August 27, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arikati:1990:LAO

[AR90] Srinivasa Rao Arikati and C. Pandu Rangan. Linear algorithm for optimal path cover problem on interval graphs. Information Processing Letters, 35 (3):149–153, July 20, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aggarwal:1991:DDS

[AR91] Alok Aggarwal and Prabhakar Raghavan. Deferred data structure for the nearest neighbor problem. Information Processing Letters, 40(3):119–122, November 8, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arvind:1992:CDS

[AR92] K. Arvind and C. Pandu Ragan. Connected domination and Steiner set on weighted permutation graphs. Information Processing Letters, 41(4):215–220, March 18, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aranha:1996:EDA

[AR96a] R. F. M. Aranha and C. Pandu Rangan. An efficient distributed algorithm for centering a spanning tree of a biconnected graph. Information Processing Letters, 59(3):145–150, August 12, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Athithan:1996:HNP

[AR96b] G. Athithan and T. Sony Roy. Hyperspherical neighbourhoods and pattern recognition using neural networks. Information Processing Letters, 59(4):223– 228, August 26, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arya:1998:FAA

[AR98] Sunil Arya and H. Ramesh. A 2.5-factor approximation algorithm for the k-MST problem. Information Processing Letters, 65(3):117–118, February 13, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arvind:1999:SMM

[AR99] A. Arvind and C. Pandu Rangan. Symmetric Min–Max heap: A simpler data structure for double-ended priority queue. Information Processing Letters, 69(4):197–199, February 26, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.

nl/gej-ng/29/28/24/47/20/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/20/22/article.
pdf.

Alonso:1993:DM

[ARS93] Laurent Alonso, Edward M.

Reingold, and René Schott. Determining the majority. Information Processing Letters, 47 (5):253–255, October 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alonso:1998:UGS

[ARS98] L. Alonso, J. L. Rémy, and R. Schott. Uniform generation of a Schröder tree. Information Processing Letters, 64(6): 305–308, January 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aggarwal:1990:CLD

[AS90] Alok Aggarwal and Subhash Suri. Computing the longest diagonal of a simple polygon. Information Processing Letters, 35 (1):13–18, June 15, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atkinson:1992:GBT

[AS92a] M. D. Atkinson and J. R. Sack.

Generating binary trees at random. Information Processing Letters, 41(1):21–23, January 21, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aurenhammer:1992:SSL

[AS92b] Franz Aurenhammer and Gerd Stöckl. Searching for segments with largest relative overlap. Information Processing Letters, 41 (2):103–108, February 14, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atkinson:1994:UGF

[AS94] M. D. Atkinson and J.-R. Sack. Uniform generation of forests of restricted height. Information Processing Letters, 50(6): 323–327, June 27, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atkinson:1995:TVR

[AS95] Simon Atkinson and David Scholefield. Transformational vs. reactive refinement in real-time systems. Information Processing Letters, 55(4):201–210, August 25, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Abbasi:1997:AFD

[AS97] Sarmad Abbasi and Anirvan Sengupta. An O(n log n) algorithm for finding dissimilar strings. Information Processing Letters, 62(3):135–139, June 4, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atkinson:1999:PSP

[AS99] M. D. Atkinson and J.-R. Sack. Pop-stacks in parallel. Information Processing Letters, 70 (2):63–67, April 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/18/
19/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/18/19/article.
pdf.

Ashir:1997:CAA

[ASA97] Yaagoub Ashir, Iain A. Stewart, and Aqeel Ahmed. Communication algorithms in k-ary ncube interconnection networks. Information Processing Letters, 61(1):43–48, January 14, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aspvall:1995:MET

[Asp95] Bengt Aspvall. Minimizing elimination tree height can increase fill more than linearly. Information Processing Letters, 56(2): 115–120, October 27, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alstrup:1997:OLD

[ASS97] Stephen Alstrup, Jens Peter Secher, and Maz Spork. Optimal on-line decremental connectivity in trees. Information Processing Letters, 64(4):161–164, December 13, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ang:1995:IT

[AT95] Chuan-Heng Ang and Kok-Phuang Tan. The interval Btree. Information Processing Letters, 53(2):85–89, January 27, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aguilera:1999:SBP

[AT99a] Marcos Kawazoe Aguilera and Sam Toueg. A simple bivalency proof that t-resilient consensus requires t + 1 rounds. Information Processing Letters, 71(3–4): 155–158, August 27, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/28/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/28/32/article.
pdf.

Arvind:1999:SSA

[AT99b] V. Arvind and Jacobo Torán. Sparse sets, approximable sets, and parallel queries to NP. Information Processing Letters, 69(4):181–188, February 26, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/20/
20/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/20/20/article.
pdf.

Alagar:1994:OAD

[AV94] Sridhar Alagar and S. Venkatesan. An optimal algorithm for distributed snapshots with causal message ordering. Information Processing Letters, 50 (6):311–316, June 27, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Albers:1995:CBT

[AvSW95] Susanne Albers, Bernhard von Stengel, and Ralph Werchner. A combined BIT and TIMES-TAMP algorithm for the list update problem. Information Processing Letters, 56(3):135–139, November 10, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atkinson:1993:EWT

[AW93] M. D. Atkinson and Louise Walker. Enumerating k-way trees. Information Processing Letters, 48(2):73–75, November 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amir:1998:OAQ

[AW98] Yair Amir and Avishai Wool.

Optimal availability quorum systems: Theory and practice. Information Processing Letters, 65(5):223–228, March 13, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ahuja:1990:FAP

[AZ90] Mohan Ahuja and Yahui Zhu.

An O(n log n) feasibility algorithm for preemptive scheduling of n independent jobs on a hypercube. Information Processing Letters, 35(1):7–11, June 15, 1990. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Acketa:1991:NLP

[AŽ91] Dragan M. Acketa and Joviša D. Žunić. On the number of linear partitions of the (m, n)-grid. Information Processing Letters, 38 (3):163–168, May 17, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barriga:1995:LUE

[BA95] Luis Barriga and Rassul Ayani. Lazy update: An efficient implementation of LRU stacks. Information Processing Letters, 54 (2):81–84, April 28, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baader:1998:CBU

[Baa98] Franz Baader. On the complexity of Boolean unification. Information Processing Letters, 67 (4):215–220, August 31, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bachmair:1992:ACR

[Bac92] Leo Bachmair. Associativecommutative reduction orderings. Information Processing Letters, 43(1):21–27, August 10, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bach:1997:CNT

[Bac97] Eric Bach. The complexity of number-theoretic constants. Information Processing Letters, 62 (3):145–152, June 4, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Backhouse:1998:PAG

[Bac98] Roland Backhouse. Pair algebras and Galois connections. Information Processing Letters, 67 (4):169–175, August 31, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bagchi:1992:SPE

[Bag92] A. Bagchi. On sorting in the presence of erroneous information. Information Processing Letters, 43(4):213–215, September 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bagchi:1997:RSM

[Bag97] Anindo Bagchi. Route selection with multiple metrics. Information Processing Letters, 64 (4):203–205, December 13, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bainbridge:1992:HMG

[Bai92] John Bainbridge. A heuristic method for generating large random expressions. Information Processing Letters, 44(3):165– 170, November 30, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Banach:1996:TTG

[Ban96] R. Banach. Transitive term graph rewriting. Information Processing Letters, 60(3):109– 114, November 11, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barbosa:1990:BVN

[Bar90] Valmir C. Barbosa. Blocking versus nonblocking interprocess communication. A note on the effect on concurrency. Information Processing Letters, 36(4): 171–175, November 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barsi:1991:MAB

[Bar91] Ferruccio Barsi. Mod m arithmetic in binary systems. Information Processing Letters, 40 (6):303–309, December 30, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barsi:1995:DRC

[Bar95]  Ferruccio Barsi.  Decoding

residue codes. Information Processing Letters, 54(4):213–222, May 26, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barth:1996:OBB

[Bar96]  Dominique Barth.  Optimal

broadcasting in the back to back d-ary trees. Information Processing Letters, 59(2):85–89, July 22, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Basart:1991:SUB

[Bas91] J. M. Basart. Some upper bounds for minimal trees. Information Processing Letters, 38 (4):209–213, May 31, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Basin:1994:TEP

[Bas94]  David A. Basin.  A term

equality problem equivalent to Graph Isomorphism. Information Processing Letters, 51(2): 61–66, July 26, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Basten:1996:BBE

[Bas96] Twan Basten. Branching bisimilarity is an equivalence indeed! Information Processing Letters, 58(3):141–147, May 13, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bax:1993:IEA

[Bax93] Eric T. Bax. Inclusion and exclusion algorithm for the Hamiltonian path problem. Information Processing Letters, 47(4): 203–207, September 27, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bax:1994:ACP

[Bax94] Eric T. Bax. Algorithms to count paths and cycles. Information Processing Letters, 52(5): 249–252, December 9, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bhattacharya:1993:FAS

[BB93] Subir Bhattacharya and A. Bagchi. A faster alternative to SSS* with extension to variable memory. Information Processing Letters, 47(4):209–214, September 27, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bhattacharya:1994:GFM

[BB94a] Subir Bhattacharya and A. Bagchi. A general framework for minimax search in game trees. Information Processing Letters, 52 (6):295–301, December 23, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bonet:1994:SDT

[BB94b] Maria Luisa Bonet and Samuel R. Buss. Size-depth tradeoffs for Boolean formulae. Information Processing Letters, 49(3):151– 155, February 11, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bianco:1993:PSM

[BBDD93] L. Bianco, J. Blazewicz, P. Dell’Olmo, and M. Drozdowski. Preemptive scheduling of multiprocessor tasks on the dedicated processor system subject to minimal lateness. Information Processing Letters, 46(3):109–113, June 11, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blazewicz:1996:SCI

[BBGT96] Jacek Błażewicz, Pascal Bouvry, Frédéric Guinand, and Denis Trystram. Scheduling complete intrees on two uniform processors with communication delays. Information Processing Letters, 58(5):255–263, June 10, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bose:1998:PMP

[BBL98] Prosenjit Bose, Jonathan F.

Buss, and Anna Lubiw. Pattern matching for permutations. Information Processing Letters, 65(5):277–283, March 13, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Biham:1999:BGD

[BBR99] Eli Biham, Dan Boneh, and Omer Reingold. Breaking generalized Diffie–Hellman modulo a composite is no easier than factoring. Information Processing Letters, 70(2): 83–87, April 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/18/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/18/22/article.
pdf.

Buchsbaum:1993:DUC

[BC93] Adam L. Buchsbaum and Martin C. Carlisle. Determining uniconnectivity in directed graphs.

Information Processing Letters, 48(1):9–12, October 29, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baldoni:1994:DAM

[BC94] R. Baldoni and B. Ciciani. Distributed algorithms for multiple entries to a critical section with priority. Information Processing Letters, 50(3):165–172, May 9, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bunke:1995:IAC

[BC95] H. Bunke and J. Csirik. An improved algorithm for computing the edit distance of runlength coded strings. Information Processing Letters, 54(2): 93–96, April 28, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertoni:1995:ELE

[BCBF95] Alberto Bertoni, Nicolò Cesa-Bianchi, and Guido Fiorino. Efficient learning with equivalence queries of conjunctions of modulo functions. Information Processing Letters, 56(1):15–17, October 13, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bilardi:1994:LBA

[BCDM94] Gianfranco Bilardi, Shiva Chaudhuri, Devdatt Dubhashi, and K. Mehlhorn. A lower bound for area-universal graphs. Information Processing Letters, 51 (2):101–105, July 26, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Breslauer:1997:TCM

[BCDM97] Dany Breslauer, Artur Czumaj, Devdatt P. Dubhashi, and Friedhelm Meyer auf der Heide. Transforming comparison model lower bounds to the parallelrandom-access-machine. Information Processing Letters, 62 (2):103–110, May 21, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brogi:1992:IDM

[BCLM92] Antonio Brogi, Anna Ciampolini, Evelina Lamma, and Paolo Mello. The implementation of a distributed model for logic programming based on multipleheaded clauses. Information Processing Letters, 42(6):331– 338, July 24, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brandt:1991:ACH

[BCM91] Jonathan Brandt, Carlos Cabrelli, and Ursula Molter. An algorithm for the computation of the Hutchinson distance. Information Processing Letters, 40(2): 113–117, October 25, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [BCM92].

Brandt:1992:CAC

[BCM92] Jonathan Brandt, Carlos Cabrelli, and Ursula Molter. Corrigendum: “An algorithm for the computation of the Hutchinson distance” [Inform. Process. Lett. 40(2), 25 October 1991, pp. 113–117]. Information Processing Letters, 41(6):347, April 17, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [BCM91].

Belogay:1997:CHD

[BCMS97] E. Belogay, C. Cabrelli, U. Molter, and R. Shonkwiler. Calculating the Hausdorff distance between curves. Information Processing Letters, 64(1):17–22, October 22, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertino:1997:TOT

[BCS97] E. Bertino, B. Catania, and B. Shidlovsky. Towards optimal two-dimensional indexing for constraint databases. Information Processing Letters, 64 (1):1–8, October 22, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Breslauer:1993:TCB

[BCT93] Dany Breslauer, Livio Colussi, and Laura Toniolo. Tight comparison bounds for the string prefix-matching problem. Information Processing Letters, 47 (1):51–57, August 9, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertino:1999:FRN

[BCW99] E. Bertino, B. Catania, and Limsoon Wong. Finitely representable nested relations. Information Processing Letters, 70 (4):165–173, May 28, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/20/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/20/18/article.
pdf.

Bajaj:1990:PNR

[BD90] Chanderjit L. Bajaj and Tamal Dey. Polygon nesting and robustness. Information Processing Letters, 35(1):23–32, June 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Burns:1996:CTP

[BD96] Alan Burns and Robert Davis.

Choosing task periods to minimise system utilisation in time triggered systems. Information Processing Letters, 58(5):223– 229, June 10, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blundo:1997:LBR

[BD97] Carlo Blundo and Alfredo De Santis. Lower bounds for robust secret sharing schemes. Information Processing Letters, 63(6): 317–321, October 8, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Burkard:1998:TSP

[BD98] Rainer E. Burkard and Vladimir G.

Deĭneko. On the traveling salesman problem with a relaxed Monge matrix. Information Processing Letters, 67(5):231– 237, September 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blazewicz:1992:SMT

[BDDS92] J. Blazewicz, P. Dell’Olmo, M. Drozdowski, and M. G. Speranza. Scheduling multiprocessor tasks on three dedicated processors. Information Processing Letters, 41(5):275–280, April 3, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [BDDS94].

Blazewicz:1994:CSM

[BDDS94] J. Błażewicz, P. Dell’Olmo, M. Drozdowski, and M. G. Speranza. Corrigendum: “Scheduling multiprocessor tasks on three dedicated processors” [Inform. Process. Lett. 41 (1992), no. 5, 275–280]. Information Processing Letters, 49(5):269–270, March 11, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [BDDS92].

Broder:1995:WCR

[BDF^{+}95] Andrei Z. Broder, Martin E. Dyer, Alan M. Frieze, Prabhakar Raghavan, and Eli Upfal. The worst-case running time of the random simplex algorithm is exponential in the height. Information Processing Letters, 56(2): 79–81, October 27, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blundo:1998:SSS

[BDV98] Carlo Blundo, Alfredo De Santis, and Ugo Vaccaro. On secret sharing schemes. Information Processing Letters, 65(1):25–32, January 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Buhrman:1999:LBQ

[BdW99] Harry Buhrman and Ronald de Wolf. A lower bound for quantum search of an ordered list. Information Processing Letters, 70(5):205–209, June 21, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/21/
23/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/21/23/article.
pdf.

Beauquier:1991:UPA

[Bea91] Danièle Beauquier. An undecidable problem about rational sets and contour words of polyominoes. Information Processing Letters, 37(5):257–263, March 14, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bern:1993:ACP

[Ber93] Marshall Bern. Approximate closest-point queries in high dimensions. Information Processing Letters, 45(2):95–99, February 26, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berkovich:1994:MIN

[Ber94] Simon Y. Berkovich. Multiprocessor interconnection network using pairwise balanced combinatorial designs. Information Processing Letters, 50(4):217– 222, May 25, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berard:1995:UTL

[Bér95]  Béatrice Bérard.  Untiming

timed languages. Information Processing Letters, 55(3):129– 135, August 11, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bernasconi:1996:SVB

[Ber96] A. Bernasconi. Sensitivity vs. block sensitivity (an averagecase study). Information Processing Letters, 59(3):151–157, August 12, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bernasconi:1999:CBB

[Ber99] A. Bernasconi. On the complexity of balanced Boolean functions. Information Processing Letters, 70(4):157–163, May 28, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/48/20/
                                        17/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/48/20/17/article.
                                        pdf.

Bax:1996:FDS

[BF96a] Eric Bax and Joel Franklin. A finite-difference sieve to count paths and cycles by length. Information Processing Letters, 60 (4):171–176, November 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Buhaceanu:1996:FDA

[BF96b] D. S. Buhăceanu and W. H. J.

Feijen. Formal derivation of an algorithm for distributed phase synchronization. Information Processing Letters, 60(4):207– 213, November 25, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bax:1997:PFM

[BF97] Eric Bax and Joel Franklin. A permanent formula with many zero-valued terms. Information Processing Letters, 63(1):33–39, July 30, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Broder:1995:BAT

[BFL^{+}95] Andrei Z. Broder, Alan Frieze, Carsten Lund, Steven Phillips, and Nick Reingold. Balanced allocations for tree-like inputs. Information Processing Letters, 55(6):329–332, September 29, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Balasubramanian:1998:IFP

[BFR98] R. Balasubramanian, Michael R.

Fellows, and Venkatesh Raman.

An improved fixed-parameter algorithm for vertex cover. Information Processing Letters, 65 (3):163–168, February 13, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barbuti:1999:AIT

[BFSV99] Roberto Barbuti, Nicoletta De Francesco, Antonella Santone, and Gigliola Vaglini. Abstract interpretation of trace semantics for concurrent calculi. Information Processing Letters, 70 (2):69–78, April 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/18/
20/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/18/20/article.
pdf.

Bartha:1991:STM

[BG91] Miklós Bartha and Éva Gombás. A structure theorem for maximum internal matchings in graphs. Information Processing Letters, 40(6):289–294, December 30, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bern:1992:DPD

[BG92] Marshall Bern and John R. Gilbert. Drawing the planar dual. Information Processing Letters, 43(1):7–13, August 10, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Borchers:1994:EQI

[BG94] Al Borchers and Prosenjit Gupta. Extending the quadrangle inequality to speed-up dynamic programming. Information Processing Letters, 49(6): 287–290, March 22, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blass:1998:VZO

[BGKL98] Andreas Blass, Yuri Gurevich, Vladik Kreinovich, and Luc Longpré. A variation on the zero-one law. Information Processing Letters, 67(1):29–30, July 16, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertoni:1990:CPA

[BGM90] A. Bertoni, M. Goldwurm, and P. Massazza. Counting problems and algebraic formal power series in noncommuting variables. Information Processing Letters, 34(3):117–121, April 9, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bokka:1995:TVO

[BGOS95] V. Bokka, H. Gurla, S. Olariu, and J. L. Schwing. Time- and VLSI-optimal convex hull computation on meshes with multiple broadcasting. Information Processing Letters, 56(5): 273–280, December 8, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baruah:1997:FLS

[BGP^{+}97] Sanjoy K. Baruah, Johannes E. Gehrk, C. Greg Plaxton, Ion Stoica, Hussein Abdel-Wahab, and Kevin Jeffay. Fair online scheduling of a dynamic set of tasks on a single resource. Information Processing Letters, 64(1):43–51, October 22, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bron:1991:SR

[BH91] Coenraad Bron and Wim H. Hesselink. Smoothsort revisited. Information Processing Letters, 39(5):269–276, September 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bonabeau:1998:SOM

[BH98] Eric Bonabeau and Florian Hénaux. Self-organizing maps for drawing large graphs. Information Processing Letters, 67 (4):177–184, August 31, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bordihn:1999:HLG

[BH99] H. Bordihn and M. Holzer. On a hierarchy of languages generated by cooperating distributed grammar systems. Information Processing Letters, 69(2): 59–62, January 29, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.

nl/gej-ng/29/28/24/47/18/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/18/18/article.
pdf.

Bhatt:1999:IWP

[Bha99]  P. C. P. Bhatt.  An in-

teresting way to partition a number. Information Processing Letters, 71(3–4):141– 148, August 27, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/28/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/28/30/article.
pdf.

Bagchi:1990:PAG

[BHMS90] A. Bagchi, S. L. Hakimi, J. Mitchem, and E. Schmeichel. Parallel algorithms for gossiping by mail. Information Processing Letters, 34(4):197–202, April 24, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Beigel:1991:PPT

[BHW91] Richard Beigel, Lane A. Hemachandra, and Gerd Wechsung. Probabilistic polynomial time is closed under parity reductions. Information Processing Letters, 37 (2):91–94, January 31, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bshouty:1999:MTR

[BHWG99] Nader H. Bshouty, Lisa Higham, and Jolanta Warpechowska-Gruca. Meeting times of random walks on graphs. Information Processing Letters, 69(5): 259–265, March 12, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/47/21/
                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/47/21/25/article.
                                        pdf.

Biedl:1998:RBS

[Bie98] Therese C. Biedl. Relating bends and size in orthogonal graph drawings. Information Processing Letters, 65(2):111– 115, January 29, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bijlsma:1991:DLP

[Bij91] A. Bijlsma. Derivation of logic programs by functional methods. Information Processing Letters, 39(6):325–332, September 27, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bijlsma:1993:CPC

[Bij93a] A. Bijlsma. Calculating with procedure calls. Information Processing Letters, 46(5):211– 217, July 9, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bijlsma:1993:QBE

[Bij93b]  A. Bijlsma.  Quasi-Boolean

equivalence. Information Processing Letters, 45(5):243–247, April 2, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Billionnet:1993:PMC

[Bil93] Alain Billionnet. Partitioning multiple-chain-like task across a host- satellite system. Information Processing Letters, 48 (5):261–266, December 10, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boreale:1992:CSA

[BIN92] M. Boreale, P. Inverardi, and M. Nesi. Complete sets of axioms for finite basic LOTOS behavioural equivalences. Information Processing Letters, 43(3): 155–160, September 14, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Biro:1990:OIR

[Bir90] Miklos Biro. Object-oriented interaction in resource constrained scheduling. Information Processing Letters, 36(2):65–67, October 15, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Birget:1992:IUR

[Bir92]  Jean-Camille Birget.  Inter-

section and union of regular languages and state complexity. Information Processing Letters, 43(4):185–190, September 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Birget:1996:SCC

[Bir96] Jean-Camille Birget. The state complexity of \overline{Σ^{∗}\overline{L}} and its connection with temporal logic. Information Processing Letters, 58 (4):185–188, May 27, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bradford:1991:ITA

[BJ91] James H. Bradford and T. A. Jenkyns. On the inadequacy of tournament algorithms for the N-SCS problem. Information Processing Letters, 38(4): 169–171, May 31, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bui:1992:FGA

[BJ92] Thang Nguyen Bui and Curt Jones. Finding good approximate vertex and edge partitions is NP-hard. Information Processing Letters, 42(3):153–159, May 25, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bouquet:1997:UOH

[BJ97] Fabrice Bouquet and Philippe Jégou. Using OBDDs to handle dynamic constraints. Information Processing Letters, 62 (3):111–120, June 4, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blokhuis:1995:ECN

[BK95] A. Blokhuis and T. Kloks. On the equivalence covering number of splitgraphs. Information Processing Letters, 54(5):301–304, June 9, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blum:1997:CAC

[BK97] Avrim Blum and David Karger.

An Õ(n^{3/14})-coloring algorithm for 3-colorable graphs. Information Processing Letters, 61(1): 49–53, January 14, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baier:1998:VQP

[BK98] Christel Baier and Marta Kwiatkowska. On the verification of qualitative properties of probabilistic processes under fairness constraints. Information Processing Letters, 66(2):71–79, April 29, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bozanis:1997:SOV

[BKMT97] Panayiotis Bozanis, Nectarios Kitsios, Christos Makris, and Athanasios Tsakalidis. The space-optimal version of a known rectangle enclosure reporting algorithm. Information Processing Letters, 61(1):37–41, January 14, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bartal:1994:BLB

[BKR94] Yair Bartal, Howard Karloff, and Yuval Rabani. A better lower bound for on-line scheduling. Information Processing Letters, 50(3):113–116, May 9, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Burris:1990:UCR

[BL90] Stanley Burris and John Lawrence. Unification in commutative rings is not finitary. Information Processing Letters, 36(1):37–38, October 1, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blair:1991:BEW

[BL91] Jean R. S. Blair and Errol L. Lloyd. The benefits of external wires in single row routing. Information Processing Letters, 40(1):43–49, October 11, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Borchert:1996:SCR

[BL96] Bernd Borchert and Antoni Lozano. Succinct circuit representations and leaf language classes are basically the same concept. Information Processing Letters, 59(4):211–215, August 26, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blackburn:1994:IRO

[Bla94] Simon R. Blackburn. Increasing the rate of output of msequences. Information Processing Letters, 51(2):73–77, July 26, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barden:1999:EDS

[BLHDW99] Benjamin Barden, Ran Libeskind-Hadas, Janet Davis, and William Williams. On edgedisjoint spanning trees in hypercubes. Information Processing Letters, 70(1):13–16, April 16, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/17/
19/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/17/19/article.
pdf.

Brogi:1990:HRL

[BLM90] Antonio Brogi, Evelina Lamma, and Paola Mello. Hypothetical reasoning in logic programming. A semantic approach. Information Processing Letters, 36 (6):285–291, December 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bottoni:1998:SVH

[BLP98] Paolo Bottoni, Stefano Levialdi, and Gheorghe Păun. Successful visual human-computer interaction is undecidable. Information Processing Letters, 67(1):13–19, July 16, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blum:1992:RDT

[Blu92] Avrim Blum. Rank-r decision trees are a subclass of rdecision lists. Information Processing Letters, 42(4):183–185, June 19, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blundo:1995:NDT

[Blu95] Carlo Blundo. A note on dynamic threshold schemes. Information Processing Letters, 55 (4):189–193, August 25, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blum:1996:ISM

[Blu96] Norbert Blum. An O(n log n) implementation of the standard method for minimizing n-state finite automata. Information Processing Letters, 57(2):65–69, January 29, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bollig:1996:ELC

[BLW96] Beate Bollig, Martin Löbbing, and Ingo Wegener. On the effect of local changes in the variable ordering of ordered decision diagrams. Information Processing Letters, 59(5):233–239, September 9, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertossi:1990:PAC

[BM90a] Alan A. Bertossi and Sabrina Moretti. Parallel algorithms on circular-arc graphs. Information Processing Letters, 33(6):275– 281, February 10, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boxer:1990:CIPb

[BM90b] L. Boxer and R. Miller. Corrigenda: “Common intersections of polygons” [Inform. Process. Lett. 33(5), 10 January 1990, pp. 249–254]. Information Processing Letters, 35(1):53, June 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [BM90c].

Boxer:1990:CIPa

[BM90c] Laurence Boxer and Russ Miller.

Common intersections of polygons. Information Processing Letters, 33(5):249–254, January 10, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See corrigenda [BM90b].

Brewster:1991:NRH

[BM91] Richard Brewster and Gary MacGillivray. A note on restricted H-colouring. Information Processing Letters, 40(3): 149–151, November 8, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bousquet-Melou:1992:NMW

[BM92] Mireille Bousquet-Mélou. The number of minimal word chains computing the Thue–Morse word. Information Processing Letters, 44(2):57–64, November 19, 1992. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Busch:1997:IRW

[BM97] Costas Busch and Marios Mavronicolas. Impossibility results for weak threshold networks. Information Processing Letters, 63(2):85–90, August 13, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bottreau:1998:SRK

[BM98] Anne Bottreau and Yves Métivier. Some remarks on the Kronecker product of graphs. Information Processing Letters, 68(2): 55–61, October 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/46/19/
                                        19/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/46/19/19/article.
                                        pdf.

Baier:1996:DLT

[BMC96] Christel Baier and Mila E. Majster-Cederbaum. Denotational linear time semantics and sequential composition. Information Processing Letters, 59 (3):135–143, August 12, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bao:1997:ACR

[BMI97] Feng Bao, Aohan Mei, and Yoshihide Igarashi. Average competitive ratios of on-line spanning trees. Information Processing Letters, 62(4):213– 216, June 11, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertoni:1994:COL

[BMP94a] A. Bertoni, Carlo Mereghetti, and Giovanni Pighizzini. Corrigendum: “An optimal lower bound for nonregular languages” [Inform. Process. Lett. 50 (27 June 1994), no. 6, 289–292; MR 95c:68126]. Information Processing Letters, 52(6):339, December 23, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [BMP94b].

Bertoni:1994:OLB

[BMP94b] Alberto Bertoni, Carlo Mereghetti, and Giovanni Pighizzini. An optimal lower bound for nonregular languages. Information Processing Letters, 50(6):289–292, June 27, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [BMP94a].

Bshouty:1997:TBA

[BMST97] Nader H. Bshouty, Yishay Mansour, Baruch Schieber, and Prasoon Tiwari. A tight bound for approximating the square root. Information Processing Letters, 63(4):211–213, September 10, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Balasubramanian:1996:PPT

[BN96] R. Balasubramanian and S. V. Nagaraj. Perfect power testing. Information Processing Letters, 58(2):59–63, May 13, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bar-Noy:1999:MAB

[BNHK99] Amotz Bar-Noy, Magnús M. Halldórsson, and Guy Kortsarz. A matched approximation bound for the sum of a greedy coloring. Information Processing Letters, 71(3–4):135– 140, August 27, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/49/28/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/49/28/29/article.
                                        pdf.

Bar-Noy:1992:GAO

[BNMN92] Amotz Bar-Noy, Rajeev Motwani, and Joseph Naor. The greedy algorithm is optimal for on-line edge coloring. Information Processing Letters, 44(5): 251–253, December 21, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bonnier:1991:SFP

[BNN91] S. Bonnier, U. Nilsson, and T. Näslund. A simple fixed point characterization of threevalued stable model semantics. Information Processing Letters, 40(2):73–78, October 25, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Balachandran:1996:CTC

[BNR96] V. Balachandran, P. Nagavamsi, and C. Pandu Rangan. Clique transversal and clique independence on comparability graphs. Information Processing Letters, 58(4):181–184, May 27, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bolognesi:1993:DGR

[Bol93] T. Bolognesi. Deriving graphical representations of process networks from algebraic expressions. Information Processing Letters, 46(6):289–294, July 26, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Book:1991:ROS

[Boo91] Ronald V. Book. On random oracle separations. Information Processing Letters, 39(1):7–10, July 12, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Book:1994:CPT

[Boo94] Ronald V. Book. On collapsing the polynomial-time hierarchy. Information Processing Letters, 52(5):235–237, December 9, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boppana:1994:DTC

[Bop94] Ravi B. Boppana. The decisiontree complexity of element distinctness. Information Processing Letters, 52(6):329–331, December 23, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boppana:1997:ASB

[Bop97] Ravi B. Boppana. The average sensitivity of bounded-depth circuits. Information Processing Letters, 63(5):257–261, September 24, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ben-Or:1996:APF

[BOR96] Michael Ben-Or and Dana Ron. Agreement in the presence of faults, on networks of bounded degree. Information Processing Letters, 57(6):329–334, March 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bataineh:1991:PBO

[BÖS91] Abdulla Bataineh, Füsun Özgüner, and Alok Sarwal. Parallel Boolean operations for information retrieval. Information Processing Letters, 39(2):99–108, July 31, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boztas:1995:RMP

[Boz95] Serdar Boztaş. A robust multipriority topology-independent transmission schedule for packet radio networks. Information Processing Letters, 55(5):291– 295, September 15, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bennaceur:1993:EAC

[BP93] Hachemi Bennaceur and Gérard Plateau. An exact algorithm for the constraint satisfaction problem: application to logical inference. Information Processing Letters, 48(3):151–158, November 19, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barsi:1994:FPA

[BP94] Ferruccio Barsi and M. Cristina Pinotti. A fully parallel algorithm for residue to binary conversion. Information Processing Letters, 50(1):1–8, April 8, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also addendum [BP95c] and erratum [BP95b].

Barbosa:1995:AFM

[BP95a] Valmir C. Barbosa and Stella C. S. Porto. An algorithm for FIFO message delivery among migrating tasks. Information Processing Letters, 53(5):261– 267, March 10, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barsi:1995:EAF

[BP95b] F. Barsi and M. C. Pinotti. Erratum: Addendum to “A fully parallel algorithm for residue to binary conversion” [Inform. Process. Lett. 50(1), 8 April 1994, pp. 1–8]. Information Processing Letters, 56(2):121, October 27, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [BP94, BP95c].

Barsi:1995:AFP

[BP95c] Ferrucio Barsi and M. Cristina Perotti. Addendum to “A fully parallel algorithm for residue to binary conversion” [Inform. Process. Lett. 50(1), 8 April 1994, pp. 1–8]. Information Processing Letters, 55(1):25–26, July 7, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [BP94]. See also erratum [BP95b].

Bergstra:1998:KTV

[BP98] Jan A. Bergstra and Alban Ponse. Kleene’s three-valued logic and process algebra. Information Processing Letters, 67 (2):95–103, July 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bapiraju:1994:EBT

[BR94a] Vinnakota Bapiraju and V. V. Bapeswara Rao. Enumeration of binary trees. Information Processing Letters, 51(3):125–127, August 10, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barth:1994:TED

[BR94b] Dominique Barth and André Raspaud. Two edge-disjoint Hamiltonian cycles in the butterfly graph. Information Processing Letters, 51(4):175–179, August 24, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blum:1994:LBS

[BR94c] Norbert Blum and Henning Rochow. A lower bound on the single-operation worst-case time complexity of the unionfind problem on intervals. Information Processing Letters, 51 (2):57–60, July 26, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bird:1997:CR

[BR97] Richard S. Bird and Jesús N.

Ravelo. On computing representatives. Information Processing Letters, 63(1):1–7, July 30, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bultena:1998:EMA

[BR98] Bette Bultena and Frank Ruskey. An Eades–McKay algorithm for well-formed parentheses strings. Information Processing Letters, 68(5):255–259, December 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bradford:1990:SMB

[Bra90] James H. Bradford. Sequence matching with binary codes. Information Processing Letters, 34 (4):193–196, April 24, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Breslauer:1992:LSS

[Bre92] Dany Breslauer. An on-line string superprimitivity test. Information Processing Letters, 44 (6):345–347, December 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Breslauer:1994:TSS

[Bre94] Dany Breslauer. Testing string superprimitivity in parallel. Information Processing Letters, 49 (5):235–241, March 11, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bright:1993:RRM

[Bri93a] Jonathan D. Bright. Rangerestricted mergeable priority queues. Information Processing Letters, 47(3):159–164, September 14, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brinksmeier:1993:PFR

[Bri93b] E. Brinksmeier. Precision fabrication research and development. Unknown journal, 46(5): 25–109, July ??, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barnard:1990:PTA

[BS90] D. T. Barnard and D. B. Skillicorn. Pipelining treestructured algorithms on SIMD architectures. Information Processing Letters, 35(2):79–84, June 29, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Belbaraka:1994:GTC

[BS94a] Mounir Belbaraka and Ivan Stojmenović. On generating B-trees with constant average delay and in lexicographic order. Information Processing Letters, 49(1): 27–32, January 14, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bergstra:1994:BTW

[BŞ94b] J. A. Bergstra and Gh. Ştefănescu.

Bisimulation is two-way simulation. Information Processing Letters, 52(6):285–287, December 23, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baclawski:1996:CIC

[BS96a] Kenneth Baclawski and Dan A.

Simovici. A characterization of the information content of a classification. Information Processing Letters, 57(4):211–214, February 26, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Binkley:1996:CSQ

[BS96b] Aaron B. Binkley and Stephen R.

Schach. A comparison of sixteen quality metrics for objectoriented design. Information Processing Letters, 58(6):271– 275, June 24, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Borchert:1997:CLL

[BS97] Bernd Borchert and Riccardo Silvestri. A characterization of the leaf language classes. Information Processing Letters, 63 (3):153–158, August 27, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baek:1998:CHS

[BS98] Nakhoon Baek and Sung-Yong Shin. On circularly-hidden surface removal. Information Processing Letters, 66(3):119–123, May 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barbosa:1999:GAA

[BS99] Valmir C. Barbosa and Jayme L. Szwarcfiter. Generating all the acyclic orientations of an undirected graph. Information Processing Letters, 72(1–2):71– 74, October 29, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/27/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/27/34/article.
pdf.

Blundo:1999:PSS

[BSG99] C. Blundo, A. De Santis, and A. Giorgio Gaggia. Probability of shares in secret sharing schemes. Information Processing Letters, 72(5–6):169– 175, December 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/60/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/60/25/28/article.
pdf.

Bshouty:1992:LBM

[Bsh92] Nader H. Bshouty. A lower bound for the multiplication of polynomials modulo a polynomial. Information Processing Letters, 41(6):321–326, April 17, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bshouty:1995:ACM

[Bsh95] Nader H. Bshouty. On the additive complexity of 2 × 2 matrix multiplication. Information Processing Letters, 56(6):329–335, December 22, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bshouty:1996:SEL

[Bsh96] Nader H. Bshouty. A subexponential exact learning algorithm for DNF using equivalence queries. Information Processing Letters, 59(1):37–39, July 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bshouty:1997:LMP

[Bsh97] Nader H. Bshouty. On learning multivariate polynomials under the uniform distribution. Information Processing Letters, 61 (6):303–309, April 24, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bodlaender:1990:BES

[BT90] Hans L. Bodlaender and Gerard Tel. Bit-optimal election in synchronous rings. Information Processing Letters, 36(1):53–56, October 1, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blondel:1997:WPM

[BT97] Vincent D. Blondel and John N. Tsitsiklis. When is a pair of matrices mortal? Information Processing Letters, 63(5): 283–286, September 24, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Buhrman:1993:TQS

[BTv93] Harry Buhrman, Leen Torenvliet, and Peter van Emde Boas. Twenty questions to a pselector. Information Processing Letters, 48(4):201–204, November 29, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bshouty:1998:LWT

[BTW98] Nader H. Bshouty, Christino Tamon, and David K. Wilson. On learning width two branching programs. Information Processing Letters, 65(4):217–222, February 27, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bodlaender:1997:IHK

[BTY97] Hans L. Bodlaender, Dimitrios M. Thilikos, and Koichi Yamazaki. It is hard to know when greedy is good for finding independent sets. Information Processing Letters, 61(2):101– 106, February 28, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boyd:1991:SPP

[BU91] Sylvia Boyd and Hasan Ural.

The synchronization problem in protocol testing and its complexity. Information Processing Letters, 40(3):131–136, November 8, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boland:1995:CSC

[BU95a] Ralph P. Boland and Jorge Urrutia. Corrigendum: “Separating collections of points in Euclidean spaces” [Inform. Process. Lett. 53(4), 24 February 1995, pp. 177–183]. Information Processing Letters, 54(6):359, June 23, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [BU95b].

Boland:1995:SCP

[BU95b] Ralph P. Boland and Jorge Urrutia. Separating collections of points in Euclidean spaces. Information Processing Letters, 53 (4):177–183, February 24, 1995.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [BU95a].

Boroujerdi:1998:EAC

[BU98] Ali Boroujerdi and Jeffrey Uhlmann. An efficient algorithm for computing least cost paths with turn constraints. Information Processing Letters, 67(6): 317–321, September 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Buning:1990:ESP

[Bün90] Hans Kleine Büning. Existence of simple propositional formulas. Information Processing Letters, 36(4):177–182, November 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Burmester:1992:ARI

[Bur92] Mike Burmester. An almostconstant round interactive zeroknowledge proof. Information Processing Letters, 42(2):81–87, May 11, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Buro:1993:MLH

[Bur93] Michael Buro. On the maximum length of Huffman codes. Information Processing Letters, 45 (5):219–223, April 2, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Buss:1992:GME

[Bus92] Samuel R. Buss. The graph of multiplication is equivalent to counting. Information Processing Letters, 41(4):199–201, March 18, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Butler:1995:EVB

[But95] Greg Butler. Easy verification of behavioural subtyping in common cases. Information Processing Letters, 55(1):57–58, July 7, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bovet:1992:RLB

[BV92] Daniel P. Bovet and Stefano Varricchio. On the regularity of languages on a binary alphabet generated by copying systems. Information Processing Letters, 44(3):119–123, November 30, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Back:1995:GWS

[Bv95] R. J. R. Back and J. von Wright. Games and winning strategies. Information Processing Letters, 53(3):165–172, February 10, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boldi:1998:MSD

[BV98] Paolo Boldi and Sebastiano Vigna. Minimal sense of direction and decision problems for Cayley graphs. Information Processing Letters, 64(6):299–303, January 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Block:1990:MEG

[BW90] Kenneth Block and Tai-Kuo Woo. A more efficient generalization of Peterson’s mutual exclusion algorithm. Information Processing Letters, 35(5): 219–222, August 28, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bollig:1998:VSF

[BW98] Beate Bollig and Ingo Wegener. A very simple function that requires exponential size read-once branching programs. Information Processing Letters, 66(2): 53–57, April 29, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baeza-Yates:1991:HBD

[BY91] Ricardo A. Baeza-Yates. Height balance distribution of search trees. Information Processing Letters, 39(6):317–324, September 27, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baeza-Yates:1996:FAS

[BYF96] Ricardo A. Baeza-Yates and Luis O. Fuentes. A framework to animate string algorithms. Information Processing Letters, 59 (5):241–244, September 9, 1996.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baeza-Yates:1996:FPA

[BYP96] Ricardo A. Baeza-Yates and Chris H. Perleberg. Fast and practical approximate string matching. Information Processing Letters, 59(1):21–27, July 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baeza-Yates:1993:FTP

[BYR93] Ricardo Baeza-Yates and Mireille Regnier. Fast two-dimensional pattern matching. Information Processing Letters, 45(1):51–57, January 25, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cadoli:1992:CMC

[Cad92] Marco Cadoli. The complexity of model checking for circumscriptive formulae. Information Processing Letters, 44 (3):113–118, November 30, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cai:1993:RUT

[Cai93] Leizhen Cai. The recognition of union trees. Information Processing Letters, 45(6):279–283, April 16, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cai:1996:FPT

[Cai96] Leizhen Cai. Fixed-parameter tractability of graph modification problems for hereditary properties. Information Processing Letters, 58(4):171–176, May 27, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cai:1999:CPP

[Cai99]  Jin-Yi Cai.  A classifica-

tion of the probabilistic polynomial time hierarchy under fault tolerant access to oracle classes. Information Processing Letters, 69(4):167–174, February 26, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/20/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/20/18/article.
pdf.

Calvert:1994:EDL

[Cal94] Ken Calvert. Eliminating disjunctions of leads-to properties (temporal logic). Information Processing Letters, 49(4):189– 194, February 25, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Campbell:1998:CMP

[Cam98] Duncan K. G. Campbell. On the CLUMPS model of parallel computation. Information Processing Letters, 66(5):231–236, June 16, 1998. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Canetti:1996:MBP

[Can96] Ran Canetti. More on BPP and the Polynomial-time Hierarchy. Information Processing Letters, 57(5):237–241, March 11, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carlsson:1991:OAD

[Car91] Svante Carlsson. An optimal algorithm for deleting the root of a heap. Information Processing Letters, 37(2):117–120, January 31, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carpi:1994:RFW

[Car94] Arturo Carpi. On repeated factors in C^{∞}-words. Information Processing Letters, 52(6): 289–294, December 23, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Caussinus:1996:NTB

[Cau96] Hervé Caussinus. A note on a theorem of Barrington, Straubing and Thérien. Information Processing Letters, 58(1):31–33, April 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Charron-Bost:1991:CSL

[CB91] Bernadette Charron-Bost. Concerning the size of logical clocks in distributed systems. Information Processing Letters, 39 (1):11–16, July 12, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carlsson:1991:OPA

[CC91] Svante Carlsson and Jingsen Chen. An optimal parallel adaptive sorting algorithm. Information Processing Letters, 39(4): 195–200, August 30, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1992:FTR

[CC92] Hsun-Wen Chang and Kuo-Liang Chung. Fault-tolerant routing in unique-path multistage Omega network. Information Processing Letters, 44 (4):201–204, December 10, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chedid:1993:NVH

[CC93] Fouad B. Chedid and Riad B. Chedid. A new variation on hypercubes with smaller diameter. Information Processing Letters, 46(6):275–280, July 26, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1994:FAE

[CC94a] Ling Chen and Henry Y. H. Chuang. A fast algorithm for Euclidean distance maps of a 2-D binary image. Information Processing Letters, 51(1):25–29, July 12, 1994. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheng:1994:FAC

[CC94b] Eddie Cheng and William H.

Cunningham. A faster algorithm for computing the strength of a network. Information Processing Letters, 49(4):209–212, February 25, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1995:CEB

[CC95a] Chui Cheng Chen and Rong Jaye Chen. Compact embedding of binary trees into hypercubes. Information Processing Letters, 54 (2):69–72, April 28, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chiang:1995:SGG

[CC95b] Wei-Kuo Chiang and Rong-Jaye Chen. The (n, k)-star graph: A generalized star graph. Information Processing Letters, 56(5): 259–264, December 8, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1997:ECI

[CC97] Hung-Yi Chang and Rong-Jaye Chen. Embedding cycles in IEH graphs. Information Processing Letters, 64(1):23–27, October 22, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chiang:1998:AG

[CC98] Wei-Kuo Chiang and Rong-Jaye Chen. On the arrangement graph. Information Processing Letters, 66(4):215–219, May 29, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1999:FMM

[CC99] Chien-Yuan Chen and Chin-Chen Chang. A fast modular multiplication algorithm for calculating the product AB modulo N. Information Processing Letters, 72(3–4):77–81, November 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/28/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/28/25/article.
pdf.

Chao:1996:MSP

[CCC96] Chia-Chiang Chao, Wen-Tsuen Chen, and Gen-Huey Chen. Multiple search problem on reconfigurable meshes. Information Processing Letters, 58(2): 65–69, May 13, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:1999:FPM

[CCG^{+}99] Maxime Crochemore, A. Czumaj, L. Gsaieniec, T. Lecroq, W. Plandowski, and W. Rytter. Fast practical multi-pattern matching. Information Processing Letters, 71(3–4):107– 113, August 27, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/28/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/28/26/article.
pdf.

Chang:1993:EBE

[CCH93a] Byeong-Mo Chang, Kwang-Moo Choe, and Taisook Han. Efficient bottom-up execution of logic programs using abstract interpretation. Information Processing Letters, 47(3):149–157, September 14, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1993:SFS

[CCH93b] Byeong-Mo Chang, Kwang-Moo Choe, and Taisook Han. Static filtering on stratified programs. Information Processing Letters, 47(5):237–244, October 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cha:1993:CMC

[CCK93] Shin Cha, In Sang Chung, and Yong Rae Kwon. Complexity measures for concurrent programs based on informationtheoretic metrics. Information Processing Letters, 46(1):43–50, April 29, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1992:BFT

[CD92] Calvin C.-Y. Chen and Sajal K.

Das. Breadth-first traversal of trees and integer sorting in parallel. Information Processing Letters, 41(1):39–49, January 21, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chakraborty:1993:URA

[CD93a] Uday Kumar Chakraborty and D. Ghosh Dastidar. Using reliability analysis to estimate the number of generations to convergence in genetic algorithms. Information Processing Letters, 46 (4):199–209, June 25, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1993:DDS

[CD93b] S. C. Chang and M. W. Du. Diamond deque: A simple data structure for priority deques. Information Processing Letters, 46 (5):231–237, July 9, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chepoi:1994:CMP

[CD94a] Victor Chepoi and Feodor Dragan. Computing a median point of a simple rectilinear polygon. Information Processing Letters, 49(6):281–285, March 22, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Collin:1994:SSD

[CD94b] Zeev Collin and Shlomi Dolev. Self-stabilizing depth-first search. Information Processing Letters, 49(6):297–301, March 22, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Colussi:1996:TSE

[CD96] Livio Colussi and Alessia De Col. A time and space efficient data structure for string searching on large texts. Information Processing Letters, 58(5): 217–222, June 10, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1998:MVP

[CD98a] Danny Z. Chen and Ovidiu Daescu. Maintaining visibility of a polygon with a moving point of view. Information Processing Letters, 65(5):269–275, March 13, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheng:1998:CSS

[CD98b] T. C. E. Cheng and Q. Ding.

The complexity of scheduling starting time dependent tasks with release times. Information Processing Letters, 65(2):75–79, January 29, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chrobak:1999:RCP

[CD99] Marek Chrobak and Christoph Dürr. Reconstructing hvconvex polyominoes from orthogonal projections. Information Processing Letters, 69(6): 283–289, March 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/22/
20/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/22/20/article.
pdf.

Chen:1991:UAP

[CDA91] Calvin C.-Y. Chen, Sajal K. Das, and Selim G. Akl. A unified approach to parallel depth-first traversals of general trees. Information Processing Letters, 38 (1):49–55, April 12, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crvenkovic:1999:NEC

[CDÉ99] S. Crvenković, I. Dolinka, and Z. Ésik. A note on equations for commutative regular languages. Information Processing Letters, 70(6):265– 267, June 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/22/
19/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/22/19/article.
pdf.

Cosnard:1990:CSX

[CDF90] Michel Cosnard, Jean Duprat, and Afonso G. Ferreira. The complexity of searching in X+Y and other multisets. Information Processing Letters, 34(2): 103–109, March 16, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cattell:1996:SLT

[CDF96] Kevin Cattell, Michael J. Dinneen, and Michael R. Fellows. A simple linear-time algorithm for finding path- decompositions of small width. Information Processing Letters, 57(4):197–203, February 26, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Clarke:1993:UAS

[CDK93a] E. M. Clarke, I. A. Draghicescu, and R. P. Kurshan. A unified approach for showing language inclusion and equivalence between various types of ω-automata. Information Processing Letters, 46 (6):301–308, July 26, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cross:1993:RNC

[CDK93b] David Cross, Reinhard Drefenstedt, and Jorg Keller. Reduction of network cost and wiring in Ranade’s butterfly routing. Information Processing Letters, 45(2):63–67, February 26, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czumaj:1996:PMI

[CDP96] Artur Czumaj, Krzysztof Diks, and Teresa M. Przytycka. Parallel maximum independent set in convex bipartite graphs. Information Processing Letters, 59(6):289–294, September 23, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cerioli:1998:HSS

[CEdK98] Marcia R. Cerioli, Hazel Everett, Celina M. H. de Figueiredo, and Sulamita Klein. The homogeneous set sandwich problem. Information Processing Letters, 67 (1):31–35, July 16, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Canetti:1995:LBS

[CEG95] Ran Canetti, Guy Even, and Oded Goldreich. Lower bounds for sampling algorithms for estimating the average. Information Processing Letters, 53(1):17–25, January 13, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Coffman:1997:AMP

[CFGG97] E. G. Coffman, Jr., L. Flatto, E. N. Gilbert, and A. G. Greenberg. An approximate model of processor communication rings under heavy load. Information Processing Letters, 64(2):61–67, November 5, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crescenzi:1991:NAM

[CFS91] P. Crescenzi, C. Fiorini, and R. Silvestri. A note on the approximation of the MAX CLIQUE problem. Information Processing Letters, 40(1): 1–5, October 11, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1991:MRI

[CG91] Chung-Kuo Chang and M. G.

Gouda. On the minimum requirements for independent recovery in distributed systems. Information Processing Letters, 37(1):1–7, January 10, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1993:TSP

[CG93] Danny Z. Chen and Sumantha Guha. Testing a simple polygon for monotonicity optimally in parallel. Information Processing Letters, 47(6):325–331, October 18, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chickering:1995:FCB

[CGH95] David M. Chickering, Dan Geiger, and David Heckerman. On finding a cycle basis with a shortest maximal cycle. Information Processing Letters, 54 (1):55–58, April 14, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Castro:1998:LNM

[CGL98] Jorge Castro, David Guijarro, and Víctor Lavín. Learning nearly monotone k-term DNF. Information Processing Letters, 67(2):75–79, July 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:1993:TDP

[CGR93] Maxime Crochemore, Leszek Gasieniec, and Wojciech Rytter. Two-dimensional pattern matching by sampling. Information Processing Letters, 46(4): 159–162, June 25, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chaudhuri:1998:NCP

[CGR98] Shiva Chaudhuri, Naveen Garg, and R. Ravi. The p-neighbor k-center problem. Information Processing Letters, 65(3):131– 134, February 13, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Campbell:1991:MVG

[CH91] Douglas Campbell and John Higgins. Minimal visibility graphs. Information Processing Letters, 37(1):49–53, January 10, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also note [SH93].

Chandru:1992:DEH

[CH92a] V. Chandru and J. N. Hooker. Detecting embedded Horn structure in propositional logic. Information Processing Letters, 42 (2):109–111, May 11, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1992:NAP

[CH92b] Rong Jaye Chen and Yu Song Hou. Non-associative parallel prefix computation. Information Processing Letters, 44(2):91–94, November 19, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cho:1992:PCC

[CH92c] Sang Cho and Dũng T. Hùynh.

The parallel complexity of coarsest set partition problems. Information Processing Letters, 42 (2):89–94, May 11, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1993:QPP

[CH93a] Gen Huey Chen and Yung Chen Hung. On the quickest path problem. Information Processing Letters, 46(3):125–128, June 11, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1993:NSN

[CH93b] Wen Tsuen Chen and Kuen Rong Hsieh. A neural sorting network with O(1) time complexity. Information Processing Letters, 45 (6):309–313, April 16, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1995:EPP

[CH95] Yung Cheng Chang and Lih Hsing Hsu. Element perturbation problems of optimum spanning trees with two-parameter objectives. Information Processing Letters, 53(1):55–59, January 13, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1997:SAG

[CH97] Chin-Chen Chang and Shin-Jia Hwang. A simple approach for generating RSA keys. Information Processing Letters, 63 (1):19–21, July 30, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1998:RGC

[CH98] Jou-Ming Chang and Chin-Wen Ho. The recognition of geodetically connected graphs. Information Processing Letters, 65 (2):81–88, January 29, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cregut:1999:PPE

[CH99] P. Crégut and B. Heyd. Progress properties for empty Unity programs. Information Processing Letters, 69(3):107–109, February 12, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/19/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/19/17/article.
pdf.

Chandrasekharan:1990:ITT

[Cha90] N. Chandrasekharan. Isomorphism testing of k-trees is in NC, for fixed k. Information Processing Letters, 34(6):283–287, May 28, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1991:SSG

[Cha91] Ruay Shiung Chang. Single step graph search problem. Information Processing Letters, 40(2): 107–111, October 25, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chandra:1992:DRH

[Cha92] Barun Chandra. Does randomization help in on-line bin packing? Information Processing Letters, 43(1):15–19, August 10, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chandra:1994:CSS

[Cha94] Barun Chandra. Constructing sparse spanners for most graphs in higher dimensions. Information Processing Letters, 51(6): 289–294, September 26, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chakraborty:1995:BPM

[Cha95a] Uday Kumar Chakraborty. A branching process model for genetic algorithms. Information Processing Letters, 56(5):281– 292, December 8, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chakraborty:1995:SDS

[Cha95b] Uday Kumar Chakraborty. A simpler derivation of schema hazard in genetic algorithms. Information Processing Letters, 56(2):77–78, October 27, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chan:1998:BAK

[Cha98a]  Timothy M. Chan.  Back-

wards analysis of the kargerklein-tarjan algorithm for minimum spanning trees. Information Processing Letters, 67(6): 303–304, September 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Charatonik:1998:UFT

[Cha98b] Witold Charatonik. An undecidable fragment of the theory of set constraints. Information Processing Letters, 68(3): 147–151, November 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/ipl/cas_sub/browse/
browse.cgi?year=1998&volume=
68&issue=3&aid=2081.

Chang:2002:CPP

[Cha02] Gerard J. Chang. Corrigendum to “The path-partition problem in block graphs”: [Information Processing Letters 52 (1994) 317–322]. Information Processing Letters, 83(5):293, September 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [YC94].

Cheng:1990:ASN

[Che90] Jingde Cheng. An algebraic semantics of notional entailment logic Cn. Information Processing Letters, 36(3):117–121, November 1, 1990. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1991:RNA

[Che91a] Zhi-Zhong Chen. A randomized NC algorithm for the maximal tree cover problem. Information Processing Letters, 40 (5):241–246, December 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheng:1991:GTA

[Che91b] Xu Cheng. A graph transformation algorithm for concurrency control in a partitioned database. Information Processing Letters, 38(1):43–48, April 12, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1992:OPT

[Che92a] Lin Chen. Optimal parallel time bounds for the maximum clique problem on intervals. Information Processing Letters, 42(4): 197–201, June 19, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1992:FEP

[Che92b] Zhi-Zhong Chen. A fast and efficient parallel algorithm for finding a satisfying truth assignment to a 2-CNF formula. Information Processing Letters, 43(4): 191–193, September 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1992:SPA

[Che92c] Zhi-Zhong Chen. A simple parallel algorithm for computing the diameters of all vertices in a tree and its application. Information Processing Letters, 42 (5):243–248, July 3, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1994:FQS

[Che94a] Y. L. Chen. Finding the k quickest simple paths in a network. Information Processing Letters, 50(2):89–92, April 22, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1994:PAF

[Che94b] Zhi-Zhong Chen. A parallel algorithm for finding a triconnected component separator with an application. Information Processing Letters, 52(5): 265–271, December 9, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chedid:1995:GTC

[Che95a] Fouad B. Chedid. On the generalized twisted cube. Information Processing Letters, 55(1):49–52, July 7, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1995:SSP

[Che95b] Lin Chen. Solving the shortestpaths problem on bipartite permutation graphs efficiently. Information Processing Letters, 55(5):259–264, September 15, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1995:SOD

[Che95c] Tung-Shou Chen. SIMPLE: An optimal disk system with two restricted heads. Information Processing Letters, 55(5):273– 277, September 15, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1995:FEN

[Che95d] Zhi-Zhong Chen. A fast and efficient NC algorithm for maximal matching. Information Processing Letters, 55(6):303– 307, September 29, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheng:1996:WES

[Che96] Siu-Wing Cheng. Widest empty L-shaped corridor. Information Processing Letters, 58(6): 277–283, June 24, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1999:ACR

[Che99] Yangjun Chen. Arc consistency revisited. Information Processing Letters, 70(4):175– 184, May 28, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/20/
19/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/20/19/article.
pdf.

Chin:1990:DCC

[Chi90] Andrew Chin. On the depth complexity of the counting functions. Information Processing Letters, 35(6):325–328, September 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chin:1993:PBP

[Chi93] Andrew Chin. Permutations on the Block PRAM. Information Processing Letters, 45(2):69–73, February 26, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chiu:1998:FTB

[Chi98] Ge-Ming Chiu. A fault-tolerant broadcasting algorithm for hypercubes. Information Processing Letters, 66(2):93–99, April 29, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chlebikova:1996:AMB

[Chl96] Janka Chlebíková. Approximating the Maximally Balanced Connected Partition Problem in graphs. Information Processing Letters, 60(5):225–230, December 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chao:1997:OEP

[CHL97] H. S. Chao, F. R. Hsu, and R. C. T. Lee. An optimal EREW parallel algorithm for computing breadth-first search trees on permutation graphs. Information Processing Letters, 61(6): 311–316, April 24, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czech:1992:OAG

[CHM92] Zbigniew J. Czech, George Havas, and Bohdan S. Majewski. An optimal algorithm for generating minimal perfect hash functions. Information Processing Letters, 43(5):257–264, October 5, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chou:1996:SPT

[Cho96] Ching-Tsun Chou. Simple proof techniques for property preservation via simulation. Information Processing Letters, 60(3): 129–134, November 11, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chaochen:1991:CD

[CHR91] Zhou Chaochen, C. A. R. Hoare, and Anders P. Ravn. A calculus of durations. Information Processing Letters, 40(5):269–276, December 13, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Christie:1996:SPB

[Chr96]  David A. Christie.  Sort-

ing permutations by blockinterchanges. Information Processing Letters, 60(4):165–169, November 25, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1990:OAC

[CHT90] Maw Shang Chang, Nen-Fu Huang, and Chuan-Yi Tang. An optimal algorithm for constructing oriented Voronoi diagrams and geographic neighborhood graphs. Information Processing Letters, 35(5):255–260, August 28, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chu:1991:ECF

[Chu91] Pai-Cheng C. Chu. Evaluating clustering factor approach to estimating block selectivities. Information Processing Letters, 40 (6):329–334, December 30, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:1992:IEH

[Chu92] Soon M. Chung. Indexed extendible hashing. Information Processing Letters, 44(1):1–6, November 9, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:1997:EHD

[Chu97a] Kuo-Liang Chung. Efficient Huffman decoding. Information Processing Letters, 61(2):97–99, February 28, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comment [TM98].

Chung:1997:FAS

[Chu97b] Kuo-Liang Chung. A fast algorithm for stereo matching. Information Processing Letters, 63 (2):57–61, August 13, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chu:1998:RO

[Chu98] Francis Chu. Reducing Ω to ♢W. Information Processing Letters, 67(6):289–293, September 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chu:1999:VCA

[Chu99] P. C. Chu. Verbs are not cases: Applying case grammar to document retrieval. Information Processing Letters, 71 (1):29–34, July 16, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/17/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/17/22/article.
pdf.

Clementi:1995:RPF

[CI95] Andrea Clementi and Russell Impagliazzo. The reachability problem for finite cellular automata. Information Processing Letters, 53(1):27–31, January 13, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cimikowski:1990:FHC

[Cim90] Robert J. Cimikowski. Finding Hamiltonian cycles in certain planar graphs. Information Processing Letters, 35(5): 249–254, August 28, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheng:1990:EDA

[CJ90] Siu Wing Cheng and Ravi Janardan. Efficient dynamic algorithms for some geometric intersection problems. Information Processing Letters, 36(5): 251–258, December 1, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Clark:1995:SRP

[CJ95] John Clark and Jeremy Jacob. On the security of recent protocols. Information Processing Letters, 56(3):151–155, November 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chmeiss:1997:GCG

[CJ97] Assef Chmeiss and Philippe Jégou. A generalization of chordal graphs and the maximum clique problem. Information Processing Letters, 62(2): 61–66, May 21, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Choi:1998:RTS

[CJKJ98] Kyunghee Choi, Gihyun Jung, Taegeun Kim, and Seunhun Jung. Real-time scheduling algorithm for minimizing maximum weighted error with O(N log N + cN) complexity. Information Processing Letters, 67(6):311– 315, September 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cooper:1990:CFA

[CK90] James Cooper and Leslie Kitchen.

CASOP: A fast algorithm for computing the n-ary composition of a binary associative operator. Information Processing Letters, 34(4):209–213, April 24, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chor:1993:CPT

[CK93a] Benny Chor and Eyal Kushilevitz. A communication-privacy tradeoff for modular addition. Information Processing Letters, 45(4):205–210, March 22, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Culik:1993:PRS

[CK93b] Karel Culik, II and Jarkko Kari.

Parametrized recurrent systems for image generation. Information Processing Letters, 48(6): 267–274, December 20, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1996:ERS

[CK96a] Danny Z. Chen and Kevin S.

Klenk. Erratum to “Rectilinear short path queries among rectangular obstacles”. Information Processing Letters 57 (6) (25 March 1996) 313–319. Information Processing Letters, 60 (2):105, October 28, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [CK96b].

Chen:1996:RSP

[CK96b] Danny Z. Chen and Kevin S. Klenk. Rectilinear short path queries among rectangular obstacles. Information Processing Letters, 57(6):313–319, March 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also errata [NAG96, CK96a].

Chen:1997:NAG

[CKK97] Jianer Chen, Saroja P. Kanchi, and Arkady Kanevsky. A note on approximating graph genus. Information Processing Letters, 61(6):317–322, April 24, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cucker:1997:CD

[CKM97] Felipe Cucker, Pascal Koiran, and Martín Matamala. Complexity and dimension. Information Processing Letters, 62(4): 209–212, June 11, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Corneil:1995:SLT

[CKN^{+}95] Derek G. Corneil, Hiryoung Kim, Sridhar Natarajan, Stephan Olariu, and Alan P. Sprague. Simple linear time recognition of unit interval graphs. Information Processing Letters, 55 (2):99–104, July 21, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czyzowicz:1998:SPR

[CKU98] Jurek Czyzowicz, Evangelos Kranakis, and Jorge Urrutia. A simple proof of the representation of bipartite planar graphs as the contact graphs of orthogonal straight line segments. Information Processing Letters, 66 (3):125–126, May 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chaudhuri:1991:SLC

[CL91a] Soma Chaudhuri and Richard E.

Ladner. Safety and liveness of ωcontext-free languages. Information Processing Letters, 37(1): 13–20, January 10, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chrobak:1991:NSP

[CL91b] Marek Chrobak and Lawrence L.

Larmore. A note on the server problem and a benevolent adversary. Information Processing Letters, 38(4):173–175, May 31, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1992:FMS

[CL92a] R. C. Chang and H. S. Lee.

Finding a maximum set of independent chords in a circle. Information Processing Letters, 41 (2):99–102, February 14, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chee:1992:ACC

[CL92b] Yeow Meng Chee and Andrew Lim. The algorithmic complexity of colour switching. Information Processing Letters, 43(2): 63–68, August 24, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheung:1992:MPR

[CL92c] Steven Cheung and Francis C. M. Lau. Mesh permutation routing with locality. Information Processing Letters, 43(2): 101–105, August 24, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1993:PAW

[CL93a] Maw Shang Chang and Yi Chang Liu. Polynomial algorithms for the weighted perfect domination problems on chordal graphs and split graphs. Information Processing Letters, 48(4):205–210, November 29, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1993:CCS

[CL93b] Cheng Chia Chen and I P’êng Lin. The computational complexity of satisfiability of temporal Horn formulas in propositional linear-time temporal logic. Information Processing Letters, 45(3):131–136, March 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheung:1993:LBP

[CL93c] S. Cheung and F. C. M. Lau.

A lower bound for permutation routing on two-dimensional bused meshes. Information Processing Letters, 45(5):225–228, April 2, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cull:1995:GTC

[CL95] Paul Cull and Shawn M. Larson.

On generalized twisted cubes. Information Processing Letters, 55(1):53–55, July 7, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1996:DSO

[CL96] Tsong Yueh Chen and Man Fai Lau. Dividing strategies for the optimization of a test suite. Information Processing Letters, 60 (3):135–141, November 11, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1997:MLS

[CL97a] Ruay-Shiung Chang and Shing-Jiuan Leu. The minimum labeling spanning trees. Information Processing Letters, 63(5): 277–282, September 24, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Choi:1997:DST

[CL97b] Y. Choi and T. W. Lam. Dynamic suffix tree and twodimensional texts management. Information Processing Letters, 61(4):213–220, March 27, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:1997:TBC

[CL97c] Maxime Crochemore and Thierry Lecroq. Tight bounds on the complexity of the Apostolico-Giancarlo algorithm. Information Processing Letters, 63(4): 195–203, September 10, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1999:ECA

[CL99] Jae-Young Chang and Sang-Goo Lee. Extended conditions for answering an aggregate query using materialized views. Information Processing Letters, 72(5–6):205–212, December 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/60/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/60/25/33/article.
pdf.

Clausen:1992:AAB

[Cla92] Michael Clausen. Almost all Boolean functions have no linear symmetries. Information Processing Letters, 41(6):291–292, April 17, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1990:OTP

[CLBW90] Wen-Huei Chen, Ching-Sung Lu, Elben R. Brozovsky, and Jin-Tuu Wang. An optimization technique for protocol conformance testing using multiple UIO sequences. Information Processing Letters, 36(1):7–11, October 1, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chrobak:1997:BLB

[CLLR97] Marek Chrobak, Lawrence L.

Larmore, Carsten Lund, and Nick Reingold. A better lower bound on the competitive ratio of the randomized 2-server problem. Information Processing Letters, 63(2):79–83, August 13, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ciampolini:1996:AIF

[CLM96] Anna Ciampolini, Evelina Lamma, and Paola Mello. An abstract interpretation framework for optimizing dynamic modular logic languages. Information Processing Letters, 58(4):163–170, May 27, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Clote:1996:NMC

[Clo96] Peter Clote. A note on the monotone complexity of 2-REF. Information Processing Letters, 57(3):117–123, February 12, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czyzowicz:1991:SFL

[CLP91] Jurek Czyzowicz, K. B. Lakshmanan, and Andrzej Pelc. Searching with a forbidden lie pattern in responses. Information Processing Letters, 37 (3):127–132, February 18, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheng:1998:MVR

[CLW98] Yi Chang Cheng, Erl Huei Lu, and Shaw Woei Wu. A modified version of the Rao– Nam algebraic-code encryption scheme. Information Processing Letters, 68(4):215–217, November 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheng:1994:ECA

[CM94] Pen Cheng and Shigeru Masuyama. On the equivalence in complexity among three computation problems on maximum number of edge-disjoint s-t paths in a probabilistic graph. Information Processing Letters, 51 (4):195–199, August 24, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:1996:SBT

[CM96] Soon M. Chung and Pyeong S. Mah. Semantics-based transaction management for multidatabase systems. Information Processing Letters, 59(2):109– 115, July 22, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cabrelli:1998:LTA

[CM98a] Carlos A. Cabrelli and Ursula M.

Molter. A linear time algorithm for a matching problem on the circle. Information Processing Letters, 66(3):161–164, May 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Collins:1998:ILB

[CM98b] Michael J. Collins and Bernard M. E. Moret. Improved lower bounds for the link length of rectilinear spanning paths in grids. Information Processing Letters, 68(6):317–319, December 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Calamoneri:1999:OLM

[CM99a] Tiziana Calamoneri and Annalisa Massini. An optimal layout of multigrid networks. Information Processing Letters, 72(3–4):137– 141, November 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/28/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/28/35/article.
pdf.

Cheriyan:1999:AHL

[CM99b] Joseph Cheriyan and Kurt Mehlhorn. An analysis of the highest-level selection rule in the preflow-push max-flow algorithm. Information Processing Letters, 69(5):239–242, March 12, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/47/21/
                                        20/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/47/21/20/article.
                                        pdf.

Codenotti:1995:ATC

[CMM95] Bruno Codenotti, Giovanni Manzini, and Luciano Margara. Algebraic techniques in communication complexity. Information Processing Letters, 56(4): 191–195, November 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:1998:AFW

[CMR98] M. Crochemore, F. Mignosi, and A. Restivo. Automata and forbidden words. Information Processing Letters, 67(3):111–117, August 17, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ching:1990:DDD

[CMS90] Yu-Tai Ching, Kurt Mehlhorn, and Michiel H. M. Smid. Dynamic deferred data structuring. Information Processing Letters, 35(1):37–40, June 15, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1994:IPL

[CN94] Wen Chin Chen and Wen Chun Ni. Internal path length of the binary representation of heapordered trees. Information Processing Letters, 51(3):129–132, August 10, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1998:WII

[CNP98] Maw-Shang Chang, P. Nagavamsi, and C. Pandu Rangan. Weighted irredundance of interval graphs. Information Processing Letters, 66(2):65–70, April 29, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carson:1992:DMT

[CNV92] S. D. Carson, V. Nirkhe, and P. Vongsathorn. A discrete-state model of the two-headed disk. Information Processing Letters, 41(6):341–345, April 17, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cohen:1994:CSG

[Coh94] Ernie Cohen. The convergence span of greedy load balancing. Information Processing Letters, 52(4):181–182, November 25, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Collette:1994:EPC

[Col94] Pierre Collette. An explanatory presentation of composition rules for assumptioncommitment specifications. Information Processing Letters, 50 (1):31–35, April 8, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Corneil:1995:LTA

[COS95] Derek G. Corneil, Stephan Olariu, and Lorna Stewart. A linear time algorithm to compute a dominating path in an AT-free graph. Information Processing Letters, 54(5):253–257, June 9, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Coupry:1997:SLA

[Cou97] Laurent Coupry. A simple linear algorithm for the edgedisjoint (s, t)-paths problem in undirected planar graphs. Information Processing Letters, 64 (2):83–86, November 5, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cook:1990:FCL

[CP90] Stephen Cook and Toniann Pitassi. A feasibly constructive lower bound for resolution proofs. Information Processing Letters, 34(2):81–85, March 16, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chrobak:1995:LTA

[CP95] M. Chrobak and T. H. Payne. A linear-time algorithm for drawing a planar graph on a grid. Information Processing Letters, 54 (4):241–246, May 26, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Calamoneri:1997:NRT

[CP97] Tiziana Calamoneri and Rossella Petreschi. A new 3D representation of trivalent Cayley networks. Information Processing Letters, 61(5):247–252, April 10, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:1990:PCM

[CR90] Maxime Crochemore and Wojciech Rytter. Parallel construction of minimal suffix and factor automata. Information Processing Letters, 35(3):121–128, July 20, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:1991:EPA

[CR91] Maxime Crochemore and Wojciech Rytter. Efficient parallel algorithms to test squarefreeness and factorize strings. Information Processing Letters, 38(2):57–60, April 26, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Caprara:1998:IFA

[CR98a] Alberto Caprara and Romeo Rizzi. Improving a family of approximation algorithms to edge color multigraphs. Information Processing Letters, 68(1):11–15, October 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chalamaiah:1998:FSP

[CR98b] Nimmagadda Chalamaiah and Badrinath Ramamurthy. Finding shortest paths in distributed loop networks. Information Processing Letters, 67(3):157–161, August 17, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Calera-Rubio:1998:CRE

[CRC98] Jorge Calera-Rubio and Rafael C. Carrasco. Computing the relative entropy between regular tree languages. Information Processing Letters, 68(6):283–289, December 30, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Creignou:1998:CVS

[Cre98] Nadia Creignou. Complexity versus stability for classes of propositional formulas. Information Processing Letters, 68 (4):161–165, November 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crippa:1991:SCD

[Cri91] Davide Crippa. A special case of the dynamization problem for least cost paths. Information Processing Letters, 39(6):297– 302, September 27, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chomicki:1990:GCW

[CS90a] Jan Chomicki and V. S.

Subrahmanian. Generalized closed world assumption is Π_{2}^{0}- complete. Information Processing Letters, 34(6):289–291, May 28, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crescenzi:1990:RCE

[CS90b] P. Crescenzi and R. Silvestri.

Relative complexity of evaluating the optimum cost and constructing the optimum for maximization problems. Information Processing Letters, 33(5): 221–226, January 10, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cherief:1991:TFA

[CS91a] F. Cherief and Ph. Schnoebelen. τ-bisimulations and full abstraction for refinement of actions. Information Processing Letters, 40(4):219–222, November 25, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Corbett:1991:UAS

[CS91b] Peter F. Corbett and Isaac D.

Scherson. A unified algorithm for sorting on multidimensional mesh-connected processors. Information Processing Letters, 37 (4):225–231, February 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cohen:1992:RDO

[CS92] D. A. Cohen and E. A. Scott. Rationality of division orderings. Information Processing Letters, 44(6):307–311, December 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chinn:1993:BSS

[CS93a] Donald D. Chinn and Rakesh K. Sinha. Bounds on sample space size for matrix product verification. Information Processing Letters, 48(2):87–91, November 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crescenzi:1993:NDC

[CS93b] Pierluigi Crescenzi and Riccardo Silvestri. A note on the descriptive complexity of maximization problems. Information Processing Letters, 47(1):11–15, August 9, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cidon:1995:MTA

[CS95] Israel Cidon and Yuval Shavitt. Message terminating algorithms for anonymous rings of unknown size. Information Processing Letters, 54(2):111–119, April 28, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cusick:1996:BNFb

[CS96] Thomas W. Cusick and Pantelimon Stănică. Bounds on the number of functions satisfying the Strict Avalanche Criterion. Information Processing Letters, 60(4):215–219, November 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [YT96].

Calamoneri:1997:SLG

[CS97a] Tiziana Calamoneri and Andrea Sterbini. 3D straight-line grid drawing of 4-colorable graphs. Information Processing Letters, 63(2):97–102, August 13, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cintioli:1997:RRK

[CS97b] Patrizio Cintioli and Riccardo Silvestri. Revisiting a result of Ko. Information Processing Letters, 61(4):189–194, March 27, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chauhan:1999:PBR

[CS99] S. R. Chauhan and I. A. Stewart. On the power of builtin relations in certain classes of program schemes. Information Processing Letters, 69(2): 77–82, January 29, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/18/
21/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/47/18/21/article.
                                        pdf.

Clark:1992:LTA

[CSS92] L. H. Clark, F. Shahrokhi, and L. A. Székely. A linear time algorithm for graph partition problems. Information Processing Letters, 42(1):19–24, April 27, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Colbourn:1992:PMI

[CST92] C. J. Colbourn, D. R. Stinson, and L. Teirlinck. A parallelization of Miller’s n^{log n} isomorphism technique. Information Processing Letters, 42(4): 223–228, June 19, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1994:EPB

[CST94] Shao Dong Chen, Hong Shen, and Rodney Topor. An efficient permutation-based parallel range-join algorithm on N-dimensional torus computers. Information Processing Letters, 52(1):35–38, October 14, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1991:COI

[CT91] Wen-Huei Chen and Chuan Yi Tang. Computing the optimal IO sequences of a protocol in polynomial time. Information Processing Letters, 40(3):

145–148, November 8, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1993:BDA

[CT93] Wen Huei Chen and Chuan Yi Tang. A 2 · |E|-bit distributed algorithm for the directed Euler trail problem. Information Processing Letters, 47(1):43–49, August 9, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cesati:1997:EPT

[CT97] Marco Cesati and Luca Trevisan. On the efficiency of polynomial time approximation schemes. Information Processing Letters, 64(4):165–171, December 13, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1998:SIV

[CTVC98] Kaijung Chang, Yan Huat Tan, Yaakov Varol, and Jiang-Hsing Chu. A study on interleaving versus segmentation. Information Processing Letters, 65(6): 319–324, March 27, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cusick:1995:CPK

[Cus95] Thomas W. Cusick. Cryptanalysis of a public key system based on Diophantine equations. Information Processing Letters, 56(2):73–75, October 27, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cusick:1996:BNFa

[Cus96] Thomas W. Cusick. Bounds on the number of functions satisfying the Strict Avalanche Criterion. Information Processing Letters, 57(5):261–263, March 11, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comment [YT96].

Chen:1994:LBR

[CvW94] Bo Chen, André van Vliet, and Gerhard J. Woeginger. A lower bound for randomized on-line scheduling algorithms. Information Processing Letters, 51(5): 219–222, September 12, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cameron:1992:NPL

[CW92a] Helen Cameron and Derick Wood. A note on the path length of red-black trees. Information Processing Letters, 42(5):287– 292, July 3, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1992:EAM

[CW92b] Maw-Shang Chang and Fu-Hsing Wang. Efficient algorithms for the maximum weight clique and maximum weight independent set problems on permutation graphs. Information Processing Letters, 43(6):293– 295, October 19, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1997:LTA

[CW97a] Hon-Chan Chen and Yue-Li Wang. A linear time algorithm for finding depth-first spanning trees on trapezoid graphs. Information Processing Letters, 63 (1):13–18, July 30, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Csirik:1997:SAL

[CW97b] János Csirik and Gerhard J.

Woeginger. Shelf algorithms for on-line strip packing. Information Processing Letters, 63(4): 171–175, September 10, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1998:IFM

[CW98a] Jiun-Liang Chen and Feng-Jian Wang. An inheritance flow model for class hierarchy analysis. Information Processing Letters, 66(6):309–315, June 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1998:TCM

[CW98b] Loon-Been Chen and I-Chen Wu. On the time complexity of minimum and maximum global snapshot problems. Information Processing Letters, 67(3): 151–156, August 17, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheng:1999:TMF

[CW99] T. C. E. Cheng and G. Wang.

Two-machine flowshop scheduling with consecutive availability constraints. Information Processing Letters, 71(2):49– 54, July 30, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/18/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/18/17/article.
pdf.

Chang:1993:MLP

[CWK93] Ee-Chien Chang, Weiguo Wang, and Mohan S. Kankanhalli. Multidimensional on-line binpacking: an algorithm and its average-case analysis. Information Processing Letters, 48(3): 121–125, November 19, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1999:MEF

[CWL99] Hong-Chung Chen, Yue-Li Wang, and Yu-Feng Lan. A memory-efficient and fast Huffman decoding algorithm. Information Processing Letters, 69(3):119–122, February 12, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/19/
20/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/19/20/article.
pdf.

Cai:1999:COB

[CXZ99] Leizhen Cai, Yinfeng Xu, and Binhai Zhu. Computing the optimal bridge between two convex polygons. Information Processing Letters, 69(3):127– 130, February 12, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/19/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/19/22/article.
pdf.

Chien:1992:WCA

[CY92] Been-Chian Chien and Wei-Pang Yang. The worst case analysis of algorithm on multiple stacks manipulation. Information Processing Letters, 44 (2):107–111, November 19, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1994:MCH

[CY94] Her-Kun Chang and Shyan-Ming Yuan. Message complexity of hierarchical quorum consensus algorithm. Information Processing Letters, 50(2):69–73, April 22, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1995:NID

[CY95] Ee-Chien Chang and Chee Yap.

A note on improved deterministic time simulation of nondeterministic space for small space.

Information Processing Letters, 55(3):155–157, August 11, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1996:MGS

[CY96] Tsong Yueh Chen and Yuen Tak Yu. A more general sufficient condition for partition testing to be better than random testing. Information Processing Letters, 57(3):145–149, February 12, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1997:EAM

[CY97] Ye-In Chang and Bi-Yen Yang. Efficient access methods for image databases. Information Processing Letters, 64(2):95–105, November 5, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1998:PAC

[CY98] Yaw-Wen Chang and Chang-Biau Yang. A parallel algorithm for circulant tridiagonal linear systems. Information Processing Letters, 65(6):331–337, March 27, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1991:SSA

[CYH91] Nian Shing Chen, Hwey Pyng Yu, and Shing Tsaan Huang. A self-stabilizing algorithm for constructing spanning trees. Information Processing Letters, 39 (3):147–151, August 16, 1991.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1992:DAO

[CYY^{+}92] Hanxiong Chen, Xu Yu, Kazunori Yamaguchi, Hiroyuki Kitagawa, Nobuo Ohbo, and Yuzuru Fujiwara. Decomposition — an approach for optimizing queries including ADT functions. Information Processing Letters, 43 (6):327–333, October 19, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Damm:1990:PCD

[Dam90] Carsten Damm. Problems complete for ⊕L. Information Processing Letters, 36(5):247–250, December 1, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dam:1992:GDB

[Dam92] Mads Dam. R-generability, and definability in branching time logics. Information Processing Letters, 41(5):281–287, April 3, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Damaschke:1995:PAN

[Dam95] Peter Damaschke. A parallel algorithm for nearly optimal edge search. Information Processing Letters, 56(4):233–236, November 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Damaschke:1998:RGT

[Dam98] Peter Damaschke. Randomized group testing for mutually obscuring defectives. Information Processing Letters, 67(3): 131–135, August 17, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Damm:1999:DES

[Dam99] Carsten Damm. Depth-efficient simulation of Boolean semiunbounded circuits by arithmetic ones. Information Processing Letters, 69(4):175–179, February 26, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/20/
19/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/20/19/article.
pdf.

Danvy:1991:SCN

[Dan91a]  Olivier Danvy.  Semantics-

directed compilation of nonlinear patterns. Information Processing Letters, 37(6):315–322, March 28, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Danzig:1991:CGA

[Dan91b] Peter B. Danzig. A cooperative game with applications to computer networks. Information Processing Letters, 38(6): 283–289, June 28, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dankelmann:1993:CAD

[Dan93] Peter Dankelmann. Computing the average distance of an interval graph. Information Processing Letters, 48(6):311–314, December 20, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dancik:1996:CBF

[Dan96] Vlado Dančík. Complexity of Boolean functions over bases with unbounded fan-in gates. Information Processing Letters, 57(1):31–34, January 15, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Das:1996:MLT

[Das96] S. K. Das. Modal logics in the theory of relational databases. Information Processing Letters, 57(1):1–7, January 15, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dasgupta:1998:AUF

[Das98] Pallab Dasgupta. Agreement under faulty interfaces. Information Processing Letters, 65 (3):125–129, February 13, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dawes:1992:SPE

[Daw92] Robin W. Dawes. Some pursuitevasion problems on grids. Information Processing Letters, 43 (5):241–247, October 5, 1992.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Diekert:1991:ICS

[DB91] Volker Diekert and Ronald V. Book. On “inherently contextsensitive” languages—an application of complexity cores. Information Processing Letters, 40 (1):21–23, October 11, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deAlmeida:1993:SMA

[dB93] Nalvo F. de Almeida, Jr. and Valmir C. Barbosa. A stringmatching algorithm for the CREW PRAM. Information Processing Letters, 47(5):257– 259, October 8, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Davis:1995:OPA

[DB95] Robert Davis and Alan Burns. Optimal priority assignment for aperiodic tasks with firm deadlines in fixed priority preemptive systems. Information Processing Letters, 53(5):249– 254, March 10, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deAzevedo:1996:STP

[dBDL96] Marcelo M. de Azevedo, Nader Bagherzadeh, Martin Dowd, and Shahram Latifi. Some topological properties of star connected cycles. Information Processing Letters, 58(2):81–85, May 13, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DiZenzo:1997:NIR

[DBM97] Silvano Di Zenzo, Paolo Bottoni, and Piero Mussio. A notion of information related to computation. Information Processing Letters, 64(4):207–215, December 13, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dey:1990:LAT

[DBT90] Pradip Dey, Barrett R. Bryant, and Tadao Takaoka. Lexical ambiguity in tree adjoining grammars. Information Processing Letters, 34(2):65–69, March 16, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dimitrov:1992:MRB

[DCD92] V. S. Dimitrov, T. V. Cooklev, and B. D. Donevsky. On the multiplication of reduced biquaternions and applications. Information Processing Letters, 43(3):161–164, September 14, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dasgupta:1995:UPP

[DCD95] Pallab Dasgupta, P. P. Chakrabarti, and S. C. De Sarkar. Utility of pathmax in partial order heuristic search. Information Processing Letters, 55(6):317– 322, September 29, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dasgupta:1996:ASU

[DCD96] Pallab Dasgupta, P. P. Chakrabarti, and S. C. DeSarkar. Agent search in uniform b-ary trees: Multiple goals and unequal costs. Information Processing Letters, 58(6):311–318, June 24, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeSantis:1999:GCP

[DCGP99] Alfredo De Santis, Giovanni Di Crescenzo, Oded Goldreich, and Giuseppe Persiano. The graph clustering problem has a perfect zero-knowledge interactive proof. Information Processing Letters, 69(4):201– 206, February 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/20/
23/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/20/23/article.
pdf.

Duh:1996:CPG

[DCH96] Dyi-Rong Duh, Gen-Huey Chen, and D. Frank Hsu. Combinatorial properties of generalized hypercube graphs. Information Processing Letters, 57(1):41–45, January 15, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dickerson:1990:FNN

[DD90] Matthew T. Dickerson and R. Scot Drysdale. Fixed-radius near neighbors search algorithms for points and segments. Information Processing Letters, 35 (5):269–273, August 28, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DePrisco:1993:BST

[DD93] Roberto De Prisco and Alfredo De Santis. On binary search trees. Information Processing Letters, 45(5):249–253, April 2, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deBerg:1997:CSC

[dDDS97] Mark de Berg, Olivier Devillers, Katrin Dobrindt, and Otfried Schwarzkopf. Computing a single cell in the overlay of two simple polygons. Information Processing Letters, 63(4):215– 219, September 10, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Diks:1998:BUH

[DDK^{+}98] Krzysztof Diks, Stefan Dobrev, Evangelos Kranakis, Andrzej Pelc, and Peter Ružička. Broadcasting in unlabeled hypercubes with a linear number of messages. Information Processing Letters, 66(4):181–186, May 29, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/23/
                                        20/44/20/25/abstract.html.

deMoor:1992:IDT

[de 92] Oege de Moor. Inductive data types for predicate transformers. Information Processing Letters, 43(3):113–117, September 14, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deLuca:1995:DPF

[de 95] Aldo de Luca. A division property of the Fibonacci word. Information Processing Letters, 54 (6):307–312, June 23, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dalal:1992:HTS

[DE92] Mukesh Dalal and David W. Etherington. A hierarchy of tractable satisfiability problems. Information Processing Letters, 44(4):173–180, December 10, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Demri:1995:CNM

[Dem95] Stéphane Demri. 3-SAT = SAT for a class of normal modal logics. Information Processing Letters, 54(5):281–287, June 9, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Deng:1990:OPA

[Den90] Xiaotie Deng. An optimal parallel algorithm for linear programming in the plane. Information Processing Letters, 35(4): 213–217, August 7, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Deng:1996:LBC

[Den96] Xiaotie Deng. A lower bound for communication on the crossbar. Information Processing Letters, 57(2):103–108, January 29, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Denis:1999:FMD

[Den99] François Denis. Finding a minimal 1-DNF consistent with a positive sample is LOGSNPcomplete. Information Processing Letters, 69(1):1–5, January 15, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/17/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/17/17/article.
pdf.

Dermouche:1995:FAS

[Der95] A. Dermouche. A fast algorithm for string matching with mismatches. Information Processing Letters, 55(2):105–110, July 21, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Devroye:1998:RCS

[Dev98] Luc Devroye. On the richness of the collection of subtrees in random binary search trees. Information Processing Letters, 65 (4):195–199, February 27, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Domingo-Ferrer:1991:DUI

[DF91] Josep Domingo-Ferrer. Distributed user identification by zero-knowledge access rights proving. Information Processing Letters, 40(5):235–239, December 13, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

dAmore:1992:OBP

[dF92] Fabrizio d’Amore and Paolo Giulio Franciosa. On the optimal binary plane partition for sets of isothetic rectangles. Information Processing Letters, 44(5):255– 259, December 21, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Danyi:1996:LEN

[DF96a] G. Dányi and Z. Fülöp. Letter to the editor: “A note on the equivalence problem of Epatterns” [Inform. Process. Lett. 57 (1996), no. 3, 125–128; MR 96m:68090]. Information Processing Letters, 58(6):327, June 24, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [DF96b].

Danyi:1996:NEP

[DF96b] G. Dányi and Z. Fülöp. A note on the equivalence problem of E-patterns. Information Processing Letters, 57(3):125–128, February 12, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See letter to the editor [DF96a].

deFigueiredo:1999:TCN

[dFMdM99] Celina M. H. de Figueiredo, João Meidanis, and Célia Picinin de Mello. Total-chromatic number and chromatic index of dually chordal graphs. Information Processing Letters, 70(3): 147–152, May 14, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/19/
21/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/19/21/article.
pdf.

Devillers:1995:IAF

[DG95] Olivier Devillers and Mordecai J. Golin. Incremental algorithms for finding the convex hulls of circles and the lower envelopes of parabolas. Information Processing Letters, 56(3): 157–164, November 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dehne:1997:EAC

[DG97] Frank Dehne and Katia Guimarães.

Exact and approximate computational geometry solutions of an unrestricted point set stereo matching problem. Information Processing Letters, 64(3):107– 114, November 29, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Deogun:1999:CRP

[DG99] Jitender S. Deogun and K. Gopalakrishnan. Consecutive retrieval property—revisited. Information Processing Letters, 69(1): 15–20, January 15, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/47/17/
                                        19/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/47/17/19/article.
                                        pdf.

Dunne:1995:MSS

[DGL95] Paul E. Dunne, Chris J. Gittings, and Paul H. Leng. Multiprocessor simulation strategies with optimal speed-up. Information Processing Letters, 54(1): 23–33, April 14, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dessmark:1998:NPC

[DGL98] Anders Dessmark, Oscar Garrido, and Andrzej Lingas. A note on parallel complexity of maximum f-matching. Information Processing Letters, 65(2): 107–109, January 29, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Demange:1994:ARM

[DGP94] Marc Demange, Pascal Grisoni, and Vangelis Th. Paschos. Approximation results for the minimum graph coloring problem. Information Processing Letters, 50(1):19–23, April 8, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DellOlmo:1997:ARD

[DGS97] P. Dell’Olmo, S. Giordani, and M. G. Speranza. An approximation result for a duo-processor task scheduling problem. Information Processing Letters, 61 (4):195–200, March 27, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ding:1998:CGO

[DH98] Cunsheng Ding and Tor Helleseth. On cyclotomic generator of order r. Information Processing Letters, 66(1):21–25, April 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dekel:1992:OAF

[DHO92] Eliezer Dekel, Jie Hu, and Wen Ouyang. An optimal algorithm for finding compact sets. Information Processing Letters, 44 (5):285–289, December 21, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Danicic:1995:PAS

[DHS95] Sebastian Danicic, Mark Harman, and Yoga Sivagurunathan. A parallel algorithm for static program slicing. Information Processing Letters, 56(6):307– 313, December 22, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:1990:MFR

[Dij90] Edsger W. Dijkstra. Making a fair roulette from a possibly biased coin. Information Processing Letters, 36(4):193, November 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:1995:HCP

[Dij95a] Edsger W. Dijkstra. Heuristics for a calculational proof. Information Processing Letters, 53 (3):141–143, February 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:1995:EUP

[Dij95b] Rutger M. Dijkstra. An experiment with the use of predicate transformers in UNITY. Information Processing Letters, 53(6):329–332, March 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:1996:EPA

[Dij96] Rutger M. Dijkstra. “everywhere” in predicate algebra and modal logic. Information Processing Letters, 58(5):237–243, June 10, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dai:1997:NPA

[DIK97] Yang Dai, Kazuo Iwano, and Naoki Katoh. A new probabilistic analysis of Karger’s randomized algorithm for minimum cut problems. Information Processing Letters, 64(5):255–261, December 23, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dillencourt:1990:RDT

[Dil90] Michael B. Dillencourt. Realizability of Delaunay triangulations. Information Processing Letters, 33(6):283–287, February 10, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dinitz:1999:AZS

[DIR99] Yefim Dinitz, Alon Itai, and Michael Rodeh. On an algorithm of Zemlyachenko for subtree isomorphism. Information Processing Letters, 70(3): 141–146, May 14, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/19/
20/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/19/20/article.
pdf.

Deo:1994:OPA

[DJM94] Narsingh Deo, Amit Jain, and Muralidhar Medidi. An optimal parallel algorithm for merging using multiselection. Information Processing Letters, 50 (2):81–87, April 22, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dimitrov:1998:AME

[DJM98] V. S. Dimitrov, G. A. Jullien, and W. C. Miller. An algorithm for modular exponentiation. Information Processing Letters, 66 (3):155–159, May 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dunten:1996:SEF

[DJS96] Brian Dunten, Julie Jones, and Jonathan Sorenson. A spaceefficient fast prime number sieve. Information Processing Letters, 59(2):79–84, July 22, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dielissen:1991:RPP

[DK91] Victor J. Dielissen and Anne Kaldewaij. Rectangular partition is polynomial in two dimensions but NP-complete in three. Information Processing Letters, 38(1):1–6, April 12, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Devroye:1994:NHS

[DK94a] Luc Devroye and Paul Kruszewski. A note on the Horton–Strahler number for random trees. Information Processing Letters, 52 (3):155–159, November 11, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dhamdhere:1994:TBR

[DK94b] Dhananjay M. Dhamdhere and Sandeep S. Kulkarni. A token based k-resilient mutual exclusion algorithm for distributed systems. Information Processing Letters, 50(3):151–157, May 9, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Duszak:1994:GND

[DK94c] Zbigniew Duszak and Waldemar W. Koczkodaj. Generalization of a new definition of consistency for pairwise comparisons. Information Processing Letters, 52(5):273–276, December 9, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Devroye:1995:NHS

[DK95] Luc Devroye and Paul Kruszewski.

A note on the Horton–Strahler number for random trees. Information Processing Letters, 56 (2):95–99, October 27, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dror:1997:SCM

[DKD97] Moshe Dror, Wieslaw Kubiak, and Paolo Dell’Olmo. Scheduling chains to minimize mean flow time. Information Processing Letters, 61(6):297–301, April 24, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dahlhaus:1992:EPA

[DKK92] Elias Dahlhaus, Marek Karpiński, and Pierre Kelsen. An efficient parallel algorithm for computing a maximal independent set in a hypergraph of dimension 3. Information Processing Letters, 42 (6):309–313, July 24, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Diks:1994:OCV

[DKK^{+}94] Krzysztof Diks, Evangelos Kranakis, Danny Krizanc, Bernard Mans, and Andrzej Pelc. Optimal coteries and voting schemes. Information Processing Letters, 51 (1):1–6, July 12, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dror:1999:TPS

[DKL99] M. Dror, W. Kubiak, and J. Y-T. Leung. Tree precedence in scheduling: The strong-weak distinction. Information Processing Letters, 71(3–4):127– 134, August 27, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/28/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/28/28/article.
pdf.

Deogun:1997:AAC

[DKS97] Jitender S. Deogun, Dieter Kratsch, and George Steiner. An approximation algorithm for clustering graphs with dominating diametral path. Information Processing Letters, 61(3): 121–127, March 14, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DellOlmo:1998:AAP

[DKST98] Paolo Dell’Olmo, Hans Kellerer, Maria Grazia Speranza, and Zsolt Tuza. A 13/12 approximation algorithm for bin packing with extendable bins. Information Processing Letters, 65(5): 229–233, March 13, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Desel:1995:SPS

[DKVW95] Jörg Desel, Ekkart Kindler, Tobias Vesper, and Rolf Walter. A simplified proof for a selfstabilizing protocol: a game of cards. Information Processing Letters, 54(6):327–328, June 23, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeIaco:1992:FAP

[DL92] Bettina De Iaco and Fabrizio Luccio. Finding all the palindromes in a binary tree in linear time and space. Information Processing Letters, 44(2):73–77, November 19, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DiBattista:1990:BGU

[DLR90] Giuseppe Di Battista, Wei-Ping Liu, and Ivan Rival. Bipartite graphs, upward drawings, and planarity. Information Processing Letters, 36(6):317–322, December 15, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

delaVega:1999:AHD

[dlVK99] W. Fernandez de la Vega and M. Karpinski. On the approximation hardness of dense TSP and other path problems. Information Processing Letters, 70 (2):53–55, April 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/18/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/18/17/article.
pdf.

Deng:1991:SPR

[DM91] Xiaotie Deng and Sanjeev Mahajan. Server problems and resistive spaces. Information Processing Letters, 37(4):193–196, February 28, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deLuca:1994:BFT

[dM94] Aldo de Luca and Licia Mione. On bispecial factors of the Thue–Morse word. Information Processing Letters, 49(4):179– 183, February 25, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Damaschke:1990:DCC

[DMK90] Peter Damaschke, Haiko Müller, and Dieter Kratsch. Domination in convex and chordal bipartite graphs. Information Processing Letters, 36(5):231–236, December 1, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dervos:1998:CSF

[DML98] D. Dervos, Y. Manolopoulos, and P. Linardis. Comparison of signature file models with superimposed coding. Information Processing Letters, 65(2): 101–106, January 29, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dahlhaus:1998:MCS

[DMM98] Elias Dahlhaus, Paul D. Manuel, and Mirka Miller. Maximum h-colourable subgraph problem in balanced graphs. Information Processing Letters, 65(6): 301–303, March 27, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dietz:1994:BRO

[DMS94] Paul F. Dietz, Ioan I. Macarie, and Joel I. Seiferas. Bits and relative order from residues, space efficiently. Information Processing Letters, 50(3):123–127, May 9, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dillenburg:1993:IED

[DN93] John F. Dillenburg and Peter C.

Nelson. Improving the efficiency of depth-first search by cycle elimination. Information Processing Letters, 45(1):5–10, January 25, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DelLungo:1998:MDS

[DNPS98] A. Del Lungo, M. Nivat, R. Pinzani, and L. Sorri. The medians of discrete sets. Information Processing Letters, 65(6): 293–299, March 27, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DomingoiFerrer:1996:NPH

[Dom96] Josep Domingo i. Ferrer. A new privacy homomorphism and applications. Information Processing Letters, 60(5):277–282, December 8, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dong:1995:IPP

[Don95] Guozhu Dong. On the index of positive programmed formal languages. Information Processing Letters, 54(2):105–110, April 28, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Davis:1993:ELS

[DP93] Robert Davis and Armand Prieditis. The expected length of a shortest path. Information Processing Letters, 46(3): 135–141, June 11, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DiCrescenzo:1994:ROP

[DP94] Giovanni Di Crescenzo and Giuseppe Persiano. Roundoptimal perfect zero-knowledge proofs. Information Processing Letters, 50(2):93–99, April 22, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DePrisco:1995:CIB

[DP95] Roberto De Prisco and Giuseppe Persiano. Characteristic inequalities for binary trees. Information Processing Letters, 53 (4):201–207, February 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dong:1997:MTC

[DP97a] Guozhu Dong and Chaoyi Pang.

Maintaining transitive closure in first order after node-set and edge-set deletions. Information Processing Letters, 62(4): 193–199, June 11, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Doring:1997:DAL

[DP97b] Andreas Döring and Wolfgang J.

Paul. Decimal adjustment of long numbers in constant time. Information Processing Letters, 62(3):161–163, June 4, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dassow:1999:MMN

[DP99] Jürgen Dassow and Gheorghe Păun. Min of Mat is not necessarily Mat. Information Processing Letters, 71(5–6):175– 177, September 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/30/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        29/28/24/49/30/25/article.
                                        pdf.

DePrisco:1996:NEP

[DPP96] Roberto De Prisco, Giuseppe Parlati, and Giuseppe Persiano. A note on the expected path length of trees with known fringe. Information Processing Letters, 59(6):309–315, September 23, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dassow:1993:UL

[DPS93] Jürgen Dassow, Gheorghe Păun, and Arto Salomaa. On the union of 0L languages. Information Processing Letters, 47(2):59–63, August 20, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dietz:1994:CUT

[DR94] Paul F. Dietz and Rajeev Raman. A constant update time finger search tree. Information Processing Letters, 52(3): 147–154, November 11, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Drewes:1993:NCC

[Dre93] Frank Drewes. NP-completeness of k-connected hyperedge-replacement languages of order k. Information Processing Letters, 45(2): 89–94, February 26, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Droubay:1995:PFW

[Dro95] Xavier Droubay. Palindromes in the Fibonacci word. Information Processing Letters, 55(4): 217–221, August 25, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Drozdowski:1996:RTS

[Dro96] Maciej Drozdowski. Real-time scheduling of linear speedup parallel tasks. Information Processing Letters, 57(1):35–40, January 15, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dickerson:1993:SAE

[DS93a] Matthew T. Dickerson and Jason Shugart. A simple algorithm for enumerating longest distances in the plane. Information Processing Letters, 45(5): 269–274, April 2, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Durand:1993:CET

[DS93b] Iréne Durand and Bruno Salinier. Constructor equivalent term rewriting systems. Information Processing Letters, 47(3):131– 137, September 14, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dey:1994:MFC

[DS94a] Tamal K. Dey and Nimish R.

Shah. Many-face complexity in incremental convex arrangements. Information Processing Letters, 51(5):227–231, September 12, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Durand:1994:CET

[DS94b] Irène Durand and Bruno Salinier. Constructor equivalent term rewriting systems are strongly sequential: A direct proof. Information Processing Letters, 52 (3):137–145, November 11, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Datta:1995:AAB

[DS95] Alak K. Datta and Ranjan K. Sen. 1-Approximation algorithm for bottleneck disjoint path matching. Information Processing Letters, 55(1):41–44, July 7, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeAgostino:1996:LVL

[DS96a] Sergio De Agostino and James A. Storer. On-line versus off-line computation in dynamic text compression. Information Processing Letters, 59(3):169–174, August 12, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dong:1996:CQC

[DS96b] Guozhu Dong and Jianwen Su. Conjunctive query containment with respect to views and constraints. Information Processing Letters, 57(2):95–102, January 29, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dannoura:1998:IYF

[DS98] Eisuke Dannoura and Kouichi Sakurai. An improvement on El-Yaniv-Fiat-Karp-Turpin’s money-making bi-directional trading strategy. Information Processing Letters, 66(1):27–33, April 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Datta:1999:ESS

[DS99] Alak Kumar Datta and Ranjan Kumar Sen. An efficient scheme to solve two problems for two-terminal series parallel graphs. Information Processing Letters, 71(1): 9–15, July 16, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/17/
19/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/17/19/article.
pdf.

Day:1992:AGC

[DT92] Khaled Day and Anand Tripathi. Arrangement graphs: A class of generalized star graphs. Information Processing Letters, 42(5):235–241, July 3, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Day:1993:USG

[DT93a] Khaled Day and Anand Tripathi. Unidirectional star graphs. Information Processing Letters, 45(3):123–129, March 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Demetrovics:1993:SPC

[DT93b] János Demetrovics and Vu Duc Thi. Some problems concerning keys for relation schemes and relations in the relational datamodel. Information Processing Letters, 46(4):179–184, June 25, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DiBattista:1992:CVR

[DTT92] Giuseppe Di Battista, Roberto Tamassia, and Ioannis G. Tollis. Constrained visibility representations of graphs. Information Processing Letters, 41(1): 1–7, January 21, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Domingo:1997:POR

[DTW97] Carlos Domingo, Tatsuie Tsukiji, and Osamu Watanabe. Partial Occam’s Razor and its applications. Information Processing Letters, 64(4):179–185, December 13, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dublish:1990:SCS

[Dub90] P. Dublish. Some comments on the subtree isomorphism problem for ordered trees. Information Processing Letters, 36(5): 273–275, December 1, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dunne:1990:CKP

[Dun90] P. E. Dunne. Comment on Kochol’s paper “Efficient monotone circuits for threshold functions” [Inform. Process. Lett. 32(3), 24 August 1989, pp. 121–122]. Information Processing Letters, 34 (5):221–222, May 7, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Koc89].

Duris:1991:SS

[DV91] Pavol Duris and Imrich Vrto.

Semelectivity is not sufficient. Information Processing Letters, 37(3):137–141, February 18, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deBerg:1994:RDL

[dv94] Mark de Berg and Marc van Kreveld. Rectilinear decompositions with low stabbing number. Information Processing Letters, 52(4):215–221, November 25, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeBonis:1998:IAG

[DV98a] Annalisa De Bonis and Ugo Vaccaro. Improved algorithms for group testing with inhibitors. Information Processing Letters, 67(2):57–64, July 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeMarco:1998:BHS

[DV98b] Gianluca De Marco and Ugo Vaccaro. Broadcasting in hypercubes and star graphs with dynamic faults. Information Processing Letters, 66(6):321–326, June 30, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dobrev:1999:OBH

[DV99] S. Dobrev and I. Vrt̆o. Optimal broadcasting in hypercubes with dynamic faults. Information Processing Letters, 71(2): 81–85, July 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/18/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/18/22/article.
pdf.

Deineko:1994:CHL

[DvR94] Vladimir G. Deĭneko, René van Dal, and Günter Rote. The convex-hull-and-line Traveling Salesman Problem: A solvable case. Information Processing Letters, 51(3):141–148, August 10, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dederichs:1990:SLM

[DW90] Frank Dederichs and Rainer Weber. Safety and liveness from a methodological point of view. Information Processing Letters, 36(1):25–30, October 1, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comments [AAA^{+}91, DW91].

Dederichs:1991:RCM

[DW91] Frank Dederichs and Rainer Weber. Reply to the comments by M. Abadi et al. Information Processing Letters, 40(3): 143, November 8, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [AAA^{+}91, DW90].

Ding:1994:CBI

[DW94] Yuzheng Ding and Mark Allen Weiss. On the complexity of building an interval heap. Information Processing Letters, 50 (3):143–144, May 9, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Deineko:1996:CHL

[DW96] Vladimir G. Deĭneko and Gerhard J. Woeginger. The convexhull-and-k-line travelling salesman problem. Information Processing Letters, 59(6):295–301, September 23, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dunne:1998:IUB

[DZ98] Paul E. Dunne and Michele Zito.

An improved upper bound on the non-3-colourability threshold. Information Processing Letters, 65(1):17–23, January 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1991:CPH

[eC91] Jian er Chen. Characterizing parallel hierarchies by reducibilities. Information Processing Letters, 39(6):303–307, September 27, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1991:NH

[eCCM91] Jian er Chen, Jim Cox, and Bud Mishra. An NL hierarchy. Information Processing Letters, 39 (1):21–26, July 12, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Estivill-Castro:1995:IPV

[ECOUX95] Vladimir Estivill-Castro, Joseph O’Rourke, Jorge Urrutia, and Dianna Xu. Illumination of polygons with vertex lights. Information Processing Letters, 56(1): 9–13, October 13, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Elmagarmid:1991:IAQ

[ED91] Ahmed K. Elmagarmid and Wei Min Du. Integrity aspects of quasi serializability. Information Processing Letters, 38(1):23–28, April 12, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eiselt:1996:OLF

[EGL96] H. A. Eiselt, Michel Gendreau, and Gilbert Laporte. Optimal location of facilities on a network with an unreliable node or link. Information Processing Letters, 58(2):71–74, May 13, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Erlingsson:1996:EMR

[EKR96] Úlfar Erlingsson, Mukkai Krishnamoorthy, and T. V. Raman. Efficient multiway radix search trees. Information Processing Letters, 60(3):115–120, November 11, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL ftp:/

/ftp.cs.princeton.edu/pub/
packages/lcc/contrib/mrst.
shar;   http://www.cs.princeton.
edu/software/lcc/doc/mrst.
pdf.

Eades:1993:FEH

[ELS93] Peter Eades, Xuemin Lin, and W. F. Smyth. A fast and effective heuristic for the feedback arc set problem. Information Processing Letters, 47(6): 319–323, October 18, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Encarnacion:1997:BBP

[Enc97] Mark J. Encarnación. Black-box polynomial resultants. Information Processing Letters, 61(4): 201–204, March 27, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Engelfriet:1992:EPD

[Eng92] Joost Engelfriet. An elementary proof of Double Greibach Normal Form. Information Processing Letters, 44(6):291–293, December 28, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eppstein:1994:ABS

[Epp94] David Eppstein. Arboricity and bipartite subgraph listing algorithms. Information Processing Letters, 51(4):207–211, August 24, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eppstein:1997:DCD

[Epp97] David Eppstein. Dynamic connectivity in digital images. Information Processing Letters, 62 (3):121–126, June 4, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ercal:1991:DEI

[Erc91] Fikret Ercal. Distributed evaluation of an iterative function for all object pairs on an SIMD hypercube. Information Processing Letters, 40(6):341–345, December 30, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Easwarakumar:1994:LAC

[ERC94] K. S. Easwarakumar, C. Pandu Rangan, and G. A. Cheston. A linear algorithm for centering a spanning tree of a biconnected graph. Information Processing Letters, 51(3):121–124, August 10, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Efrat:1998:SSL

[ES98] Alon Efrat and Otfried Schwarzkopf.

Separating and shattering long line segments. Information Processing Letters, 64(6):309–314, January 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Esparza:1992:SCP

[Esp92] Javier Esparza. A solution to the covering problem for 1-bounded conflict-free Petri nets using linear programming. Information Processing Letters, 41(6):313– 319, April 17, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Esselink:1992:OAA

[Ess92] Klaas Esselink. Order of Appel’s algorithm. Information Processing Letters, 41(3):141–147, March 6, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Emerald:1999:NIU

[EST99] J. D. Emerald, K. G. Subramanian, and D. G. Thomas. A note on inferring uniquely terminating code languages. Information Processing Letters, 70(5): 217–222, June 21, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/21/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/21/17/article.
pdf.

Eades:1996:TKB

[ESW96] Peter Eades, Charles Stirk, and Sue Whitesides. The techniques of Komolgorov and Bardzin for three-dimensional orthogonal graph drawings. Information Processing Letters, 60(2): 97–103, October 28, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Edalat:1997:BAI

[ESW97] Abbas Edalat, David Sharp, and Lyndon While. Bounding the attractor of an IFS. Information Processing Letters, 64(4): 197–202, December 13, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Erbas:1992:LCE

[ETA92] Cengiz Erbas, Murat M. Tanik, and Zekeriya Aliyazicioglu. Linear congruence equations for the solutions of the N-queens problem. Information Processing Letters, 41(6):301–306, April 17, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ebert:1997:SGC

[EV97] Jürgen Ebert and Gottfried Vossen. I-serializability: Generalized correctness for transactionbased environments. Information Processing Letters, 63(4): 221–227, September 10, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

El-Yaniv:1995:GTS

[EYK95] Ran El-Yaniv and Jon Kleinberg. Geometric two-server algorithms. Information Processing Letters, 53(6):355–358, March 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fagerberg:1996:GBQ

[Fag96] Rolf Fagerberg. A generalization of binomial queues. Information Processing Letters, 57(2): 109–114, January 29, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Falkowski:1990:PR

[Fal90] Bernd-Jürgen Falkowski. Perceptrons revisited. Information Processing Letters, 36(4):207– 213, November 15, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fernandez-Baca:1991:MHG

[FBW91] David Fernández-Baca and Mark A. Williams. On matroids and hierarchical graphs. Information Processing Letters, 38 (3):117–121, May 17, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feige:1992:CFR

[Fei92] Uriel Feige. On the complexity of finite random functions. Information Processing Letters, 44 (6):295–296, December 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ferragina:1994:SDP

[Fer94] P. Ferragina. Static and dynamic parallel computation of connected components. Information Processing Letters, 50 (2):63–68, April 22, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fernau:1995:NUL

[Fer95] Henning Fernau. A note on uniformly limited ET0L systems with unique interpretation. Information Processing Letters, 54 (4):199–204, May 26, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fields:1990:FAM

[FF90] Malcolm C. Fields and Greg N. Frederickson. A faster algorithm for the maximum weighted tardiness problem. Information Processing Letters, 36(1):39–44, October 1, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fulop:1993:IDT

[FG93] Z. Fülöp and P. Gyenizse. On injectivity of deterministic topdown tree transducers. Information Processing Letters, 48 (4):183–188, November 29, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fischer:1993:AMP

[FGHP93] Ted Fischer, Andrew V. Goldberg, David J. Haglin, and Serge Plotkin. Approximating matchings in parallel. Information Processing Letters, 46(3):115– 118, June 11, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Franciosa:1997:IMD

[FGN97] Paolo G. Franciosa, Giorgio Gambosi, and Umberto Nanni. The incremental maintenance of a Depth-First-Search tree in directed acyclic graphs. Information Processing Letters, 61 (2):113–120, February 28, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fantechi:1991:CBN

[FGR91] Alessandro Fantechi, Stefania Gnesi, and Gioia Ristori. Compositionality and bisimulation. A negative result. Information Processing Letters, 39(2): 109–114, July 31, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Frank:1998:AFS

[FGW98] Jeremy Frank, Ian P. Gent, and Toby Walsh. Asymptotic and finite size parameters for phase transitions: Hamiltonian circuit as a case study. Information Processing Letters, 65(5): 241–245, March 13, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fischer:1994:CMA

[FH94] Paul Fischer and Klaus-Uwe Höffgen. Computing a maximum axis-aligned rectangle in a convex polygon. Information Processing Letters, 51(4):189–193, August 24, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fris:1997:DCC

[FHL97] Ivan Friš, Ivan Havel, and Petr Liebl. The diameter of the cube-connected cycles. Information Processing Letters, 61(3): 157–160, March 14, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Frederickson:1993:SPC

[FHT93] Greg N. Frederickson, Susanne E. Hambrusch, and Hung Yi Tu. Shortest path computations in source-deplanarized graphs. Information Processing Letters, 47(2):71–75, August 20, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fidge:1998:LVT

[Fid98] C. J. Fidge. A limitation of vector timestamps for reconstructing distributed computations. Information Processing Letters, 68(2):87–91, October 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/46/19/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/46/19/18/article.
pdf.

Fischer:1995:NCL

[Fis95] S. T. Fischer. A note on the complexity of local search problems. Information Processing Letters, 53(2):69–75, January 27, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fromentin:1995:FAD

[FJJR95] Eddy Fromentin, Claude Jard, Guy-Vincent Jourdan, and Michel Raynal. On-the-fly analysis of distributed computations. Information Processing Letters, 54(5):267–274, June 9, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fraigniaud:1993:FTS

[FKP93] Pierre Fraigniaud, Claire Kenyon, and Andrzej Pelc. Finding a target subnetwork in sparse networks with random faults. Information Processing Letters, 48 (6):297–303, December 20, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Flouret:1997:NMA

[FL97] M. Flouret and E. Laugerotte.

Noncommutative minimization algorithms. Information Processing Letters, 64(3):123–126, November 29, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Flammini:1994:LMD

[Fla94] Michele Flammini. On the learnability of monotone kμ-DNF formulae under product distributions. Information Processing Letters, 52(3):167–173, November 11, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feuerstein:1998:ETB

[FLMSS98] Esteban Feuerstein, Stefano Leonardi, Alberto Marchetti-Spaccamela, and Nicola Santoro. Efficient token-based control in rings. Information Processing Letters, 66(4):175–180, May 29, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Floren:1991:NFA

[Flo91]  Rolf Floren.  A note on:

“A faster approximation algorithm for the Steiner problem in graphs” [Inform. Process. Lett. 27 (1988), no. 3, 125–128; MR 89d:68031] by K. Mehlhorn. Information Processing Letters, 38 (4):177–178, May 31, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Meh88].

Flocchini:1997:MSD

[Flo97] Paola Flocchini. Minimal sense of direction in regular networks. Information Processing Letters, 61(6):331–338, April 24, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fischetti:1993:CIP

[FLSS93a] Vincent A. Fischetti, Gad M. Landau, Jeanette P. Schmidt, and Peter H. Sellers. Corrigendum: “Identifying periodic occurrences of a template with applications to protein structure”. Information Processing Letters, 46(3):157, June 11, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [FLSS93b].

Fischetti:1993:IPO

[FLSS93b] Vincent A. Fischetti, Gad M.

Landau, Peter H. Sellers, and Jeanette P. Schmidt. Identifying periodic occurrences of a template with applications to protein structure. Information Processing Letters, 45(1):11–18, January 25, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [FLSS93a].

Farach:1997:OPR

[FM97a] Martin Farach and S. Muthukrishnan. Optimal parallel randomized renaming. Information Processing Letters, 61(1):7–10, January 14, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Formica:1997:VAI

[FM97b] Anna Formica and Michele Missikoff. A verification algorithm for inheritance hierarchies in object-oriented databases. Information Processing Letters, 62 (5):269–279, July 2, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fujiwara:1995:OPA

[FMF95] Akihiro Fujiwara, Toshimitsu Masuzawa, and Hideo Fujiwara. An optimal parallel algorithm for the Euclidean distance maps of 2-D binary images. Information Processing Letters, 54(5): 295–300, June 9, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Flocchini:1997:ISD

[FMS97] Paola Flocchini, Bernard Mans, and Nicola Santoro. On the impact of sense of direction on message complexity. Information Processing Letters, 63(1):23–31, July 30, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fischer:1993:SAC

[FMT93] Michael J. Fischer, Shlomo Moran, and Gadi Taubenfeld. Space-efficient asynchronous consensus without shared memory initialization. Information Processing Letters, 45(2):101– 105, February 26, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fung:1990:SLT

[FNTV90] Khun Yee Fung, Tina M. Nicholl, Robert E. Tarjan, and Christopher J. Van Wyk. Simplified linear-time Jordan sorting and polygon clipping. Information Processing Letters, 35 (2):85–92, June 29, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fokkink:1994:CEA

[Fok94] Wan Fokkink. A complete equational axiomatization for prefix iteration. Information Processing Letters, 52(6):333–337, December 23, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fokkink:1997:UIS

[Fok97] Wan Fokkink. Unification for infinite sets of equations between finite terms. Information Processing Letters, 62(4):183–188, June 11, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Foster:1993:GOH

[Fos93] James A. Foster. The generic oracle hypothesis is false. Information Processing Letters, 45(2): 59–62, February 26, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fotouhi:1991:PON

[FP91a] Farshad Fotouhi and Sakti Pramanik. Problem of optimizing the number of block accesses in performing relational join is NP-hard. Information Processing Letters, 38(5):271–275, June 14, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fraigniaud:1991:BHW

[FP91b] Pierre Fraigniaud and Claudine Peyrat. Broadcasting in a hypercube when some calls fail. Information Processing Letters, 39 (3):115–119, August 16, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Frazier:1994:PGA

[FP94] Michael Frazier and C. David Page, Jr. Prefix grammars: An alternative characterization of the regular languages. Information Processing Letters, 51 (2):67–71, July 26, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fekete:1998:TBM

[FP98] Sándor P. Fekete and William R. Pulleyblank. Traveling the boundary of Minkowski sums. Information Processing Letters, 66(4):171–174, May 29, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fudos:1996:POR

[FPS96] Ioannis Fudos, Evaggelia Pitoura, and Wojciech Szpankowski. On pattern occurrences in a random text. Information Processing Letters, 57(6):307–312, March 25, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Farach:1995:AMT

[FPT95] Martin Farach, Teresa M. Przytycka, and Mikkel Thorup. On the agreement of many trees. Information Processing Letters, 55(6):297–301, September 29, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fouquet:1997:TAM

[FPT97] J. L. Fouquet, I. Parfenoff, and H. Thuillier. An O(n) time algorithm for maximum matching in P_{4}-tidy graphs. Information Processing Letters, 62(6): 281–287, July 16, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fraigniaud:1998:HBN

[Fra98] Pierre Fraigniaud. Hierarchical broadcast networks. Information Processing Letters, 68 (6):303–305, December 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Franzblau:1999:EDB

[Fra99]  D. S. Franzblau.  Ear de-

composition with bounds on ear length. Information Processing Letters, 70(5):245–249, June 21, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/21/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/21/22/article.
pdf.

Frederickson:1983:SUT

[Fre83] Greg N. Frederickson. Scheduling unit-time tasks with integer release times and deadlines. Information Processing Letters, 16 (4):171–173, May 13, 1983. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See note [SY93].

Fraenkel:1994:EMD

[FRS94] Aviezri S. Fraenkel, Edward M. Reingold, and Prashant Saxena. Efficient management of dynamic tables. Information Processing Letters, 50(1):25–30, April 8, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fortnow:1992:PTL

[FS92] Lance Fortnow and Márió Szegedy. On the power of twolocal random reductions. Information Processing Letters, 44 (6):303–306, December 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Frandsen:1997:DMM

[FS97] Gudmund S. Frandsen and Sven Skyum. Dynamic maintenance of majority information in constant time per update. Information Processing Letters, 63(2): 75–78, August 13, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feder:1998:OCA

[FS98] Tomás Feder and Sunil M. Shende. Online channel allocation in FDMA networks with reuse constraints. Information Processing Letters, 67(6):295– 302, September 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fujita:1995:NSM

[Fuj95] Satoshi Fujita. A note on the size of a multicast tree in hypercubes. Information Processing Letters, 54(4):223–227, May 26, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fujito:1996:NAV

[Fuj96] Toshihiro Fujito. A note on approximation of the vertex cover and feedback vertex set problems—unified approach. Information Processing Letters, 59 (2):59–63, July 22, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fujita:1998:QBM

[Fuj98] Satoshi Fujita. A quorum based k-mutual exclusion by weighted k-quorum systems. Information Processing Letters, 67(4): 191–197, August 31, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fulop:1990:EPU

[FV90] Z. Fülöp and S. Vágvölgyi.

The emptiness problem is undecidable for domains of partial monadic 2-modular tree transformations. Information Processing Letters, 36(6):293–296, December 15, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feijs:1997:ADT

[Fv97] L. M. G. Feijs and R. C. van Ommering. Abstract derivation of transitive closure algorithms. Information Processing Letters, 63(3):159–164, August 27, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Franzle:1995:GNS

[FvW95] Martin Fränzle, Bernhard von Stengel, and Arne Wittmüss. A generalized notion of semantic independence. Information Processing Letters, 53(1):5–9, January 13, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fleischner:1992:DCT

[FW92] Herbert Fleischner and Gerhard J. Woeginger. Detecting cycles through three fixed vertices in a graph. Information Processing Letters, 42(1):29–33, April 27, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fujita:1993:FGS

[FY93] Satoshi Fujita and Masafumi Yamashita. Fast gossiping on square mesh computers. Information Processing Letters, 48 (3):127–130, November 19, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fraczak:1998:NSO

[FZ98] Wojciech Fraczak and Marek B. Zaremba. A non-SOS operational semantics for a process algebra. Information Processing Letters, 68(1):47–54, October 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garg:1991:FPN

[GAG91] M. L. Garg, S. I. Ahson, and P. V. Gupta. A fuzzy Petri net for knowledge representation and reasoning. Information Processing Letters, 39(3):165–171, August 16, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gaines:1997:PME

[Gai97] J. G. Gaines. Partitions with minimum entropy of regions in R^{2}. Information Processing Letters, 63(2):69–73, August 13, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Galvez:1994:NPL

[Gál94] José Fortes Gálvez. A note on a proposed LALR parser for extended context-free grammars. Information Processing Letters, 50(6):303–305, June 27, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gal:1997:SFR

[Gál97] Anna Gál. A simple function that requires exponential size read-once branching programs. Information Processing Letters, 62(1):13–16, May 7, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garg:1992:SOA

[Gar92] Vijay K. Garg. Some optimal algorithms for decomposed partially ordered sets. Information Processing Letters, 44(1):39–43, November 9, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gartner:1993:RRL

[Gär93] Stefan Gärtner. A remark on the regulation of klETOL systems. Information Processing Letters, 48(2):83–85, November 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gasarch:1991:SLR

[Gas91] William I. Gasarch. On selecting the k largest with restricted quadratic queries. Information Processing Letters, 38 (4):193–195, May 31, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gavalda:1993:PRP

[Gav93a] Ricard Gavaldà. A positive relativization of polynomial time versus polylog space. Information Processing Letters, 46(3): 119–123, June 11, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gavril:1993:ESG

[Gav93b] Fănică Gavril. An efficiently solvable graph partition problem to which many problems are reducible. Information Processing Letters, 45(6):285–290, April 16, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gawiejnowicz:1996:NSS

[Gaw96] Stanisław Gawiejnowicz. A note on scheduling on a single processor with speed dependent on a number of executed jobs. Information Processing Letters, 57 (6):297–300, March 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gonnet:1990:AKS

[GBY90] Gaston H. Gonnet and Ricardo A. Baeza-Yates. An analysis of the Karp-Rabin string matching algorithm. Information Processing Letters, 34(5): 271–274, May 7, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Geist:1997:PBM

[Gei97]  Robert Geist.  Performance

bounds for modeling NUMA architectures. Information Processing Letters, 63(2):113–117, August 13, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gemignani:1992:FIH

[Gem92] Luca Gemignani. Fast inversion of Hankel and Toeplitz matrices. Information Processing Letters, 41(3):119–123, March 6, 1992.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gergov:1994:TST

[Ger94] Jordan Gergov. Time-space tradeoffs for integer, multiplication on various types of input oblivious sequential machines. Information Processing Letters, 51(5):265–269, September 12, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gergatsoulis:1997:UFT

[Ger97] Manolis Gergatsoulis. Unfold/ fold transformations for disjunctive logic programs. Information Processing Letters, 62(1):23–29, May 7, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ghosh:1996:EFC

[GG96] Sukumar Ghosh and Arobinda Gupta. An exercise in faultcontainment: self-stabilizing leader election. Information Processing Letters, 59(5):281– 288, September 9, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldmann:1994:ATH

[GGH94] Mikael Goldmann, Per Grape, and Johan Håstad. On average time hierarchies (turing machines). Information Processing Letters, 49(1):15–20, January 14, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goralcik:1991:APF

[GGK91] P. Goralčík, A. Goralčíková, and V. Koubek. Alternation with a pebble. Information Processing Letters, 38(1):7–13, April 12, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gilbert:1998:ASR

[GGOQ98] Henri Gilbert, Dipankar Gupta, Andrew Odlyzko, and Jean-Jacques Quisquater. Attacks on Shamir’s ‘RSA for paranoids’. Information Processing Letters, 68(4):197–199, November 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.research.att.com/~amo/
doc/rsa.for.paranoids.pdf;
 http://www.research.att.
com/~amo/doc/rsa.for.paranoids.
ps;   http://www.research.
att.com/~amo/doc/rsa.for.
paranoids.tex.

Gouda:1990:SUC

[GH90] Mohamed G. Gouda and Ted Herman. Stabilizing unison. Information Processing Letters, 35 (4):171–175, August 7, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldmann:1992:SLB

[GH92] Mikael Goldmann and Johan Håstad. A simple lower bound for monotone clique using a communication game. Information Processing Letters, 41(4):

221–226, March 18, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gemmell:1994:TBE

[GH94] Peter Gemmell and Mor Harchol. Tight bounds on expected time to add correctly and add mostly correctly. Information Processing Letters, 49(2):77–83, January 28, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldsmith:1996:SIP

[GH96] Judy Goldsmith and Steven Homer. Scalability and the Isomorphism Problem. Information Processing Letters, 57(3):137– 143, February 12, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldreich:1998:CIP

[GH98] Oded Goldreich and Johan Håstad. On the complexity of interactive proofs with bounded communication. Information Processing Letters, 67(4):205– 214, August 31, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ghosh:1991:BSD

[Gho91] Sukumar Ghosh. Binary selfstabilization in distributed systems. Information Processing Letters, 40(3):153–159, November 8, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gopalakrishnan:1993:NCC

[GHS93] K. Gopalakrishnan, D. G. Hoffman, and D. R. Stinson. A note on a conjecture concerning symmetric resilient functions. Information Processing Letters, 47(3):139–143, September 14, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldschmidt:1993:MGH

[GHY93] Olivier Goldschmidt, Dorit S.

Hochbaum, and Gang Yu. A modified greedy heuristic for the Set Covering problem with improved worst case bound. Information Processing Letters, 48 (6):305–310, December 20, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Galil:1991:NSU

[GI91] Zvi Galil and Giuseppe F. Italiano. A note on set union with arbitrary deunions. Information Processing Letters, 37(6): 331–335, March 28, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gianuzzi:1994:DTD

[Gia94] Vittoria Gianuzzi. Distributed termination detection in reducible communication graphs. Information Processing Letters, 49(1):1–8, January 14, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Giakoumakis:1996:LGN

[Gia96] Vassilis Giakoumakis. P_{4}-laden graphs: A new class of brittle graphs. Information Processing Letters, 60(1):29–36, October 14, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ghandeharizadeh:1996:LAO

[GIZ96] Shahram Ghandeharizadeh, Doug Ierardi, and Roger Zimmermann. An on-line algorithm to optimize file layout in a dynamic environment. Information Processing Letters, 57(2):75–81, January 29, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Greenberg:1992:DMC

[GJK92] Ronald Greenberg, Joseph JáJá, and Sridhar Krishnamurthy. On the difficulty of Manhattan channel routing. Information Processing Letters, 44(5):281– 284, December 21, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garrido:1996:SRP

[GJLR96] Oscar Garrido, Stefan Jarominek, Andrzej Lingas, and Wojciech Rytter. A simple randomized parallel algorithm for maximal f-matchings. Information Processing Letters, 57(2):83–87, January 29, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gupta:1997:TAR

[GJS97] Prosenjit Gupta, Ravi Janardan, and Michiel Smid. A technique for adding range restrictions to generalized searching problems. Information Processing Letters, 64(5):263–269, December 23, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gupta:1999:EAC

[GJS99] Prosenjit Gupta, Ravi Janardan, and Michiel Smid. Efficient algorithms for counting and reporting pairwise intersections between convex polygons. Information Processing Letters, 69 (1):7–13, January 15, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/17/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/17/18/article.
pdf.

Garcia:1998:EAL

[GJT98] Alfredo García, Pedro Jodrá, and Javier Tejel. An efficient algorithm for on-line searching of minima in Monge pathdecomposable tridimensional arrays. Information Processing Letters, 68(1):3–9, October 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gluck:1997:RSG

[GK97] Robert Glück and Andrei Klimov. A regeneration scheme for generating extensions. Information Processing Letters, 62 (3):127–134, June 4, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gordon:1998:SMS

[GK98a] Valery Gordon and Wieslaw Kubiak. Single machine scheduling with release and due date assignment to minimize the weighted number of late jobs. Information Processing Letters, 68(3): 153–159, November 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/ipl/cas_sub/browse/
browse.cgi?year=1998&volume=
68&issue=3&aid=2079.

Graf:1998:SDQ

[GK98b] T. Graf and V. Kamakoti. Sparse dominance queries for many points in optimal time and space. Information Processing Letters, 64(6):287–291, January 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garrido:1996:SAM

[GKL96] Oscar Garrido, Pierre Kelsen, and Andrzej Lingas. A simple NC-algorithm for a maximal independent set in a hypergraph of poly-log arboricity. Information Processing Letters, 58(2):55–58, May 13, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gens:1998:ABS

[GL98] George Gens and Eugene Levner. An approximate binary search algorithm for the multiple-choice knapsack problem. Information Processing Letters, 67(5):261–265, September 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guijarro:1999:ELW

[GLR99] David Guijarro, Víctor Lavín, and Vijay Raghavan. Exact learning when irrelevant variables abound. Information Processing Letters, 70(5):233– 239, June 21, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/21/
20/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/21/20/article.
pdf.

Gusfield:1992:EAA

[GLS92] Dan Gusfield, Gad M. Landau, and Baruch Schieber. An efficient algorithm for the All Pairs Suffix — Prefix Problem. Information Processing Letters, 41 (4):181–185, March 18, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Govindan:1998:APO

[GLY98] Rajeev Govindan, Michael A.

Langston, and Xudong Yan. Approximating the pathwidth of outerplanar graphs. Information Processing Letters, 68(1):17–23, October 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ghosh:1990:OAC

[GM90] Subir Kumar Ghosh and Anil Maheshwari. An optimal algorithm for computing a minimum nested nonconvex polygon. Information Processing Letters, 36 (6):277–280, December 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ghosh:1991:BHS

[GM91] Subrata Ghosh and Ambuj Mahanti. Bidirectional heuristic search with limited resources. Information Processing Letters, 40(6):335–340, December 30, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ghosh:1992:OPA

[GM92a] Subir Kumar Ghosh and Anil Maheshwari. An optimal parallel algorithm for computing furthest neighbors in a tree. Information Processing Letters, 44 (3):155–160, November 30, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goles:1992:ANO

[GM92b] Eric Goles and Servet Martínez. Automata networks and optimization. Information Processing Letters, 42(6):339–343, July 24, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Greenberg:1992:MSS

[GM92c] Ronald I. Greenberg and F. Miller Maley. Minimum separation for single-layer channel routing. Information Processing Letters, 43(4):201–205, September 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gergov:1994:CAM

[GM94a] Jordan Gergov and Christoph Meinel. On the complexity of analysis and manipulation of Boolean functions in terms of decision graphs. Information Processing Letters, 50(6):317–322, June 27, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gil:1994:DAE

[GM94b] Joseph Gil and Yossi Matias.

Designing algorithms by expectations. Information Processing Letters, 51(1):31–34, July 12, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garg:1997:DCG

[GM97] Vijay K. Garg and J. Roger Mitchell. Detecting conjunctions of global predicates. Information Processing Letters, 63(6): 295–302, October 8, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gabarro:1998:PDL

[GM98] Joaquim Gabarró and Xavier Messeguer. Parallel dictionaries with local rules on AVL and brother trees. Information Processing Letters, 68(2): 79–85, October 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/46/19/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/46/19/17/article.
pdf.

Galbiati:1994:SNA

[GMM94] G. Galbiati, F. Maffioli, and A. Morzenti. A short note on the approximability of the Maximum Leaves Spanning Tree Problem. Information Processing Letters, 52(1):45–49, October 14, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guaiana:1998:EWP

[GMPW98] G. Guaiana, R. Meyer, A. Petit, and P. Weil. An extension of the wreath product principle for finite Mazurkiewicz traces. Information Processing Letters, 67(6):277–282, September 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grigg:1990:SMP

[GMR90] Ken Grigg, Serge Miguet, and Yves Robert. Symmetric matrixvector product on a ring of processors. Information Processing Letters, 35(5):239–248, August 28, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldreich:1999:ASL

[GMSS99] O. Goldreich, D. Micciancio, S. Safra, and J.-P. Seifert. Approximating shortest lattice vectors is not harder than approximating closest lattice vectors. Information Processing Letters, 71(2):55–61, July 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/18/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/18/18/article.
pdf.

Gobel:1999:RTU

[Göb99] Manfred Göbel. A rewriting technique for universal polynomial invariants. Information Processing Letters, 69(6):271– 273, March 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/22/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/22/18/article.
pdf.

Goeman:1990:TTS

[Goe90] Henk Goeman. Towards a theory of (self) applicative communicating processes. A short note. Information Processing Letters, 34(3):139–142, April 9, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldreich:1990:NCI

[Gol90] Oded Goldreich. A note on computational indistinguishability. Information Processing Letters, 34(6):277–281, May 28, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldberg:1991:PIM

[Gol91] Andrew V. Goldberg. Processorefficient implementation of a maximum flow algorithm. Information Processing Letters, 38 (4):179–185, May 31, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldberg:1992:NRS

[Gol92] Andrew V. Goldberg. A natural randomization strategy for multicommodity flow and related algorithms. Information Processing Letters, 42(5):249–256, July 3, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldmann:1995:NPM

[Gol95a] Mikael Goldmann. A note on the power of majority gates and modular gates. Information Processing Letters, 53(6):321–327, March 24, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldwurm:1995:RGW

[Gol95b] Massimiliano Goldwurm. Random generation of words in an algebraic language in linear binary space. Information Processing Letters, 54(4):229–233, May 26, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldmann:1997:PTG

[Gol97] Mikael Goldmann. On the power of a threshold gate at the top. Information Processing Letters, 63(6):287–293, October 8, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Golic:1999:SCE

[Gol99]  J. Dj. Golic´ .  Stream ci-

pher encryption of random access files. Information Processing Letters, 69(3):145–148, February 12, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/19/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/19/25/article.
pdf.

Gonzalez:1991:CSP

[Gon91] Teofilo F. Gonzalez. Covering a set of points in multidimensional space. Information Processing Letters, 40(4):181–188, November 25, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gong:1995:CKH

[Gon95a] Li Gong. Collisionful keyed hash functions with selectable collisions. Information Processing Letters, 55(3):167–170, August 11, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gonzalez:1995:SLF

[Gon95b] Teofilo F. Gonzalez. A simple LP-free approximation algorithm for the minimum weight vertex cover problem. Information Processing Letters, 54(3): 129–131, May 12, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gonzalez:1998:IAA

[Gon98] Teofilo F. Gonzalez. Improved approximation algorithms for embedding hyperedges in a cycle. Information Processing Letters, 67(5):267–271, September 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gastin:1992:DSP

[GOPR92] Paul Gastin, Edward Ochmański, Antoine Petit, and Brigitte Rozoy. Decidability of the Star Problem in A ∗ ×{b}∗. Information Processing Letters, 44(2): 65–71, November 19, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gouda:1994:SOD

[Gou94] Mohamed G. Gouda. Stabilizing observers. Information Processing Letters, 52(2):99–103, October 28, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Galil:1990:LAC

[GP90a] Zvi Galil and Kunsoo Park. A linear-time algorithm for concave one-dimensional dynamic programming. Information Processing Letters, 33(6):309–311, February 10, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldreich:1990:BBW

[GP90b] Oded Goldreich and Erez Petrank. The best of both worlds: guaranteeing termination in fast randomized Byzantine Agreement protocols. Information Processing Letters, 36(1):45–49, October 1, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gawiejnowicz:1995:SJV

[GP95a] Stanislaw Gawiejnowicz and Lidia Pankowska. Scheduling jobs with varying processing times. Information Processing Letters, 54(3):175–178, May 12, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gu:1995:NNC

[GP95b] Qian Ping Gu and Shietung Peng. Node-to-node cluster fault tolerant routing in star graphs. Information Processing Letters, 56(1):29–35, October 13, 1995.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Germain:1996:TSP

[GP96] Christian Germain and Jean Pallo. Two shortest path metrics on well-formed parentheses strings. Information Processing Letters, 60(6):283–287, December 23, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gu:1997:NSD

[GP97] Qian-Ping Gu and Shietung Peng. Node-to-set disjoint paths problem in star graphs. Information Processing Letters, 62(4): 201–207, June 11, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grabner:1998:ASR

[GP98a] Peter J. Grabner and Helmut Prodinger. An asymptotic study of a recursion occurring in the analysis of an algorithm on broadcast communication. Information Processing Letters, 65 (2):89–93, January 29, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gu:1998:EAP

[GP98b] Qian-Ping Gu and Shietung Peng. An efficient algorithm for k-pairwise disjoint paths in star graphs. Information Processing Letters, 67(6):283–287, September 30, 1998. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Galiev:1992:DDR

[GPS92] V. I. Galiev, A. F. Polupanov, and I. E. Shparlinski. Distances from differences of roots of polynomials to the nearest integers. Information Processing Letters, 43(3):143–146, September 14, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gyssens:1990:HCN

[GPV90] Marc Gyssens, Jan Paredaens, and Dirk Van Gucht. On a hierarchy of classes for nested databases. Information Processing Letters, 36(5):259–266, December 1, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldreich:1997:ULA

[GR97] Oded Goldreich and Dana Ron.

On universal learning algorithms. Information Processing Letters, 63(3):131–136, August 27, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guruswami:1998:NFO

[GR98] Venkatesan Guruswami and C. Pandu Rangan. A natural family of optimization problems with arbitrarily small approximation thresholds. Information Processing Letters, 68(5):241– 248, December 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gradel:1990:SIA

[Grä90] Erich Grädel. Simple interpretations among complicated theories. Information Processing Letters, 35(5):235–238, August 28, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grable:1996:NPH

[Gra96a] David A. Grable. Nearly-perfect hypergraph packing is in NC. Information Processing Letters, 60 (6):295–299, December 23, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grandi:1996:INP

[Gra96b] Piercarlo Grandi. Implementing (nondeterministic) parallel assignments. Information Processing Letters, 58(4):177–179, May 27, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gray:1997:CJV

[Gra97] James W. Gray, III. On the Clark-Jacob version of SPLICE/ AS. Information Processing Letters, 62(5):251–254, July 2, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Green:1991:OS

[Gre91] Frederic Green. Oracle separating ⊕P from P P^{P H}. Information Processing Letters, 37 (3):149–153, February 18, 1991.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Greenwood:1995:EME

[Gre95] Garrison W. Greenwood. On the equity of mutual exclusion algorithms in distributed systems. Information Processing Letters, 56(1):19–22, October 13, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Greco:1998:ETF

[Gre98] G. Greco. Embeddings and the trace of finite sets. Information Processing Letters, 67(4): 199–203, August 31, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grossi:1991:FCS

[Gro91a] Roberto Grossi. Further comments on the subtree isomorphism for ordered trees: “On the subtree isomorphism problem for ordered trees” [Inform. Process. Lett. 32 (1989), no. 5, 271–273; MR 90k:68139] by E. Mäkinen. Information Processing Letters, 40(5):255–256, December 13, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Mäk89].

Grossi:1991:NSI

[Gro91b] Roberto Grossi. A note on the subtree isomorphism for ordered trees and related problems. Information Processing Letters, 39 (2):81–84, July 31, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Groote:1992:SPD

[Gro92] Jan Friso Groote. A short proof of the decidability of bisimulation for normed BPA-processes. Information Processing Letters, 42(3):167–171, May 25, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grolmusz:1998:LBD

[Gro98] Vince Grolmusz. A lower bound for depth-3 circuits with MOD m gates. Information Processing Letters, 67(2):87–90, July 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gunther:1990:PIE

[GS90] Neil J. Gunther and John G. Shaw. Path integral evaluation of ALOHA network transients. Information Processing Letters, 33(6):289–295, February 10, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guha:1991:ETA

[GS91] Sumanta Guha and Arunabha Sen. Expected time analysis of interpolation merge — A simple new merging algorithm. Information Processing Letters, 40 (5):277–281, December 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gabrani:1992:NRB

[GS92a] Naveen Gabrani and Priti Shankar. A note on the reconstruction of a binary tree from its traversals. Information Processing Letters, 42(2):117–119, May 11, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gemmell:1992:HRC

[GS92b] Peter Gemmell and Madhu Sudan. Highly resilient correctors for polynomials. Information Processing Letters, 43(4): 169–174, September 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Golin:1993:Q

[GS93] Mordecia J. Golin and Robert Sedgewick. Queue-mergesort. Information Processing Letters, 48(5):253–259, December 10, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gries:1995:EPL

[GS95] David Gries and Fred B. Schneider. Equational propositional logic. Information Processing Letters, 53(3):145–152, February 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Glaister:1996:LBT

[GS96a] Ian Glaister and Jeffrey Shallit. A lower bound technique for the size of nondeterministic finite automata. Information Processing Letters, 59(2):75–77, July 22, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gopalakrishnan:1996:SAE

[GS96b] K. Gopalakrishnan and D. R. Stinson. A simple analysis of the error probability of two-point based sampling. Information Processing Letters, 60(2):91–96, October 28, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gartner:1998:EPS

[GS98a] Bernd Gärtner and Sven Schönherr. Exact primitives for smallest enclosing ellipses. Information Processing Letters, 68(1):33–38, October 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gavril:1998:IGA

[GS98b] Fǎnicǎ Gavril and Oded Shmueli. Intersection graphs of k-acyclic families of subtrees and relational database query processing. Information Processing Letters, 66(1):1–6, April 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Golic:1997:FCA

[GSSD97] J. Dj. Golić, M. Salmasizadeh, L. Simpson, and E. Dawson. Fast correlation attacks on nonlinear filter generators. Information Processing Letters, 64(1):

37–42, October 22, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grinstead:1993:EED

[GSSH93] Dana L. Grinstead, Peter J.

Slater, Naveed A. Sherwani, and Nancy D. Holmes. Efficient edge domination problems in graphs. Information Processing Letters, 48(5):221–228, December 10, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garcia:1996:UTM

[GT96] Alfredo García and Javier Tejel.

Using total monotonicity for two optimization problems on the plane. Information Processing Letters, 60(1):13–17, October 14, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guan:1998:OMR

[Gua98] D. J. Guan. An optimal message routing algorithm for doubleloop networks. Information Processing Letters, 65(5):255–260, March 13, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guenther:1990:EEF

[Gue90] G. R. Guenther. Efficient expansion of factored expressions. Information Processing Letters, 35 (2):69–72, June 29, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gumm:1993:AGA

[Gum93] H. Peter Gumm. Another glance at the Alpern-Schneider characterization of safety and liveness in concurrent executions. Information Processing Letters, 47 (6):291–294, October 18, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gupta:1992:DPE

[Gup92a] Gopal Gupta. Dynamic parallel evaluation of the cross-product set using time-stamps. Information Processing Letters, 44 (5):273–280, December 21, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gupta:1992:CCF

[Gup92b] Sanjay Gupta. On the closure of certain function classes under integer division by polynomiallybounded functions. Information Processing Letters, 44(4): 205–210, December 10, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gupta:1993:BPO

[Gup93] Sanjay Gupta. On boundedprobability operators and C_{=}P. Information Processing Letters, 48(2):93–98, November 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gupta:1995:GD

[Gup95] Rajiv Gupta. Generalized dominators. Information Processing Letters, 53(4):193–200, February 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gurla:1993:LOC

[Gur93] Himabindu Gurla. Leftmost one computation on meshes with row broadcasting. Information Processing Letters, 47(5):261–266, October 8, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gurla:1994:LOC

[Gur94] Himabindu Gurla. Leftmost one computation on meshes with row broadcasting. Information Processing Letters, 50(2):111, April 22, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gusfield:1994:FIS

[Gus94] Dan Gusfield. Faster implementation of a shortest superstring approximation. Information Processing Letters, 51(5): 271–274, September 12, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garcia-Villalba:1999:GCN

[GVFS99] L. J. García-Villalba and A. Fúster-Sabater. On the general classification of nonlinear filters of m-sequences. Information Processing Letters, 69(5):

227–232, March 12, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/21/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/21/18/article.
pdf.

Gargano:1993:FTR

[GVV93] L. Gargano, U. Vaccaro, and A. Vozella. Fault tolerant routing in the star and pancake interconnection networks. Information Processing Letters, 45(6): 315–320, April 16, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goltz:1996:MCA

[GW96] Ursula Goltz and Heike Wehrheim. Modelling causality via action dependencies in branching time semantics. Information Processing Letters, 59(4):179–184, August 26, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gibbons:1991:BBT

[GZ91] Alan Gibbons and Ridha Ziani. The balanced binary tree technique on mesh-connected computers. Information Processing Letters, 37(2):101–109, January 31, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guan:1997:CPW

[GZ97] D. J. Guan and Xuding Zhu. A coloring problem for weighted graphs. Information Processing Letters, 61(2):77–81, February 28, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hachem:1993:AAP

[Hac93]  Nabil I. Hachem.  An ap-

proximate analysis of the performance of extendible hashing with elastic buckets. Information Processing Letters, 48(1): 13–20, October 29, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hagerup:1992:CTR

[Hag92] Torben Hagerup. On a compaction theorem of Ragde. Information Processing Letters, 43 (6):335–340, October 19, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hagerup:1995:PCI

[Hag95] Torben Hagerup. The parallel complexity of integer prefix summation. Information Processing Letters, 56(1):59–64, October 13, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hall:1993:UOE

[Hal93a] C. V. Hall. Using overloading to express distinctions between evaluators. Information Processing Letters, 48(1):1–8, October 29, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Halldorsson:1993:AMM

[Hal93b] Magnús M. Halldórsson. Approximating the minimum maximal independence number. Information Processing Letters, 46 (4):169–172, June 25, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Halldorsson:1993:SBP

[Hal93c] Magnús M. Halldórsson. A still better performance guarantee for approximate graph coloring. Information Processing Letters, 45(1):19–23, January 25, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hamdi:1995:TPD

[Ham95] Mounir Hamdi. Topological properties of the directional hypercube. Information Processing Letters, 53(5):277–286, March 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Honma:2011:EAS

[HAM11] Hirotoshi Honma, Kodai Abe, and Shigeru Masuyama. Erratum and addendum to “A linear time algorithm for finding all hinge vertices of a permutation graph” [Information Processing Letters 59 (2) (1996) 103–107]. Information Processing Letters, 111(18):891–894, September 30, 2011. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

sciencedirect.com/science/

article/pii/S0020019011000238.

See [HWJ96].

Handley:1991:SED

[Han91] C. C. Handley. A space efficient distributive sort. Information Processing Letters, 37(2):75–78, January 31, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Han:1992:BPB

[Han92] Jiawei Han. Binding propagation beyond the reach of rule/ goal graphs. Information Processing Letters, 42(5):263–268, July 3, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hancart:1993:SSS

[Han93] Christophe Hancart. On Simon’s string searching algorithm. Information Processing Letters, 47 (2):95–99, August 20, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Han:1995:IPC

[Han95] Yijie Han. An improvement on parallel computation of a maximal matching. Information Processing Letters, 56(6):343–348, December 22, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Harper:1994:SAP

[Har94] Robert Harper. A simplified account of polymorphic references. Information Processing Letters, 51(4):201–206, August 24, 1994.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also note [Har96].

Harper:1996:NSA

[Har96] Robert Harper. A note on: “A simplified account of polymorphic references” [Inform. Process. Lett. 51 (1994), no. 4, 201– 206; MR 95f:68142]. Information Processing Letters, 57(1): 15–16, January 15, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Har94].

Harms:1998:SOA

[Har98] Janelle J. Harms. A simple optimal algorithm for scheduling variable-sized requests. Information Processing Letters, 68 (6):291–293, December 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hur:1990:SS

[HC90] Jin Ho Hur and Kilnam Chon. Self and selftype. Information Processing Letters, 36(5): 225–230, December 1, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:1992:SSA

[HC92] Shing Tsaan Huang and Nian Shing Chen. A self-stabilizing algorithm for constructing breadthfirst trees. Information Processing Letters, 41(2):109–117, February 14, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hwang:1995:SSA

[HC95] Tzonelih Hwang and Yung-Hsiang Chen. On the security of SPLICE/AS — the authentication system in WIDE Internet. Information Processing Letters, 53(2):97–101, January 27, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1998:TSG

[hC98] Wen huei Chen. Test sequence generation from the protocol data portion based on the Selecting Chinese Postman algorithm. Information Processing Letters, 65(5):261–268, March 13, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ho:1999:SAP

[HC99a] Chin-Wen Ho and Jou-Ming Chang. Solving the allpairs-shortest-length problem on chordal bipartite graphs. Information Processing Letters, 69(2):87–93, January 29, 1999. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/18/
23/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/18/23/article.
pdf.

Hwang:1999:NCC

[HC99b] Shien-Ching Hwang and Gen-Huey Chen. A note on cycliccubes. Information Processing Letters, 72(3–4):131–135, November 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/28/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/28/34/article.
pdf.

Horng:1991:OSA

[HCF91] Shi-Jinn Horng, Wen-Tsuen Chen, and Ming-Yi Fang. Optimal speed-up algorithms for template matching on SIMD hypercube multiprocessors with restricted local memory. Information Processing Letters, 38 (1):29–37, April 12, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hu:1994:NCA

[HCH94] X. D. Hu, P. D. Chen, and F. K. Hwang. A new competitive algorithm for the counterfeit coin problem. Information Processing Letters, 51(4):213–218, August 24, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hwang:1996:AES

[HCY96] Shin-Jia Hwang, Chin-Chen Chang, and Wei-Pang Yang. Authenticated encryption schemes with message linkage. Information Processing Letters, 58(4): 189–194, May 27, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huizing:1991:IDC

[Hd91] C. Huizing and W. P. de Roever.

Introduction to design choices in the semantics of Statecharts. Information Processing Letters, 37 (4):205–213, February 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

He:1991:EPA

[He91] Xin He. An efficient parallel algorithm for finding minimum weight matching for points on a convex polygon. Information Processing Letters, 37(2): 111–116, January 31, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Heath:1990:CSA

[Hea90] Lenwood S. Heath. Covering a set with arithmetic progressions is NP-complete. Information Processing Letters, 34(6): 293–298, May 28, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hebrard:1995:UHR

[Héb95] Jean-Jacques Hébrard. Unique Horn renaming and Unique 2-Satisfiability. Information Processing Letters, 54(4):235–239, May 26, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Heckmann:1991:LUP

[Hec91] Reinhold Heckmann. Lower and upper power domain constructions commute on all cpos. Information Processing Letters, 40 (1):7–11, October 11, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Heffernan:1991:TSM

[Hef91] Paul J. Heffernan. The translation square map and approximate congruence. Information Processing Letters, 39(3): 153–159, August 16, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hemaspaandra:1994:CTC

[Hem94] Edith Hemaspaandra. Census techniques collapse space classes. Information Processing Letters, 51(2):79–84, July 26, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Henzinger:1992:SST

[Hen92] Thomas A. Henzinger. Sooner is safer than later (reactive systems). Information Processing Letters, 43(3):135–141, September 14, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Henk:1997:NSN

[Hen97] Martin Henk. Note on shortest and nearest lattice vectors. Information Processing Letters, 61(4):183–188, March 27, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Herman:1990:PS

[Her90]  Ted Herman.  Probabilistic

self-stabilization. Information Processing Letters, 35(2):63–67, June 29, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Herley:1991:NTD

[Her91a] Kieran T. Herley. A note on the token distribution problem. Information Processing Letters, 38 (6):329–334, June 28, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hershberger:1991:NDS

[Her91b]  John Hershberger.  A new

data structure for shortest path queries in a simple polygon. Information Processing Letters, 38 (5):231–235, June 14, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hershberger:1993:FAT

[Her93] John Hershberger. A faster algorithm for the two-center decision problem. Information Processing Letters, 47(1):23–29, August 9, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

HernandezBarrera:1996:ADC

[Her96] Antonio Hernández Barrera. Algorithms for deciding the containment of polygons. Information Processing Letters, 59 (5):261–265, September 9, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hermo:1997:CUC

[Her97] Montserrat Hermo. Compressibility and uniform complexity. Information Processing Letters, 62(5):259–264, July 2, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Herrmann:1998:TAR

[Her98] Philippe Herrmann. Timed automata and recognizability. Information Processing Letters, 65 (6):313–318, March 27, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hesselink:1991:RKU

[Hes91] Wim H. Hesselink. Repetitions, known or unknown? Information Processing Letters, 40(1): 51–57, October 11, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hesselink:1998:ICH

[Hes98] Wim H. Hesselink. Invariants for the construction of a handshake register. Information Processing Letters, 68(4):173–177, November 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Herman:1995:SPC

[HG95] Ted Herman and Sukumar Ghosh. Stabilizing phase-clocks. Information Processing Letters, 54(5):259–265, June 9, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:1991:CRA

[HH91] Nen Fu Huang and Ching Ho Huang. Complexity of the repeaters allocating problem. Information Processing Letters, 40 (1):13–20, October 11, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hsu:1992:SSA

[HH92] Su Chu Hsu and Shing Tsaan Huang. A self-stabilizing algorithm for maximal matching. Information Processing Letters, 43 (2):77–81, August 24, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hassin:1994:MNU

[HH94] Refael Hassin and Shlomo Lahav (Haddad). Maximizing the number of unused colors in the vertex coloring problem. Information Processing Letters, 52(2):87–90, October 28, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Halava:1997:GPZ

[HHI97] Vesa Halava, Tero Harju, and Lucian Ilie. On a geometric problem of zigzags. Information Processing Letters, 62(1):1– 4, May 7, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hofmeister:1991:SNT

[HHK91] Thomas Hofmeister, Walter Hohberg, and Susanne Köhling.

Some notes on threshold circuits, and multiplication in depth 4. Information Processing Letters, 39(4):219–225, August 30, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:1995:EBP

[HHL95] Chung-Ming Huang, Jenq-Muh Hsu, and Shiun-Wei Lee. ECFSM-based probabilistic protocol verification. Information Processing Letters, 55(1):1–9, July 7, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:1999:OEC

[HHL99] Chien-Hung Huang, Ju-Yuan Hsiao, and R. C. T. Lee. An optimal embedding of cycles into incomplete hypercubes. Information Processing Letters, 72(5–6):213–218, December 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/60/25/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/60/25/34/article.
pdf.

Hung:1999:CTV

[HHS99] Chun-Nan Hung, Lih-Hsing Hsu, and Ting-Yi Sung. Christmas tree: A versatile 1fault-tolerant design for token rings. Information Processing Letters, 72(1–2):55–63, October 29, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/27/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/27/32/article.
pdf.

Huang:1994:SAS

[HHW94] Nen-Fu F. Huang, Ching-Ho H.

Huang, and Yue-Li L. Wang. A sweepline algorithm to solve the two-center problem. Information Processing Letters, 49 (4):171–177, February 25, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Han:1990:TLB

[HI90] Yijie Han and Yoshihide Igarashi.

Time lower bounds for sorting on multi-dimensional meshconnected processor arrays. Information Processing Letters, 33 (5):233–238, January 10, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hirose:1997:CKD

[HI97] Shouichi Hirose and Katsuo Ikeda. A conference key distribution system for the star configuration based on the discrete logarithm problem. Information Processing Letters, 62(4): 189–192, June 11, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hill:1995:LB

[Hil95] Steve Hill. The lazy z-buffer. Information Processing Letters, 55 (2):65–70, July 21, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hirata:1996:ULT

[Hir96] Tomio Hirata. A unified lineartime algorithm for computing distance maps. Information Processing Letters, 58(3):129–133, May 13, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hirata:1997:CSR

[Hir97] Kouichi Hirata. Constructing simply recursive programs from a finite set of good examples. Information Processing Letters, 64 (5):225–230, December 23, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hsu:1991:FMV

[HJL^{+}91] Lih-Hsing Hsu, Rong-Hong Jan, Yu-Che Lee, Chun-Nan Hung, and Maw-Sheng Chern. Finding the most vital edge with respect to minimum spanning tree in weighted graphs. Information Processing Letters, 39(5): 277–281, September 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Harn:1990:EPE

[HK90] Lein Harn and Thomas Kiesler. An efficient probabilistic encryption scheme. Information Processing Letters, 34(3):123–129, April 9, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hammer:1992:HFT

[HK92] Peter L. Hammer and Alexander Kogan. Horn functions and their DNFs. Information Processing Letters, 44(1):23–29, November 9, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Harikumar:1996:IDM

[HK96a] S. Harikumar and Shashi Kumar. Iterative deepening multiobjective A^{∗}. Information Processing Letters, 58(1):11–15, April 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Herman:1996:MCS

[HK96b] Michael W. Herman and Waldemar W. Koczkodaj. A Monte Carlo study of pairwise comparison. Information Processing Letters, 57(1):25–29, January 15, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Holsztynski:1996:CIA

[HK96c] Włodzimierz Holsztyński and Waldemar W. Koczkodaj. Convergence of inconsistency algorithms for the pairwise comparisons. Information Processing Letters, 59(4):197–202, August 26, 1996. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Ho:1997:BFC

[HK97a] Jan-Ming Ho and Ming-Tat Ko. Bounded fan-out m-center problem. Information Processing Letters, 63(2):103–108, August 13, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hong:1997:RDC

[HK97b] Seok Hee Hong and Myoung Ho Kim. Resolving data conflicts with multiple versions and precedence relationships in realtime databases. Information Processing Letters, 61(3):149– 156, March 14, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huttenlocher:1992:VDR

[HKK92] Daniel P. Huttenlocher, Klara Kedem, and Jon M. Kleinberg. Voronoĭ diagrams of rigidly moving sets of points. Information Processing Letters, 43(4): 217–223, September 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hellerstein:1990:TCR

[HKW90] Lisa Hellerstein, Philip Klein, and Robert Wilber. On the time-space complexity of reachability queries for preprocessed graphs. Information Processing Letters, 35(5):261–267, August 28, 1990. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Hambrusch:1991:PAC

[HL91a] Susanne Hambrusch and Michael Luby. Parallel asynchronous connected components in a mesh. Information Processing Letters, 38(5):257–263, June 14, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:1991:SSM

[HL91b] Bing-Chao Huang and Michael A.

Langston. Stable set and multiset operations in optimal time and space. Information Processing Letters, 39(3):131–136, August 16, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Han:1992:SRT

[HL92] Ching-Chih Han and Kwei-Jay Lin. Scheduling real-time computations with separation constraints. Information Processing Letters, 42(2):61–66, May 11, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hong:1995:PAD

[HL95] Dawei Hong and Joseph Y.-T.

Leung. Probabilistic analysis of k-dimensional packing algorithms. Information Processing Letters, 55(1):17–24, July 7, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hofmeister:1996:ISG

[HL96] Thomas Hofmeister and Hanno Lefmann. Independent sets in graphs with triangles. Information Processing Letters, 58(5): 207–210, June 10, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:1998:FSS

[HL98] Shing-Tsaan Huang and Tzong-Jye Liu. Four-state stabilizing phase clock for unidirectional rings of odd size. Information Processing Letters, 65(6): 325–329, March 27, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hsu:1992:SSW

[HLC92] F. R. Hsu, R. C. T. Lee, and R. C. Chang. Special subgraphs of weighted visibility graphs. Information Processing Letters, 43 (2):69–75, August 24, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hwang:1995:TAN

[HLL^{+}95] Tzonelih Hwang, Narn-Yih Lee, Chuan-Ming Li, Ming-Yung Ko, and Yung-Hsiang Chen. Two attacks on Neuman-Stubblebine authentication protocols. Information Processing Letters, 53 (2):103–107, January 27, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hanus:1998:SSI

[HLM98] Michael Hanus, Salvador Lucas, and Aart Middeldorp. Strongly sequential and inductively sequential term rewriting systems. Information Processing Letters, 67(1):1–8, July 16, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huy:1992:WFM

[HLV92] Phan Trung Huy, Igor Livotsky, and – Dõ Long Vân. Which finite monoids are syntactic monoids of rational ω-languages. Information Processing Letters, 42 (3):127–132, May 25, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:1997:ERM

[HLV97] Shou-Hsuan S. Huang, Hongfei Liu, and Rakesh M. Verma. On embedding rectangular meshes into rectangular meshes of smaller aspect ratio. Information Processing Letters, 63(3): 123–129, August 27, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ho:1993:CST

[HLW93] Kevin I.-J. Ho, Joseph Y.-T. Leung, and W.-D. Wei. Complexity of scheduling tasks with timedependent execution times. Information Processing Letters, 48 (6):315–320, December 20, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huynh:1995:ICD

[HM95] Tien Huynh and Kim Marriott.

Incremental constraint deletion in systems of linear constraints.

Information Processing Letters, 55(2):111–115, July 21, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hui:1996:ASA

[HM96] Lucas Chi Kwong Hui and Charles U. Martel. Analyzing self-adjusting linear list algorithms with deletions and unsuccessful searches. Information Processing Letters, 58(5): 231–236, June 10, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ha:1998:CMM

[HM98] Jae-Cheol Ha and Sang-Jae Moon. A common-multiplicand method to the Montgomery algorithm for speeding up exponentiation. Information Processing Letters, 66(2):105–107, April 29, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Harper:1999:PVG

[HM99] Robert Harper and John C. Mitchell. Parametricity and variants of Girard’s J operator. Information Processing Letters, 70(1):1–5, April 16, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/17/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/17/17/article.
pdf.

HerreradeFigueiredo:1995:LTA

[HMP95] Celina M. Herrera de Figueiredo, João Meidanis, and Célia Picinin de Mello. A linear-time algorithm for proper interval graph recognition. Information Processing Letters, 56(3):179–184, November 10, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hunt:1996:EAC

[HMPS96] Galen C. Hunt, Maged M.

Michael, Srinivasan Parthasarathy, and Michael L. Scott. An efficient algorithm for concurrent priority queue heaps. Information Processing Letters, 60(3): 151–157, November 11, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Han:1992:MCT

[HNC92] Y. Han, B. Narahari, and H.- A. Choi. Mapping a chain task to chained processors. Information Processing Letters, 44 (3):141–148, November 30, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoare:1990:FPI

[Hoa90] C. A. R. Hoare. Fixed points of increasing functions. Information Processing Letters, 34(3): 111–112, April 9, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoffgen:1993:CLT

[Höf93] Klaus-Uwe Höffgen. Computational limitations on training sigmoid neural networks. Information Processing Letters, 46 (6):269–274, July 26, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hofri:1994:TGD

[Hof94]  Micha Hofri.  On timeout

for global deadlock detection in decentralized database systems. Information Processing Letters, 51(6):295–302, September 26, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoogerwoord:1990:CDC

[Hoo90] Rob R. Hoogerwoord. A calculational derivation of the CASOP algorithm. Information Processing Letters, 36(6):297–299, December 15, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Horng:1995:PAU

[Hor95] Gwoboa Horng. Password authentication without using a password table. Information Processing Letters, 55(5):247– 250, September 15, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Horng:1998:AAP

[Hor98] Gwoboa Horng. An active attack on protocols for serveraided RSA signature computation. Information Processing Letters, 65(2):71–73, January 29, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Howell:1991:CPI

[How91] Rodney R. Howell. The complexity of problems involving structurally bounded and conservative Petri nets. Information Processing Letters, 39(6): 309–315, September 27, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hower:1995:CSA

[How95] Walter Hower. Constraint satisfaction — algorithms and complexity analysis. Information Processing Letters, 55(3):171– 178, August 11, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hower:1998:RGC

[How98] Walter Hower. Revisiting global constraint satisfaction. Information Processing Letters, 66(1): 41–48, April 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Haldar:1990:OCC

[HP90] Chandan Haldar and L. M. Patnaik. Oracle complexities for computational geometry of semialgebraic sets and Voronoi diagrams. Information Processing Letters, 36(2):95–102, October 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoover:1992:DSS

[HP92] Debra Hoover and Joseph Poole. A distributed self-stabilizing solution to the dining philosophers problem. Information Processing Letters, 41(4):209–213, March 18, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hung:1994:FRD

[HP94] Ching Yu Hung and Behrooz Parhami. Fast RNS division algorithms for fixed divisors with application to RSA encryption. Information Processing Letters, 51(4):163–169, August 24, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Higham:1996:SEA

[HP96] Lisa Higham and Teresa Przytycka. A simple, efficient algorithm for maximum finding on rings. Information Processing Letters, 58(6):319–324, June 24, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Harbane:1997:SBK

[HP97a] Rabah Harbane and Carles Padró. Spanners of de Bruijn and Kautz graphs. Information Processing Letters, 62(5):231– 236, July 2, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Harbane:1997:SUG

[HP97b] Rabah Harbane and Carles Padró. Spanners of underlying graphs of iterated line digraphs. Information Processing Letters, 62(5):237–244, July 2, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Haastad:1993:WCA

[HPS93] Johan Håstad, Steven Phillips, and Shmuel Safra. A wellcharacterized approximation problem. Information Processing Letters, 47(6):301–305, October 18, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hagerup:1990:GTC

[HR90] Torben Hagerup and Christine Rüb. A guided tour of Chernoff bounds. Information Processing Letters, 33(6):305–308, February 10, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hassin:1994:AMA

[HR94] Refael Hassin and Shlomi Rubinstein. Approximations for the maximum acyclic subgraph problem. Information Processing Letters, 51(3):133–140, August 10, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hassin:1997:AAM

[HR97] Refael Hassin and Shlomi Rubinstein. An approximation algorithm for maximum packing of 3-edge paths. Information Processing Letters, 63(2):63–67, August 13, 1997. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Harary:1998:BSC

[HR98a] Frank Harary and Desh Ranjan. Breaking symmetry in complete graphs by orienting edges: asymptotic bounds. Information Processing Letters, 67(5): 227–230, September 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hassin:1998:AAM

[HR98b] Refael Hassin and Shlomi Rubinstein. An approximation algorithm for the maximum traveling salesman problem. Information Processing Letters, 67(3): 125–130, August 17, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hemaspaandra:1998:RWG

[HR98c] Edith Hemaspaandra and Jörg Rothe. Recognizing when greed can approximate maximum independent sets is complete for parallel access to NP. Information Processing Letters, 65 (3):151–156, February 13, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hagerup:1990:INS

[HS90] Torben Hagerup and Hong Shen. Improved nonconservative sequential and parallel integer sorting. Information Processing Letters, 36(2):57–63, October 15, 1990. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Halperin:1991:DCC

[HS91a] Dan Halperin and Micha Sharir.

On disjoint concave chains in arrangements of (pseudo) lines. Information Processing Letters, 40 (4):189–192, November 25, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [HS94].

Hofri:1991:OCS

[HS91b] Micha Hofri and Hadas Shachnai.

On the optimality of the counter scheme for dynamic linear lists. Information Processing Letters, 37(3):175–179$, February 18, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hirschberg:1993:BTT

[HS93a] D. S. Hirschberg and S. S. Seiden. A bounded-space tree traversal algorithm. Information Processing Letters, 47(4): 215–219, September 27, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hsu:1993:OBS

[HS93b] Suchen H. Hsu and Richard Snodgrass. Optimal block size for set-valued attributes. Information Processing Letters, 45 (3):153–158, March 8, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Halperin:1994:CDC

[HS94] Dan Halperin and Micha Sharir. Corrigendum: “On disjoint concave chains in arrangements of (pseudo) lines” [Inform. Process. Lett. 40 (1991), no. 4, 189–192; MR 92m:52029]. Information Processing Letters, 51(1):53–56, July 12, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [HS91a].

Heydemann:1995:NRP

[HS95] M. C. Heydemann and D. Sotteau. A note on recursive properties of the de Bruijn, Kautz, and FFT digraphs. Information Processing Letters, 53(5): 255–259, March 10, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hasunuma:1997:CBN

[HS97] Toru Hasunuma and Yukio Shibata. Containment of butterflies in networks constructed by the line digraph operation. Information Processing Letters, 61(1): 25–30, January 14, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hensgen:1993:FBA

[HSC93] Debra A. Hensgen, David L. Sims, and David Charley. A fair Banker’s Algorithm for read and write locks. Information Processing Letters, 48(3):131–137, November 19, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hakimi:1997:OGO

[HSY97] S. Louis Hakimi, Edward F.

Schmeichel, and Neal E. Young. Orienting graphs to optimize reachability. Information Processing Letters, 63(5):229–235, September 24, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ho-Stuart:1993:CWB

[HSZF93] Chris Ho-Stuart, H. S. M.

Zedan, and Ming Fang. Congruent weak bisimulation with dense real-time. Information Processing Letters, 46(2):55–61, May 17, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hsu:1991:LTA

[HT91] Wen-Lian Hsu and Kuo-Hui Tsai. Linear time algorithms on circular-arc graphs. Information Processing Letters, 40(3): 123–129, November 8, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hassin:1995:MDS

[HT95] Refael Hassin and Arie Tamir.

On the minimum diameter spanning tree problem. Information Processing Letters, 53(2): 109–111, January 27, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hu:1998:OAT

[HT98] T. C. Hu and P. A. Tucker. Optimal alphabetic trees for binary search. Information Processing Letters, 67(3):137–140, August 17, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hsiao:1991:SSS

[HTC91] Ju Yuan Hsiao, Chuan Yi Tang, and Ruay Shiung Chang. Solving the single step graph searching problem by solving the maximum two-independent set problem. Information Processing Letters, 40(5):283–287, December 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hsiao:1992:EAF

[HTC92] Ju Yuan Hsiao, Chuan Yi Tang, and Ruay Shiung Chang. An efficient algorithm for finding a maximum weight 2-independent set on interval graphs. Information Processing Letters, 43(5): 229–235, October 5, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huhne:1993:LSD

[Hüh93] Martin Hühne. Linear speedup does not hold on Turing machines with tree storages. Information Processing Letters, 47 (6):313–318, October 18, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huynh:1991:EES

[Hùy91] Dũng T. Hùynh. The effective entropies of some extensions of context-free languages. Information Processing Letters, 37 (3):165–169, February 18, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Howard:1996:PLI

[HV96] Paul G. Howard and Jeffrey Scott Vitter. Parallel lossless image compression using Huffman and arithmetic coding. Information Processing Letters, 59(2):65–73, July 22, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hingston:1990:DJA

[HW90] Philips Hingston and Ross Wilkinson. A distributed join algorithm. Information Processing Letters, 33(6):313–317, February 10, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hanschke:1993:SSB

[HW93a] Philipp Hanschke and Jörg Würtz. Satisfiability of the smallest binary program. Information Processing Letters, 45 (5):237–241, April 2, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoory:1993:UTS

[HW93b] Shlomo Hoory and Avi Wigderson. Universal traversal sequences for expander graphs. Information Processing Letters, 46 (2):67–69, May 17, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Henzinger:1996:NSC

[HW96] Monika Henzinger and David P. Williamson. On the number of small cuts in a graph. Information Processing Letters, 59(1): 41–44, July 8, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hwang:1998:AUS

[HW98] Tzonelih Hwang and Chih-Hung Wang. Arbitrated unconditionally secure authentication scheme with multi-senders. Information Processing Letters, 65 (4):189–193, February 27, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hwang:1999:PRL

[HW99] W.-P. Hwang and C.-L. Wang. In-place random list permutations. Information Processing Letters, 70(6):255–257, June 30, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/22/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/22/17/article.
pdf.

Hwang:1992:AOT

[Hwa92a] Tzonelih Hwang. Attacks on Okamoto and Tanaka’s oneway ID-based key distribution system. Information Processing Letters, 43(2):83–86, August 24, 1992. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Hwang:1992:EIB

[Hwa92b] Tzonelih Hwang. Efficient IDbased key distribution with tamperfree devices. Information Processing Letters, 44(1):31–34, November 9, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hwang:1992:PGO

[Hwa92c] Tzonelih Hwang. Protocols for group oriented secret sharing. Information Processing Letters, 42(4):179–182, June 19, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hwang:1993:SSD

[Hwa93] Tzonelih Hwang. Scheme for secure digital mobile communications based on symmetric key cryptography. Information Processing Letters, 48(1):35–37, October 29, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ho:1996:LTA

[HWJ96] Ting-Yem Ho, Yue-Li Wang, and Ming-Tsan Juan. A linear time algorithm for finding all hinge vertices of a permutation graph. Information Processing Letters, 59(2):103–107, July 22, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See errata [HAM11].

Huang:1999:TDP

[HWT99] Kuo-Chan Huang, Feng-Jian Wang, and Jyun-Hwei Tsai. Two design patterns for dataparallel computation based on master-slave model. Information Processing Letters, 70(4): 197–204, May 28, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/20/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/20/22/article.
pdf.

Hsieh:1998:H

[HYHW98] Tsu-Miin Hsieh, Yi-Shiung Yeh, Yung-Cheng Hsieh, and Chan-Chi Wang. A homophonic DES. Information Processing Letters, 66(6):317–320, June 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hu:1999:RG

[HZS99] Qing Hu, Yixin Zhang, and Xiaojun Shen. Rearrangeable graphs. Information Processing Letters, 71(1):23– 27, July 16, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/17/
21/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/17/21/article.
pdf.

Ibarra:1997:FPM

[Iba97] Louis Ibarra. Finding pattern matchings for permutations. Information Processing Letters, 61 (6):293–295, April 24, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iyengar:1998:EFS

[IC98] Vikram Iyengar and Krishnendu Chakrabarty. An efficient finitestate machine implementation of Huffman decoders. Information Processing Letters, 64(6): 271–275, January 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [TM98].

Iwama:1997:FPA

[IIO97] K. Iwama, C. Iwamoto, and T. Ohsawa. A faster parallel algorithm for k-connectivity. Information Processing Letters, 61 (5):265–269, April 10, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Inoue:1992:RBN

[IIT92] Katsushi Inoue, Akira Itō, and Itsuo Takanami. A relationship between nondeterministic Turing machines and 1-inkdot Turing machines with small space. Information Processing Letters, 43(4):225–227, September 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Inoue:1993:NMI

[IITY93] Katsushi Inoue, Akira Itō, Itsuo Takanami, and Tsunehiro Yoshinaga. A note on multiinkdot nondeterministic Turing machines with small space. Information Processing Letters, 48 (6):285–288, December 20, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iwano:1993:EAF

[IK93] Kazuo Iwano and Naoki Katoh. Efficient algorithms for finding the most vital edge of a minimum spanning tree. Information Processing Letters, 48(5): 211–213, December 10, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Imrich:1997:RHG

[IK97] Wilfried Imrich and Sandi Klavžar. Recognizing Hamming graphs in linear time and space. Information Processing Letters, 63(2):91–95, August 13, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iwasaki:1999:IST

[IKOI99] Yukihiro Iwasaki, Yuka Kajiwara, Koji Obokata, and Yoshihide Igarashi. Independent spanning trees of chordal rings. Information Processing Letters, 69(3):155–160, February 12, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/19/
27/abstract.html;   http:

//www.elsevier.nl/gej-ng/
10/23/20/47/19/27/article.
pdf.

Ivkovic:1996:FRF

[IL96] Zoran Ivković and Errol L.

Lloyd. A fundamental restriction on fully dynamic maintenance of bin packing. Information Processing Letters, 59(4): 229–232, August 26, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ivkovic:1997:PDP

[IL97] Zoran Ivković and Errol L.

Lloyd. Partially dynamic bin packing can be solved within 1+ϵ in (amortized) polylogarithmic time. Information Processing Letters, 63(1):45–50, July 30, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Inverardi:1990:RSV

[IN90] P. Inverardi and M. Nesi. A rewriting strategy to verify observational congruence. Information Processing Letters, 35 (4):191–199, August 7, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nakayama:1997:PAS

[iNM97] Shin ichi Nakayama and Shigeru Masuyama. A parallel algorithm for solving the coloring problem on trapezoid graphs. Information Processing Letters, 62(6): 323–327, July 16, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nakano:1997:LTA

[iNRN97] Shin ichi Nakano, Md. Saidur Rahman, and Takao Nishizeki. A linear-time algorithm for fourpartitioning four-connected planar graphs. Information Processing Letters, 62(6):315–322, July 16, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Igarashi:1994:ABS

[IOU94] Yoshihide Igarashi, Shingo Osawa, and Walter Unger. Automorphisms of broadcasting schemes with respect to start rounds. Information Processing Letters, 51(1):35–41, July 12, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iwama:1995:ELB

[IP95] Kazuo Iwama and Toniann Pitassi. Exponential lower bounds for the tree-like Hajós calculus. Information Processing Letters, 54(5):289–294, June 9, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Irani:1991:CSA

[IR91] Sandy Irani and Ronitt Rubinfeld. A competitive 2-server algorithm. Information Processing Letters, 39(2):85–91, July 31, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Irani:1991:TRL

[Ira91] Sandy Irani. Two results on the list update problem. Information Processing Letters, 38(6): 301–306, June 28, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Irving:1991:AMI

[Irv91] Robert W. Irving. On approximating the minimum independent dominating set. Information Processing Letters, 37 (4):197–200, February 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Israeli:1998:TCU

[IS98] Amos Israeli and Asaf Shirazi.

The time complexity of updating snapshot memories. Information Processing Letters, 65(1):33–40, January 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Istrate:1997:SEE

[Ist97] Gabriel Istrate. The strong equivalence of ET0L grammars. Information Processing Letters, 62(4):171–176, June 11, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iwamoto:1994:FHC

[IT94] Chuzo Iwamoto and Godfried T.

Toussaint. Finding Hamiltonian circuits in arrangements of Jordan curves is NP-complete. Information Processing Letters, 52 (4):183–189, November 25, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tokunaga:1996:INT

[iT96] Shin ichi Tokunaga. Intersection number of two connected geometric graphs. Information Processing Letters, 59(6):331– 333, September 23, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Itoh:1991:CFI

[Ito91] Toshiya Itoh. Characterization for a family of infinitely many irreducible equally spaced polynomials. Information Processing Letters, 37(5):273–277, March 14, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ito:1996:SRI

[ITY96] Masayuki Ito, Naofumi Takagi, and Shuzo Yajima. Square rooting by iterative multiplyadditions. Information Processing Letters, 60(5):267–269, December 8, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ihler:1993:MHG

[IWW93] Edmund Ihler, Dorothea Wagner, and Frank Wagner. Modeling hypergraphs by graphs with the same mincut properties. Information Processing Letters, 45 (4):171–175, March 22, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ito:1998:LTA

[IY98] Hiro Ito and Mitsuo Yokoyama.

Linear time algorithms for graph search and connectivity determination on complement graphs. Information Processing Letters, 66(4):209–213, May 29, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Istrail:1994:BWP

[IZ94] Sorin Istrail and Dejan Zivkovic.

Bounded-width polynomial-size Boolean formulas compute exactly those functions in AC^{0}. Information Processing Letters, 50 (4):211–216, May 25, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jacob:1990:MRC

[Jac90a] Jeremy Jacob. A model of reconfiguration in communicating sequential processes. Information Processing Letters, 35(1):19–22, June 15, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jacob:1990:SDH

[Jac90b] Jeremy Jacob. Separability and the detection of hidden channels. Information Processing Letters, 34(1):27–29, February 22, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jacobs:1991:PCA

[Jac91] D. P. Jacobs. Probabilistic checking of associativity in algebras. Information Processing Letters, 37(4):187–191, February 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jacob:1992:MRC

[Jac92] Jeremy Jacob. A model of reconfiguration in communicating sequential processes with a notion of transactions. Information Processing Letters, 41(1): 9–12, January 21, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jagadish:1997:AHC

[Jag97] H. V. Jagadish. Analysis of the Hilbert curve for representing two-dimensional space. Information Processing Letters, 62 (1):17–22, May 7, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jain:1995:QAL

[Jai95] Sanjay Jain. On a question about learning nearly minimal programs. Information Processing Letters, 53(1):1–4, January 13, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jain:1999:QNM

[Jai99] Sanjay Jain. On a question of nearly minimal identification of functions. Information Processing Letters, 70(3):

113–117, May 14, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/19/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/19/18/article.
pdf.

Janardan:1991:DMM

[Jan91a] Ravi Janardan. On the dynamic maintenance of maximal points in the plane. Information Processing Letters, 40(2):59–64, October 25, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Janssen:1991:OPR

[Jan91b] A. J. E. M. Janssen. An optimization problem related to neural networks. Information Processing Letters, 37(3):155–157, February 18, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jansen:1992:AAG

[Jan92] Klaus Jansen. An approximation algorithm for the general routing problem. Information Processing Letters, 41(6): 333–339, April 17, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jancar:1999:NWQ

[Jan99]  Petr Jančar.  A note on

well quasi-orderings for powersets. Information Processing Letters, 72(5–6):155–160, December 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/60/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/60/25/26/article.
pdf.

Jiang:1999:OVO

[JB99] Xiaoyi Jiang and Horst Bunke. Optimal vertex ordering of graphs. Information Processing Letters, 72(5–6):149–154, December 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/60/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/60/25/25/article.
pdf.

Jung:1994:ECR

[JCH94] Min-Soo Jung, Kwang-Moo Choe, and Taisook Han. An efficient computation of right context for LR-based error repair. Information Processing Letters, 49(2):63–71, January 28, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jayanti:1999:CGD

[JCT99] Prasad Jayanti, Tushar Deepak Chandra, and Sam Toueg. The cost of graceful degradation for omission failures. Information Processing Letters, 71(3–4):167– 172, August 27, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/28/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/28/34/article.
pdf.

Jedrzejowicz:1990:IHS

[Jed90] Joanna Jedrzejowicz. Infinite hierarchy of shuffle expressions over a finite alphabet. Information Processing Letters, 36(1): 13–17, October 1, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jeffs:1991:OIN

[Jef91] Janice Jeffs. Order independent NCE grammars recognized in polynomial time. Information Processing Letters, 39(3): 161–164, August 16, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jenner:1995:KPN

[Jen95] Birgit Jenner. Knapsack problems for NL. Information Processing Letters, 54(3):169–174, May 12, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jerrum:1994:CTG

[Jer94] Mark Jerrum. Counting trees in a graph is #P-complete. Information Processing Letters, 51 (3):111–116, August 10, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jifeng:1993:AOS

[JH93a] He Jifeng and C. A. R. Hoare. From algebra to operational semantics. Information Processing Letters, 45(2):75–80, February 26, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Johnson:1993:SCP

[JH93b] Theodore Johnson and Krishna Harathi. A simple correctness proof of the MCS contentionfree lock. Information Processing Letters, 48(5):215–220, December 10, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:1991:DTG

[Jia91a]  Bin Jiang.  DFS-traversing

graphs in a paging environment, LRU or MRU? Information Processing Letters, 40(4):193–196, November 25, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:1991:TGP

[Jia91b] Bin Jiang. Traversing graphs in a paging environment, BFS or DFS? Information Processing Letters, 37(3):143–147, February 18, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:1993:COR

[Jia93] Bin Jiang. I/O- and CPUoptimal recognition of strongly connected components. Information Processing Letters, 45 (3):111–115, March 8, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:1992:NSS

[JLD92] Tao Jiang, Ming Li, and Ding Zhu Du. A note on shortest superstrings with flipping. Information Processing Letters, 44 (4):195–199, December 10, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Julia:1996:CCG

[JLP96] S. Julia, I. Litovsky, and B. Patrou. On codes, ω-codes and ω-generators. Information Processing Letters, 60(1):1–5, October 14, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jimbo:1994:RBB

[JM94a] Shuji Jimbo and Akira Maruoka.

On the relationship between ϵbiased random variables and ϵdependent random variables. Information Processing Letters, 51 (1):17–23, July 12, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jimbo:1994:RBD

[JM94b] Shuji Jimbo and Akira Maruoka.

On the relationship between the diameter and the size of a boundary of a directed graph. Information Processing Letters, 50(5):277–282, June 10, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jovanovic:1994:FTS

[JM94c] Zoran Jovanović and Jelena Mišić. Fault tolerance of the star graph interconnection network. Information Processing Letters, 49(3):145–150, February 11, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Johannsen:1998:LBM

[Joh98] Jan Johannsen. Lower bounds for monotone real circuit depth and formula size and tree-like Cutting Planes. Information Processing Letters, 67(1):37–41, July 16, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Johansson:1999:SDD

[Joh99]  Öjvind Johansson.  Simple

distributed Δ + 1-coloring of graphs. Information Processing Letters, 70(5):229–232, June 21, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/48/21/
                                        19/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/48/21/19/article.
                                        pdf.

Jones:1995:PFL

[Jon95] C. B. Jones. Partial functions and logics: A warning. Information Processing Letters, 54 (2):65–67, April 28, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jonsson:1997:NRF

[Jon97] P. Jonsson. A nonapproximability result for finite function generation. Information Processing Letters, 63(3):143–145, August 27, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jonsson:1998:NON

[Jon98] P. Jonsson. Near-optimal nonapproximability results for some Npo PB-complete problems. Information Processing Letters, 68 (5):249–253, December 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Justin:1997:SFI

[JP97] J. Justin and G. Pirillo. On some factorizations of infinite words by elements of codes. Information Processing Letters, 62(6): 289–294, July 16, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:1991:NSC

[JR91] Tao Jiang and B. Ravikumar.

A note on the space complexity of some decision problems for finite automata. Information Processing Letters, 40(1):25–31, October 11, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jain:1993:NMI

[JS93] Sanjay Jain and Arun Sharma.

On the non-existence of maximal inference degrees for language identification. Information Processing Letters, 47(2): 81–88, August 20, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jha:1995:SCD

[JS95] Pranava K. Jha and Giora Slutzki. A scheme to construct distance-three codes using Latin squares, with applications to the n-cube. Information Processing Letters, 55(3):123–127, August 11, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jerebic:1992:ORT

[JT92] Izidor Jerebic and Roman Trobec. Optimal routing in toroidal networks. Information Processing Letters, 43(6):285– 291, October 19, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jwo:1994:TDD

[JT94] Jung Sing Jwo and Tai Ching Tuan. On transmitting delay in a distance-transitive strongly antipodal graph. Information Processing Letters, 51(5):233– 235, September 12, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jan:1996:SIE

[JT96] Jinn-Ke Jan and Yuh-Min Tseng. On the security of image encryption method. Information Processing Letters, 60(5): 261–265, December 8, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jukna:1995:CTF

[Juk95]  Stasys Jukna.  Computing

threshold functions by depth-3 threshold circuits with smaller thresholds of their gates. Information Processing Letters, 56 (3):147–150, November 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jukna:1999:LCH

[Juk99] Stasys Jukna. Linear codes are hard for oblivious readonce parity branching programs. Information Processing Letters, 69(6):267–269, March 26, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/22/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/22/17/article.
pdf.

Jurdzinski:1998:DWP

[Jur98] Marcin Jurdziński. Deciding the winner in parity games is in UP ∩ co-UP. Information Processing Letters, 68(3):119– 124, November 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/ipl/cas_sub/browse/
browse.cgi?year=1998&volume=
68&issue=3&aid=2076.

Jain:1995:AAA

[JW95] Ravi Jain and John Werth. Analysis of approximate algorithms for edge-coloring bipartite graphs. Information Processing Letters, 54(3):163–168, May 12, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Juurlink:1996:CPB

[JW96] Ben H. H. Juurlink and Harry A. G. Wijshoff. Communication primitives for BSP computers. Information Processing Letters, 58(6):303–310, June 24, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kabadi:1997:ESP

[KA97a] S. N. Kabadi and Y. P. Aneja. An efficient, strongly polynomial, ε-approximation parametric optimization scheme. Information Processing Letters, 64 (4):173–177, December 13, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khan:1997:ACM

[KA97b] Md. Mozammel Huq Azad Khan and Md. Shamsul Alam. Algorithms for conversion of minterms to positive polarity Reed–Muller coefficients and vice versa. Information Processing Letters, 62(5):223–230, July 2, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kulkarni:1997:MBS

[KA97c] Sandeep S. Kulkarni and Anish Arora. Multitolerant barrier synchronization. Information Processing Letters, 64(1): 29–36, October 22, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kaivola:1995:MSS

[Kai95] Roope Kaivola. On modal mu-calculus and Büchi tree automata. Information Processing Letters, 54(1):17–22, April 14, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kale:1990:APH

[Kal90] L. V. Kale. An almost perfect heuristic for the N-nonattacking queens problem. Information Processing Letters, 34(4):173– 178, April 24, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1993:AAC

[KAL93] Myung Soo Kim, Jae-Woo Ahn, and Soon-Bum Lim. An algebraic algorithm to compute the exact general sweep boundary of a 2D curved object. Information Processing Letters, 47(5): 221–229, October 8, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kamper:1990:RRD

[Käm90] Jürgen Kämper. A result relating disjunctive self-reducibility to P -immunity. Information Processing Letters, 33(5):239– 242, January 10, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kamin:1992:HSM

[Kam92] Samuel Kamin. Head-strictness is not a monotonic abstract property. Information Processing Letters, 41(4):195–198, March 18, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kann:1991:MBD

[Kan91] Viggo Kann. Maximum bounded 3-dimensional matching is MAX SNP-complete. Information Processing Letters, 37(1):27–35, January 10, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kanchanasut:1994:SPA

[Kan94a]  Kanchana Kanchanasut.  A

shortest-path algorithm for Manhattan graphs. Information Processing Letters, 49(1):21–25, January 14, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kann:1994:MBM

[Kan94b] Viggo Kann. Maximum bounded H-matching is Max SNPcomplete. Information Processing Letters, 49(6):309–318, March 22, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kao:1995:LTO

[Kao95] Ming-Yang Kao. Linear-time optimal augmentation for componentwise bipartite-completeness of graphs. Information Processing Letters, 54(1):59–63, April 14, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Karner:1990:NTB

[Kar90] George Karner. A note on terminal balancing of algebraic systems. Information Processing Letters, 35(6):291–293, September 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Karloff:1993:FAA

[Kar93] Howard Karloff. Fast algorithms for approximately counting mismatches. Information Processing Letters, 48(2):53–60, November 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kocsis:1990:RCR

[KB90] F. Kocsis and J. F. Böhme.

Rotation-based computations for ray-tracing second-order surfaces and curves. Information Processing Letters, 36(6):281– 283, December 15, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kapur:1992:ULL

[KB92] Shyam Kapur and Gianfranco Bilardi. On uniform learnability of language families. Information Processing Letters, 44(1): 35–38, November 9, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kumar:1994:ONP

[KB94] G. Prem Kumar and G. Phanendra Babu. Optimal network partitioning for fault-tolerant network management using evolutionary programming. Information Processing Letters, 50(3): 145–149, May 9, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kulasinghe:1995:MTH

[KB95] Priyalal Kulasinghe and Said Bettayeb. Multiply-twisted hypercube with five or more dimensions is not vertex-transitive. Information Processing Letters, 53(1):33–36, January 13, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kannan:1990:FHT

[KC90] C. S. Kannan and Henry Y. H. Chuang. Fast Hough transform on a mesh connected processor array. Information Processing Letters, 33(5):243–248, January 10, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1991:YAE

[KC91a] Do-Hyung Kim and Kwang-Moo Choe. Yet another efficient backward execution algorithm in the AND/OR process model. Information Processing Letters, 40 (4):201–211, November 25, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigenda [KC92].

Kumar:1991:HAH

[KC91b] Akhil Kumar and Shun Yan Cheung. A high availability \sqrt{N} hierarchical grid algorithm for replicated data. Information Processing Letters, 40(6): 311–316, December 30, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1992:CYA

[KC92] Do-Hyung Kim and Kwang-Moo Choe. Corrigenda: “Yet another efficient backward execution algorithm in the AND/OR Process Model” [Inform. Process. Lett. 40(4), 25 November 1991, pp. 201–211]. Information Processing Letters, 43(3): 165, September 14, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [KC91a].

Kocberber:1996:PEQ

[KC96] Seyit Kocberber and Fazli Can.

Partial evaluation of queries for bit-sliced signature files. Information Processing Letters, 60 (6):305–311, December 23, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1998:ASL

[KC98] Seongyeol Kim and Ilyong Chung. Application of the special Latin square to a parallel routing algorithm on a recursive circulant network. Information Processing Letters, 66(3): 141–147, May 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1993:MPM

[KCK93] Jeong Uk Kim, Ho Chang, and Tag Gon Kim. Multidisk partial match file design with known access pattern. Information Processing Letters, 45(1):33–39, January 25, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Karpovsky:1999:CVF

[KCLA99] Mark G. Karpovsky, Krishnendu Chakrabarty, Lev B. Levitin, and Dimiter R. Avresky. On the covering of vertices for fault diagnosis in hypercubes. Information Processing Letters, 69(2): 99–103, January 29, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/47/18/
                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/47/18/25/article.
                                        pdf.

Keil:1991:SAD

[Kei91] M. Keil. A simple algorithm for determining the envelope of a set of lines. Information Processing Letters, 39(3):121–124, August 16, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kelsen:1994:OPA

[Kel94] Pierre Kelsen. An optimal parallel algorithm for maximal matching. Information Processing Letters, 52(4):223–228, November 25, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kemp:1999:OOC

[Kem99]  Rainer Kemp.  A one-to-

one correspondence between a class of leftist trees and binary trees. Information Processing Letters, 71(3–4):97–105, August 27, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/28/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/28/25/article.
pdf.

Kennedy:1991:PCS

[Ken91] Robert Kennedy. Parallel cardinality stacks and an application. Information Processing Letters, 38(6):295–299, June 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kern:1992:LCB

[Ker92] W. Kern. Learning convex bodies under uniform distribution. Information Processing Letters, 43(1):35–39, August 10, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kakugawa:1994:DME

[KFYA94] Hirotsugu Kakugawa, Satoshi Fujita, Masafumi Yamashita, and Tadashi Ae. A distributed k-mutual exclusion algorithm using k-coterie. Information Processing Letters, 49 (4):213–218, February 25, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krishnamurti:1999:AAN

[KG99] Ramesh Krishnamurti and Daya Ram Gaur. An approximation algorithm for nonpreemptive scheduling on hypercube parallel task systems. Information Processing Letters, 72(5–6):183–188, December 30, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/60/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/60/25/30/article.
pdf.

Kuo:1996:SSC

[KH96] Yu-Chen Kuo and Shing-Tsaan Huang. A simple scheme to construct k-coteries with O(\sqrt{N}) uniform quorum sizes. Information Processing Letters, 59(1): 31–36, July 8, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khan:1994:AHF

[Kha94a] Md. Mozammel Huq Azad Khan. An algorithm for hazardfree minimization of incompletely specified switching function. Information Processing Letters, 52(1):23–29, October 14, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khardon:1994:UFT

[Kha94b] Roni Khardon. On using the Fourier transform to learn disjoint DNF. Information Processing Letters, 49(5):219–222, March 11, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1998:RBS

[KHJ98] Jin-Soo Kim, Soonhoi Ha, and Chu Shik Jhon. Relaxed barrier synchronization for the BSP model of computation on message-passing architectures. Information Processing Letters, 66(5):247–253, June 16, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khoong:1993:OPC

[Kho93] C. M. Khoong. Optimal parallel construction of heaps. Information Processing Letters, 48 (4):159–161, November 29, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khuller:1990:CAK

[Khu90] Samir Khuller. Coloring algorithms for K_{5}-minor free graphs. Information Processing Letters, 34(4):203–208, April 24, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khuller:1999:VAS

[Khu99] Samir Khuller. An O(|V |^{2}) algorithm for single connectedness. Information Processing Letters, 72(3–4):105–107, November 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/28/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/28/29/article.

pdf. See addendum [Khu00].

Khuller:2000:ASV

[Khu00] Samir Khuller. Addendum to “An O(|V |^{2}) algorithm for single connectedness”, [Information Processing Letters 72(3– 4) (1999) 105–107]. Information Processing Letters, 74(5– 6):263, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/28/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/28/37/article.

pdf. See [Khu99].

Kim:1990:FMI

[Kim90a] Haklin Kim. Finding a maximum independent set in a permutation graph. Information Processing Letters, 36(1):19–23, October 1, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1990:PAF

[Kim90b] Sung Kwon Kim. A parallel algorithm for finding a maximum clique of a set of circular arcs of a circle. Information Processing Letters, 34(5):235–241, May 7, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1990:PAS

[Kim90c] Sung Kwon Kim. Parallel algorithms for the segment dragging problem. Information Processing Letters, 36(6):323–327, December 15, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1994:RCM

[Kim94a] Sung Kwon Kim. The range cominima problem. Information Processing Letters, 49(3):117– 121, February 11, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1994:OSA

[Kim94b] Yong-Seok Kim. An optimal scheduling algorithm for preemptable real-time tasks. Information Processing Letters, 50 (1):43–48, April 8, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1996:UDC

[Kim96a] Changwook Kim. Unambiguous description of chain code picture languages. Information Processing Letters, 58(2):75–79, May 13, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1996:NFC

[Kim96b] Sung Kwon Kim. A note on finding compact sets in graphs represented by an adjacency list. Information Processing Letters, 57(6):335–338, March 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1997:LWL

[Kim97] Sung Kwon Kim. Logarithmic width, linear area upward drawing of AVL trees. Information Processing Letters, 63(6): 303–307, October 8, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kinnersley:1992:VSN

[Kin92] Nancy G. Kinnersley. The vertex separation number of a graph equals its path-width (NP-complete problem). Information Processing Letters, 42 (6):345–350, July 24, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kirchherr:1992:KCR

[Kir92] W. W. Kirchherr. Kolmogorov complexity and random graphs. Information Processing Letters, 41(3):125–130, March 6, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kirrinnis:1994:OBP

[Kir94] Peter Kirrinnis. An optimal bound for path weights in Huffman trees. Information Processing Letters, 51(2):107–110, July 26, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kumar:1990:CDD

[KIS90] D. Kumar, S. S. Iyengar, and M. B. Sharma. Corrections to a distributed depth-first search algorithm. Information Processing Letters, 35(1):55–56, June 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kurihara:1990:MTR

[KK90] Masahito Kurihara and Ikuo Kaji. Modular term rewriting systems and the termination. Information Processing Letters, 34 (1):1–4, February 22, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1992:MVA

[KK92a] Cheeha Kim and Jong-Sung Kim. A mean value analysis of the Ethernet throughput. Information Processing Letters, 43 (6):315–320, October 19, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kolountzakis:1992:FCE

[KK92b] Mihail N. Kolountzakis and Kiriakos N. Kutulakos. Fast computation of the Euclidian distance maps for binary images. Information Processing Letters, 43(4):181–184, September 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1994:NAM

[KK94] Dongseung Kim and Seung-Hoon Kim. O(log n) numerical algorithms on a mesh with wormhole routing. Information Processing Letters, 50(3): 129–136, May 9, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kavraki:1995:PPA

[KK95a] Lydia E. Kavraki and Mihail N. Kolountzakis. Partitioning a planar assembly into two connected parts is NP-complete. Information Processing Letters, 55 (3):159–165, August 11, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kloks:1995:CPE

[KK95b] T. Kloks and D. Kratsch. Computing a perfect edge without vertex elimination ordering of a chordal bipartite graph. Information Processing Letters, 55 (1):11–16, July 7, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1998:SFS

[KK98] Heung-Taek Kim and Myoung Ho Kim. Starvation-free secure multiversion concurrency control. Information Processing Letters, 65(5):247–253, March 13, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kang:1997:AOD

[KKK97] Hee-Joong Kang, Kawon Kim, and Jin H. Kim. Approximating optimally discrete probability distribution with kth-order dependency for combining multiple decisions. Information Processing Letters, 62(2):67–75, May 21, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1999:TAP

[KKK99] Jong Soo Kim, Sung Tak Kang, and Myoung Ho Kim. On temporal aggregate processing based on time points. Information Processing Letters, 71(5–6):213– 220, September 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/30/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/30/32/article.
pdf.

Kloks:1998:BCG

[KKM98] Ton Kloks, Dieter Kratsch, and Haiko Müller. Bandwidth of chain graphs. Information Processing Letters, 68(6):313–315, December 30, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Klavzar:1999:GWL

[KKM99] S. Klavžar, J. Koolen, and H. M. Mulder. Graphs which locally mirror the hypercube structure. Information Processing Letters, 71(2):87–90, July 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/49/18/
                                        23/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/49/18/23/article.
                                        pdf.

Kavadias:1991:CRC

[KKS91] Dimitris Kavadias, Lefteris M. Kirousis, and Paul Spirakis. The complexity of the reliable connectivity problem. Information Processing Letters, 39(5):245– 252, September 13, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1990:PSP

[KL90] Hyoung Joong Kim and Jang Gyu Lee. Partial sum problem mapping into a hypercube. Information Processing Letters, 36(5): 221–224, December 1, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Korsh:1998:SLG

[KL98] James F. Korsh and Seymour Lipschutz. Shifts and loopless generation of k-ary trees. Information Processing Letters, 65 (5):235–240, March 13, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Korsh:1999:LGG

[KL99] James F. Korsh and Paul LaFollette. Loopless generation of Gray codes for k-ary trees. Information Processing Letters, 70 (1):7–11, April 16, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/17/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/17/18/article.
pdf.

Klein:1994:DSB

[Kle94a] Philip N. Klein. A data structure for bicategories, with application to speeding up an approximation algorithm. Information Processing Letters, 52(6):303–307, December 23, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kleinberg:1994:LBT

[Kle94b] Jon M. Kleinberg. A lower bound for two-server balancing algorithms. Information Processing Letters, 52(1):39–43, October 14, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1997:ESD

[KLK97] Seung-Hoon Kim, Kyungshik Lim, and Cheeha Kim. Efficient stream distribution algorithm for heterogeneous multimedia multicast with link capacity constraint. Information Processing Letters, 63(6):309–315, October 8, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1993:AFO

[KLL93] Myoung Ho Kim, Joon Ho Lee, and Yoon Joon Lee. Analysis of fuzzy operators for high quality information retrieval. Information Processing Letters, 46 (5):251–256, July 9, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kloks:1996:FFG

[Klo96] T. Kloks. K_{1,3}-free and W_{4}-free graphs. Information Processing Letters, 60(4):221–223, November 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kann:1996:AMS

[KLP96] Viggo Kann, Jens Lagergren, and Alessandro Panconesi. Approximability of maximum splitting of k-sets and some other Apx-complete problems. Information Processing Letters, 58 (3):105–110, May 13, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kann:1998:AMC

[KLP98] Viggo Kann, Jens Lagergren, and Alessandro Panconesi. Approximate Max k-Cut with subgraph guarantee. Information Processing Letters, 65(3):145– 150, February 13, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Konana:1997:UTI

[KLR97] Prabhudev Konana, Juhnyoung Lee, and Sudha Ram. Updating timestamp interval for dynamic adjustment of serialization order in Optimistic Concurrency Control-Time Interval (OCCTI) protocol. Information Processing Letters, 63(4): 189–193, September 10, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kant:1997:ARV

[KLTT97] Goos Kant, Giuseppe Liotta, Roberto Tamassia, and Ioannis G. Tollis. Area requirement of visibility representations of trees. Information Processing Letters, 62(2):81–88, May 21, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kavraki:1993:CAP

[KLW93] Lydia Kavraki, Jean-Claude Latombe, and Randall H. Wilson. On the complexity of assembly partitioning. Information Processing Letters, 48(5):229– 235, December 10, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khuller:1990:TCP

[KM90] Samir Khuller and Joseph S. B. Mitchell. On a triangle counting problem. Information Processing Letters, 33(6):319–321, February 10, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kalamboukis:1991:TOD

[KM91] T. Z. Kalamboukis and S. L. Mantzaris. Towards optimal distributed election on chordal rings. Information Processing Letters, 38(5):265–270, June 14, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krishnamachari:1996:DAG

[KM96] Bagirath R. Krishnamachari and Ravi Mittal. Design and analysis of a generalized multi-ring architecture. Information Processing Letters, 58(1):17–21, April 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khuller:1999:BMC

[KMN99] Samir Khuller, Anna Moss, and Joseph (Seffi) Naor. The budgeted maximum coverage problem. Information Processing Letters, 70(1):39–45, April 16, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/48/17/

23/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/17/23/article.
pdf.

Kollias:1990:OEO

[KMP90] John G. Kollias, Yannis Manolopoulos, and Christos H. Papadimitriou. The optimum execution order of queries in linear storage. Information Processing Letters, 36(3):141–145, November 1, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Klop:1994:MCS

[KMTd94] Jan Willem Klop, Aart Middeldorp, Yoshihito Toyama, and Roel de Vrijer. Modularity of confluence: A simplified proof. Information Processing Letters, 49(2):101–109, January 28, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kloks:1998:VRA

[KMW98] Ton Kloks, Haiko Müller, and C. K. Wong. Vertex ranking of asteroidal triple-free graphs. Information Processing Letters, 68 (4):201–206, November 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Knapp:1990:PTP

[Kna90] Edgar Knapp. A predicate transformer for progress. Information Processing Letters, 33 (6):323–330, February 10, 1990.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kurosawa:1996:CLB

[KO96] Kaoru Kurosawa and Koji Okada. Combinatorial lower bounds for secret sharing schemes. Information Processing Letters, 60(6):301–304, December 23, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kobuchi:1992:OSF

[Kob92] Youichi Kobuchi. Order of state functions and logic functions. Information Processing Letters, 41(2):61–66, February 14, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kobayashi:1993:CPP

[Kob93] Kojiro Kobayashi. Σ_{n}^{0}-complete properties of programs and Martin-Löf randomness. Information Processing Letters, 46 (1):37–42, April 29, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kochol:1989:EMC

[Koc89] Martin Kochol. Efficient monotone circuits for threshold functions. Information Processing Letters, 32(3):121–122, August 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comment [Dun90].

Kolountzakis:1994:SLS

[Kol94] Mihail N. Kolountzakis. Selection of a large sum-free subset in polynomial time. Information Processing Letters, 49(5): 255–256, March 11, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Korsh:1993:CRG

[Kor93] James F. Korsh. Counting and randomly generating binary trees. Information Processing Letters, 45(6):291–294, April 16, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Korsh:1994:LGA

[Kor94] James F. Korsh. Loopless generation of k-ary tree sequences. Information Processing Letters, 52 (5):243–247, December 9, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kornerup:1995:MFN

[Kor95] Jacob Kornerup. Mapping a functional notation for parallel programs onto hypercubes. Information Processing Letters, 53 (3):153–158, February 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kornerup:1997:OES

[Kor97] Jacob Kornerup. Odd–even sort in powerlists. Information Processing Letters, 61(1):15–24, January 14, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kosub:1999:NUF

[Kos99]  Sven Kosub.  A note on

unambiguous function classes. Information Processing Letters, 72(5–6):197–203, December 30, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/60/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/60/25/32/article.
pdf.

Kurosawa:1995:LEA

[KOT95] Kaoru Kurosawa, Koji Okada, and Shigeo Tsujii. Low exponent attack against elliptic curve RSA. Information Processing Letters, 53(2):77–83, January 27, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1991:FDM

[KP91] Myoung Ho Kim and Sakti Pramanik. The FX distribution method for parallel processing of partial match queries. Information Processing Letters, 38 (5):243–252, June 14, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Koutsoupias:1992:GAS

[KP92] Elias Koutsoupias and Christos H. Papadimitriou. On the greedy algorithm for satisfiability. Information Processing Letters, 43(1):53–55, August 10, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1994:TWD

[KP94] Dongseung Kim and Joonyoung Park. Two-way dominant sequence clustering for processor scheduling. Information Processing Letters, 49(4):203–208, February 25, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Koutsoupias:1996:EP

[KP96] Elias Koutsoupias and Christos Papadimitriou. The 2evader problem. Information Processing Letters, 57(5):249– 252, March 11, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Knapik:1998:FQC

[KP98a] Teodor Knapik and Étienne Payet. The full quotient and its closure property for regular languages. Information Processing Letters, 65(2):57–62, January 29, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kwai:1998:PTD

[KP98b] Ding-Ming Kwai and Behrooz Parhami. Pruned threedimensional toroidal networks. Information Processing Letters, 68(4):179–183, November 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Katajainen:1999:PSF

[KP99] Jyrki Katajainen and Tomi A. Pasanen. In-place sorting with fewer moves. Information Processing Letters, 70(1): 31–37, April 16, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/17/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/17/22/article.
pdf.

Kim:1995:OAF

[KPC^{+}95] Sung-Ho Kim, Jung-Heum Park, Seung-Hak Choi, Sung Yong Shin, and Kyung-Yong Chwa. An optimal algorithm for finding the edge visibility polygon under limited visibility. Information Processing Letters, 53(6): 359–365, March 24, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

King:1997:OEP

[KPRS97a] Valerie King, Chung Keung Poon, Vijaya Ramachandran, and Santanu Sinha. An optimal EREW PRAM algorithm for minimum spanning tree verification. Information Processing Letters, 62(3):153–159, June 4, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kumar:1997:NOR

[KPRS97b] S. Ravi Kumar, Rina Panigrahy, Alexander Russell, and Ravi Sundaram. A note on optical routing on trees. Information Processing Letters, 62(6): 295–300, July 16, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kamakoti:1992:OAR

[KR92a] V. Kamakoti and C. Pandu Rangan. An optimal algorithm for reconstructing a binary tree. Information Processing Letters, 42 (2):113–115, May 11, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ko:1992:GPB

[KR92b] Chun Wa Ko and Frank Ruskey.

Generating permutations of a bag by interchanges. Information Processing Letters, 41(5): 263–269, April 3, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kucherov:1995:UGR

[KR95] Gregory Kucherov and Michaël Rusinowitch. Undecidability of ground reducibility for word rewriting systems with variables. Information Processing Letters, 53(4):209–215, February 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krizanc:1993:ISM

[Kri93] Danny Krizanc. Integer sorting on a mesh-connected array of processors. Information Processing Letters, 47(6):283–289, October 18, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krumke:1995:GCP

[Kru95] S. O. Krumke. On a generalization of the p-Center Problem. Information Processing Letters, 56(2):67–71, October 27, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kruszewski:1999:NHS

[Kru99]  Paul Kruszewski.  A note

on the Horton–Strahler number for random binary search trees. Information Processing Letters, 69(1):47–51, January 15, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/17/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/17/24/article.
pdf.

Khuller:1994:DMC

[KRY94] Samir Khuller, Balaji Raghavachari, and Neal Young. Designing multi-commodity flow trees. Information Processing Letters, 50 (1):49–55, April 8, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Klein:1990:PAE

[KS90] Philip Klein and Clifford Stein. A parallel algorithm for eliminating cycles in undirected graphs. Information Processing Letters, 34(6):307–312, May 28, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kaldewaij:1991:DTB

[KS91] Anne Kaldewaij and Berry Schoenmakers. The derivation of a tighter bound for top-down skew heaps. Information Processing Letters, 37(5):265–271, March 14, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khuller:1992:IST

[KS92] Samir Khuller and Baruch Schieber. On independent spanning trees. Information Processing Letters, 42(6):321–323, July 24, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Katz:1993:OSS

[KS93a] Matthew J. Katz and Micha Sharir. Optimal slope selection via expanders. Information Processing Letters, 47(3): 115–122, September 14, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kimbrel:1993:PAV

[KS93b] Tracy Kimbrel and Rakesh Kumar Sinha. A probabilistic algorithm for verifying matrix products using O(n^{2}) time and log_{2} n + O(1) random bits. Information Processing Letters, 45 (2):107–110, February 26, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kaplan:1994:DNP

[KS94] Haim Kaplan and Ron Shamir. The domatic number problem on some perfect graph families. Information Processing Letters, 49(1):51–56, January 14, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kovalyov:1997:BSD

[KS97a] Mikhail Y. Kovalyov and Yakov M. Shafransky. Batch scheduling with deadlines on parallel machines: An NP-hard case. Information Processing Letters, 64(2):69–74, November 5, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kratsch:1997:TDT

[KS97b] Dieter Kratsch and Lorna Stewart. Total domination and transformation. Information Processing Letters, 63(3):167–170, August 27, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kumar:1997:CEA

[KS97c] Devendra Kumar and Abraham Silberschatz. A counter-example to an algorithm for the generalized input-output construct of CSP. Information Processing Letters, 61(6):287, April 24, 1997. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Kanth:1998:EDR

[KS98a] K. V. Ravi Kanth and Ambuj Singh. Efficient dynamic range searching using data replication. Information Processing Letters, 68(2):97–105, October 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/46/19/
23/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/46/19/23/article.
pdf.

Krivelevich:1998:CRG

[KS98b] Michael Krivelevich and Benny Sudakov. Coloring random graphs. Information Processing Letters, 67(2):71–74, July 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kirousis:1996:SAS

[KST96] Lefteris M. Kirousis, Paul Spirakis, and Philippas Tsigas. Simple atomic snapshots: A linear complexity solution with unbounded time-stamps. Information Processing Letters, 58 (1):47–53, April 8, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kuo:1990:NSC

[KT90] R. T. Kuo and S. S. Tseng. The necessary and sufficient condition for the worst-case male optimal stable matching. Information Processing Letters, 34(5): 261–263, May 7, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kao:1991:OMB

[KT91] Ming-Yang Kao and Stephen R. Tate. Online matching with blocked input. Information Processing Letters, 38(3):113–116, May 17, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kitsios:1996:SOH

[KT96] Nectarios Kitsios and Athanasios Tsakalidis. Space-optimal hidden line elimination for rectangles. Information Processing Letters, 60(4):195–200, November 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

King:1997:LAD

[KT97] Gow-Hsing King and Wen-Guey Tzeng. On-line algorithms for the dominating set problem. Information Processing Letters, 61 (1):11–14, January 14, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kimura:1996:ESA

[KTS96] Shigetomo Kimura, Atsushi Togashi, and Norio Shiratori. Extension of synthesis algorithm of recursive processes to μ-calculus. Information Processing Letters, 58(2):97–104, May 13, 1996.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kucera:1997:CRF

[Kuč97] Luděk Kučera. Computing OR on a randomized fixed adversary CRCW PRAM. Information Processing Letters, 63(3): 165–166, August 27, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kucera:1999:FRI

[Kuč99a]  Antonín Kučera.  On fi-

nite representations of infinitestate behaviours. Information Processing Letters, 70(1): 23–30, April 16, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/17/
21/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/17/21/article.
pdf.

Kucera:1999:RNP

[Kuč99b] Antonín Kučera. Regularity of normed PA processes. Information Processing Letters, 72(1–2): 9–17, October 29, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/27/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/27/26/article.
pdf.

Kuhn:1998:LCV

[Küh98] Ulrich Kühn. Local calculation of Voronoi diagrams. Information Processing Letters, 68 (6):307–312, December 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kuich:1995:RCS

[Kui95] Werner Kuich. Representations and complete semiring morphisms. Information Processing Letters, 56(6):293–298, December 22, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kulasinghe:1997:CCC

[Kul97] Priyalal D. Kulasinghe. Connectivity of the crossed cube. Information Processing Letters, 61(4):221–226, March 27, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kummer:1995:LTC

[Kum95] Martin Kummer. A learningtheoretic characterization of classes of recursive functions. Information Processing Letters, 54 (4):205–211, May 26, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kuncheva:1993:GAF

[Kun93] Ludmila Kuncheva. Genetic algorithm for feature selection for parallel classifiers. Information Processing Letters, 46(4):

163–168, June 25, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kushilevitz:1997:SAL

[Kus97] Eyal Kushilevitz. A simple algorithm for learning O(log n)- term DNF. Information Processing Letters, 61(6):289–292, April 24, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kutylowski:1990:RSO

[Kut90] Mirosław Kutyłowski. Remarks on sorting and one-way multihead finite automata. Information Processing Letters, 36(4): 215–218, November 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kuzjurin:1994:MPS

[Kuz94] Nikolai N. Kuzjurin. Multiprocessor scheduling and expanders. Information Processing Letters, 51(6):315–319, September 26, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khuller:1994:PCD

[KV94] Samir Khuller and Uzi Vishkin.

On the parallel complexity of digraph reachability. Information Processing Letters, 52(5): 239–241, December 9, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kindler:1999:LFR

[KvdA99] E. Kindler and W. van der Aalst. Liveness, fairness, and recurrence in Petri nets. Information Processing Letters, 70(6): 269–274, June 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/22/
20/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/22/20/article.
pdf.

Kao:1992:APD

[KW92] Ming-Yang Kao and Fang Wan. Not all planar digraphs have small cycle separators. Information Processing Letters, 44(2): 79–83, November 19, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Keller:1995:NIC

[KW95] Jörg Keller and Thomas Walle. A note on implementing combining networks. Information Processing Letters, 55(4):195–200, August 25, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kindler:1997:MNF

[KW97] Ekkart Kindler and Rolf Walter. Mutex needs fairness. Information Processing Letters, 62 (1):31–39, May 7, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krumke:1998:MLS

[KW98] Sven O. Krumke and Hans-Christoph Wirth. On the minimum label spanning tree problem. Information Processing Letters, 66(2):81–85, April 29, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kwiatkowska:1990:MT

[Kwi90] Marta Z. Kwiatkowska. A metric for traces. Information Processing Letters, 35(3):129–135, July 20, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lacoste:1996:FPL

[Lac96]  Thierry Lacoste.  Finitistic

proofs of 0-1 laws for fragments of second-order logic. Information Processing Letters, 58(1):1– 4, April 8, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lam:1995:NCH

[Lam95] Nguyen Huong Lam. A note on codes having no finite completions. Information Processing Letters, 55(4):185–188, August 25, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lambert:1997:OAR

[Lam97] Timothy Lambert. An optimal algorithm for realizing a Delaunay triangulation. Information Processing Letters, 62(5):

245–250, July 2, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Langemyr:1991:CCG

[Lan91] Lars Langemyr. Circuits for computing the GCD of two polynomials over an algebraic number field. Information Processing Letters, 38(3):129–134, May 17, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Larmore:1990:OAU

[Lar90] Lawrence L. Larmore. An optimal algorithm with unknown time complexity for convex matrix searching. Information Processing Letters, 36(3):147–151, November 1, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Laroussinie:1995:AEP

[Lar95] F. Laroussinie. About the expressive power of CTL combinators. Information Processing Letters, 54(6):343–345, June 23, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Larsen:1998:REN

[Lar98] Kim S. Larsen. Regular expressions with nested levels of back referencing form a hierarchy. Information Processing Letters, 65 (4):169–172, February 27, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Latifi:1991:DSI

[Lat91] Shahram Latifi. Distributed subcube identification algorithms for reliable hypercubes. Information Processing Letters, 38 (6):315–321, June 28, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Latifi:1993:FDS

[Lat93] Shahram Latifi. On the faultdiameter of the star graph. Information Processing Letters, 46 (3):143–150, June 11, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Laumond:1990:CPT

[Lau90a] Jean-Paul Laumond. Connectivity of plane triangulations. Information Processing Letters, 34 (2):87–96, March 16, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Launchbury:1990:SAA

[Lau90b] John Launchbury. Strictness analysis aids inductive proofs. Information Processing Letters, 35(3):155–159, July 20, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liang:1995:CCB

[LB95] Y. Daniel Liang and Norbert Blum. Circular convex bipartite graphs: maximum matching and Hamiltonian circuits. Information Processing Letters, 56 (4):215–219, November 24, 1995.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luke:1998:OEM

[LBL98] Edward A. Luke, Ioana Banicescu, and Jin Li. The optimal effectiveness metric for parallel application analysis. Information Processing Letters, 66 (5):223–229, June 16, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1991:CCG

[LC91a] Myung-Joon Lee and Kwang-Moo Choe. Corrigenda: “SLR(k) covering for LR(k) grammars”. Information Processing Letters, 38(5):281, June 14, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [LC91b].

Lee:1991:CG

[LC91b] Myung-Joon Lee and Kwang-Moo Choe. SLR(k) covering for LR(k) grammars. Information Processing Letters, 37(6): 337–347, March 28, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigenda [LC91a].

Lin:1993:MVE

[LC93] Kao Chêng Lin and Maw Sheng Chern. The most vital edges in the minimum spanning tree problem. Information Processing Letters, 45(1):25–31, January 25, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1994:GCD

[LC94a] Eun-Jung Lee and Kwang-Moo Choe. Grammar coverings of a deterministic parser with action conflicts. Information Processing Letters, 51(2):85–92, July 26, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:1994:AEF

[LC94b] Tze Fen Li and Sung Wu Chang.

An algorithm to estimate the fraction defective and the exponential mean life using unlabeled samples. Information Processing Letters, 49(3):129–133, February 11, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1996:IAP

[LC96] Wei-Bin Lee and Chin-Chen Chang. Integrating authentication in public key distribution system. Information Processing Letters, 57(1):49–52, January 15, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1997:AES

[LC97a] Wei-Bin Lee and Chin-Chen Chang. Authenticated encryption schemes with linkage between message blocks. Information Processing Letters, 63(5): 247–250, September 24, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:1997:CCR

[LC97b] Min-Sheng Lin and Deng-Jyi Chen. The computational complexity of the reliability problem on distributed systems. Information Processing Letters, 64 (3):143–147, November 29, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1999:LPP

[LC99a] Gyung-Ok Lee and Kwang-Moo Choe. An LR parser with pre-determined reduction goals. Information Processing Letters, 72(5–6):189–196, December 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/60/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/60/25/31/article.
pdf.

Lee:1999:TAS

[LC99b] Jae-Ha Lee and Kyung-Yong Chwa. Tight analysis of a self-approaching strategy for the online kernel-search problem. Information Processing Letters, 69(1):39–45, January 15, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/17/
23/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/17/23/article.
pdf.

Lim:1993:SJM

[LCC93] Andrew Lim, Yeow Meng Chee, and Siu Wing Cheng. Single jog minimum area joining of compacted cells. Information Processing Letters, 47(4): 167–172, September 27, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1999:MNN

[LCC99] Ju-Hong Lee, Guang-Ho Cha, and Chin-Wan Chung. A model for k-nearest neighbor query processing cost in multidimensional data space. Information Processing Letters, 69(2): 69–76, January 29, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/18/
20/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/18/20/article.
pdf.

Lee:1993:DDB

[LCH93] DongGill Lee, Kwang-Moo Choe, and Taisook Han. A description of dynamic behavior for compilers based on object oriented modeling. Information Processing Letters, 48(4):163– 170, November 29, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:1993:ACO

[LCL93] Lung-Tien Liu, Gen-Huey Chen, and Kwei-Jay Lin. An algorithm for coalescing operations with precedence constraints in realtime systems. Information Processing Letters, 46(3):129–133, June 11, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lu:1994:CDA

[LCW94a] Erl Huei Lu, Yi Chang Chen, and Hsiao Peng Wuu. A complete decoding algorithm for double-error-correcting primitive binary BCH codes of odd m. Information Processing Letters, 51(3):117–120, August 10, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [LCW94b].

Lu:1994:CCD

[LCW94b] Erl Huei Lu, Yi Chang Cheng, and Hsiao Peng Wuu. Corrigendum: “A complete decoding algorithm for double-error-correcting primitive binary BCH codes of odd m” [Inform. Process. Lett. 51 (1994), no. 3, 117–120; 1290203]. Information Processing Letters, 52(6): 341, December 23, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [LCW94a].

Le:1997:NSC

[Lê97] Ngo.c-Minh Lê. On non-smooth convex distance functions. Information Processing Letters, 63 (6):323–329, October 8, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1990:EAG

[Lee90] De-Lei Lee. Efficient address generation in a parallel processor. Information Processing Letters, 36(3):111–116, November 1, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leermakers:1992:RAE

[Lee92] Rene Leermakers. A recursive ascent Earley parser. Information Processing Letters, 41(2): 87–91, February 14, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1996:SDM

[Lee96] Dong-Tsan Lee. Semantic data modelling using linear logic. Information Processing Letters, 60 (1):19–27, October 14, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leino:1995:CPE

[Lei95] K. Rustan M. Leino. Constructing a program with exceptions. Information Processing Letters, 53(3):159–163, February 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leung:1992:NFA

[Leu92]  Hing Leung.  A note on

finitely ambiguous distance automata. Information Processing Letters, 44(6):329–331, December 28, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lu:1992:CGR

[LH92] James J. Lu and Lawrence J. Henschen. The completeness of gp-resolution for annotated logics. Information Processing Letters, 44(3):135–140, November 30, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1993:SPE

[LH93] Sang Ho Lee and Lawrence J. Henschen. Semantics and properties of existential quantifiers in deductive databases. Information Processing Letters, 47(2): 101–108, August 20, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:1995:FRS

[LH95] Hung-Yu Lin and Lein Harn. Fair reconstruction of a secret. Information Processing Letters, 55(1):45–47, July 7, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1999:STI

[LHW99] Narn-Yih Lee, Tzonelih Hwang, and Chih-Hung Wang. The security of two ID-based multisignature protocols for sequential and broadcasting architectures. Information Processing Letters, 70 (2):79–81, April 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/18/
21/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/18/21/article.
pdf.

Li:1990:FUM

[Li90] Shuo-Yen Robert Li. On full utilization of multi-channel capacity with priority protocol. Information Processing Letters, 36 (3):129–133, November 1, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:1994:TPS

[Li94] Huade Li. Trajectory planning in H-space. Information Processing Letters, 50(5):253–258, June 10, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:1999:CBA

[Li99] C. M. Li. A constraint-based approach to narrow search trees for satisfiability. Information Processing Letters, 71(2): 75–80, July 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/18/
21/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/18/21/article.
pdf.

Liang:1994:DTG

[Lia94a] Y. Daniel Liang. Dominations in trapezoid graphs. Information Processing Letters, 52(6): 309–315, December 23, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liang:1994:FVS

[Lia94b] Y. Daniel Liang. On the feedback vertex set problem in permutation graphs. Information Processing Letters, 52(3):123– 129, November 11, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liang:1995:FPA

[Lia95a] Weifa Liang. Fast parallel algorithms for the approximate edge-coloring problem. Information Processing Letters, 55(6): 333–338, September 29, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liang:1995:SSC

[Lia95b] Y. Daniel Liang. Steiner set and connected domination in trapezoid graphs. Information Processing Letters, 56(2):101–108, October 27, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liberatore:1998:UMR

[Lib98] Vincenzo Liberatore. Uniform multipaging reduces to paging. Information Processing Letters, 67(1):9–12, July 16, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lim:1996:MAJ

[Lim96] Andrew Lim. Minimum area joining of k compacted cells. Information Processing Letters, 58 (6):265–269, June 24, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:1990:PAG

[Lin90] Chau-Jy Lin. Parallel algorithm for generating permutations on linear array. Information Processing Letters, 35(4):167–170, August 7, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:1991:FAL

[Lin91a] Rong Lin. Fast algorithms for lowest common ancestors on a processor array with reconfigurable buses. Information Processing Letters, 40(4):223–230, November 25, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lingas:1991:BCM

[Lin91b] Andrzej Lingas. Bit complexity of matrix products. Information Processing Letters, 38(5): 237–242, June 14, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lins:1992:CRC

[Lin92] Rafael D. Lins. Cyclic reference counting with lazy markscan. Information Processing Letters, 44(4):215–220, December 10, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:1993:PTS

[Lin93a] Yi-Bing Lin. Parallel tracedriven simulation of packetswitched multiplexer under priority scheduling policy. Information Processing Letters, 47(4): 197–201, September 27, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lins:1993:GCR

[Lin93b] Rafael D. Lins. Generational cyclic reference counting. Information Processing Letters, 46 (1):19–20, April 29, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:1994:DGP

[Lin94] Yi-Bing Lin. Determining the global progress of parallel simulation with FIFO communication property. Information Processing Letters, 50(1):13–17, April 8, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:1996:POM

[Lin96] Yen-Chun Lin. Perfectly overlapped merging and sorting on a two-way linear array. Information Processing Letters, 60 (4):183–187, November 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lipton:1996:PGN

[Lip96] Richard J. Lipton. On proving that a graph has no large clique:

A connection with Ramsey theory. Information Processing Letters, 58(1):39–42, April 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liskiewicz:1993:RBD

[Liś93] Maciej Liśkiewicz. On the relationship between deterministic time and deterministic reversal. Information Processing Letters, 45(3):143–146, March 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Litovsky:1991:PFL

[Lit91] Igor Litovsky. Prefix-free languages as ω-generators. Information Processing Letters, 37 (2):61–65, January 31, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Litow:1992:IIP

[Lit92] Bruce Litow. On iterated integer product. Information Processing Letters, 42(5):269–272, July 3, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Litow:1995:IGS

[Lit95] B. Litow. The influence of graph structure on generalized dimension exchange. Information Processing Letters, 54(6):347–353, June 23, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:1990:NGB

[Liu90] Zhen Liu. A note on Graham’s bound. Information Processing Letters, 36(1):1–5, October 1, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:1992:EPS

[LK92] Xiao Qing Liu and Junguk L. Kim. An efficient parallel sorting algorithm. Information Processing Letters, 43(3):129–133, September 14, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1995:RAD

[LK95] Soojung Lee and Junguk L. Kim. Resolving all deadlocks in distributed systems. Information Processing Letters, 55(5): 265–271, September 15, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Levcopoulos:1996:TLB

[LK96] Christos Levcopoulos and Drago Krznaric. Tight lower bounds for minimum weight triangulation heuristics. Information Processing Letters, 57(3):129–135, February 12, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1997:CEG

[LK97] Gyung-Ok Lee and Do-Hyung Kim. Characterization of extended LR(k) grammars. Information Processing Letters, 64 (2):75–82, November 5, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1994:PPE

[LKC94] Su-Hyun Lee, Do-Hyung Kim, and Kwang-Moo Choe. Path for AND-parallel execution of logic programs. Information Processing Letters, 52(4):191–199, November 25, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luo:1991:ODR

[LKCNW91] Kenneth Luo, William Klostermeyer, Yuan-Chieh Chow, and Richard Newman-Wolfe. Optimal deadlock resolutions in edge-disjoint reducible wait-for graphs. Information Processing Letters, 38(6):307–313, June 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liaw:1994:OAS

[LL94a] Bern Cherng Liaw and R. C. T.

Lee. An optimal algorithm to solve the minimum weakly cooperative guards problem for 1spiral polygons. Information Processing Letters, 52(2):69–75, October 28, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:1994:MFA

[LL94b] Chiu-Chuan Lin and Ferng-Ching Lin. Minimal fully adaptive wormhole routing on hypercubes. Information Processing Letters, 50(6):297–301, June 27, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lim:1995:SPP

[LL95] Chae Hoon Lim and Pil Joong Lee. Several practical protocols for authentication and key exchange. Information Processing Letters, 53(2):91–96, January 27, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lenhart:1996:DOM

[LL96] William Lenhart and Giuseppe Liotta. Drawing outerplanar minimum weight triangulations. Information Processing Letters, 57(5):253–260, March 11, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lakhal:1998:PAE

[LL98] J. Lakhal and L. Litzler. A polynomial algorithm for the extendability problem in bipartite graphs. Information Processing Letters, 65(1):11–16, January 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Levene:1999:HPI

[LL99a] Mark Levene and George Loizou. How to prevent interaction of functional and inclusion dependencies. Information Processing Letters, 71(3–4): 115–125, August 27, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/28/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/28/27/article.
pdf.

Lin:1999:FOP

[LL99b] Yen-Chun Lin and Chun-Keng Liu. Finding optimal parallel prefix circuits with fan-out 2 in constant time. Information Processing Letters, 70(4): 191–195, May 28, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/20/
21/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/20/21/article.
pdf.

Lecroq:1997:FLS

[LLM97] Thierry Lecroq, Guillaume Luce, and Jean Frédéric Myoupo. A faster linear systolic algorithm for recovering a longest common subsequence. Information Processing Letters, 61(3): 129–136, March 14, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lodi:1990:RTS

[LLP90] E. Lodi, F. Luccio, and L. Pagli.

Routing in times square mode. Information Processing Letters, 35(1):41–48, June 15, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Longpre:1993:SIO

[LM93] Luc Longpré and Sarah Mocas. Symmetry of information and one-way functions. Information Processing Letters, 46(2): 95–100, May 17, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Litow:1998:NAC

[LM98] B. Litow and B. Mans. A note on the Ádám conjecture for double loops. Information Processing Letters, 66(3):149–153, May 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leuchner:1994:NES

[LMS94] John H. Leuchner, Les Miller, and Giora Slutzki. A note on the equivalence of a set of egds to a set of FDs. Information Processing Letters, 49(4):185–188, February 25, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liang:1996:NAD

[LMS96] Weifa Liang, Brendan D. McKay, and Hong Shen. NC algorithms for dynamically solving the all pairs shortest paths problem and related problems. Information Processing Letters, 58 (3):149–155, May 13, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lagogiannis:1999:NAR

[LMT99] George Lagogiannis, Christos Makris, and Athanasios Tsakalidis. A new algorithm for rectangle enclosure reporting. Information Processing Letters, 72(5–6):177–182, December 30, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/60/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/60/25/29/article.
pdf.

Lambrichts:1990:CFC

[LNP^{+}90] Erik Lambrichts, Peter Nees, Jan Paredaens, Peter Peelman, and Letizia Tanca. Checking functional consistency in deductive databases. Information Processing Letters, 35(3):115–120, July 20, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:1990:FPA

[LO90] R. Lin and S. Olariu. A fast parallel algorithm to recognize partitionable graphs. Information Processing Letters, 36(3): 153–157, November 1, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lopez-Ortiz:1994:NLB

[LO94] Alejandro López-Ortiz. New lower bounds for element distinctness on a one-tape Turing machine. Information Processing Letters, 51(6):311–314, September 26, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Loescher:1997:OUF

[Loe97] Bernd Loescher. One unary function says less than two in existential second order logic. Information Processing Letters, 61 (2):69–75, February 28, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Long:1994:HLL

[Lon94] Philip M. Long. Halfspace learning, linear programming, and nonmalicious distributions. Information Processing Letters, 51(5):245–250, September 12, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lowe:1995:ANS

[Low95] Gavin Lowe. An attack on the Needham–Schroeder publickey authentication protocol. Information Processing Letters, 56 (3):131–133, November 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Low:1998:FSA

[Low98a] Chor Ping Low. A fast search algorithm for the quorumcast routing problem. Information Processing Letters, 66(2):87–92, April 29, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Low:1998:PTS

[Low98b] Chor-Ping Low. A polynomial time solvable instance of the feasible minimum cover problem.

Information Processing Letters, 68(3):141–146, November 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/ipl/cas_sub/browse/
browse.cgi?year=1998&volume=
68&issue=3&aid=2075.

Lozinskii:1992:CPM

[Loz92] Eliezer L. Lozinskii. Counting propositional models. Information Processing Letters, 41(6): 327–332, April 17, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lipton:1990:PIP

[LP90] Richard J. Lipton and Arvin Park. The processor identity problem. Information Processing Letters, 36(2):91–94, October 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Levcopoulos:1991:SAS

[LP91a] Christos Levcopoulos and Ola Petersson. Splitsort—an adaptive sorting algorithm. Information Processing Letters, 39(4): 205–211, August 30, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luccio:1991:EAS

[LP91b] Fabrizio Luccio and Linda Pagli.

An efficient algorithm for some tree matching problems. Information Processing Letters, 39 (1):51–57, July 12, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1993:APQ

[LP93a] D. T. Lee and E. Papadopoulou. The all-pairs quickest path problem. Information Processing Letters, 45(5):261–267, April 2, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luchetti:1993:SCB

[LP93b] Carlo Luchetti and M. Cristina Pinotti. Some comments on building heaps in parallel. Information Processing Letters, 47(3):145–148, September 14, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ladermann:1994:NLD

[LP94a] M. Ladermann and H. Petersen. Notes on looping deterministic two-way pushdown automata. Information Processing Letters, 49(3):123–127, February 11, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luccio:1994:PLU

[LP94b] F. Luccio and A. Pedrotti. A parallel list update problem. Information Processing Letters, 52 (5):277–284, December 9, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lalgudi:1997:CSS

[LP97a] Kumar N. Lalgudi and Marios C. Papaefthymiou. Computing strictly-second shortest paths. Information Processing Letters, 63(4):177–181, September 10, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luccio:1997:IPC

[LP97b] Fabrizio Luccio and Linda Pagli.

An insight on PRAM computational bounds. Information Processing Letters, 63(6):331–336, October 8, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1990:UBB

[LPK90] Jae Moon Lee, Jong Soo Park, and Myunghwan Kim. An upper bound on buffer size for join operation using nonclustered indexes. Information Processing Letters, 36(2):85–90, October 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luccio:1991:APU

[LPP91] Fabrizio Luccio, Andrea Pietracaprina, and Geppino Pucci. Analysis of parallel uniform hashing. Information Processing Letters, 37(2):67–69, January 31, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lotker:1999:NRM

[LPS99] Zvi Lotker and Boaz Patt-Shamir. A note on randomized mutual search. Information Processing Letters, 71(5–6):187– 191, September 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/49/30/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/49/30/28/article.
                                        pdf.

Leiss:1991:ECB

[LR91] Ernst L. Leiss and Hari N. Reddy. Embedding complete binary trees into hypercubes. Information Processing Letters, 38 (4):197–199, May 31, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liang:1993:FMM

[LR93] Y. Daniel Liang and Chongkye Rhee. Finding a maximum matching in a circular-arc graph. Information Processing Letters, 45(4):185–190, March 22, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Larmore:1994:OST

[LR94] Lawrence L. Larmore and Wojciech Rytter. An optimal sublinear time parallel algorithm for some dynamic programming problems. Information Processing Letters, 52(1):31–34, October 14, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liang:1996:FBC

[LR96] Y. Daniel Liang and Chongkye Rhee. Finding biconnected components in O(n) time for a class of graphs. Information Processing Letters, 60(3):159–163, November 11, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liang:1991:NAD

[LRDL91] Y. Liang, C. Rhee, S. K. Dhall, and S. Lakshmivarahan. A new approach for the domination problem on permutation graphs. Information Processing Letters, 37(4):219–224, February 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leung:1995:SPD

[LRH^{+}95] H. Leung, D. Ranjan, H. J.

Hernández, D. Tang, and A. González. A simple proof on the decidability of equivalence between recursive and nonrecursive Datalog programs. Information Processing Letters, 55(5): 279–282, September 15, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Loui:1990:ALB

[LS90] Michael C. Loui and Milind A.

Sohoni. An algorithm for load balancing in multiprocessor systems. Information Processing Letters, 35(5):223–228, August 28, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lobo:1992:RMM

[LS92] Jorge Lobo and V. S. Subrahmanian. Relating minimal models and pre-requisite-free normal defaults. Information Processing Letters, 44(3):129–133, November 30, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lawler:1995:AUG

[LS95a] Eugene L. Lawler and Sergei Sarkissian. An algorithm for “Ulam’s Game” and its application to error correcting codes. Information Processing Letters, 56(2):89–93, October 27, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Likas:1995:PAM

[LS95b] Aristidis Likas and Andreas Stafylopatis. A parallel algorithm for the minimum weighted vertex cover problem. Information Processing Letters, 53 (4):229–234, February 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lavault:1999:WCA

[LS99a] Christian Lavault and S. Mohamed Sedjelmaci. Worst-case analysis of Weber’s GCD algorithm. Information Processing Letters, 72(3–4):125– 130, November 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/28/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/

29/28/24/60/28/33/article.
pdf.

Leino:1999:VCS

[LS99b] K. Rustan M. Leino and Raymie Stata. Virginity: A contribution to the specification of object-oriented software. Information Processing Letters, 70 (2):99–105, April 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/18/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/18/25/article.
pdf.

Larsen:1992:NFR

[LSS92] Kim S. Larsen, Michael I.

Schwartzbach, and Erik M. Schmidt. A new formalism for relational algebra. Information Processing Letters, 41(3):163– 168, March 6, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lobbing:1998:POC

[LSW98] Martin Löbbing, Detlef Sieling, and Ingo Wegener. Parity OB-DDs cannot be handled efficiently enough. Information Processing Letters, 67(4):163– 168, August 31, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luby:1993:OSV

[LSZ93] Michael Luby, Alistair Sinclair, and David Zuckerman. Optimal speedup of Las Vegas algorithms. Information Processing Letters, 47(4):173–180, September 27, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lu:1997:LTA

[LT97] Chin Lung Lu and Chuan Yi Tang. A linear-time algorithm for the weighted feedback vertex problem on interval graphs. Information Processing Letters, 61 (2):107–111, February 28, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Laih:1995:SLF

[LTT95] Chi Sung Laih, Fu Kuan Tu, and Wen Chung Tai. On the security of the Lucas function. Information Processing Letters, 53(5): 243–247, March 10, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lu:1993:PMA

[Lu93] Jian Lu. Parallelizing Mallat algorithm for 2-D wavelet transforms. Information Processing Letters, 45(5):255–259, April 2, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lubiw:1990:CCS

[Lub90] Anna Lubiw. Counterexample to a conjecture of Szymanski on hypercube routing. Information Processing Letters, 35(2):57–61, June 29, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lucas:1998:RNA

[Luc98a] Salvador Lucas. Root-neededness and approximations of neededness. Information Processing Letters, 67(5):245–254, September 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luccio:1998:AEM

[Luc98b] Flaminia L. Luccio. Almost exact minimum feedback vertex set in meshes and butterflies. Information Processing Letters, 66 (2):59–64, April 29, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lutz:1990:SPD

[Lut90] Earlin Lutz. Some proofs of data refinement. Information Processing Letters, 34(4):179–185, April 24, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:1992:ACC

[LV92a] Ming Li and P. M. B. Vitanyi.

Average case complexity under the universal distribution equals worst-case complexity. Information Processing Letters, 42(3): 145–149, May 25, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:1992:OWF

[LV92b] Ming Li and Paul M. B. Vitányi.

Optimality of wait-free atomic multiwriter variables. Information Processing Letters, 43(2):

107–112, August 24, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:1992:AAG

[LV92c] Jyh-Han Lin and Jeffrey Scott Vitter. Approximation algorithms for geometric median problems. Information Processing Letters, 44(5):245–249, December 21, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luo:1992:FET

[LvBDW92] Gang Luo, Gregor v. Bochmann, Anindya Das, and Cheng Wu. Failure-equivalent transformation of transition systems to avoid internal actions. Information Processing Letters, 44 (6):333–343, December 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lai:1993:EHP

[LW93] Ten Hwang Lai and Shu-Shang Wei. The Edge Hamiltonian Path Problem is NP-complete for bipartite graphs. Information Processing Letters, 46(1): 21–26, April 29, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Libkin:1994:CNR

[LW94] Leonid Libkin and Limsoon Wong. Conservativity of nested relational calculi with internal generic functions. Information Processing Letters, 49(6):273– 280, March 22, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leung:1995:TBH

[LW95a] Joseph Y.-T.-T. Leung and W.- D.-D. Wei. Tighter bounds on a heuristic for a partition problem. Information Processing Letters, 56(1):51–57, October 13, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Libkin:1995:RQI

[LW95b] Leonid Libkin and Limsoon Wong. On representation and querying incomplete information in databases with bags. Information Processing Letters, 56 (4):209–214, November 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lan:1999:LTA

[LWS99] Yu-Feng Lan, Yue-Li Wang, and Hitoshi Suzuki. A linear-time algorithm for solving the center problem on weighted cactus graphs. Information Processing Letters, 71(5–6):205– 212, September 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/30/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/30/31/article.
pdf.

Lin:1999:STP

[LX99] Guo-Hui Lin and Guoliang Xue. Steiner tree problem with minimum number of Steiner points and bounded edge-length. Information Processing Letters, 69(2):53–57, January 29, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/47/18/
                                        17/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/47/18/17/article.
                                        pdf.

Leung:1990:PSM

[LY90] Joseph Y.-T. Leung and Gilbert H. Young. Preemptive scheduling to minimize mean weighted flow time. Information Processing Letters, 34(1):47–50, February 22, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:1992:INS

[LY92] Zhiyong Liu and Jia-Huai You. An implementation of a nonlinear skewing scheme. Information Processing Letters, 42(4): 209–215, June 19, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1998:DSA

[LY98] Kyung Oh Lee and Heon Y. Yeom. A dynamic scheduling algorithm for large scale multimedia servers. Information Processing Letters, 68(5):235–240, December 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:1999:EPP

[LY99] Y.-C. Lin and C.-S. Yeh. Efficient parallel prefix algorithms on multiport messagepassing systems. Information Processing Letters, 71(2): 91–95, July 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/18/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/18/24/article.
pdf.

Liu:1999:SGU

[LYC99] Zhaohui Liu, Wenci Yu, and T. C. Edwin Cheng. Scheduling groups of unit length jobs on two identical parallel machines. Information Processing Letters, 69(6):275– 281, March 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/22/
19/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/22/19/article.
pdf.

Lysne:1994:EBM

[Lys94] Olav Lysne. Extending Bachmair’s method for proof by consistency to the final algebra. Information Processing Letters, 51(6):303–310, September 26, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morrison:1990:CBM

[MABD90] R. Morrison, M. P. Atkinson, A. L. Brown, and A. Dearle. On the classification of binding mechanisms. Information Processing Letters, 34(1):51–55, February 22, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Macarie:1995:DBT

[Mac95] Ioan I. Macarie. Decreasing the bandwidth of a transition matrix. Information Processing Letters, 53(6):315–320, March 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maes:1990:CSC

[Mae90] Maurice Maes. On a cyclic string-to-string correction problem. Information Processing Letters, 35(2):73–78, June 29, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mahmoud:1991:MEB

[Mah91] Sabri A. Mahmoud. Motion estimation based on modified Fourier spectrum. Information Processing Letters, 37(6): 311–313, March 28, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mahmoud:1998:RFB

[Mah98] Hosam M. Mahmoud. On rotations in fringe-balanced binary trees. Information Processing Letters, 65(1):41–46, January 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maire:1994:PPG

[Mai94a] Frédéric Maire. Polyominos and perfect graphs. Information Processing Letters, 50(2):57–61, April 22, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mairson:1994:GWC

[Mai94b] Harry G. Mairson. Generating words in a context-free language uniformly at random. Information Processing Letters, 49(2): 95–99, January 28, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makinen:1989:SIP

[Mäk89] Erkki Mäkinen. On the subtree isomorphism problem for ordered trees. Information Processing Letters, 32(5):271–273, September 22, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also comments [Gro91a].

Makarov:1990:SFA

[Mak90a] O. M. Makarov. On the synthesis of fast algorithms for signal processing. Information Processing Letters, 36(5):267–272, December 1, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makinen:1990:GIP

[Mak90b] Erkki Makinen. The grammatical inference problem for the Szilard languages of linear grammars. Information Processing Letters, 36(4):203–206, November 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makinen:1992:SGI

[Mäk92a] Erkki Mäkinen. On the structural grammatical inference problem for some classes of context-free grammars. Information Processing Letters, 42 (1):1–5, April 27, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makinen:1992:RSG

[Mäk92b] Erkki Mäkinen. Remarks on the structural grammatical inference problem for context-free grammars. Information Processing Letters, 44(3):125–127, November 30, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makinen:1997:IUT

[Mäk97]  Erkki Mäkinen.  Inferring

uniquely terminating regular languages from positive data. Information Processing Letters, 62(2):57–60, May 21, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Malvestuto:1998:CAF

[Mal98] F. M. Malvestuto. A complete axiomatization of full acyclic join dependencies. Information Processing Letters, 68(3): 133–139, November 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/ipl/cas_sub/browse/
browse.cgi?year=1998&volume=
68&issue=3&aid=2078.

Manber:1990:RBS

[Man90]  Udi Manber.  Recognizing

breadth-first search trees in linear time. Information Processing Letters, 34(4):167–171, April 24, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manzini:1991:RSH

[Man91] Giovanni Manzini. Radix sort on the hypercube. Information Processing Letters, 38(2):77–81, April 26, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manzini:1994:SMV

[Man94] Giovanni Manzini. Sparse matrix vector multiplication on distributed architectures: Lower bounds and average complexity results. Information Processing Letters, 50(5):231–238, June 10, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manku:1996:LTA

[Man96] Gurmeet Singh Manku. A linear time algorithm for the bottleneck biconnected spanning subgraph problem. Information Processing Letters, 59(1):1– 7, July 8, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mandoiu:1998:OEP

[Măn98] Ion I. Măndoiu. Optimum extensions of prefix codes. Information Processing Letters, 66 (1):35–40, April 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manos:1999:CH

[Man99] H. Manos. Construction of halvers. Information Processing Letters, 69(6):303–307, March 26, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/47/22/
                                        23/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/47/22/23/article.
                                        pdf.

Martin:1990:NDO

[Mar90] Ursula Martin. A note on division orderings on strings. Information Processing Letters, 36 (5):237–240, December 1, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martel:1991:SAM

[Mar91] Charles Martel. Self-adjusting multi-way search trees. Information Processing Letters, 38(3): 135–141, May 17, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martel:1994:MFM

[Mar94] Charles U. Martel. Maximum finding on a multiple access broadcast network. Information Processing Letters, 52(1):7–13, October 14, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martinelli:1998:IAS

[Mar98] Fabio Martinelli. An improvement of algorithms for solving interface equations. Information Processing Letters, 67(4): 185–190, August 31, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Massart:1991:ACS

[Mas91] T. Massart. An agent calculus with simple actions where the enabling and disabling are derived operators. Information Processing Letters, 40(4):213– 218, November 25, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Matsliach:1990:PAF

[Mat90] Gabriel Matsliach. Performance analysis of file organizations that use multi-bucket data leaves. Information Processing Letters, 36 (6):301–310, December 15, 1990.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Matousek:1991:CD

[Mat91a] Jiri Matousek. Computing dominances in E^{n}. Information Processing Letters, 38(5):277–278, June 14, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Matousek:1991:ROA

[Mat91b] Jiří Matoušek. Randomized optimal algorithm for slope selection. Information Processing Letters, 39(4):183–187, August 30, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

MPCG:1995:FPC

[Mat95a] Mathematics of Program Construction Group. Fixed-point calculus. Information Processing Letters, 53(3):131–136, February 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Matousek:1995:EPC

[Mat95b] Jiří Matoušek. On enclosing k points by a circle. Information Processing Letters, 53 (4):217–221, February 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Miyadera:1998:DFL

[MAUY98] Youzou Miyadera, Koushi Anzai, Hiroshi Unno, and Takeo Yaku. Depth-first layout algorithm for trees. Information Processing Letters, 66(4): 187–194, May 29, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mazurkiewicz:1993:DDM

[Maz93]  Antoni Mazurkiewicz.  Dis-

tributed disassembly of mosaics. Information Processing Letters, 46(4):173–178, June 25, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mazzanti:1995:SIC

[Maz95] S. Mazzanti. Succinct iterative characterizations of primitive computable unary functions. Information Processing Letters, 56(6):315–319, December 22, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mazurkiewicz:1997:DE

[Maz97]  Antoni Mazurkiewicz.  Dis-

tributed enumeration. Information Processing Letters, 61(5): 233–239, April 10, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manber:1991:ASM

[MBY91] Udi Manber and Ricardo Baeza-Yates. An algorithm for string matching with a sequence of don’t cares. Information Processing Letters, 37(3):133–136, February 18, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Majster-Cederbaum:1992:EEB

[MC92] Mila E. Majster-Cederbaum. Ensuring the existence of a BCNF-decomposition that preserves functional dependencies in O(N^{2}) time. Information Processing Letters, 43(2):95– 100, August 24, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

McCune:1998:APC

[McC98] William McCune. Automatic proofs and counterexamples for some ortholattice identities. Information Processing Letters, 65 (6):285–291, March 27, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Majster-Cederbaum:1998:TSE

[MCR98] Mila Majster-Cederbaum and Markus Roggenbach. Transition systems from event structures revisited. Information Processing Letters, 67(3):119–124, August 17, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mok:1997:IDI

[ME97] Wai Yin Mok and David W. Embley. On improving dependency implication algorithms. Information Processing Letters, 64 (3):135–141, November 29, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Megiddo:1992:NAL

[Meg92] Nimrod Megiddo. A note on approximate linear programming. Information Processing Letters, 42(1):53, April 27, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Megson:1993:SPA

[Meg93] G. M. Megson. Systolic partitioning algorithms. Information Processing Letters, 46(1):13–18, April 29, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mehlhorn:1988:FAA

[Meh88] Kurt Mehlhorn. A faster approximation algorithm for the Steiner problem in graphs. Information Processing Letters, 27 (3):125–128, March 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See note [Flo91].

Meinel:1990:LVC

[Mei90] Christoph Meinel. Logic vs.

complexity theoretic properties of the graph accessibility problem for directed graphs of bounded degree. Information Processing Letters, 34(3):143– 146, April 9, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Meidanis:1991:LBA

[Mei91] João Meidânis. Lower bounds for arithmetic problems. Information Processing Letters, 38 (2):83–87, April 26, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Melancon:1996:VFI

[Mel96] Guy Melançon. Viennot factorization of infinite words. Information Processing Letters, 60 (2):53–57, October 28, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Messeguer:1991:DBU

[Mes91] Xavier Messeguer. Dynamic behaviour in updating process over BST of size two with probabilistic deletion algorithms. Information Processing Letters, 38(2): 89–100, April 26, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Megson:1990:IMT

[MG90] G. M. Megson and F. M. F. Gaston. Improved matrix triangularisation using a double pipeline systolic array. Information Processing Letters, 36(2):103–109, October 15, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Misra:1992:CPV

[MG92] J. Misra and David Gries. A constructive proof of Vizing’s theorem. Information Processing Letters, 41(3):131–133, March 6, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makki:1996:DAD

[MH96] S. A. M. Makki and George Havas. Distributed algorithms for depth-first search. Information Processing Letters, 60(1): 7–12, October 14, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mei:1994:ESR

[MI94] Aohan Mei and Yoshihide Igarashi. An efficient strategy for robot navigation in unknown environment. Information Processing Letters, 52(1):51–56, October 14, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Micheneau:1997:DHC

[Mic97] C. Micheneau. Disjoint Hamiltonian cycles in recursive circulant graphs. Information Processing Letters, 61(5):259–264, April 10, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Middendorf:1993:MBT

[Mid93] Martin Middendorf. Minimum broadcast time is NP-complete for 3-regular planar graphs and deadline 2. Information Processing Letters, 46(6):281–287, July 26, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Miller:1992:GHA

[Mil92a] L. L. Miller. Generating hinges from arbitrary subhypergraphs. Information Processing Letters, 41(6):307–312, April 17, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Miltersen:1992:CDR

[Mil92b] Peter Bro Miltersen. Circuit depth relative to a random oracle. Information Processing Letters, 42(6):295–298, July 24, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Minsker:1991:TAP

[Min91] Steven Minsker. The Towers of Antwerpen problem. Information Processing Letters, 38(2): 107–111, April 26, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Misra:1991:PS

[Mis91] Jayadev Misra. Phase synchronization. Information Processing Letters, 38(2):101–105, April 26, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigenda [Mis92].

Misra:1992:CPS

[Mis92] Jayadev Misra. Corrigenda: “Phase synchronization” [Inform. Process. Lett. 38(2), 26 April 1991, pp. 101–105]. Information Processing Letters, 41 (1):59, January 21, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Mis91].

Mitrana:1998:PM

[Mit98] Victor Mitrana. Primitive morphisms. Information Processing Letters, 64(6):277–281, January 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mizuno:1999:SAD

[Miz99] Masaaki Mizuno. A structured approach for developing concurrent programs in Java. Information Processing Letters, 69(5): 233–238, March 12, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/21/
19/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/21/19/article.
pdf.

Morii:1992:PSP

[MK92] Masakatu Morii and Masao Kasahara. Perfect staircase profile of linear complexity for finite sequences. Information Processing Letters, 44(2):85–89, November 19, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mayo:1994:DTD

[MK94] Jean Mayo and Phil Kearns.

Distributed termination detection with roughly synchronized clocks. Information Processing Letters, 52(2):105–108, October 28, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Murshed:1996:SDT

[MK96] M. Manzur Murshed and M. Kaykobad. Seek distances in two-headed disk systems. Information Processing Letters, 57 (4):205–209, February 26, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

May:1995:NSD

[ML95] J. H. R. May and A. D. Lunn. New statistics for demand-based software testing. Information Processing Letters, 53(6):307– 314, March 24, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manacher:1992:INV

[MM92] Glenn K. Manacher and Terrance A. Mankus. Incorporating negative-weight vertices in certain vertex-search graph algorithms. Information Processing Letters, 42(6):293–294, July 24, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morgan:1996:U

[MM96] Carroll Morgan and Annabelle McIver. Unifying wp and wlp. Information Processing Letters, 59(3):159–163, August 12, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Matamala:1999:CSC

[MM99] Martin Matamala and Klaus Meer. On the computational structure of the connected components of a hard problem. Information Processing Letters, 72(3–4):83–90, November 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/28/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/28/26/article.
pdf.

Mehta:1991:CIA

[MMN91] Shashank Mehta, Maharaj Mukherjee, and George Nagy. Constrained integer approximation to planar line intersection. Information Processing Letters, 40(3):137–139, November 8, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manacher:1990:OAF

[MMS90] Glenn K. Manacher, Terrance A.

Mankus, and Carol Joan Smith. An optimum Θ(n log n) algorithm for finding a canonical Hamiltonian path and a canonical Hamiltonian circuit in a set of intervals. Information Processing Letters, 35(4):205–211, August 7, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mehlhorn:1990:BOD

[MN90] Kurt Mehlhorn and Stefan Näher. Bounded ordered dictionaries in O(log log n) time and O(n) space. Information Processing Letters, 35(4):183–189, August 7, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Masuyama:1992:DWG

[MN92] Shigeru Masuyama and Shozo Naito. Deciding whether graph G has page number one is in NC. Information Processing Letters, 44(1):7–10, November 9, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Meskes:1993:GSM

[MN93a] Michael Meskes and Jörg Noack. The generalized supplementary magic-sets transformation for stratified Datalog. Information Processing Letters, 47(1):31–41, August 9, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Muller:1993:PTA

[MN93b] Haiko Müller and Falk Nicolai. Polynomial time algorithms for Hamiltonian problems on bipartite distance-hereditary graphs. Information Processing Letters, 46(5):225–230, July 9, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mizuno:1998:TSS

[MN98] Masaaki Mizuno and Mikhail Nesterenko. A transformation of self-stabilizing serial model programs for asynchronous parallel computing environments. Information Processing Letters, 66 (6):285–290, June 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ma:1992:OED

[MNT92] Eva Ma, Bhagirath Narahari, and Lixin Tao. Optimal embedding of 2-D torus into ring. Information Processing Letters, 41(4):227–231, March 18, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mehlhorn:1990:HLE

[MNU90] K. Mehlhorn, S. Näher, and C. Uhrig. Hidden line elimination for isooriented rectangles. Information Processing Letters, 35(3):137–143, July 20, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mnuk:1991:DDB

[Mnu91] Michal Mnuk. A div(n) depth Boolean circuit for smooth modular inverse. Information Processing Letters, 38(3):153–156, May 17, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mogensen:1994:WEC

[Mog94] Torben Æ. Mogensen. WORM-2DPDAs: An extension to 2DP-DAs that can be simulated in linear time. Information Processing Letters, 52(1):15–22, October 14, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mok:1997:KNF

[Mok97] Wai Yin Mok. On keys and normal forms. Information Processing Letters, 62(5):255–258, July 2, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Monti:1996:CCG

[Mon96] Angelo Monti. On the computational complexity of graph closures. Information Processing Letters, 57(6):291–295, March 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morita:1992:CUO

[Mor92] Kenichi Morita. Computationuniversality of one-dimensional one-way reversible cellular automata. Information Processing Letters, 42(6):325–329, July 24, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moriya:1994:TWT

[Mor94] Etsuro Moriya. On two-way tree automata. Information Processing Letters, 50(3):117–121, May 9, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morris:1997:NDE

[Mor97]  Joseph M. Morris.  Non-

deterministic expressions and predicate transformers. Information Processing Letters, 61 (5):241–246, April 10, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mostowski:1990:AAS

[Mos90] A. W. Mostowski. Alternating automata with start formulas. Information Processing Letters, 35(6):287–290, September 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mosca:1997:PAM

[Mos97] Raffaele Mosca. Polynomial algorithms for the maximum stable set problem on particular classes of P_{5}-free graphs. Information Processing Letters, 61 (3):137–143, March 14, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Motoki:1991:IIA

[Mot91] Tatsuya Motoki. Inductive inference from all positive and some negative data. Information Processing Letters, 39(4):177–182, August 30, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mount:1999:ERF

[Mou99] John A. Mount. Estimating the range of a function in an online setting. Information Processing Letters, 72(1–2):31– 35, October 29, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/27/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/27/29/article.
pdf.

Mowbray:1992:FLS

[Mow92] Miranda Mowbray. Finitary logics for some CCS observational bisimulations. Information Processing Letters, 42(3):161–165, May 25, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mahapatra:1991:MFP

[MP91] Rabi N. Mahapatra and Harish Pareek. Modelling a fast parallel thinning algorithm for shared memory SIMD computers. Information Processing Letters, 40 (5):257–261, December 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mereghetti:1995:RMS

[MP95] Carlo Mereghetti and Giovanni Pighizzini. A remark on middle space bounded alternating Turing machines. Information Processing Letters, 56(4):229–232, November 24, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Muscholl:1996:NCC

[MP96] Anca Muscholl and Holger Petersen. A note on the commutative closure of star-free languages. Information Processing Letters, 57(2):71–74, January 29, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ma:1997:ILB

[MP97] Yuan Ma and Serge Plotkin. An improved lower bound for load balancing of tasks with unknown duration. Information Processing Letters, 62(6):301–303, July 16, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mansour:1998:LCN

[MP98a] Yishay Mansour and Michal Parnas. Learning conjunctions with noise under product distributions. Information Processing Letters, 68(4):189–196, November 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Montana:1998:KCR

[MP98b] J. L. Montaña and Luis M.

Pardo. On Kolmogorov complexity in the real Turing machine setting. Information Processing Letters, 67(2):81–86, July 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morillo:1999:WTS

[MPSV99] P. Morillo, C. Padró, G. Sáez, and J. L. Villar. Weighted threshold secret sharing schemes. Information Processing Letters, 70(5):211–216, June 21, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/23/
20/48/21/24/abstract.html;
 http://www.elsevier.nl/gej-
ng/10/23/20/48/21/24/article.
pdf.

Meijer:1992:CMW

[MR92] Henk Meijer and David Rappaport. Computing the minimum weight triangulation of a set of linearly ordered points. Information Processing Letters, 42 (1):35–38, April 27, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Monti:1995:CRC

[MR95] Angelo Monti and Alessandro Roncato. Completeness results concerning systolic tree automata and E0L languages. Information Processing Letters, 53 (1):11–16, January 13, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Marathe:1996:AAM

[MR96a] M. V. Marathe and S. S. Ravi. On approximation algorithms for the minimum satisfiability problem. Information Processing Letters, 58(1):23–29, April 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Monti:1996:GTA

[MR96b] Angelo Monti and Alessandro Roncato. A gap theorem for the anonymous torus. Information Processing Letters, 57(5): 279–285, March 11, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Muller:1996:CES

[MRR96] Michael Müller, Christine Rüb, and Wolfgang Rülling. A circuit for exact summation of floating-point numbers. Information Processing Letters, 57 (3):159–163, February 12, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mitchell:1995:CCP

[MRSW95] Joseph S. B. Mitchell, Günter Rote, Gopalakrishnan Sundaram, and Gerhard Woeginger. Counting convex polygons in planar point sets. Information Processing Letters, 56(1):45–49, October 13, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Matsliach:1990:DTL

[MS90a] Gabriel Matsliach and Oded Shmueli. Distributing a B^{+}- tree in a loosely coupled environment. Information Processing Letters, 34(6):313–321, May 28, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Murphy:1990:FNN

[MS90b] O. J. Murphy and S. M. Selkow.

Finding nearest neighbors with Voronoi tessellations. Information Processing Letters, 34(1): 37–41, February 22, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mukhopadhyaya:1992:HGM

[MS92] Krishnendu Mukhopadhyaya and Bhabani P. Sinha. Hamiltonian graphs with minimum number of edges for fault-tolerant topologies. Information Processing Letters, 44(2):95–99, November 19, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moore:1994:OAC

[MS94] Dennis Moore and W. F. Smyth. An optimal algorithm to compute all the covers of a string. Information Processing Letters, 50 (5):239–246, June 10, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moore:1995:COA

[MS95] Dennis Moore and W. F. Smyth. A correction to “An optimal algorithm to compute all the covers of a string”. Information Processing Letters, 54(2): 101–103, April 28, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Melkman:1996:APC

[MS96a] Avraham A. Melkman and Solomon E. Shimony. Algorithms for parsimonious complete sets in directed graphs. Information Processing Letters, 59(6):335–339, September 23, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Metivier:1996:MCP

[MS96b] Yves Métivier and Nasser Saheb. Medians and centres of polyominoes. Information Processing Letters, 57(4):175–181, February 26, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maggs:1998:RTE

[MS98] Bruce M. Maggs and Eric J.

Schwabe. Real-time emulations of bounded-degree networks. Information Processing Letters, 66 (5):269–276, June 16, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Macarie:1999:ASP

[MS99a] Ioan I. Macarie and Joel I.

Seiferas. Amplification of slight probabilistic advantage at absolutely no cost in space. Information Processing Letters, 72(3–4):113–118, November 26, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/28/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/28/31/article.
pdf.

Maitra:1999:HWC

[MS99b] Subhamoy Maitra and Palash Sarkar. Hamming weights of correlation immune Boolean functions. Information Processing Letters, 71(3–4):149– 153, August 27, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/28/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/23/20/49/28/31/article.
                                        pdf.

Malhotra:1996:SED

[MSK96] Vishv M. Malhotra, Bala Srinivasan, and Santosh Kulkarni. Storage-efficient data structure for large lookup dictionaries. Information Processing Letters, 58 (4):201–206, May 27, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Marchetti-Spaccamela:1996:MTO

[MSNR96] Alberto Marchetti-Spaccamela, Umberto Nanni, and Hans Rohnert. Maintaining a topological order under edge insertions. Information Processing Letters, 59(1):53–58, July 8, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Merlini:1996:UMS

[MSV96] Donatella Merlini, Renzo Sprugnoli, and M. Cecilia Verri. A uniform model for the storage utilization of B-tree-like structures. Information Processing Letters, 57(1):53–58, January 15, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makedon:1993:SLT

[MSW93] Fillia Makedon, Dafna Sheinwald, and Yaron Wolfsthal. A simple linear-time algorithm for the recognition of bandwidth-2 biconnected graphs. Information Processing Letters, 46(2): 103–107, May 17, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martelli:1992:NS

[MT92a] Maurizio Martelli and Chiara Tricomi. A new SLDNF-tree. Information Processing Letters, 43 (2):57–62, August 24, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martello:1992:NGA

[MT92b] Silvano Martello and Paolo Toth. A note on 0.5-bounded greedy algorithms for the 0/1 knapsack problem. Information Processing Letters, 44(4):221– 222, December 10, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [SCGD92b].

Merritt:1993:SLF

[MT93] Michael Merritt and Gadi Taubenfeld. Speeding Lamport’s fast mutual exclusion algorithm. Information Processing Letters, 45(3):137–142, March 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makris:1998:ATD

[MT98a] Christos Makris and Athanasios Tsakalidis. Algorithms for threedimensional dominance searching in linear space. Information Processing Letters, 66(6): 277–283, June 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Molva:1998:SSA

[MT98b] R. Molva and G. Tsudik. Secret sets and applications. Information Processing Letters, 65(1): 47–55, January 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mahajan:1994:NSF

[MTV94] Meena Mahajan, Thomas Thierauf, and N. V. Vinodchandran. A note on SpanP functions. Information Processing Letters, 51 (1):7–10, July 12, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mullin:1991:CUC

[Mul91] James K. Mullin. A caution on universal classes of hash functions. Information Processing Letters, 37(5):247–256, March 14, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Muller:1992:CLC

[Mül92] Fritz Müller. Confluence of the lambda calculus with left-linear algebraic rewriting. Information Processing Letters, 41(6): 293–299, April 17, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Murphy:1990:UFT

[Mur90] Owen Murphy. A unifying framework for trie design heuristics. Information Processing Letters, 34(5):243–249, May 7, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manolopoulos:1991:SDD

[MV91] Yannis Manolopoulos and Athina Vakali. Seek distances in disks with two independent heads per surface. Information Processing Letters, 37(1):37–42, January 10, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mahajan:1995:NMG

[MV95] Meena Mahajan and N. V. Vinodchandran. A note on Mod and generalised Mod classes. Information Processing Letters, 55 (1):27–31, July 7, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Madanlal:1996:TSI

[MVR96] M. S. Madanlal, G. Venkatesan, and C. Pandu Rangan. Tree 3-spanners on interval, permutation and regular bipartite graphs. Information Processing Letters, 59(2):97–102, July 22, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morris:1991:SCI

[MW91] Robert J. T. Morris and Wing Shing Wong. Systematic choice of initial points in local search. extensions and application to neural networks. Information Processing Letters, 39 (4):213–217, August 30, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manber:1994:AAM

[MW94] Udi Manber and Sun Wu. An algorithm for approximate membership checking with application to password security. Information Processing Letters, 50 (4):191–197, May 25, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Myoupo:1997:ILS

[MW97] Jean-Frédéric Myoupo and Ahmad Wabbi. Improved linear systolic algorithms for substring statistics. Information Processing Letters, 61(5):253–258, April 10, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martinez:1990:CRC

[MWL90] Alejandro D. Martinez, Rosita Wachenchauzer, and Rafael D. Lins. Cyclic reference counting with local mark-scan. Information Processing Letters, 34(1): 31–35, February 22, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moriya:1996:LSL

[MY96] Tetsuo Moriya and Hideki Yamasaki. Literal shuffle on ωlanguages. Information Processing Letters, 59(3):165–168, August 12, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Myers:1995:AMC

[Mye95] Gene Myers. Approximately matching context-free languages.

Information Processing Letters, 54(2):85–92, April 28, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Megson:1999:HTH

[MYL99] G. M. Megson, Xiaofan Yang, and Xiaoping Liu. Honeycomb tori are Hamiltonian. Information Processing Letters, 72(3–4): 99–103, November 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/28/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/28/28/article.
pdf.

Myoupo:1991:DPL

[Myo91] Jean Frédéric Myoupo. Dynamic programming on linear pipelines. Information Processing Letters, 39(6):333–341, September 27, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigenda [Myo92].

Myoupo:1992:CDP

[Myo92] Jean Frédéric Myoupo. Corrigenda: “Dynamic programming on linear pipelines”. Information Processing Letters, 41 (5):289, April 3, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Myo91].

Moffat:1993:SRA

[MZ93] Alistair Moffat and Justin Zobel. Supporting random access in files of variable length records. Information Processing Letters, 46(2):71–77, May 17, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ma:1999:ADR

[MZM99] Z. M. Ma, W. J. Zhang, and W. Y. Ma. Assessment of data redundancy in fuzzy relational databases based on semantic inclusion degree. Information Processing Letters, 72(1–2):25– 29, October 29, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/27/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/27/28/article.
pdf.

Nayak:1995:NIC

[NAG95] Amiya Nayak, Vincenzo Accia, and Paolo Gissi. A note on isomorphic chordal rings. Information Processing Letters, 55(6): 339–341, September 29, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [NAG96].

Nayak:1996:ENI

[NAG96] Amiya Nayak, Vincenzo Acciaro, and Paolo Gissi. Erratum: “A note on isomorphic chordal rings” [Inform. Process. Lett. 55 (1995), no. 6, 339– 341; MR 96d:68016]. Information Processing Letters, 57(2): 115, January 29, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [NAG95, CK96b].

Nassar:1995:MMM

[Nas95] Hamed Nassar. A Markov model for multibus multiprocessor systems under asynchronous operation. Information Processing Letters, 54(1):11–16, April 14, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nash:1998:SPP

[Nas98] Jonathan Nash. Scalable and predictable performance for irregular problems using the WPRAM computational model. Information Processing Letters, 66(5):237–246, June 16, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Naumann:1994:RTP

[Nau94] David A. Naumann. A recursion theorem for predicate transformers on inductive data types. Information Processing Letters, 50 (6):329–336, June 27, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nawrocki:1991:CDR

[Naw91] Jerzy R. Nawrocki. Conflict detection and resolution in a lexical analyzer generator. Information Processing Letters, 38(6): 323–328, June 28, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nassif:1996:GES

[NB96] Nayla Nassif and Nader Bagherzadeh. A grid embedding into the star graph for image analysis solutions. Information Processing Letters, 60(5):255–260, December 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Navarro:1999:VFS

[NBY99] Gonzalo Navarro and Ricardo Baeza-Yates. Very fast and simple approximate string matching. Information Processing Letters, 72(1–2):65–70, October 29, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/27/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/27/33/article.
pdf.

Nawrocki:1998:SCT

[NCC98] Jerzy R. Nawrocki, Adam Czajka, and Wojciech Complak. Scheduling cyclic tasks with binary periods. Information Processing Letters, 65(4):173–178, February 27, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nath:1999:AEL

[NCK99] Suman Kumar Nath, Rezaul Alam Chowdhury, and M. Kaykobad. On average edge length of minimum spanning trees. Information Processing Letters, 70(5):

241–243, June 21, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/21/
21/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/21/21/article.
pdf.

Nievergelt:1995:MGE

[ND95] J. Nievergelt and Narsingh Deo.

Metric graphs elastically embeddable in the plane. Information Processing Letters, 55(6): 309–315, September 29, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nievergelt:1999:MEE

[NDM99] Jurg Nievergelt, Narsingh Deo, and Ambros Marzetta. Memoryefficient enumeration of constrained spanning trees. Information Processing Letters, 72(1– 2):47–53, October 29, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/27/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/27/31/article.
pdf.

Neiger:1994:DCR

[Nei94] Gil Neiger. Distributed Consensus revisited. Information Processing Letters, 49(4):195–201, February 25, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nepomniaschy:1999:SVM

[Nep99] V. A. Nepomniaschy. Symbolic verification method for definite iteration over data structures. Information Processing Letters, 69(4):207–213, February 26, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/47/20/
                                        24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/47/20/24/article.
                                        pdf.

Neufeld:1996:PEP

[Neu96] Stewart W. Neufeld. A pursuitevasion problem on a grid. Information Processing Letters, 58 (1):5–9, April 8, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Newman:1991:PVC

[New91] Ilan Newman. Private vs. common random bits in communication complexity. Information Processing Letters, 39(2):67–71, July 31, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Narayanaswami:1992:EIH

[NF92] Chandrasekhar Narayanaswami and William Randolph Franklin. Edge intersection on the hypercube computer. Information Processing Letters, 41(5): 257–262, April 3, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Natu:1995:PPC

[NF95] Madan Natu and Shu Cherng Fang. On the point-to-point connection problem. Information Processing Letters, 53(6): 333–336, March 24, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nagamochi:1998:NMS

[NI98] Hiroshi Nagamochi and Toshihide Ibaraki. A note on minimizing submodular functions. Information Processing Letters, 67(5):239–244, September 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nieuwenhuis:1993:SLC

[Nie93] Robert Nieuwenhuis. Simple LPO constraint solving methods. Information Processing Letters, 47(2):65–69, August 20, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nielsen:1996:OSP

[Nie96]  Franck Nielsen.  Output-

sensitive peeling of convex and maximal layers. Information Processing Letters, 59(5):255– 259, September 9, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nikolik:1995:CPT

[Nik95a] Borislav Nikolik. Constraint preservation through loops. Information Processing Letters, 55 (3):143–148, August 11, 1995.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nikolopoulos:1995:CTP

[Nik95b] Stavros D. Nikolopoulos. Constanttime parallel recognition of split graphs. Information Processing Letters, 54(1):1–8, April 14, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Narayanaswami:1994:AE

[NL94] Chandrasekhar Narayanaswami and William Luken. Approximating x^{n} efficiently. Information Processing Letters, 50(4): 205–210, May 25, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nagumo:1996:LLF

[NLW96] Hideo Nagumo, Mi Lu, and Karan Watson. On-line longest fragment first parsing algorithm. Information Processing Letters, 59(2):91–96, July 22, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Neilsen:1994:NCM

[NM94] Mitchell L. Neilsen and Masaaki Mizuno. Nondominated kcoteries for multiple mutual exclusion. Information Processing Letters, 50(5):247–252, June 10, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [NM96].

Neilsen:1996:ENC

[NM96] Mitchell N. Neilsen and Masaaki Mizuno. Erratum to “Nondominated k-coteries for multiple mutual exclusion”. Information Processing Letters 50 (5) (10 June 1994) 247–252. Information Processing Letters, 60 (6):319, December 23, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [NM94].

Nardelli:1997:CPR

[NMS97a] Enrico Nardelli, Vincenzo Mastrobuoni, and Alesiano Santomo. Computing a poset from its realizer. Information Processing Letters, 64(3):149–154, November 29, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nardelli:1997:BTR

[NMS97b] Enrico Nardelli, Vincenzo Mastrobuoni, and Alesiano Santomo. On building the transitive reduction of a two-dimensional poset. Information Processing Letters, 63(1):9–12, July 30, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nieuwenhuis:1991:EDE

[NN91] Robert Nieuwenhuis and Pilar Nivela. Efficient deduction in equality Horn logic by Horncompletion. Information Processing Letters, 39(1):1–6, July 12, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nitsche:1996:ASP

[NO96] Ulrich Nitsche and Peter Ochsenschläger. Approximately satisfied properties of systems and simple language homomorphisms. Information Processing Letters, 60(4):201–206, November 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nardelli:1998:FDC

[NPW98] Enrico Nardelli, Guido Proietti, and Peter Widmayer. Finding the detour-critical edge of a shortest path between two nodes. Information Processing Letters, 67(1):51–54, July 16, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nikolopoulos:1998:NST

[NR98] S. D. Nikolopoulos and P. Rondogiannis. On the number of spanning trees of multi-star related graphs. Information Processing Letters, 65(4):183–188, February 27, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nourine:1999:FAB

[NR99] Lhouari Nourine and Olivier Raynaud. A fast algorithm for building lattices. Information Processing Letters, 71(5–6):199– 204, September 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/49/30/

30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/30/30/article.
pdf.

Natschlager:1996:EEV

[NS96a] Thomas Natschläger and Michael Schmitt. Erratum to “Exact VC-dimension of Boolean monomials” Information Processing Letters 59 (1) (8 July 1996) 19–20. Information Processing Letters, 60(2):107, October 28, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [NS96b].

Natschlager:1996:EVD

[NS96b] Thomas Natschläger and Michael Schmitt. Exact VC-dimension of Boolean monomials. Information Processing Letters, 59(1): 19–20, July 8, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [NS96a].

Nuutila:1994:FSC

[NSS94] Esko Nuutila and Eljas Soisalon-Soininen. On finding the strongly connected components in a directed graph. Information Processing Letters, 49(1): 9–14, January 14, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ntafos:1990:RRP

[Nta90] Simeon Ntafos. The robber route problem. Information Processing Letters, 34(2):59–63, March 16, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nykanen:1994:FLC

[NU94] Matti Nykänen and Esko Ukkonen. Finding lowest common ancestors in arbitrarily directed trees. Information Processing Letters, 50(6):307–310, June 27, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nuutila:1994:ETC

[Nuu94] Esko Nuutila. An efficient transitive closure algorithm for cyclic digraphs. Information Processing Letters, 52(4):207–213, November 25, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nystrom:1996:TNF

[Nys96] Sven-Olof Nyström. There is no fully abstract fixpoint semantics for non-deterministic languages with infinite computations. Information Processing Letters, 60 (6):289–293, December 23, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Oberoi:1995:CGR

[Obe95] Sundeep Oberoi. λ_{β′} — A λcalculus with a generalized βreduction rule. Information Processing Letters, 54(1):45–53, April 14, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

OConnor:1994:UBN

[O’C94]  Luke O’Connor.  An up-

per bound on the number of functions satisfying the Strict Avalanche Criterion. Information Processing Letters, 52(6): 325–327, December 23, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

OConnor:1995:NLB

[O’C95] Luke O’Connor. A new lower bound on the expected size of irredundant forms for Boolean functions. Information Processing Letters, 53(6):347–353, March 24, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Orozco:1997:RMS

[OCSF97] J. Orozco, R. Cayssials, J. Santos, and E. Ferro. 802.4 Rate Monotonic scheduling in hard real-time environments: Setting the medium access control parameters. Information Processing Letters, 62(1):47–55, May 7, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ogihara:1995:HPL

[Ogi95] Mitsunori Ogihara. On helping by parity-like languages. Information Processing Letters, 54 (1):41–43, April 14, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ogihara:1996:FCL

[Ogi96] Mitsunori Ogihara. Functions computable with limited access to NP. Information Processing Letters, 58(1):35–38, April 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Olsson:1991:ASE

[OH91] Ronald A. Olsson and Daniel T. Huang. Axiomatic semantics for ‘escape’ statements. Information Processing Letters, 39(1):27–33, July 12, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Oh:1993:AEK

[Oh93] Sang Bong Oh. An analytical evidence for Kalé’s heuristic for the N queens problem. Information Processing Letters, 46(1):51–54, April 29, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ohlebusch:1995:TMC

[Ohl95] Enno Ohlebusch. Termination is not modular for confluent variable-preserving term rewriting systems. Information Processing Letters, 53(4):223–228, February 24, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ohtsuka:1993:PSL

[Oht93] Hiroshi Ohtsuka. A proof of the substitution lemma in de Bruijn’s notation. Information Processing Letters, 46(2):63–66, May 17, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Okuno:1998:PCS

[OiMI98] Hiroshi G. Okuno, Shin ichi Minato, and Hideki Isozaki. On the properties of combination set operations. Information Processing Letters, 66(4):195–199, May 29, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Obana:1996:VIS

[OK96] Satoshi Obana and Kaoru Kurosawa. Veto is impossible in secret sharing schemes. Information Processing Letters, 58(6): 293–295, June 24, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Olariu:1990:GCS

[Ola90a] Stephan Olariu. A generalization of Chvatal’s star-cutset lemma. Information Processing Letters, 33(6):301–303, February 10, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Olariu:1990:CTG

[Ola90b] Stephan Olariu. On the closure of triangle-free graphs under substitution. Information Processing Letters, 34(2):97–101, March 16, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Olariu:1991:OGH

[Ola91] Stephan Olariu. An optimal greedy heuristic to color interval graphs. Information Processing Letters, 37(1):21–25, January 10, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Olson:1997:AAL

[Ols97] Clark F. Olson. An approximation algorithm for least median of squares regression. Information Processing Letters, 63(5): 237–241, September 24, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Orda:1997:ETC

[OM97] Ariel Orda and Michael Merritt. Efficient test&set constructions for faulty shared memory. Information Processing Letters, 62 (1):41–46, May 7, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Omondi:1991:FOM

[Omo91] Amos Omondi. Fast one’scomplement multiplication. Information Processing Letters, 39 (2):73–79, July 31, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Obokata:1996:POE

[ONI96] Koji Obokata, Yasuaki Nishitani, and Yoshihide Igarashi. A probably optimal embedding of hyper-rings in hypercubes. Information Processing Letters, 58 (3):117–122, May 13, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Oellermann:1995:PAT

[OS95] Ortrud Oellermann and Jeremy P.

Spinrad. A polynomial algorithm for testing whether a graph is 3-Steiner distance hereditary. Information Processing Letters, 55(3):149–154, August 11, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ogata:1996:MMV

[OSK96] Wakaha Ogata, Keiichi Sakano, and Kaoru Kurosawa. Multisymbol majority vote and hard core. Information Processing Letters, 58(6):285–292, June 24, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Otto:1995:SWE

[Ott95a] Friedrich Otto. Solvability of word equations modulo finite special and confluent stringrewriting systems is undecidable in general. Information Processing Letters, 53(5):237–242, March 10, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Otto:1995:NNM

[Ott95b] Martin Otto. A note on the number of monadic quantifiers in monadic Σ_{1}^{1}. Information Processing Letters, 53(6):337–339, March 24, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Otto:1996:FOQ

[OV96] Martin Otto and Jan Van den Bussche. First-order queries on databases embedded in an infinite structure. Information Processing Letters, 60(1):37–41, October 14, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Odijk:1998:ISS

[Ov98] Michiel A. Odijk and Hans van Maaren. Improved solutions to the Steiner triple covering problem. Information Processing Letters, 65(2):67–69, January 29, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Overmars:1992:PLF

[Ove92] Mark H. Overmars. Point location in fat subdivisions. Information Processing Letters, 44 (5):261–265, December 21, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Owolabi:1993:EPS

[Owo93] Olumide Owolabi. Efficient pattern searching over large dictionaries. Information Processing Letters, 47(1):17–21, August 9, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Oyang:1995:TUB

[Oya95] Yen-Jen Oyang. A tight upper bound of the lumped disk seek time for the Scan disk scheduling policy. Information Processing Letters, 54(6):355–358, June 23, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pachl:1992:CSP

[Pac92a] J. Pachl. Corrigendum: “A simple proof of a completeness result for leads-to in the UNITY logic” [Inform. Process. Lett. 41(1), 21 January 1992, pp. 35–38]. Information Processing Letters, 44(3):171, November 30, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Pac92b].

Pachl:1992:SPC

[Pac92b] Jan Pachl. A simple proof of a completeness result for leads-to in the UNITY logic. Information Processing Letters, 41(1):35–38, January 21, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [Pac92a].

Padro:1998:RVS

[Pad98] Carles Padró. Robust vector space secret sharing schemes. Information Processing Letters, 68 (3):107–111, November 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Powell:1999:VDC

[PAD99] David R. Powell, Lloyd Allison, and Trevor I. Dix. A versatile divide and conquer technique for optimal string alignment. Information Processing Letters, 70 (3):127–139, May 14, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/19/
19/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/19/19/article.
pdf.

Pallo:1990:DMB

[Pal90] J. M. Pallo. A distance metric on binary trees using latticetheoretic measures. Information Processing Letters, 34(3): 113–116, April 9, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Palsberg:1993:NFP

[Pal93] Jens Palsberg. Normal forms have partial types. Information Processing Letters, 45(1): 1–3, January 25, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Panaite:1996:HPR

[Pan96a] Petrişor Panaite. Hypercube permutations routable under all dimension orderings. Information Processing Letters, 59(4): 185–189, August 26, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Panda:1996:NLT

[Pan96b] B. S. Panda. New linear time algorithms for generating perfect elimination orderings of chordal graphs. Information Processing Letters, 58(3):111–115, May 13, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Parekh:1991:AGH

[Par91a] Abhay K. Parekh. Analysis of a greedy heuristic for finding small dominating sets in graphs. Information Processing Letters, 39(5):237–240, September 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Park:1991:SCV

[Par91b] James K. Park. A special case of the n-vertex traveling-salesman problem that can be solved in O(n) time. Information Processing Letters, 40(5):247–254, December 13, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Parberry:1995:RTA

[Par95] Ian Parberry. A real-time algorithm for the (n^{2} − 1)-puzzle. Information Processing Letters, 56(1):23–28, October 13, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Paschos:1992:AAM

[Pas92a] V. Th. Paschos. A (Δ/2)- approximation algorithm for the maximum independent set problem. Information Processing Letters, 44(1):11–13, November 9, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pascoletti:1992:OAP

[Pas92b] Adriano Pascoletti. An optimal algorithm for the period of a strongly connected digraph. Information Processing Letters, 44 (6):313–316, December 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Paulik:1990:WAG

[Pau90] A. Paulik. Worst-case analysis of a generalized heapsort algorithm. Information Processing Letters, 36(3):159–165, November 1, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Paul:1994:NBS

[Pau94] Wolfgang J. Paul. A note on bitonic sorting. Information Processing Letters, 49(5): 223–225, March 11, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Paun:1996:SST

[Pău96] Gheorghe Păun. Splicing systems with targets are computationally universal. Information Processing Letters, 59(3): 129–133, August 12, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pavlu:1992:EIL

[Pav92] Petr Pavlu. On efficient implementation of LR-attributed grammars. Information Processing Letters, 41(2):67–75, February 14, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peng:1992:SLT

[PC92] Shen Lung Peng and Maw Shang Chang. A simple linear time algorithm for the domatic partition problem on strongly chordal graphs. Information Processing Letters, 43(6):297–300, October 19, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Park:1995:UPR

[PCK95] Ung Kyu Park, Hwang Kyu Choi, and Tag Gon Kim. Uniform partitioning of relations using histogram equalization framework: An efficient parallel hash-based join. Information Processing Letters, 55(5):283– 289, September 15, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pramanik:1991:DCK

[PDBF91] P. Pramanik, P. K. Das, A. K.

Bandyopadhyay, and D. Q. M. Fay. A deadlock-free communication kernel for loop architecture. Information Processing Letters, 38(3):157–161, May 17, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pedrotti:1994:ALU

[Ped94] A. Pedrotti. Analysis of a listupdate strategy. Information Processing Letters, 52(3):115– 121, November 11, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pelc:1991:BCN

[Pel91] Andrzej Pelc. Broadcasting in complete networks with faulty nodes using unreliable calls. Information Processing Letters, 40 (3):169–174, November 8, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Penczek:1992:UPT

[Pen92] Wojciech Penczek. On undecidability of propositional temporal logics on trace systems. Information Processing Letters, 43(3):147–153, September 14, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Perleberg:1994:SCS

[Per94] Chris H. Perleberg. Single character searching methods and the shift-or pattern-matching algorithm. Information Processing Letters, 50(5):269–275, June 10, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pettersson:1991:LEC

[Pet91] Jimmi S. Pettersson. Letter to the editor: Comments on “Always-true is not invariant”:

Assertional reasoning about invariance [Inform. Process. Lett. 35(6), 15 September 1990, pp. 277–279]. Information Processing Letters, 40(5):231–233, December 13, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Ver87, VT90].

Petersen:1994:DPT

[Pet94a] H. Petersen. On the determinacy problem for two-way pushdown automata. Information Processing Letters, 52(6):323–324, December 23, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Petersen:1994:RSM

[Pet94b] H. Petersen. Refined simulation of multihead automata. Information Processing Letters, 52 (5):229–233, December 9, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Petersen:1995:SFF

[Pet95] H. Petersen. On space functions fully constructed by twodimensional Turing machines. Information Processing Letters, 54(1):9–10, April 14, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Park:1995:SAO

[PG95a] Young Park and Benjamin Goldberg. Static analysis for optimizing reference counting. Information Processing Letters, 55(4):

229–234, August 25, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Park:1995:ODA

[PG95b] Young G. Park and Benjamin Goldberg. Order-of-demand analysis for lazy languages. Information Processing Letters, 55(6):343–348, September 29, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Parker:1992:OVN

[PH92] A. Parker and J. O. Hamblen. Optimal value for the Newton– Raphson division algorithm. Information Processing Letters, 42 (3):141–144, May 25, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pin:1995:NAQ

[Pin95] Jean-Eric Pin. A negative answer to a question of Wilke on varieties of ω-languages. Information Processing Letters, 56 (4):197–200, November 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Piotrowski:1990:FMS

[Pio90] Jerzy A. Piotrowski. A functional model of a simplified sequential machine. Information Processing Letters, 35(3):161– 166, July 20, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pirillo:1994:IWB

[Pir94] Giuseppe Pirillo. Infinite words and biprefix codes. Information Processing Letters, 50(6): 293–295, June 27, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pitsch:1995:CCF

[Pit95] Gisela Pitsch. LR(k)-coupled-context-free grammars. Information Processing Letters, 55(6):349–358, September 29, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Park:1990:DSS

[PK90] Moon Hwa Park and Myunghwan Kim. A distributed synchronization scheme for fair multiprocess handshakes. Information Processing Letters, 34(3): 131–138, April 9, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pilarski:1993:SBC

[PK93] Sławomir Pilarski and Tiko Kameda. Simple bounds on the convergence rate of an ergodic Markov chain. Information Processing Letters, 45(2): 81–87, February 26, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Plaisted:1999:CSC

[PK99] David A. Plaisted and Gregory Kucherov. The complexity of some complementation problems. Information Processing Letters, 71(3–4):159– 165, August 27, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/28/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/28/33/article.
pdf.

Park:1993:TOA

[PKP93] Heonchul Park, Hyoung Joong Kim, and Viktor K. Prasanna. An O(1) time optimal algorithm for multiplying matrices on reconfigurable mesh. Information Processing Letters, 47(2): 109–113, August 20, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Park:1993:RP

[PLC93] Woo-Jun Park, Myung-Joon Lee, and Kwang-Moo Choe. On the reduction of LR(k) parsers. Information Processing Letters, 47(5):245–251, October 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Plehn:1990:PSI

[Ple90]  Jürgen Plehn.  Preemptive

scheduling of independent jobs with release times and deadlines on a hypercube. Information Processing Letters, 34(3): 161–166, April 9, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Park:1997:IBN

[PLSK97] Sung Mo Park, Sangjin Lee, Soo Hak Sung, and Kwangjo Kim. Improving bounds for the number of correlation immune Boolean functions. Information Processing Letters, 61(4): 209–212, March 27, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Panda:1994:RAI

[PM94] B. S. Panda and S. P. Mohanty.

Recognition algorithm for intersection graphs of edge disjoint paths in a tree. Information Processing Letters, 49(3):139–143, February 11, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Park:1997:WSA

[PMC97] Kunsoo Park, Sang Lyul Min, and Yookun Cho. The working set algorithm has competitive ratio less than two. Information Processing Letters, 63(4): 183–188, September 10, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Punnen:1994:FSA

[PN94] Abraham P. Punnen and K. P. K.

Nair. A fast and simple algorithm for the bottleneck biconnected spanning subgraph problem. Information Processing Letters, 50(5):283–286, June 10, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pombortsis:1990:SSP

[Pom90] A. Pombortsis. Sharing special purpose resources in a multiprocessor environment. Information Processing Letters, 34(5): 255–260, May 7, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pombortsis:1992:AHB

[Pom92] A. S. Pombortsis. Analysis of hierarchical bus-based multicomputer architectures. Information Processing Letters, 43(5): 265–270, October 5, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Purdy:1990:ATC

[PP90] Carla Neaderhouser Purdy and George B. Purdy. The area-time complexity of the greatest common divisor problem: a lower bound. Information Processing Letters, 34(1):43–46, February 22, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pinotti:1991:PPQ

[PP91] Maria Cristina Pinotti and Geppino Pucci. Parallel priority queues. Information Processing Letters, 40(1):33–40, October 11, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pagli:1994:CNF

[PP94] Linda Pagli and Geppino Pucci. Counting the number of fault patterns in redundant VLSI arrays. Information Processing Letters, 50(6):337–342, June 27, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Phillips:1998:LTA

[PPK98] Jeffrey Mark Phillips, Abraham P. Punnen, and S. N. Kabadi. A linear time algorithm for the bottleneck traveling salesman problem on a Halin graph. Information Processing Letters, 67(2):105–110, July 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pan:1991:PCM

[PR91] V. Pan and J. Reif. The parallel computation of minimum cost paths in graphs by stream contraction. Information Processing Letters, 40(2):79–83, October 25, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pillaipakkamnatt:1993:LTE

[PR93] Krishnan Pillaipakkamnatt and Vijay Raghavan. A linear time equivalence test for read-twice DNF formulas. Information Processing Letters, 48(6):289–295, December 20, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pedreschi:1998:WPP

[PR98] Dino Pedreschi and Salvatore Ruggieri. Weakest preconditions for pure Prolog programs. Information Processing Letters, 67 (3):145–150, August 17, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Prather:1996:SPS

[Pra96] Ronald E. Prather. The subprogram problem for software metric design. Information Processing Letters, 60(3):143–149, November 11, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Preparata:1991:IVM

[Pre91] Franco P. Preparata. Inverting a Vandermonde matrix in minimum parallel time. Information Processing Letters, 38(6): 291–294, June 28, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pretolani:1993:LTA

[Pre93] Daniele Pretolani. A linear time algorithm for unique Horn satisfiability. Information Processing Letters, 48(2):61–66, November 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pritchard:1995:SSQ

[Pri95] Paul Pritchard. A simple subquadratic algorithm for computing the subset partial order. Information Processing Letters, 56 (6):337–341, December 22, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0020019095001654.

Pritchard:1997:OSQ

[Pri97] Paul Pritchard. An old subquadratic algorithm for finding extremal sets. Information Processing Letters, 62(6):329–334, July 16, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Prodinger:1994:ACP

[Pro94] Helmut Prodinger. An asymptotic comment on a paper by A. Analyti and S. Pramanik: “Performance analysis of a main memory multi-directory hashing technique” [Inform. Process. Lett. 45 (1993), no. 4, 191– 197; MR 93k:68026]. Information Processing Letters, 49(6): 327–328, March 22, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [AP93].

Prodinger:1995:MQH

[Pro95] Helmut Prodinger. Multiple Quickselect — Hoare’s Find algorithm for several elements. Information Processing Letters, 56 (3):123–129, November 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Plandowski:1996:PTC

[PRS96] Wojciech Plandowski, Wojciech Rytter, and Tomasz Szymacha. Parallel tree-contraction and Fibonacci numbers. Information Processing Letters, 59(5):267– 271, September 9, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pruhs:1994:ACS

[Pru94] Kirk R. Pruhs. Average-case scalable on-line algorithms for fault replacement. Information Processing Letters, 52(3):131– 136, November 11, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Palsberg:1992:SAV

[PS92] Jens Palsberg and Michael I. Schwartzbach. Safety analysis versus type inference for partial types. Information Processing Letters, 43(4):175–180, September 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Prodinger:1993:NBR

[PS93] Helmut Prodinger and Wojciech Szpankowski. A note on binomial recurrences arising in the analysis of algorithms. Information Processing Letters, 46(6): 309–311, July 26, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pemmaraju:1994:AWC

[PS94] Sriram V. Pemmaraju and Clifford A. Shaffer. Analysis of the worst case space complexity of a PR quadtree. Information Processing Letters, 49(5):263–267, March 11, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Petreschi:1995:ERS

[PS95a] Rossella Petreschi and Andrea Sterbini. Erratum: “Recognizing strict 2-threshold graphs in O(m) time”. Information Processing Letters, 56(1):65, October 13, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [PS95b].

Petreschi:1995:RSR

[PS95b] Rossella Petreschi and Andrea Sterbini. Recognizing strict 2threshold graphs in O(m) time. Information Processing Letters, 54(4):193–198, May 26, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [PS95a].

Posenato:1998:NLB

[PS98] Roberto Posenato and Massimo Santini. A new lower bound on approximability of the ground state problem for tridimensional Ising spin glasses. Information Processing Letters, 68(4):167– 171, November 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pan:1992:PDR

[PSL92] Victor Pan, Akimou Sadikou, and Elliott Landowne. Polynomial division with a remainder by means of evaluation and interpolation. Information Processing Letters, 44(3):149–153, November 30, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Paterson:1993:SPD

[PSSV93] Michael S. Paterson, Heiko Schröder, Ondrej Sýkora, and Imrich Vř’to. A short proof of the dilation of a toroidal mesh in a path. Information Processing Letters, 48(4):197–199, November 29, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Plateau:1992:OTE

[PT92] Brigitte Plateau and Denis Trystam. Optimal total exchange for a 3-D torus of processors. Information Processing Letters, 42 (2):95–102, May 11, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pugh:1990:SOB

[Pug90] William Pugh. Slow optimally balanced search strategies vs. cached fast uniformly balanced search strategies. Information Processing Letters, 34(5):251– 254, May 7, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peled:1997:SIT

[PW97] Doron Peled and Thomas Wilke. Stutter-invariant temporal properties are expressible without the next-time operator. Information Processing Letters, 63(5): 243–246, September 24, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peng:1994:RSP

[PWL94] C. H. Peng, J. S. Wang, and R. C. T. Lee. Recognizing shortest-path trees in linear time. Information Processing Letters, 52(2):77–85, October 28, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Poon:1998:PTS

[PZC98] Chung Keung Poon, Binhai Zhu, and Francis Chin. A polynomial time solution for labeling a rectilinear map. Information Processing Letters, 65(4):201–207, February 27, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Quaglia:1999:EIO

[QB99] Francesco Quaglia and Roberto Baldoni. Exploiting intraobject dependencies in parallel simulation. Information Processing Letters, 70(3):119– 125, May 14, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/19/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/19/22/article.
pdf.

Qian:1990:ASD

[Qia90] Xiaolei Qian. An axiom system for database transactions. Information Processing Letters, 36 (4):183–189, November 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Qiu:1990:NDA

[QM90] Ke Qiu and Henk Meijer. A note on diameter of acyclic directed hypercubes. Information Processing Letters, 36(1):51–52, October 1, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Qiu:1991:DSG

[QMA91] K. Qiu, H. Meijer, and S. Akl. Decomposing a star graph into disjoint cycles. Information Processing Letters, 39(3):125–129, August 16, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Queinnec:1998:FCD

[Que98]  Christian Queinnec.  Fast

and compact dispatching for dynamic object-oriented languages. Information Processing Letters, 64(6):315–321, January 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rabinovich:1998:NEL

[Rab98] A. Rabinovich. Non-elementary lower bound for Propositional Duration Calculus. Information Processing Letters, 66(1):7–11, April 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raczunas:1993:RES

[Rac93] Marek Raczunas. Remarks on the equivalence of c-e structures and Petri nets. Information Processing Letters, 45(4):165–169, March 22, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Radhakrishnan:1992:IBC

[Rad92] Jaikumar Radhakrishnan. Improved bounds for covering complete uniform hypergraphs. Information Processing Letters, 41 (4):203–207, March 18, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ragde:1991:AAP

[Rag91] Prabhakar Ragde. Analysis of an asynchronous PRAM algorithm. Information Processing Letters, 39(5):253–256, September 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Roy:1994:NMS

[RAGR94] T. Sony Roy, G. Athithan, M. S.

Ganagi, and A. Sivasankara Reddy. A new method to solve non-linear equations. Information Processing Letters, 50(2): 75–79, April 22, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rajcani:1992:OPC

[Raj92] Peter Rajcani. Optimal parallel 3-coloring algorithm for rooted trees and its applications. Information Processing Letters, 41 (3):153–156, March 6, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rajsbaum:1994:ULB

[Raj94] Sergio Rajsbaum. Upper and lower bounds for stochastic marked graphs. Information Processing Letters, 49(6):291– 295, March 22, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rajasekaran:1998:OPA

[Raj98] Sanguthevar Rajasekaran. An optimal parallel algorithm for sorting multisets. Information Processing Letters, 67(3):141– 143, August 17, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramesh:1990:CMP

[Ram90] S. Ramesh. On the completeness of modular proof systems. Information Processing Letters, 36 (4):195–201, November 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramachandran:1995:MCL

[Ram95]  J. Ramachandran.  Modulo

classes and logarithmic advice. Information Processing Letters, 55(5):241–245, September 15, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

RamirezAlfonsin:1998:SAM

[Ram98] J. L. Ramírez Alfonsín. A special arrangement with minimal number of triangles. Information Processing Letters, 67(6): 273–276, September 30, 1998.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rao:1995:NSS

[Rao95] Rajesh P. N. Rao. A note on P - selective sets and closeness. Information Processing Letters, 54 (3):179–185, May 12, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ravi:1994:PDA

[Rav94] R. Ravi. A primal-dual approximation algorithm for the Steiner forest problem. Information Processing Letters, 50(4): 185–189, May 25, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raymond:1989:DAM

[Ray89] Kerry Raymond. A distributed algorithm for multiple entries to a critical section. Information Processing Letters, 30(4):189– 193, February 27, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See note [SM92].

Raychaudhuri:1995:TIN

[Ray95] Arundhati Raychaudhuri. The total interval number of a tree and the Hamiltonian completion number of its line graph. Information Processing Letters, 56 (6):299–306, December 22, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raz:1994:SCO

[Raz94] Yoav Raz. Serializability by commitment ordering. Information Processing Letters, 51(5): 257–264, September 12, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Resende:1994:STN

[RE94] R. F. Resende and A. El Abbadi. On the serializability theorem for nested transactions. Information Processing Letters, 50 (4):177–183, May 25, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rehof:1996:SNN

[Reh96] Jakob Rehof. Strong normalization for non-structural subtyping via saturated sets. Information Processing Letters, 58 (4):157–162, May 27, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ressler:1992:RLP

[Res92] Eugene K. Ressler. Random list permutations in place. Information Processing Letters, 43(5): 271–275, October 5, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rescigno:1996:CCP

[Res96] Adele A. Rescigno. On the communication complexity of polling. Information Processing Letters, 59(6):317–323, September 23, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reichwein:1993:MSA

[RF93] Georg Reichwein and José Luiz Fiadeiro. Models for the substitution axiom of UNITY logic. Information Processing Letters, 48(4):171–176, November 29, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ravindran:1993:DED

[RG93] Somasundaram Ravindran and Alan Gibbons. Dense edgedisjoint embedding of complete binary trees in the hypercube. Information Processing Letters, 45(6):321–325, April 16, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rhee:1998:CIG

[Rhe98] Chongkye Rhee. On the chromatic index of graphs with 2m+ 1 vertices and 2m^{2} edges. Information Processing Letters, 66 (3):115–118, May 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rajanarayanan:1991:NOD

[RI91] Subbiah Rajanarayanan and Sitharama S. Iyengar. A new optimal distributed algorithm for the set intersection problem. Information Processing Letters, 38 (3):143–148, May 17, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rifa:1998:NAA

[Rif98] Josep Rifà. A new algebraic algorithm to decode the ternary Golay code. Information Processing Letters, 68(6):271–274, December 30, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rittri:1994:SUT

[Rit94] Mikael Rittri. Semi-unification of two terms in Abelian groups. Information Processing Letters, 52(2):61–68, October 28, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [Rit95].

Rittri:1995:CSU

[Rit95] Mikael Rittri. Corrigendum: “Semi-unification of two terms in Abelian groups” [Inform. Process. Lett. 52(2), 28 October 1994, pp. 61–68]. Information Processing Letters, 53(4): 235, February 24, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Rit94].

Rivin:1993:PFT

[Riv93] Eugene Rivin. Precision fabrication technology in the former Soviet Union and other East European countries. Unknown journal, 46(5):111–131, July ??, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ra:1996:PPA

[RK96] Dong-Yul Ra and Jong-Hyun Kim. A parallel parsing algorithm for arbitrary context-free grammars. Information Processing Letters, 58(2):87–96, May 13, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rao:1993:SPC

[RL93] Nageswara S. V. Rao and Ching Luo. On similarity of polynomial configurations. Information Processing Letters, 47(5): 231–236, October 8, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reddy:1994:ASL

[RL94] H. N. Reddy and E. L. Leiss. An O(log N) algorithm to solve linear recurrences on hypercubes. Information Processing Letters, 49(6):319–325, March 22, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rhee:1996:ACC

[RL96] Chongkye Rhee and Y. Daniel Liang. An NC algorithm for the clique cover problem in cocomparability graphs and its application. Information Processing Letters, 57(5):287–290, March 11, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rhee:1994:EAF

[RLDL94] C. Rhee, Y. Daniel Liang, S. K. Dhall, and S. Lakshmivarahan. Efficient algorithms for finding depth-first and breadthfirst search trees in permutation graphs. Information Processing Letters, 49(1):45–50, January 14, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Robson:1990:RAM

[Rob90]  J. M. Robson.  Random

access machines with multidimensional memories. Information Processing Letters, 34 (5):265–266, May 7, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ronn:1991:LER

[Ron91] Stefan Ronn. On the logarithmic evaluation of recurrence relations. Information Processing Letters, 40(4):197–199, November 25, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rosenblueth:1993:EMN

[Ros93] David A. Rosenblueth. An execution mechanism for nondeterministic, state-oriented programs based on a chart parser. Information Processing Letters, 45(4):211–217, March 22, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rosenbloom:1997:MRC

[Ros97] Arnold Rosenbloom. Monotone real circuits are more powerful than monotone Boolean circuits. Information Processing Letters, 61(3):161–164, March 14, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rosenberg:1998:FNR

[Ros98] Burton Rosenberg. Fast nondeterministic recognition of context-free languages using two queues. Information Processing Letters, 67(2):91–93, July 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rosenbloom:1999:SPM

[Ros99] Arnold Rosenbloom. On the sets of perfect matchings for two bipartite graphs. Information Processing Letters, 70(2): 95–97, April 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/18/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/18/24/article.
pdf.

Rote:1991:CMH

[Rot91] Günter Rote. Computing the minimum Hausdorff distance between two-point sets on a line under translation. Information Processing Letters, 38(3):123– 127, May 17, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Roupin:1997:AMC

[Rou97] Frédéric Roupin. On approximating the memory-Constrained Module Allocation Problem. Information Processing Letters, 61 (4):205–208, March 27, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rao:1991:IAR

[RP91] D. Sreenivasa Rao and John D. Provence. An integrated approach to routing and via minimization. Information Processing Letters, 39(5):257–263, September 13, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Radhakrishnan:1999:PAD

[RP99] D. Radhakrishnan and A. P. Preethy. A parallel approach to direct analog-to-residue conversion. Information Processing Letters, 69(5):249–252, March 12, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/21/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/21/22/article.
pdf.

Ranjan:1998:EAT

[RPG98] D. Ranjan, E. Pontelli, and G. Gupta. Efficient algorithms for the temporal precedence problem. Information Processing Letters, 68(2):71– 78, October 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/46/19/
21/abstract.html;   http:

//www.elsevier.nl/gej-ng/
10/23/20/46/19/21/article.
pdf.

Ramalingam:1990:NSP

[RR90] G. Ramalingam and C. Pandu Rangan. New sequential and parallel algorithms for interval graph recognition. Information Processing Letters, 34(4): 215–219, April 24, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rossmanith:1992:OTP

[RR92] Peter Rossmanith and Wojciech Rytter. Observations on log(n) time parallel recognition of unambiguous CFL’s. Information Processing Letters, 44(5):267– 272, December 21, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ranjan:1993:TAM

[RR93] Desh Ranjan and Daniela Rus.

A tool for the analysis of manipulation. Information Processing Letters, 45(3):117–121, March 8, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramalingam:1994:CLA

[RR94] G. Ramalingam and Thomas Reps. On competitive online algorithms for the dynamic priority-ordering problem. Information Processing Letters, 51 (3):155–161, August 10, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramnath:1999:SSR

[RR99a] Sarnath Ramnath and Venkatesh Raman. Selecting small ranks in EREW PRAM. Information Processing Letters, 71(5–6):183– 186, September 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/30/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/30/27/article.
pdf.

Roussel:1999:LAC

[RR99b] F. Roussel and I. Rusu. A linear algorithm to color itriangulated graphs. Information Processing Letters, 70(2): 57–62, April 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/18/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/18/18/article.
pdf.

Raman:1998:SNC

[RRS98] Venkatesh Raman, B. Ravikumar, and S. Srinivasa Rao. A simplified NP-complete MAXSAT problem. Information Processing Letters, 65(1):1–6, January 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Robins:1999:DSR

[RRS99] Gabriel Robins, Brian L. Robinson, and Bhupinder S. Sethi.

On detecting spatial regularity in noisy images. Information Processing Letters, 69(4):189– 195, February 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/20/
21/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/20/21/article.
pdf.

Rao:1994:USR

[RRW94] Rajesh P. N. Rao, Jörg Rothe, and Osamu Watanabe. Upward separation for FewP and related classes. Information Processing Letters, 52(4):175–180, November 25, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See corrigendum [RRW00].

Rao:2000:CSS

[RRW00] Rajesh P. N. Rao, Jörg Rothe, and Osamu Watanabe. Corrigendum to “Upward separation for FewP and related classes”, Information Processing Letters 52(4) (1994) 175–180. Information Processing Letters, 74 (1–2):89, April 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/25/37/article.

pdf. See [RRW94].

Ronn:1990:DTD

[RS90] Stefan Ronn and Heikki Saikkonen. Distributed termination detection with counters. Information Processing Letters, 34(5): 223–227, May 7, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rytter:1991:CRP

[RS91] W. Rytter and A. Saoudi. On the complexity of the recognition of parallel 2D-image languages. Information Processing Letters, 38(5):225–229, June 14, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rouskov:1993:FDS

[RS93] Yordan Rouskov and Pradip K. Srimani. Fault diameter of star graphs. Information Processing Letters, 48(5):243–251, December 10, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Radhakrishnan:1994:DMC

[RS94a] Jaikumar Radhakrishnan and K. V. Subrahmanyam. Directed monotone contact networks for threshold functions. Information Processing Letters, 50(4): 199–203, May 25, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raspaud:1994:GSS

[RS94b] André Raspaud and Eric Sopena. Good and semi-strong colorings of oriented planar graphs. Information Processing Letters, 51 (4):171–174, August 24, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rubinstein:1995:SMF

[RS95a] Roy S. Rubinstein and John N.

Shutt. Self-modifying finite automata: An introduction. Information Processing Letters, 56 (4):185–190, November 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Russell:1995:RRB

[RS95b] Alexander Russell and Ravi Sundaram. The relativized relationship between probabilistically checkable debate systems, IP and PSPACE. Information Processing Letters, 53(2):61–68, January 27, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramnath:1996:TPS

[RS96] Sarnath Ramnath and Sivaprakasam Sunder. On two-processor scheduling and maximum matching in permutation graphs. Information Processing Letters, 57 (6):321–327, March 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rabi:1997:OAO

[RS97] Muhammad Rabi and Alan T.

Sherman. An observation on associative one-way functions in complexity theory. Information Processing Letters, 64(5):239– 244, December 23, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ryan:1998:ARA

[RS98] P. Y. A. Ryan and S. A. Schneider. An attack on a recursive authentication protocol — A cautionary tale. Information Processing Letters, 65(1):7–10, January 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ra:1999:NOP

[RS99] Dong-Yul Ra and George C. Stockman. A new one pass algorithm for estimating stochastic context-free grammars. Information Processing Letters, 72(1– 2):37–45, October 29, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/27/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/27/30/article.
pdf.

Raynal:1991:COA

[RST91] Michel Raynal, André Schiper, and Sam Toueg. The causal ordering abstraction and a simple way to implement it. Information Processing Letters, 39(6): 343–350, September 27, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raynal:1999:RFD

[RT99] Michel Raynal and Frédéric Tronel. Restricted failure detectors: Definition and reduction protocols. Information Processing Letters, 72(3–4):91– 97, November 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/28/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/28/27/article.
pdf.

Rodrigues:1993:ESC

[RU93] Marion Rodrigues and Hasan Ural. Exact solutions for the construction of optimal length test sequences. Information Processing Letters, 48(6):275–280, December 20, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rubinfeld:1990:CTR

[Rub90] Ronitt Rubinfeld. The cover time of a regular expander is O(n log n). Information Processing Letters, 35(1):49–51, June 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rubinfeld:1992:BCA

[Rub92] Ronitt Rubinfeld. Batch checking with applications to linear functions. Information Processing Letters, 42(2):77–80, May 11, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reingold:1997:KMP

[RUG97] Edward M. Reingold, Kenneth J. Urban, and David Gries. K-M-P string matching revisited. Information Processing Letters, 64(5):217–223, December 23, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rusu:1995:QPP

[Rus95] Irena Rusu. Quasi-parity and perfect graphs. Information Processing Letters, 54(1):35–39, April 14, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramarao:1993:LBD

[RV93] K. V. S. Ramarao and S. Venkatesan. The lower bounds on distributed shortest paths. Information Processing Letters, 48 (3):145–149, November 19, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Romijn:1996:NFA

[RV96] Judi Romijn and Frits Vaandrager. A note on fairness in I/O automata. Information Processing Letters, 59(5):245–250, September 9, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ragde:1991:LSC

[RW91] Prabhakar Ragde and Avi Wigderson. Linear-size constantdepth polylog-threshold circuits. Information Processing Letters, 39(3):143–146, August 16, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rote:1992:CCG

[RW92] Günter Rote and Gerhard Woeginger. Counting convex k-gons in planar point sets. Information Processing Letters, 41(4): 191–194, March 18, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Razborov:1993:LBS

[RW93] Alexander Razborov and Avi Wigderson. n^{Ω(log n)} lower bounds on the size of depth-3 threshold circuits with AND gates at the bottom. Information Processing Letters, 45(6): 303–307, April 16, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Roos:1994:VLP

[RW94] Thomas Roos and Peter Widmayer. k-violation linear programming. Information Processing Letters, 52(2):109–114, October 28, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reingold:1996:LAL

[RW96] Nick Reingold and Jeffery Westbrook. Off-line algorithms for the list update problem. Information Processing Letters, 60 (2):75–80, October 28, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rhee:1997:TBS

[RW97] Injong Rhee and Jennifer L. Welch. Time bounds on synchronization in a periodic distributed system. Information Processing Letters, 64(2):87–93, November 5, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rote:1991:CSC

[RWZW91] Günter Rote, Gerhard Woeginger, Binhai Zhu, and Zhengyan Wang. Counting k-subsets and convex k-gons in the plane. Information Processing Letters, 38 (3):149–151, May 17, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rao:1991:BHP

[RZ91] Nageswara S. V. Rao and Weixiong Zhang. Building heaps in parallel. Information Processing Letters, 37(6):355–358, March 28, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rivin:1992:DPS

[RZ92] Igor Rivin and Ramin Zabih. A dynamic programming solution to the n-queens problem. Information Processing Letters, 41 (5):253–256, April 3, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Saab:1995:IIV

[Saa95] Youssef Saab. Iterative improvement of vertex covers. Information Processing Letters, 55 (2):95–98, July 21, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Saab:1997:IAI

[Saa97]  Youssef G. Saab.  An im-

proved algorithm for intersecting convex polygons. Information Processing Letters, 61(2):89–90, February 28, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sacca:1996:MTS

[Sac96] Domenico Saccà. Multiple total stable models are definitely needed to solve unique solution problems. Information Processing Letters, 58(5):249–254, June 10, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schlipf:1995:FSE

[SAFS95] John S. Schlipf, Fred S. Annexstein, John V. Franco, and R. P. Swaminathan. On finding solutions for extended Horn formulas. Information Processing Letters, 54(3):133–137, May 12, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sakakibara:1991:LQC

[Sak91a] Yasubumi Sakakibara. On learning from queries and counterexamples in the presence of noise. Information Processing Letters, 37(5):279–284, March 14, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sakkinen:1991:SSC

[Sak91b] Markku Sakkinen. Selftype is a special case. Information Processing Letters, 38(4):221–224, May 31, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Salowe:1992:NLB

[Sal92a] J. S. Salowe. A note on lower bounds for rectilinear Steiner trees. Information Processing Letters, 42(3):151–152, May 25, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Saluja:1992:NPV

[Sal92b] Sanjeev Saluja. A note on the permanent value problem. Information Processing Letters, 43 (1):1–5, August 10, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Salomaa:1996:YLT

[Sal96] Kai Salomaa. Yield-languages of two-way pushdown tree automata. Information Processing Letters, 58(4):195–199, May 27, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sanchis:1990:CTC

[San90] Laura A. Sanchis. On the complexity of test case generation for NP-hard problems. Information Processing Letters, 36(3): 135–140, November 1, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Santos:1996:ISB

[San96] Francisco Santos. Inscribing a symmetric body in an ellipse. Information Processing Letters, 59(4):175–178, August 26, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sanders:1998:RPD

[San98]  P. Sanders.  Random per-

mutations on distributed, external and hierarchical memory. Information Processing Letters, 67(6):305–309, September 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sasaki:1991:EDS

[Sas91] Galen Sasaki. The effect of the density of states on the Metropolis algorithm. Information Processing Letters, 37(3):159–163, February 18, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sato:1993:DSP

[Sat93] Tadakazu Sato. Decidability for some problems of linear cellular automata over finite commutative rings. Information Processing Letters, 46(3):151–155, June 11, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sato:1997:ELC

[Sat97] Tadakazu Sato. Ergodicity of linear cellular automata over Z_{m}. Information Processing Letters, 61(3):169–172, March 14, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sato:1998:SLC

[Sat98] Tadakazu Sato. Surjective linear cellular automata over Z_{m}. Information Processing Letters, 66 (2):101–104, April 29, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Saxena:1994:TCL

[Sax94] Sanjeev Saxena. Two-coloring linked lists is NC^{1}-complete for logarithmic space. Information Processing Letters, 49(2):73–76, January 28, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sengupta:1990:DFR

[SB90] A. Sengupta and S. Bandyopadhyay. Deadlock-free routing in kary hypercube network in presence of processor failures. Information Processing Letters, 34 (6):323–328, May 28, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Suraweera:1993:PAM

[SB93] Francis Suraweera and Prabir Bhattacharya. An O(log m) parallel algorithm for the minimum spanning tree problem. Information Processing Letters, 45(3): 159–163, March 8, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schmeiser:1995:PTP

[SB95] James P. Schmeiser and David T.

Barnard. Producing a topdown parse order with bottomup parsing. Information Processing Letters, 54(6):323–326, June 23, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shidlovsky:1996:NDO

[SB96] Boris Shidlovsky and Elisa Bertino. On the number of descendants in an object DAG. Information Processing Letters, 58 (5):211–216, June 10, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sheu:1990:EAC

[SC90] Jang-Ping Sheu and Zen-Fu Chiang. Efficient allocation of chainlike task on chain-like network computers. Information Processing Letters, 36(5):241–245, December 1, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shin:1993:IPG

[SC93] Heung-Chul Shin and Kwang-Moo Choe. An improved LALR(k) parser generation for regular right part grammars. Information Processing Letters, 47(3):123–129, September 14, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szwarcfiter:1994:EKD

[SC94] Jayme L. Szwarcfiter and Guy Chaty. Enumerating the kernels of a directed graph with no odd circuits. Information Processing Letters, 51(3):149–153, August 10, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sarkar:1991:MSB

[SCGD91] U. K. Sarkar, P. P. Chakrabarti, S. Ghose, and S. C. De Sarkar. Multiple stack branch and bound. Information Processing Letters, 37(1):43–48, January 10, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sarkar:1992:EUM

[SCGD92a] U. K. Sarkar, P. P. Chakrabarti, S. Ghose, and S. C. De Sarkar. Effective use of memory in iterative deepening search. Information Processing Letters, 42 (1):47–52, April 27, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sarkar:1992:SGA

[SCGD92b] U. K. Sarkar, P. P. Chakrabarti, S. Ghose, and S. C. De Sarkar. A simple 0.5-bounded greedy algorithm for the 0/1 knapsack problem. Information Processing Letters, 42(3):173–177, May 25, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See note [MT92b].

Schweikard:1991:TPS

[Sch91] Achim Schweikard. Trigonometric polynomials with simple roots. Information Processing Letters, 39(5):231–236, September 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schirra:1992:ADA

[Sch92a] Stefan Schirra. Approximate decision algorithms for approximate congruence. Information Processing Letters, 43(1):29–34, August 10, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schwabe:1992:EMT

[Sch92b] Eric J. Schwabe. Embedding meshes of trees into de Bruijn graphs. Information Processing Letters, 43(5):237–240, October 5, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schoenmakers:1993:SAS

[Sch93a] Berry Schoenmakers. A systematic analysis of splaying. Information Processing Letters, 45 (1):41–50, January 25, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schoning:1993:RRS

[Sch93b] Uwe Schoning. On random reductions from sparse sets to tally sets. Information Processing Letters, 46(5):239–241, July 9, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schwabe:1993:CSS

[Sch93c] Eric J. Schwabe. Constantslowdown simulations of normal hypercube algorithms on the butterfly network. Information Processing Letters, 45(6): 295–301, April 16, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schuler:1995:SPS

[Sch95a] Rainer Schuler. Some properties of sets tractable under every polynomial-time computable distribution. Information Processing Letters, 55(4):179–184, August 25, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Spiliopoulou:1995:QPM

[SCH95b] Myra Spiliopoulou, Yannis Cotronis, and Michael Hatzopoulos. Query processing for multimedia applications on optical media. Information Processing Letters, 53(6):301–306, March 24, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schoenmakers:1997:TLB

[Sch97] Berry Schoenmakers. A tight lower bound for top-down skew heaps. Information Processing Letters, 61(5):279–284, April 10, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sung:1999:GAC

[SCP99] Soo Hak Sung, Seongtaek Chee, and Choonsik Park.

Global avalanche characteristics and propagation criterion of balanced Boolean functions. Information Processing Letters, 69(1):21–24, January 15, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/17/
20/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/17/20/article.
pdf.

Subramaniyam:1991:CSE

[SD91] R. V. Subramaniyam and A. A.

Diwan. A counterexample for the sufficiency of edge guards in star polygons. Information Processing Letters, 40(2):97–99, October 25, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schittenkopf:1997:FAM

[SDB97] Christian Schittenkopf, Gustavo Deco, and Wilfried Brauer. Finite automata-models for the investigation of dynamical systems. Information Processing Letters, 63(3):137–141, August 27, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sen:1992:GPP

[SDG92] Arunabha Sen, Haiyong Deng, and Sumanta Guha. On a graph partition problem with application to VLSI layout. Information Processing Letters, 43(2):87–94, August 24, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Seidl:1994:HOD

[Sei94] Helmut Seidl. Haskell overloading is DEXPTIME-complete. Information Processing Letters, 52(2):57–60, October 28, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Seidl:1996:FSN

[Sei96] Helmut Seidl. Fast and simple nested fixpoints. Information Processing Letters, 59(6): 303–308, September 23, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sen:1990:FAM

[Sen90] Sandeep Sen. Finding an approximate median with high probability in constant parallel time. Information Processing Letters, 34(2):77–80, March 16, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sen:1991:SOS

[Sen91] Sandeep Sen. Some observations on skip-lists. Information Processing Letters, 39(4):173–176, August 30, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sengupta:1997:CEP

[Sen97] Rimli Sengupta. Cancellation is exponentially powerful for computing the determinant. Information Processing Letters, 62 (4):177–181, June 11, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sengupta:1998:REH

[Sen98] Abhijit Sengupta. On ring embedding in hypercubes with faulty nodes and links. Information Processing Letters, 68 (4):207–214, November 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Serna:1991:ALP

[Ser91] Maria Serna. Approximating linear programming is log-space complete for P. Information Processing Letters, 37(4):233– 236, February 28, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sere:1996:PAR

[Ser96] K. Sere. Procedures and atomicity refinement. Information Processing Letters, 60(2):67–74, October 28, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sereno:1998:BSE

[Ser98] Matteo Sereno. Binary search with errors and variable cost queries. Information Processing Letters, 68(5):261–270, December 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stotts:1992:PTN

[SG92] P. David Stotts and Parke Godfrey. Place/transition nets with debit arcs. Information Processing Letters, 41(1):25–33, January 21, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sgall:1997:LBR

[Sga97] Jiří Sgall. A lower bound for randomized on-line multiprocessor scheduling. Information Processing Letters, 63(1):51–55, July 30, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sen:1992:RPF

[SGB92] Arunabha Sen, Abhijit Sen Gupta, and Subir Bandyopadhyay. On the routing problem in faulty supercubes. Information Processing Letters, 42(1):39–46, April 27, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Swaminathan:1995:PCB

[SGB95] R. P. Swaminathan, D. Giriraj, and D. K. Bhatia. The pagenumber of the class of bandwidthk graphs is k − 1. Information Processing Letters, 55(2):71–74, July 21, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Salhi:1998:PIG

[SGD98] A. Salhi, H. Glaser, and D. De Roure. Parallel implementation of a genetic-programming based tool for symbolic regression. Information Processing Letters, 66 (6):299–307, June 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Simpson:1999:FCA

[SGSD99] L. Simpson, J. Golić, M. Salmasizadeh, and E. Dawson. A fast correlation attack on multiplexer generators. Information Processing Letters, 70(2): 89–93, April 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/18/
23/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/18/23/article.
pdf.

Shen:1993:NMV

[SH93] Xiaojun Shen and Qing Hu.

A note on: “Minimal visibility graphs” [Inform. Process. Lett. 37 (1991), no. 1, 49–53; MR 91m:68160] by D. Campbell and J. C. Higgins. Information Processing Letters, 46(2):101, May 17, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [CH91].

Seiden:1994:FSO

[SH94] Steven S. Seiden and Daniel S.

Hirschberg. Finding succinct ordered minimal perfect hash functions. Information Processing Letters, 51(6):283–288, September 26, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shah:1993:PAC

[Sha93] Nimish R. Shah. A parallel algorithm for constructing projection polyhedra. Information Processing Letters, 48(3):113– 119, November 19, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sharir:1996:EAS

[Sha96] Micha Sharir. Excess in arrangements of segments. Information Processing Letters, 58(5): 245–247, June 10, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shen:1993:SGS

[SHC^{+}93] Xiaojun Shen, Qing Hu, Bin Cong, Hal Sudborough, Mike Girou, and Said Bettayeb. The 4-star graph is not a subgraph of any hypercube. Information Processing Letters, 45(4): 199–203, March 22, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shermer:1991:CAD

[She91a] Thomas Shermer. A counterexample to the algorithms for determining opaque minimal forests. Information Processing Letters, 40(1):41–42, October 11, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sheu:1991:FTP

[She91b] Jang-Ping Sheu. Fault-tolerant parallel k selection algorithm in n-cube networks. Information Processing Letters, 39(2):93–97, July 31, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shermer:1992:LAB

[She92] Thomas C. Shermer. A linear algorithm for bisecting a polygon. Information Processing Letters, 41(3):135–140, March 6, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shen:1993:SBA

[She93] Zhizhang Shen. Static behavior analysis of a mesh system. Information Processing Letters, 47 (6):295–299, October 18, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shen:1994:RAP

[SHL94] Xiao Jun Shen, Qing Hu, and Weifa Liang. Realization of an arbitrary permutation on a hypercube. Information Processing Letters, 51(5):237–243, September 12, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shoup:1990:DCF

[Sho90] Victor Shoup. On the deterministic complexity of factoring polynomials over finite fields. Information Processing Letters, 33 (5):261–267, January 10, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shonkwiler:1991:CHS

[Sho91a] R. Shonkwiler. Computing the Hausdorff set distance in linear time for any L_{p} point distance. Information Processing Letters, 38(4):201–207, May 31, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shoup:1991:SFP

[Sho91b] Victor Shoup. Smoothness and factoring polynomials over finite fields. Information Processing Letters, 38(1):39–42, April 12, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shoudai:1992:CLD

[Sho92]  Takayoshi Shoudai.  A P -

complete language describable with iterated shuffle. Information Processing Letters, 41(5): 233–238, April 3, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shimozono:1998:HAM

[SHS98] Shinichi Shimozono, Kouichi Hirata, and Ayumi Shinohara. On the hardness of approximating the minimum consistent acyclic DFA and decision diagram. Information Processing Letters, 66 (4):165–170, May 29, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shvartsman:1991:AOC

[Shv91] Alex A. Shvartsman. Achieving optimal CRCW PRAM faulttolerance. Information Processing Letters, 39(2):59–66, July 31, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shvartsman:1992:EWA

[Shv92] Alex A. Shvartsman. An efficient Write-All algorithm for fail-stop PRAM without initialized memory. Information Processing Letters, 44(4):223–231, December 10, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sugihara:1995:WDT

[SI95] Kokichi Sugihara and Hiroshi Inagaki. Why is the 3D Delaunay triangulation difficult to construct? Information Processing Letters, 54(5):275–280, June 9, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sugimoto:1999:GLD

[SI99] Noriko Sugimoto and Hiroki Ishizaka. Generating languages by a derivation procedure for elementary formal systems. Information Processing Letters, 69(4):161–166, February 26, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/20/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/20/17/article.
pdf.

Sideris:1990:AGA

[Sid90] Martha Sideris. On attribute grammars without attribute synthesis. Information Processing Letters, 35(2):103–109, June 29, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Simonson:1990:RCP

[Sim90] Shai Simonson. Routing with critical paths. Information Processing Letters, 34(1):13–19, February 22, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Simon:1991:VCD

[Sim91]  Hans Ulrich Simon.  The

Vapnik–Chervonenkis dimension of decision trees with bounded rank. Information Processing Letters, 39(3):137–141, August 16, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Simon:1995:NLB

[Sim95] Klaus Simon. A note on lexicographic breadth first search for chordal graphs. Information Processing Letters, 54(5): 249–251, June 9, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Singh:1992:TUU

[Sin92a] Ambuj K. Singh. Towards an understanding of unbounded variables in asynchronous systems. Information Processing Letters, 42(1):7–17, April 27, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Singh:1992:ECL

[Sin92b] S. Singh. Expected connectivity and leader election in unreliable networks. Information Processing Letters, 42(5):283–285, July 3, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sinachopoulos:1994:LDL

[Sin94a] Antonia Sinachopoulos. Logics and decidability for labelled preand partially-ordered Kripke structures. Information Processing Letters, 51(1):43–52, July 12, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Singh:1994:RTL

[Sin94b] Gurdip Singh. Real-time leader election. Information Processing Letters, 49(2):57–61, January 28, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sinha:1998:SPS

[Sin98] Rakesh K. Sinha. Simulation of PRAMs with scan primitives by unbounded fan-in circuits. Information Processing Letters, 68 (6):275–282, December 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sistla:1991:PCR

[Sis91] A. P. Sistla. Proving correctness with respect to nondeterministic safety specifications. Information Processing Letters, 39 (1):45–49, July 12, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Snelick:1995:USP

[SJKL95] Robert Snelick, Joseph JáJá, Raghu Kacker, and Gordon Lyon. Using synthetic perturbations and statistical screening to assay shared-memory programs. Information Processing Letters, 54(3):147–153, May 12, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Srikant:1991:FPA

[SK91] R. Srikant and Kamala Krithivasan. Fastest path across constrained moving rectilinear obstacles. Information Processing Letters, 37 (6):349–353, March 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Singhal:1992:EIV

[SK92a] Mukesh Singhal and Ajay Kshemkalyani. An efficient implementation of vector clocks. Information Processing Letters, 43(1):47–52, August 10, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sprague:1992:OPA

[SK92b] Alan P. Sprague and K. H.

Kulkarni. Optimal parallel algorithms for finding cut vertices and bridges of interval graphs. Information Processing Letters, 42(4):229–234, June 19, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schwarz:1998:BCF

[SK98a] S. Schwarz and S. O. Krumke.

On budget-constrained flow improvement. Information Processing Letters, 66(6):291–297, June 30, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Segal:1998:EPS

[SK98b] Michael Segal and Klara Kedem. Enclosing k points in the smallest axis parallel rectangle. Information Processing Letters, 65(2):95–99, January 29, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Skillicorn:1996:PTD

[Ski96] D. B. Skillicorn. A parallel tree difference algorithm. Information Processing Letters, 60(5): 231–235, December 8, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shin:1998:ADB

[SKKC98] Chan-Su Shin, Sung Kwon Kim, Sung-Ho Kim, and Kyung-Yong Chwa. Algorithms for drawing binary trees in the plane. Information Processing Letters, 66 (3):133–139, May 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Skyum:1991:SAC

[Sky91] Sven Skyum. A simple algorithm for computing the smallest enclosing circle. Information Processing Letters, 37(3):121–125, February 18, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sempere:1995:MPN

[SL95] Jose M. Sempere and Damián López. A McCulloch-Pitts neural net to characterize even linear languages. Information Processing Letters, 56(4):201–208, November 24, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Seo:1996:AGA

[SL96] Sang Koo Seo and Yoon Joon Lee. Applicability of genetic algorithms to optimal evaluation of path predicates in objectoriented queries. Information Processing Letters, 58(3):123– 128, May 13, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sedjelmaci:1997:IAI

[SL97] Mohamed S. Sedjelmaci and Christian Lavault. Improvements on the accelerated integer GCD algorithm. Information Processing Letters, 61(1):31–36, January 14, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Slavik:1997:IPG

[Sla97] Petr Slavík. Improved performance of the greedy algorithm for partial cover. Information Processing Letters, 64(5):251– 254, December 23, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sheu:1993:BAS

[SLC93] Jang Ping Sheu, Wen Hwa Liaw, and Tzung Shi Chen. A broadcasting algorithm in star graph interconnection networks. Information Processing Letters, 48 (5):237–241, December 10, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sung:1998:FTT

[SLCH98] Ting-Yi Sung, Chun-Yuan Lin, Yen-Chu Chuang, and Lih-Hsing Hsu. Fault tolerant token ring embedding in double loop networks. Information Processing Letters, 66(4):201–207, May 29, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Son:1998:GAC

[SLCS98] Jung Je Son, Jong In Lim, Seongtaek Chee, and Soo Hak Sung. Global avalanche characteristics and nonlinearity of balanced Boolean functions. Information Processing Letters, 65 (3):139–144, February 13, 1998.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Slobodova:1990:OGD

[Slo90] Anna Slobodova. One-way globally deterministic synchronized alternating finite automata recognize exactly deterministic context-sensitive languages. Information Processing Letters, 36 (2):69–72, October 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sloan:1995:FTN

[Slo95] Robert H. Sloan. Four types of noise in data for PAC learning. Information Processing Letters, 54(3):157–162, May 12, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Siromoney:1990:PKC

[SM90] Rani Siromoney and Lisa Mathew. A public key cryptosystem based on Lyndon words. Information Processing Letters, 35(1):33–36, June 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Satyanarayanan:1992:NRT

[SM92] R. Satyanarayanan and D. R. Muthukrishnan. A note on Raymond’s tree based algorithm for distributed mutual exclusion. Information Processing Letters, 43(5):249–255, October 5, 1992.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Ray89].

Simha:1994:LLU

[SM94] Rahul Simha and Amitava Majumdar. On lookahead in the list update problem. Information Processing Letters, 50(2): 105–110, April 22, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Singhal:1995:OPA

[SM95] Mukesh Singhal and Friedemann Mattern. An optimality proof for asynchronous recovery algorithms in distributed systems. Information Processing Letters, 55(3):117–121, August 11, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Siromoney:1994:ILW

[SMDS94] Rani Siromoney, Lisa Mathew, V. R. Dare, and K. G. Subramanian. Infinite Lyndon words. Information Processing Letters, 50 (2):101–104, April 22, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Srinivasan:1995:EDB

[SMN^{+}95] Anand Srinivasan, K. Madhukar, P. Nagavamsi, C. Pandu Rangan, and Maw-Shang Chang. Edge domination on bipartite permutation graphs and cotriangulated graphs. Information Processing Letters, 56(3):165– 171, November 10, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Snyder:1991:CLB

[Sny91a] Timothy Law Snyder. Corrigendum: “Lower bounds for rectilinear Steiner trees in bounded space” [Inform. Process. Lett. 37(2), 31 January 1991, pp. 71– 74]. Information Processing Letters, 38(5):279, June 14, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Sny91b].

Snyder:1991:LBR

[Sny91b] Timothy Law Snyder. Lower bounds for rectilinear Steiner trees in bounded space. Information Processing Letters, 37(2): 71–74, January 31, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [Sny91a].

Snyder:1993:CRP

[Sny93] Wayne Snyder. On the complexity of recursive path orderings. Information Processing Letters, 46(5):257–262, July 9, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Santos:1993:RMS

[SO93] J. Santos and J. Orozco. Rate monotonic scheduling in hard real-time systems. Information Processing Letters, 48(1):39–45, October 29, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sassa:1997:SDM

[SO97] Masataka Sassa and Takuya Ookubo. Systematic debugging method for attribute grammar description. Information Processing Letters, 62(6):305–313, July 16, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Spinrad:1991:FLH

[Spi91] Jeremy P. Spinrad. Finding large holes. Information Processing Letters, 39(4):227–229, August 30, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Spinrad:1993:DLO

[Spi93] Jeremy P. Spinrad. Doubly lexical ordering of dense 0-1 matrices. Information Processing Letters, 45(5):229–235, April 2, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shen:1991:SH

[SR91] Xiaojun Shen and Edward M.

Reingold. Scheduling on a hypercube. Information Processing Letters, 40(6):323–328, December 30, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Srimani:1992:ADA

[SR92] Pradip K. Srimani and Rachamallu L. N. Reddy. Another distributed algorithm for multiple entries to a critical section. Information Processing Letters, 41(1):51–57, January 21, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Saxena:1995:PAC

[SR95] Sanjeev Saxena and N. Malahal Rao. Parallel algorithms for connectivity problems on interval graphs. Information Processing Letters, 56(1):37–44, October 13, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Seshadri:1996:THD

[SR96a] S. Seshadri and D. Rotem. The two headed disk: Stochastic dominance of the greedy policy. Information Processing Letters, 57(5):273–277, March 11, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shen:1996:OPS

[SR96b] Hong Shen and Sarnath Ramnath. Optimal parallel selection in sorted matrices. Information Processing Letters, 59(3): 117–122, August 12, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sterbini:1998:TAR

[SR98] Andrea Sterbini and Thomas Raschle. An O(n^{3}) time algorithm for recognizing threshold dimension 2 graphs. Information Processing Letters, 67(5): 255–259, September 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schobbens:1999:LSS

[SR99] P.-Y. Schobbens and J.-F.

Raskin. The logic of “initially” and “next”: Complete axiomatization and complexity. Information Processing Letters, 69(5): 221–225, March 12, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/21/
23/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/21/23/article.
pdf.

Sridharan:1999:CTP

[Sri99] K. Sridharan. Computing two penetration measures for curved 2D objects. Information Processing Letters, 72(3–4):143– 148, November 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/28/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/28/36/article.
pdf.

Shi:1995:SBT

[SRW95] Hongchi Shi, Gerhard X. Ritter, and Joseph N. Wilson. Simulations between two reconfigurable mesh models. Information Processing Letters, 55(3): 137–142, August 11, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Saluja:1992:PEC

[SS92a] Sanjeev Saluja and K. V. Subrahmanyam. On the power of enumerative counting. Information Processing Letters, 43(3): 119–125, September 14, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Simonson:1992:CTE

[SS92b] Shai Simonson and I. Hal Sudborough. On the complexity of tree embedding problems. Information Processing Letters, 44 (6):323–328, December 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szymanski:1992:CCV

[SS92c] Boleslaw K. Szymanski and Balaram Sinharoy. Complexity of the closest vector problem in a lattice generated by (0, 1)- matrix. Information Processing Letters, 42(3):121–126, May 25, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigenda [SS92d].

Szymanski:1992:CCC

[SS92d] Boleslaw K. Szymanski and Balaram Sinharoy. Corrigenda: “Complexity of the closest vector problem in a lattice generated by (0, 1)-matrix” [Inform. Process. Lett. 42 (1992), no. 3, 121–126; MR 94c:68096]. Information Processing Letters, 43 (3):167, September 14, 1992.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [SS92c].

Sajith:1994:OPA

[SS94a] G. Sajith and Sanjeev Saxena.

Optimal parallel algorithms for coloring bounded degree graphs and finding maximal independent sets in rooted trees. Information Processing Letters, 49 (6):303–308, March 22, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [SS95a].

Sun:1994:DTS

[SS94b] Hung Min Sun and Shiuh Pyng Shieh. On dynamic threshold schemes. Information Processing Letters, 52(4):201–206, November 25, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sajith:1995:COP

[SS95a] G. Sajith and Sanjeev Saxena.

Corrigendum: “Optimal parallel algorithms for coloring bounded degree graphs and finding maximal independent sets in rooted trees” [Inform. Process. Lett. 49 (22 March 1994), no. 6, 303– 308; MR 95a:68058]. Information Processing Letters, 54 (5):305, June 9, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [SS94a].

Steiger:1995:PAS

[SS95b] William Steiger and Ileana Streinu. A pseudo-algorithmic separation of lines from pseudolines. Information Processing Letters, 53(5):295–299, March 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sanchez:1998:ICP

[SS98a] Pedro P. Sanchez and Refik Soyer. Information concepts and pairwise comparison matrices. Information Processing Letters, 68(4):185–188, November 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shmueli:1998:DSQ

[SS98b] Oded Shmueli and Kurt Shoens. Data sufficiency for queries on cache. Information Processing Letters, 65(4):209–216, February 27, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shin:1998:WDC

[SSC98] Chan-Su Shin, Sung Yong Shin, and Kyung-Yong Chwa. The widest k-dense corridor problems. Information Processing Letters, 68(1):25–31, October 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sridharan:1994:DMI

[SSCK94] K. Sridharan, H. E. Stephanou, K. C. Craig, and S. S. Keerthi.

Distance measures on intersecting objects and their applications. Information Processing Letters, 51(4):181–188, August 24, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shibata:1994:CCW

[SSO94] Yukio Shibata, Miyuki Shirahata, and Shingo Osawa. Counting closed walks in generalized de Bruijn graphs. Information Processing Letters, 49(3):135– 138, February 11, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shokrollahi:1998:RMR

[SSS98] M. A. Shokrollahi, D. A. Spielman, and V. Stemann. A remark on matrix rigidity. Information Processing Letters, 64(6): 283–285, January 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sheu:1990:EPK

[ST90] Jang-Ping Sheu and Jyh-Shyan Tang. Efficient parallel k selection algorithm. Information Processing Letters, 35(6):313– 316, September 15, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Siklossy:1991:SRM

[ST91] Laurent Siklossy and Eduard Tulp. The space reduction method: a method to reduce the size of search spaces. Information Processing Letters, 38 (4):187–192, May 31, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stamoulis:1993:EAM

[ST93] George D. Stamoulis and John N. Tsitsiklis. An efficient algorithm for multiple simultaneous broadcasts in the hypercube. Information Processing Letters, 46(5):219–224, July 9, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shachnai:1999:MMT

[ST99] Hadas Shachnai and John J. Turek. Multiresource malleable task scheduling to minimize response time. Information Processing Letters, 70(1): 47–52, April 16, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/17/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/17/24/article.
pdf.

Staiger:1996:CSW

[Sta96] Ludwig Staiger. Codes, simplifying words, and open set condition. Information Processing Letters, 58(6):297–301, June 24, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stewart:1991:CPS

[Ste91] Iain A. Stewart. Complete problems for symmetric logspace involving free groups. Information Processing Letters, 40(5): 263–267, December 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Steinbach:1994:GPO

[Ste94] Joachim Steinbach. Generating polynomial orderings. Information Processing Letters, 49(2): 85–93, January 28, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shikishima-Tsuji:1997:TCO

[STKK97] Kayoko Shikishima-Tsuji, Masashi Katsura, and Yuji Kobayashi. On termination of confluent onerule string-rewriting systems. Information Processing Letters, 61(2):91–96, February 28, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Suzuki:1990:LAB

[STN90] Hitoshi Suzuki, Naomi Takahashi, and Takao Nishizeki. A linear algorithm for bipartition of biconnected graphs. Information Processing Letters, 33(5): 227–231, January 10, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stohr:1991:BBN

[Stö91a] Elena Stöhr. Broadcasting in the Butterfly network. Information Processing Letters, 39(1):41–43, July 12, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stojmenovic:1991:BHS

[Sto91b] Ivan Stojmenović. Bisections and ham-sandwich cuts of convex polygons and polyhedra. Information Processing Letters, 38 (1):15–21, April 12, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Subramanian:1995:EPB

[Sub95a] Ashok Subramanian. Erratum: “A polynomial bound on the number of light cycles in an undirected graph” [Inform. Process. Lett. 53 (1995), no. 4, 173–176; MR 95k:05096]. Information Processing Letters, 54 (4):247, May 26, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Sub95b].

Subramanian:1995:PBN

[Sub95b] Ashok Subramanian. A polynomial bound on the number of light cycles in an undirected graph. Information Processing Letters, 53(4):173–176, February 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [Sub95a].

Sugihara:1994:SPR

[Sug94] Kōkichi Sugihara. Simpler proof of a realizability theorem on Delaunay triangulations. Information Processing Letters, 50(4): 173–176, May 25, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sun:1996:TRH

[Sun96] Yong Sun. Term rewriting and Hoare logic — Coded rewriting. Information Processing Letters, 60(5):237–242, December 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sengupta:1994:FTF

[SV94] Abhijit Sengupta and Suresh Viswanathan. On fault-tolerant fixed routing in hypercubes. Information Processing Letters, 51 (2):93–99, July 26, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schwarzkopf:1996:RSL

[SV96] Otfried Schwarzkopf and Jules Vleugels. Range searching in low-density environments. Information Processing Letters, 60 (3):121–127, November 11, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Saouter:1999:CAN

[SV99] Yannick Saouter and Hervé Le Verge. Computability of affine non-conditional recurrent systems. Information Processing Letters, 69(6):291– 295, March 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/22/
21/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/22/21/article.
pdf.

Strijk:1999:LRM

[SvK99] Tycho Strijk and Marc van Kreveld. Labeling a rectilinear map more efficiently. Information Processing Letters, 69(1): 25–30, January 15, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/47/17/
                                        21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/47/17/21/article.
                                        pdf.

Shmoys:1990:AHK

[SW90] David B. Shmoys and David P. Williamson. Analyzing the Held-Karp TSP bound. A monotonicity property with application. Information Processing Letters, 35(6):281–285, September 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stein:1992:AMC

[SW92] Clifford Stein and Joel Wein. Approximating the minimumcost maximum flow is P - complete. Information Processing Letters, 42(6):315–319, July 24, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sieling:1993:ROL

[SW93a] Detlef Sieling and Ingo Wegener. Reduction of OBDDs in linear time. Information Processing Letters, 48(3):139–144, November 19, 1993. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Steel:1993:KTT

[SW93b] Mike Steel and Tandy Warnow.

Kaikoura tree theorems: Computing the maximum agreement subtree. Information Processing Letters, 48(2):77–82, November 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Simon:1998:LDA

[SW98] J. Simon and J.-M. Wierum.

The Latency-of-Data-Access model for analyzing parallel computation. Information Processing Letters, 66(5):255–261, June 16, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stinson:1999:ARS

[SW99] D. R. Stinson and R. Wei. An application of ramp schemes to broadcast encryption. Information Processing Letters, 69 (3):131–135, February 12, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/23/
20/47/19/23/abstract.html;
http://www.elsevier.nl/gej-
ng/10/23/20/47/19/23/article.
pdf.

Salomaa:1990:IPL

[SY90] Kai Salomaa and Sheng Yu. The immortality problem for Lag systems. Information Processing Letters, 36(6):311–315, December 15, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Steiner:1993:NSU

[SY93] George Steiner and Scott Yeomans. A note on: “Scheduling unit-time tasks with integer release times and deadlines” [Inform. Process. Lett. 16 (1983), no. 4, 171–173; MR 84m:68030] by G. Frederickson. Information Processing Letters, 47(3): 165–166, September 14, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Fre83].

Shiau:1996:FMF

[SY96] Shyue-Horng Shiau and Chang-Biau Yang. A fast maximum finding algorithm on broadcast communication. Information Processing Letters, 60(2):81–89, October 28, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Symvonis:1996:RT

[Sym96] Antonios Symvonis. Routing on trees. Information Processing Letters, 57(4):215–223, February 26, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Symvonis:1999:NDW

[Sym99] Antonios Symvonis. A note on deflection worm routing on meshes. Information Processing Letters, 71(5–6):235–239, September 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/30/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/30/35/article.
pdf.

Suzuki:1998:BTW

[SYUK98] Ichiro Suzuki, Masafumi Yamashita, Hideki Umemoto, and Tsunehiko Kameda. Bushiness and a tight worst-case upper bound on the search number of a simple polygon. Information Processing Letters, 66(1):49–52, April 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Scholefield:1992:WPS

[SZ92a] D. Scholefield and H. S. M.

Zedan. Weakest precondition semantics for time and concurrency. Information Processing Letters, 43(6):301–308, October 19, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sharary:1992:SOP

[SZ92b] Ahmad Sharary and Nejib Zaguia. On a setup optimization problem for interval orders. Information Processing Letters, 44 (1):51–55, November 9, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szalas:1992:AFL

[Sza92] Andrzej Szałas. Axiomatizing fixpoint logics. Information Processing Letters, 41(4):175–180, March 18, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szepietowski:1996:EDP

[Sze96] Andrzej Szepietowski. The element distinctness problem on one-tape Turing machines. Information Processing Letters, 59 (4):203–206, August 26, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szepietowski:1998:WSO

[Sze98] Andrzej Szepietowski. Weak and strong one-way space complexity classes. Information Processing Letters, 68(6):299–302, December 30, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Seberry:1994:ISA

[SZZ94] Jennifer Seberry, Xian Mo Zhang, and Yuliang Zheng. Improving the strict avalanche characteristics of cryptographic functions. Information Processing Letters, 50(1):37–41, April 8, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Takaoka:1992:NUB

[Tak92] Tadao Takaoka. A new upper bound on the complexity of the all pairs shortest path problem. Information Processing Letters, 43(4):195–199, September 28, 1992. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Tamassia:1991:IRM

[Tam91] Roberto Tamassia. An incremental reconstruction method for dynamic planar point location. Information Processing Letters, 37(2):79–83, January 31, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tanaka:1991:DAS

[Tan91] Yoshihiro Tanaka. A dual algorithm for the satisfiability problem. Information Processing Letters, 37(2):85–89, January 31, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Taubenfeld:1991:NRC

[Tau91a] Gadi Taubenfeld. On the nonexistence of resilient consensus protocols. Information Processing Letters, 37(5):285–289, March 14, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Taubner:1991:NNR

[Tau91b] Dirk Taubner. A note on the notation of recursion in process algebras. Information Processing Letters, 37(6):299–303, March 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Thilikos:1997:FPA

[TB97] Dimitrios M. Thilikos and Hans L. Bodlaender. Fast partitioning l-apex graphs with applications to approximating maximum induced-subgraph problems. Information Processing Letters, 61(5):227–232, April 10, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Terekhov:1999:TED

[TC99] I. Terekhov and T. Camp. Time efficient deadlock resolution algorithms. Information Processing Letters, 69(3):149– 154, February 12, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/19/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/19/26/article.
pdf.

Tucci:1991:RNP

[TCC91] Maurizio Tucci, Gennaro Costagliola, and Shi-Kuo Chang. A remark on NP-completeness of picture matching. Information Processing Letters, 39(5):241– 243, September 13, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tscha:1993:RNC

[TCL93] Yeonghwan Tscha, Yanghee Choi, and Kyoon Ha Lee. Rearrangeable nonblocking condition for multi-log_{2} N multiconnection networks. Information Processing Letters, 48(3):107– 112, November 19, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Teia:1993:LBR

[Tei93] Boris Teia. A lower bound for randomized list update algorithms. Information Processing Letters, 47(1):5–9, August 9, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tel:1994:MMS

[Tel94] Gerard Tel. Maximal matching stabilizes in quadratic time. Information Processing Letters, 49(6):271–272, March 22, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Teng:1990:SEP

[Ten90] Shang-Hua Teng. Space efficient processor identity protocol. Information Processing Letters, 34 (3):147–154, April 9, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tzeng:1999:IPI

[TH99] Wen-Guey Tzeng and Chi-Ming Hu. Inter-protocol interleaving attacks on some authentication and key distribution protocols. Information Processing Letters, 69(6):297– 302, March 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/22/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        29/28/24/47/22/22/article.
                                        pdf.

ETAC:1995:CGA

[The95] The Eindhoven Tuesday Afternoon Club. Constructing the Galois adjoint. Information Processing Letters, 53(3):137–139, February 10, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsai:1999:CDT

[THHC99] Chang-Hsiung Tsai, Chun-Nan Hung, Lih-Hsing Hsu, and Chung-Haw Chang. The correct diameter of trivalent Cayley graphs. Information Processing Letters, 72(3–4):109– 111, November 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/60/28/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/60/28/30/article.
                                        pdf.

Tan:1991:ISP

[THI91] Xue-Hou Tan, Tomio Hirata, and Yasuyoshi Inagaki. The intersection searching problem for c-oriented polygons. Information Processing Letters, 37 (4):201–204, February 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Thirunarayan:1993:LIN

[Thi93] Krishnaprasad Thirunarayan. Locality in inheritance networks.

Information Processing Letters, 46(6):263–268, July 26, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Thomas:1995:GCS

[Tho95] Stephen Thomas. Garbage collection in shared-environment closure reducers: Space-efficient depth first copying using a tailored approach. Information Processing Letters, 56(1):1–7, October 13, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Thomborson:1998:ELM

[Tho98] Clark D. Thomborson. The economics of large-memory computations. Information Processing Letters, 66(5):263–268, June 16, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Taylor:1997:AHN

[THS97] Stephen Taylor, Nabil Hachem, and Stanley Selkow. The average height of a node in the BANG abstract directory tree. Information Processing Letters, 61(1): 55–61, January 14, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsuchiya:1999:CBQ

[TIK99] Tatsuhiro Tsuchiya, Nobuhiko Ido, and Tohru Kikuno. Constructing Byzantine Quorum systems from combinatorial designs. Information Processing Letters, 71(1):35–42, July 16, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/17/
23/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/17/23/article.
pdf.

Tipper:1990:SIA

[Tip90] John C. Tipper. A straightforward iterative algorithm for the planar Voronoi diagram. Information Processing Letters, 34 (3):155–160, April 9, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tseng:1999:ATS

[TJ99] Yuh-Min Tseng and Jinn-Ke Jan. Attacks on threshold signature schemes with traceable signers. Information Processing Letters, 71(1): 1–4, July 16, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/17/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/17/17/article.
pdf.

Torres:1999:SIS

[TKI99] Marco A. Torres, Susumu Kuroyanagi, and Akira Iwata. The Self-Indexed Search Algorithm: A bit-level approach to minimal perfect hashing. Information Processing Letters, 69(5):253–258, March 12, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/21/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/21/24/article.
pdf.

Tan:1993:RSM

[TL93] Kian-Lee Tan and Hongium Lu.

On resource scheduling of multijoin queries in parallel database systems. Information Processing Letters, 48(4):189–195, November 29, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tan:1995:WSM

[TL95] Kian-Lee Tan and Hongjun Lu.

Workload scheduling for multiple query processing. Information Processing Letters, 55(5): 251–257, September 15, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Thangavel:1993:PAA

[TM93] P. Thangavel and V. P. Muthuswamy.

Parallel algorithms for addition and multiplication on processor arrays with reconfigurable bus systems. Information Processing Letters, 46(2):89–94, May 17, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Turpin:1998:CEH

[TM98] Andrew Turpin and Alistair Moffat. Comment on “Efficient Huffman decoding” and “An efficient finite-state machine implementation of Huffman decoders”. Information Processing Letters, 68(1):1–2, October 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Chu97a, IC98].

Tanaka:1996:NLC

[TNB96] Keisuke Tanaka, Tetsuro Nishino, and Robert Beals. Negationlimited circuit complexity of symmetric functions. Information Processing Letters, 59(5): 273–279, September 9, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Toda:1990:CTS

[Tod90] Seinosuke Toda. On the complexity of topological sorting. Information Processing Letters, 35 (5):229–233, August 28, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Toledo:1993:APS

[Tol93] Sivan Toledo. Approximate parametric searching. Information Processing Letters, 47(1):1– 4, August 9, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tomkins:1995:LBT

[Tom95] Andrew Tomkins. Lower bounds for two call control problems. Information Processing Letters, 56 (3):173–178, November 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Thirusangu:1995:NCM

[TR95] K. Thirusangu and K. Rangarajan. A note on the construction of marked graphs. Information Processing Letters, 55(4): 211–215, August 25, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Trevisan:1996:NMA

[Tre96] Luca Trevisan. A note on minimum-area upward drawing of complete and Fibonacci trees. Information Processing Letters, 57(5):231–236, March 11, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Trier:1992:AWS

[Tri92] Uwe Trier. Additive weights of a special class of nonuniformly distributed backtrack trees. Information Processing Letters, 42 (2):67–76, May 11, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tan:1995:HTL

[TS95a] Xuehou Tan and Xiaoyu Song.

Hexagonal three-layer channel routing. Information Processing Letters, 55(4):223–228, August 25, 1995. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Tromp:1995:SCG

[TS95b] John Tromp and Jeffrey Shallit. Subword complexity of a generalized Thue–Morse word. Information Processing Letters, 54 (6):313–316, June 23, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tu:1995:TD

[TS95c] Hsieh-Chang Tu and Carl H. Smith. Training digraphs. Information Processing Letters, 53 (4):185–192, February 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ta-Shma:1996:NPV

[TS96] Amnon Ta-Shma. A note on PCP vs. MIP. Information Processing Letters, 58(3):135–140, May 13, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsalgatidou:1990:MAI

[Tsa90] Aphrodite Tsalgatidou. Modelling and animating information systems dynamics. Information Processing Letters, 36(3): 123–127, November 1, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsai:1992:AFT

[Tsa92] Li-Hui Tsai. An algorithm for flow time minimization and its asymptotic makespan properties. Information Processing Letters, 43(1):41–46, August 10, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsay:1993:EIP

[Tsa93] Jyh Jong Tsay. An efficient implementation of priority queues using fixed-sized systolic coprocessors. Information Processing Letters, 46(4):193–198, June 25, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tseng:1996:ERH

[Tse96] Yu-Chee Tseng. Embedding a ring in a hypercube with both faulty links and faulty nodes. Information Processing Letters, 59 (4):217–222, August 26, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Teo:1990:PAG

[TT90] Kim-Heng Teo and Tai-Ching Tuan. Performance analysis of greedy heuristic to find a minimum total-jogs layout for river routing. Information Processing Letters, 36(4):167–170, November 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Teo:1991:IUB

[TT91] Kim-Heng Teo and Tai-Ching Tuan. An improved upper bound on the number of intersections between two rectangular paths. Information Processing Letters, 37(6):305–309, March 28, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tamura:1992:DCT

[TT92a] Akihisa Tamura and Yoshiko Tamura. Degree constrained tree embedding into points in the plane. Information Processing Letters, 44(4):211–214, December 10, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Teo:1992:EOS

[TT92b] Kim-Heng Teo and Tai-Ching Tuan. An efficient one-side height minimization algorithm for routing around a rectangle. Information Processing Letters, 44(2):101–105, November 19, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsai:1993:CRA

[TT93] Ying Teh Tsai and Chuan Yi Tang. The competitiveness of randomized algorithms for online Steiner tree and on-line spanning tree problems. Information Processing Letters, 48 (4):177–182, November 29, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tiwari:1994:DVS

[TT94] Prasoon Tiwari and Martin Tompa. A direct version of Shamir and Snir’s lower bounds on monotone circuit depth. Information Processing Letters, 49 (5):243–248, March 11, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsai:1994:APG

[TTC94] Ying Teh Tsai, Chuan Yi Tang, and Yunn Yen Chen. Average performance of a greedy algorithm for the on-line minimum matching problem on Euclidean space. Information Processing Letters, 51(6):275–282, September 26, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tamassia:1991:LBP

[TTV91] Roberto Tamassia, Ioannis G.

Tollis, and Jeffrey Scott Vitter. Lower bounds for planar orthogonal drawings of graphs. Information Processing Letters, 39 (1):35–40, July 12, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Torng:1999:TLB

[TU99] Eric Torng and Patchrawat Uthaisombut. A tight lower bound for the best-α algorithm. Information Processing Letters, 71(1):17–22, July 16, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/17/
20/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/49/17/20/article.
pdf.

Turau:1991:FNN

[Tur91a] Volker Turau. Fixed-radius near neighbors search. Information Processing Letters, 39(4):201– 203, August 30, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Turski:1991:SSR

[Tur91b]  Wladyslaw M. Turski.  On

starvation and some related issues. Information Processing Letters, 37(3):171–174, February 18, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tuza:1990:PSD

[Tuz90] Zsolt Tuza. Periodic string division generated by deterministic L systems. Information Processing Letters, 34(5):229–234, May 7, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tardos:1993:IBM

[TV93] Éva Tardos and Vijay V. Vazirani. Improved bounds for the max-flow min-multicut ratio for planar and K_{r,r}-free graphs. Information Processing Letters, 47 (2):77–80, August 20, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Turan:1996:SDT

[TV96] György Turán and Farrokh Vatan. A size-depth trade-off for the analog computation of Boolean functions. Information Processing Letters, 59(5):251– 254, September 9, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tokuda:1994:AEC

[TW94] Takehiro Tokuda and Yoshimichi Watanabe. An attribute evaluation of context-free languages. Information Processing Letters, 52(2):91–98, October 28, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tokuda:1995:ESE

[TW95] Takehiro Tokuda and Yoshimichi Watanabe. An efficient semantic evaluator for warped LC(1) attributed grammars. Information Processing Letters, 53(5): 269–276, March 10, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsai:1999:EDF

[TWK99] Jichiang Tsai, Yi-Min Wang, and Sy-Yen Kuo. Evaluations of domino-free communicationinduced checkpointing protocols. Information Processing Letters, 69(1):31–37, January 15, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/17/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/17/22/article.
pdf.

Ting:1994:RAF

[TY94] Hing F. Ting and Andrew C. Yao. A randomized algorithm for finding maximum with O((log n)^{2}) polynomial tests. Information Processing Letters, 49 (1):39–43, January 14, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tyszkiewicz:1997:NKD

[Tys97] Jerzy Tyszkiewicz. A note on the Kolmogorov data complexity and nonuniform logical definitions. Information Processing Letters, 64(4):187–195, December 13, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tzeng:1996:PEN

[Tze96] Wen-Guey Tzeng. On path equivalence of nondeterministic finite automata. Information Processing Letters, 58(1):43–46, April 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tzeng:1999:CMC

[Tze99]  Wen-Guey Tzeng.  Com-

mon modulus and chosenmessage attacks on publickey schemes with linear recurrence relations. Information Processing Letters, 70(3): 153–156, May 14, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/48/19/

23/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/19/23/article.
pdf.

Uehara:1995:ESB

[Ueh95] Ryuhei Uehara. Efficient simulations by a biased coin. Information Processing Letters, 56(5): 245–248, December 8, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Uehara:1997:CPS

[Ueh97] Ryuhei Uehara. Collapse of PP with a semi-random source to BPP. Information Processing Letters, 61(2):83–87, February 28, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Urrutia:1992:AFO

[UG92] J. Urrutia and F. Gavril. An algorithm for fraternal orientation of graphs. Information Processing Letters, 41(5):271–274, April 3, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Utard:1995:DFA

[UH95] Gil Utard and Gaétan Hains.

Deadlock-free absorption of barrier synchronisations. Information Processing Letters, 56(4): 221–227, November 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Uhlmann:1991:SGP

[Uhl91] Jeffrey K. Uhlmann. Satisfying general proximity/similarity queries with metric trees. Information Processing Letters, 40 (4):175–179, November 25, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Udink:1993:UPS

[UK93] R. T. Udink and J. N. Kok. Unity properties and sequences of states, some observations. Information Processing Letters, 47 (6):307–311, October 18, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ustimenko:1996:ATL

[Ust96]  A. P. Ustimenko.  Algebra

of two-level cause-effect structures. Information Processing Letters, 59(6):325–330, September 23, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ustimenko:1998:CCE

[Ust98] A. P. Ustimenko. Colored cause– effect structures. Information Processing Letters, 68(5):219– 225, December 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Varadarajan:1997:LAS

[VA97a] Kasturi R. Varadarajan and Pankaj K. Agarwal. Linear approximation of simple objects. Information Processing Letters, 62(2):89–94, May 21, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Verhaegh:1997:PTA

[VA97b] W. F. J. Verhaegh and E. H. L.

Aarts. A polynomial-time algorithm for Knapsack with divisible item sizes. Information Processing Letters, 62(4):217–221, June 11, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vaidyanathan:1992:SPR

[Vai92] Ramachandran Vaidyanathan.

Sorting on PRAMs with reconfigurable buses. Information Processing Letters, 42(4):203– 208, June 19, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vajnovszki:1998:LGB

[Vaj98] Vincent Vajnovszki. On the loopless generation of binary tree sequences. Information Processing Letters, 68(3):113– 117, November 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/46/18/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/46/18/24/article.
pdf.

Valmari:1995:WDP

[Val95a] Antti Valmari. The weakest deadlock-preserving congruence. Information Processing Letters, 53(6):341–346, March 24, 1995.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Valois:1995:VWP

[Val95b] John D. Valois. A 3-valued wakeup protocol. Information Processing Letters, 55(2):89–93, July 21, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanBaronaigien:1991:LAG

[van91a] D. Roelants van Baronaigien. A loopless algorithm for generating binary tree sequences. Information Processing Letters, 39(4): 189–194, August 30, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vandeSnepscheut:1991:IRT

[van91b] Jan L. A. van de Snepscheut. Inversion of a recursive tree traversal. Information Processing Letters, 39(5):265–267, September 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

VanDeSnepscheut:1992:LPE

[Van92a] Jan L. A. Van De Snepscheut. A LISP programming exercise. Information Processing Letters, 42 (2):103–108, May 11, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanVliet:1992:ILB

[van92b] André van Vliet. An improved lower bound for on-line bin packing algorithms. Information Processing Letters, 43(5):277– 284, October 5, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanKreveld:1993:PPP

[van93] Marc van Kreveld. The power of parallel projection (computational geometry). Information Processing Letters, 46(4): 185–191, June 25, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanDijk:1997:MIT

[van97] Marten van Dijk. More information theoretical inequalities to be used in secret sharing? Information Processing Letters, 63 (1):41–44, July 30, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vardi:1990:ECM

[Var90] Moshe Y. Vardi. Endmarkers can make a difference. Information Processing Letters, 35(3): 145–148, July 20, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vasconcelos:1995:UKI

[Vas95] Vasco Thudichum Vasconcelos.

Unification of kinded infinite trees. Information Processing Letters, 55(6):323–328, September 29, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vassiliev:1997:DTO

[Vas97] V. A. Vassiliev. On decision trees for orthants. Information Processing Letters, 62(5): 265–268, July 2, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vavasis:1990:QPN

[Vav90] Stephen A. Vavasis. Quadratic programming is in NP. Information Processing Letters, 36(2): 73–77, October 15, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Veanes:1996:CNC

[VB96a] Margus Veanes and Jonas Barklund. Construction of natural cycletrees. Information Processing Letters, 60(6):313–318, December 23, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Veanes:1996:NEC

[VB96b] Margus Veanes and Jonas Barklund. On the number of edges in cycletrees. Information Processing Letters, 57(4):225–229, February 26, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vonKarger:1997:CKD

[vB97] Burghard von Karger and Rudolf Berghammer. Computing kernels in directed bichromatic graphs. Information Processing Letters, 62(1):5–11, May 7, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baronaigien:1999:MSM

[vB99] Dominique Roelants van Baronaigien. A multi-stack method for the fast generation of permutations with minimal length increasing subsequences. Information Processing Letters, 69(3):123–126, February 12, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/19/
21/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/19/21/article.
pdf.

Viswanathan:1996:FTE

[VCS96] Suresh Viswanathan, Éva Czabarka, and Abhijit Sengupta. On fault-tolerant embedding of Hamiltonian circuits in line digraph interconnection networks. Information Processing Letters, 57(5):265–271, March 11, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Varghese:1995:DGV

[VCW95] George Varghese, Roger Chamberlain, and William E. Weihl. Deriving global virtual time algorithms from conservative simulation protocols. Information Processing Letters, 54(2):121– 126, April 28, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Veith:1997:LRB

[Vei97] Helmut Veith. Languages represented by Boolean formulas. Information Processing Letters, 63(5):251–256, September 24, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Veljan:1994:CVP

[Vel94] D. Veljan. Computing values of a polynomial with only few multiplications. Information Processing Letters, 49(1):33–37, January 14, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Veloso:1996:PCM

[Vel96] Paulo A. S. Veloso. On pushout consistency, modularity and interpolation for logical specifications. Information Processing Letters, 60(2):59–66, October 28, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Venkatesh:1998:PAB

[Ven98] S. Venkatesh. Pseudo-average block sensitivity equals average sensitivity. Information Processing Letters, 68(2):93– 95, October 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/46/19/
                                        22/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/46/19/22/article.
                                        pdf.

Verjus:1987:PDA

[Ver87] J. Pierre Verjus. On the proof of a distributed algorithm. Information Processing Letters, 25 (3):145–147, May 29, 1987. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comments [VT90, Pet91].

Verma:1992:STP

[Ver92] Rakesh M. Verma. Strings, trees, and patterns. Information Processing Letters, 41(3): 157–161, March 6, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Verbarg:1997:ACP

[Ver97] Knut Verbarg. Approximate center points in dense point sets. Information Processing Letters, 61(5):271–278, April 10, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Verhoeven:1999:RSU

[Ver99]  Yann Verhoeven.  Random

2-SAT and unsatisfiability. Information Processing Letters, 72(3–4):119–123, November 26, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/28/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/28/32/article.
pdf.

vonKarger:1995:SC

[vH95] Burghard von Karger and C. A. R. Hoare. Sequential calculus. Information Processing Letters, 53(3):123–130, February 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vaidyanathan:1992:PVW

[VHV92] Ramachandran Vaidyanathan, Carlos R. P. Hartmann, and Pramod K. Varshney. PRAMs with variable word-size. Information Processing Letters, 42 (4):217–222, June 19, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vaidyanathan:1993:RAD

[VHV93] Ramachandran Vaidyanathan, Carlos R. P. Hartmann, and Pramod K. Varshney. Running ASCEND, DESCEND and PIPELINE algorithms in parallel using small processors. Information Processing Letters, 46 (1):31–36, April 29, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vedantham:1998:BAP

[VI98] Sundararajan Vedantham and S. S. Iyengar. The Bandwidth Allocation Problem in the ATM network model is NPcomplete. Information Processing Letters, 65(4):179–182, February 27, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vidyasankar:1991:VSC

[Vid91] K. Vidyasankar. A very simple construction of 1-writer multireader multivalued atomic variable. Information Processing Letters, 37(6):323–326, March 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vilar:1995:ROA

[Vil95] Juan Miguel Vilar. Reducing the overhead of the AESA metricspace nearest neighbour searching algorithm. Information Processing Letters, 56(5):265–271, December 8, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vishwanathan:1992:AAA

[Vis92] Sundar Vishwanathan. An approximation algorithm for the asymmetric travelling salesman problem with distances one and two. Information Processing Letters, 44(6):297–302, December 28, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Veeraraghavulu:1992:LAI

[VKM92] N. Ch. Veeraraghavulu, P. Sreenivasa Kumar, and C. E. Veni Madhavan. A linear-time algorithm for isomorphism of a subclass of chordal graphs. Information Processing Letters, 44(1): 45–49, November 9, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Veloso:1995:MTL

[VM95] Paulo A. S. Veloso and Thomas S. E. Maibaum. On the Modularization Theorem for logical specifications. Information Processing Letters, 53(5):287–293, March 10, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vakali:1997:EAE

[VM97] Athena Vakali and Yannis Manolopoulos. An exact analysis on expected seeks in shadowed disks. Information Processing Letters, 61(6):323–329, April 24, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vitini:1999:LCO

[VMD99] F. Montoya Vitini, J. Muñoz Masqué, and A. Peinado Domínguez. Linear complexity of the x^{2} mod p orbits. Information Processing Letters, 72(1–2):3–7, October 29, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/27/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/27/25/article.
pdf.

Vogler:1995:FPO

[Vog95] Walter Vogler. Fairness and partial order semantics. Information Processing Letters, 55(1): 33–39, July 7, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Volpano:1996:LBT

[Vol96] Dennis M. Volpano. Lower bounds on type checking overloading. Information Processing Letters, 57(1):9–13, January 15, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vonStengel:1991:ACS

[von91a] Bernhard von Stengel. An algebraic characterization of semantic independence. Information Processing Letters, 39(6): 291–296, September 27, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vonWright:1991:PIR

[von91b] J. von Wright. Program inversion in the refinement calculus. Information Processing Letters, 37(2):95–100, January 31, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

VonZurGathen:1992:PEE

[Von92] Joachim Von Zur Gathen.

Processor-efficient exponentiation in finite fields. Information Processing Letters, 41(2):81–86, February 14, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Voss:1993:WCP

[Vos93] Stefan Voss. Worst-case performance of some heuristics for Steiner’s problem in directed graphs. Information Processing Letters, 48(2):99–105, November 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Venkatesan:1996:ATC

[VR96] G. Venkatesan and C. Pandu Rangan. Approximate triclique coloring for register allocation. Information Processing Letters, 60(5):249–253, December 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vasconcelos:1999:CEC

[VR99] Vasco T. Vasconcelos and António Ravara. Communication errors in the π-calculus are undecidable. Information Processing Letters, 71(5–6):229– 233, September 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/30/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/30/34/article.
pdf.

Vincent:1993:NRS

[VS93] Millist W. Vincent and Bala Srinivasan. A note on relation schemes which are in 3NF but not in BCNF. Information Processing Letters, 48(6):281–283, December 20, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vadapalli:1995:TCG

[VS95a] Premkumar Vadapalli and Pradip K.

Srimani. Trivalent Cayley graphs for interconnection networks. Information Processing Letters, 54(6):329–335, June 23, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Volpano:1995:TSP

[VS95b] Dennis Volpano and Geoffrey Smith. A type soundness proof for variables in LCF ML. Information Processing Letters, 56 (3):141–146, November 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vadapalli:1996:SRT

[VS96] Premkumar Vadapalli and Pradip K.

Srimani. Shortest routing in trivalent Cayley graph network. Information Processing Letters, 57(4):183–188, February 26, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Venema:1996:NAS

[VSS96] Sven Venema, Hong Shen, and Francis Suraweera. NC algorithms for the Single Most Vital Edge problem with respect to shortest paths. Information Processing Letters, 60(5):243–248, December 8, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

VanGasteren:1990:CPD

[VT90] A. J. M. Van Gasteren and G. Tel. Comments on “On the proof of a distributed algorithm”: always-true is not invariant. Information Processing Letters, 35(6):277–279, September 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Ver87, Pet91].

Vaidyanathan:1993:OSM

[VT93] Ramachandran Vaidyanathan and Jerry L. Trahan. Optimal simulation of multidimensional reconfigurable meshes by two-dimensional reconfigurable meshes. Information Processing Letters, 47(5):267–273, October 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanWezel:1994:APR

[VT94] Michiel C. Van Wezel and Gerard Tel. An assertional proof of Rana’s algorithm. Information Processing Letters, 49(5): 227–233, March 11, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Veloso:1997:MSI

[VV97] Paulo A. S. Veloso and Sheila R. M. Veloso. On methods for safe introduction of operations. Information Processing Letters, 64(5):231–238, December 23, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vyskoc:1990:MBR

[Vys90] Jozef Vyskoc. Making bubblesort recursive. Information Processing Letters, 36(4):219–220, November 15, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wagner:1994:AML

[Wag94] Frank Wagner. Approximate Map Labeling is in Ω(n log n). Information Processing Letters, 52(3):161–165, November 11, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wagner:1998:NPQ

[Wag98] Klaus W. Wagner. A note on parallel queries and the symmetric-difference hierarchy. Information Processing Letters, 66(1):13–20, April 15, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Waldmann:1998:ERO

[Wal98] Uwe Waldmann. Extending reduction orderings to ACUcompatible reduction orderings. Information Processing Letters, 67(1):43–49, July 16, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wand:1990:SPL

[Wan90a] Mitchell Wand. A short proof of the lexical addressing algorithm. Information Processing Letters, 35(1):1–5, June 15, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1990:OLE

[Wan90b] Ching-Lin Wang. Obtaining lazy evaluation with continuations in Scheme. Information Processing Letters, 35(2):93–97, June 29, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1991:CEP

[Wan91] Xiaodong Wang. On the complexity of the extreme points decision problem. Information Processing Letters, 40(2):101– 106, October 25, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1992:IPA

[Wan92a] Guorong Wang. An improved parallel algorithm for computing the generalized inverse A^{+}. Information Processing Letters, 41 (5):243–251, April 3, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1992:NTW

[Wan92b] Jie Wang. A note on two-way probabilistic automata. Information Processing Letters, 43 (6):321–326, October 19, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1994:NRA

[Wan94] Chien-Min M. Wang. A new routing algorithm for cyclic shifts on BRGC hypercubes. Information Processing Letters, 49 (4):165–169, February 25, 1994.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1995:SRS

[Wan95] Jie Wang. Some results on selectivity and self-reducibility. Information Processing Letters, 55 (2):81–87, July 21, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1997:LCP

[Wan97a] M. Z. Wang. Linear complexity profiles and jump complexity. Information Processing Letters, 61(3):165–168, March 14, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1997:DRF

[Wan97b] Sying-Jyan Wang. Distributed routing in a fault-tolerant multistage interconnection network. Information Processing Letters, 63(4):205–210, September 10, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1997:NHS

[Wan97c] Yongge Wang. NP-hard sets are superterse unless NP is small. Information Processing Letters, 61(1):1–6, January 14, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1998:SNR

[Wan98]  Huaxiong Wang.  On syn-

tactic nuclei of rational languages. Information Processing Letters, 67(5):221–226, September 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1999:STR

[Wan99a] Yongge Wang. A separation of two randomness concepts. Information Processing Letters, 69(3):115–118, February 12, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/19/
19/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/19/19/article.
pdf.

Wang:1999:CQS

[Wan99b] Zheng Wang. On the complexity of quality of service routing. Information Processing Letters, 69(3):111–114, February 12, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/19/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/19/18/article.
pdf.

Warners:1998:LTT

[War98] Joost P. Warners. A lineartime transformation of linear inequalities into conjunctive normal form. Information Processing Letters, 68(2): 63–69, October 30, 1998. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/46/19/
20/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/46/19/20/article.
pdf.

Wang:1993:HIS

[WB93] Jie Wang and Jay Belanger.

Honest iteration schemes of randomizing algorithms. Information Processing Letters, 45(6): 275–278, April 16, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1990:TDP

[WC90] Biing-Feng Wang and Gen-Huey Chen. Two-dimensional processor array with a reconfigurable bus system is at least as powerful as CRCW model. Information Processing Letters, 36(1):31–36, October 1, 1990. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1991:LSP

[WC91] Qingzhou Wang and Kam Hoi Cheng. List scheduling of parallel tasks. Information Processing Letters, 37(5):291–297, March 14, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wrightson:1992:TTC

[WC92a] G. Wrightson and J. Coldwell. A truncation technique for clausal analytic tableaux. Information Processing Letters, 42(5):273– 281, July 3, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:1992:THP

[WC92b] Jer Shyan Wu and Rong Jaye Chen. The Towers of Hanoi problem with parallel moves. Information Processing Letters, 44 (5):241–243, December 21, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:1993:THP

[WC93] Jer Shyan Wu and Rong Jaye Chen. The Towers of Hanoi problem with cyclic parallel moves. Information Processing Letters, 46(1):1–6, April 29, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wei:1997:BSP

[WC97] Li-Chiun Wei and Ruay-Shiung Chang. Broadcast scheduling in packet radio networks by Hopfield neural networks. Information Processing Letters, 63(5): 271–276, September 24, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1991:SAI

[WCC91] Biing-Feng Wang, Chuen-Liang Chen, and Gen-Huey Chen. A simple approach to implementing multiplication with small tables. Information Processing Letters, 37(6):327–329, March 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1990:CTS

[WCL90] Biing-Feng Wang, Gen-Huey Chen, and Ferng-Ching Lin. Constant time sorting on a processor array with a reconfigurable bus system. Information Processing Letters, 34(4): 187–192, April 24, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1995:PAC

[WCL95] Yue Li Wang, Hon Chan Chen, and Chen Yu Lee. An O(log n) parallel algorithm for constructing a spanning tree on permutation graphs. Information Processing Letters, 56(2):83–87, October 27, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:1998:NCL

[WCL98] Q. S. Wu, K. M. Chao, and R. C. T. Lee. The NPOcompleteness of the longest Hamiltonian cycle problem. Information Processing Letters, 65 (3):119–123, February 13, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:1999:EAL

[WCT99] Bang Ye Wu, Kun-Mao Chao, and Chuan Yi Tang. An efficient algorithm for the lengthconstrained heaviest path problem on a tree. Information Processing Letters, 69(2):63– 67, January 29, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/18/
19/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/18/19/article.
pdf.

Wee:1990:ARN

[WCW90] Young C. Wee, Seth Chaiken, and Dan E. Willard. On the angle restricted nearest neighbor problem. Information Processing Letters, 34(2):71–76, March 16, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1999:MWT

[WCY99] Cao An Wang, Francis Y. Chin, and Bo Ting Yang. Maximum weight triangulation and graph drawing. Information Processing Letters, 70(1):17– 22, April 16, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/17/
20/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/17/20/article.
pdf.

Wegener:1990:ESC

[Weg90] Ingo Wegener. Efficient simulation of circuits by EREW PRAMs. Information Processing Letters, 35(2):99–102, June 29, 1990. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Wegener:1993:OLB

[Weg93]  Ingo Wegener.  Optimal

lower bounds on the depth of polynomial-size threshold circuits for some arithmetic functions. Information Processing Letters, 46(2):85–87, May 17, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wegener:1996:CEA

[Weg96] Ingo Wegener. On the complexity of encoding in analog circuits. Information Processing Letters, 60(1):49–52, October 14, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wein:1991:VRA

[Wei91] Joel Wein. Las Vegas RNC algorithms for unary weighted perfect matching and T -join problems. Information Processing Letters, 40(3):161–167, November 8, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Weiss:1994:LTC

[Wei94] Mark Allen Weiss. Lineartime construction of treaps and Cartesian trees. Information Processing Letters, 52(5):253– 257, December 9, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Weiss:1995:NCT

[Wei95] M. A. Weiss. A note on construction of treaps and Cartesian trees. Information Processing Letters, 54(2):127, April 28, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wen:1991:PMS

[Wen91] Zhaofang Wen. Parallel multiple search. Information Processing Letters, 37(4):181–186, February 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wen:1994:NAL

[Wen94]  Zhaofang Wen.  New algo-

rithms for the LCA problem and the binary tree reconstruction problem. Information Processing Letters, 51(1):11–16, July 12, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1996:FGD

[WF96] Xiaodong Wang and Qingxiang Fu. A frame for general divideand-conquer recurrences 1. Information Processing Letters, 59 (1):45–51, July 8, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wuu:1994:IAU

[WH94] Lih-Chyau Wuu and Shing-Tsaan Huang. Identity assignment in uniform synchronous rings. Information Processing Letters, 49(5):257–262, March 11, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Whaley:1992:ADC

[Wha92] Tom Whaley. Alternative developments of cyclic-permutation algorithms. Information Processing Letters, 41(5):239–241, April 3, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1998:OHG

[WHH98] Jeng-Jung Wang, Chun-Nan Hung, and Lih-Hsing Hsu. Optimal 1-Hamiltonian graphs. Information Processing Letters, 65 (3):157–161, February 13, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1999:CTG

[WHL99] Chih-Hung Wang, Tzonelih Hwang, and Narn-Yih Lee. Comments on two group signatures. Information Processing Letters, 69(2):95–97, January 29, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/18/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/18/24/article.
pdf.

Wang:1997:FEF

[WHS97] Shih-Yih Wang, Lih-Hsing Hsu, and Ting-Yi Sung. Faithful 1edge fault tolerant graphs. Information Processing Letters, 61 (4):173–181, March 27, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wierzbowska:1997:CDS

[Wie97] Izabela Wierzbowska. Checker for data structures which sort elements. Information Processing Letters, 64(5):245–249, December 23, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1999:NSM

[WIIO99] Y. Wang, K. Inoue, A. Ito, and T. Okazaki. A note on self-modifying finite automata. Information Processing Letters, 72(1–2):19–24, October 29, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/60/27/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/60/27/27/article.
                                        pdf.

Wismath:1992:CFV

[Wis92] S. K. Wismath. Computing the full visibility graph of a set of line segments. Information Processing Letters, 42(5):257–261, July 3, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wise:1993:SOR

[Wis93] David S. Wise. Stop-and-copy and one-bit reference counting.

Information Processing Letters, 46(5):243–249, July 9, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wolfson:1995:ADD

[WJ95] Ouri Wolfson and Sushil Jajodia. An algorithm for dynamic data allocation in distributed systems. Information Processing Letters, 53(2):113–119, January 27, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1993:CMS

[WL93] Feng Hsu Wang and Ferng-Ching Lin. On constructing multiple spanning trees in a hypercube. Information Processing Letters, 45(4):177–183, March 22, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1994:CGC

[WLF94] Yi-Min Wang, Andy Lowry, and W. Kent Fuchs. Consistent global checkpoints based on direct dependency tracking. Information Processing Letters, 50 (4):223–230, May 25, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wada:1992:OFT

[WLK92] Koichi Wada, Yupin Luo, and Kimio Kawaguchi. Optimal fault-tolerant routings for connected graphs. Information Processing Letters, 41(3):169–174, March 6, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wagh:1996:HCT

[WM96] Meghanad D. Wagh and Jiancheng Mo. Hamilton cycles in Trivalent Cayley graphs. Information Processing Letters, 60(4):177–181, November 25, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:1990:SCA

[WMMM90] Sun Wu, Udi Manber, Gene Myers, and Webb Miller. O(NP) sequence comparison algorithm. Information Processing Letters, 35(6):317–323, September 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Woeginger:1990:SST

[Woe90] Gerhard Woeginger. A simple solution to the two paths problem in planar graphs. Information Processing Letters, 36 (4):191–192, November 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Woeginger:1991:MST

[Woe91] Gerhard Woeginger. On minimizing the sum of k tardinesses. Information Processing Letters, 38(5):253–256, June 14, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Woeginger:1992:CFA

[Woe92a] Gerhard J. Woeginger. The complexity of finding arborescences in hypergraphs. Information Processing Letters, 44 (3):161–164, November 30, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Woeginger:1992:FCE

[Woe92b] Gerhard J. Woeginger. Finding the closest extreme vertex to a fixed point. Information Processing Letters, 43(3):127– 128, September 14, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Woeginger:1995:STD

[Woe95] Gerhard J. Woeginger. Scheduling with time-dependent execution times. Information Processing Letters, 54(3):155–156, May 12, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Woeginger:1998:TNA

[Woe98] Gerhard J. Woeginger. There is no asymptotic PTAS for twodimensional vector packing. Information Processing Letters, 64 (6):293–297, January 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wojna:1999:CM

[Woj99] Arkadiusz Wojna. Counter machines. Information Processing Letters, 71(5–6):193–197, September 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/30/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/30/29/article.
pdf.

Wright:1999:RRS

[Wri99] William E. Wright. A refinement of replacement selection. Information Processing Letters, 70 (3):107–111, May 14, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/48/19/
17/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/48/19/17/article.
pdf.

Weiss:1990:MSI

[WS90] Mark Allen Weiss and Robert Sedgewick. More on Shellsort increment sequences. Information Processing Letters, 34(5): 267–270, May 7, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Weems:1994:ATD

[WS94] Bob P. Weems and Lloyd C. Swayze. Allocation techniques for distributed reduction data elements. Information Processing Letters, 50(3):137–142, May 9, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Watjen:1996:DRC

[WS96] Dietmar Wätjen and Heike Spilker. Decidability results concerning k-limited ED0L systems. Information Processing Letters, 59(1):13–17, July 8, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:1992:SSP

[WT92] L. C. Wu and C. Y. Tang. Solving the satisfiability problem by using randomized approach. Information Processing Letters, 41 (4):187–190, March 18, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:1997:AFO

[WT97] B. Y. Wu and C. Y. Tang.

An O(n) algorithm for finding an optimal position with relative distances in an evolutionary tree. Information Processing Letters, 63(5):263–269, September 24, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:1990:AEH

[WWT90] Chwan-Hwa Wu, Chia-Jiu Wang, and Heng-Ming Tai. Application of exponential heteroassociative memory on frequency classifier. Information Processing Letters, 35(6):305– 311, September 15, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1992:PIC

[WY92a] Ke Wang and Li Yan Yuan. Preservation of integrity constraints in definite DATALOG programs. Information Processing Letters, 44(4):185–193, December 10, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Woeginger:1992:ESS

[WY92b] Gerhard J. Woeginger and Zhong Liang Yu. On the equalsubset-sum problem. Information Processing Letters, 42(6): 299–302, July 24, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Westbrook:1995:LBL

[WY95] Jeffery Westbrook and Dicky Yan. Linear bounds for on-line Steiner problems. Information Processing Letters, 55(2):59–63, July 21, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xu:1993:OPL

[XL93] C.-Z. Xu and F. C. M. Lau. Optimal parameters for load balancing using the diffusion method in k-ary n-cube networks. Information Processing Letters, 47(4):181–187, September 27, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xue:1998:FDT

[XSR98] Guoliang Xue, Shangzhi Sun, and J. Ben Rosen. Fast data transmission and maximal dynamic flow. Information Processing Letters, 66(3):127–132, May 15, 1998. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xiang:1998:APB

[XU98] Limin Xiang and Kazuo Ushijima. ANSV problem on BSRs. Information Processing Letters, 65(3):135–138, February 13, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xiang:1999:TRB

[XU99a] L. Xiang and K. Ushijima. A theorem on the relation between BSR_{k} and BSR^{+}. Information Processing Letters, 71(2): 71–73, July 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/18/
20/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/18/20/article.
pdf.

Xiang:1999:RSI

[XU99b] Limin Xiang and Kazuo Ushijima. Rearranging scattered information on BSR. Information Processing Letters, 71(1): 43–47, July 16, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/49/17/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/23/20/49/17/24/article.
                                        pdf.

Xue:1996:CSF

[Xue96] Qingbo Xue. On a class of square-free graphs. Information Processing Letters, 57(1):47–48, January 15, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:1991:RCR

[Yan91a] Chang-Biau B. Yang. Reducing conflict resolution time for solving graph problems in broadcast communications. Information Processing Letters, 40(6): 295–302, December 30, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:1991:NER

[Yan91b] Yi Xian Yang. New enumeration results about the optical orthogonal codes. Information Processing Letters, 40(2):85–87, October 25, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:1993:ILP

[Yan93a] Wuu Yang. An incremental LL(1) parsing algorithm. Information Processing Letters, 48 (2):67–72, November 8, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:1993:NBS

[Yan93b] Yi Xian Yang. New binary sequences with perfect staircase profile of linear complexity. Information Processing Letters, 46 (1):27–29, April 29, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:1995:BSM

[Yan95] Bo-Ting Yang. A better subgraph of the minimum weight triangulation. Information Processing Letters, 56(5):255–258, December 8, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:1998:DPA

[Yan98] Wuu Yang. A data-parallel algorithm for minimum-width tree layout. Information Processing Letters, 67(1):21–28, July 16, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yao:1992:FAS

[Yao92] Xin Yao. Finding approximate solutions to NP-hard problems by neural networks is hard. Information Processing Letters, 41 (2):93–98, February 14, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yao:1995:NNS

[Yao95] Xin Yao. A note on neural sorting networks with O(1) time complexity. Information Processing Letters, 56(5):253–254, December 8, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cai:1990:LBC

[yC90] Jin yi Cai. Lower bounds for constant-depth circuits in the presence of help bits. Information Processing Letters, 36(2): 79–83, October 15, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yan:1994:PPP

[YC94] Jing Ho Yan and Gerard J. Chang. The path-partition problem in block graphs. Information Processing Letters, 52(6):317– 322, December 23, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See corrigendum [Cha02].

Chou:1999:DFT

[yCGlW99a] C. y. Chou, D. J. Guan, and K. l. Wang. A dynamic fault-tolerant message routing algorithm for doubleloop networks. Information Processing Letters, 70(6):259– 264, June 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/48/22/
18/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/48/22/18/article.

pdf. See erratum [yCGlW99b].

Chou:1999:ESD

[yCGlW99b] Chun yen Chou, D. J. Guan, and Kuei lin Wang. Erratum to “A dynamic fault-tolerant message routing algorithm for double-loop networks”, Information Processing Letters 70(6)

(1999) 259–264. Information Processing Letters, 71(5–6):241, September 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/49/30/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/49/30/36/article.

pdf. See [yCGlW99a].

Cai:1991:NEC

[yCH91] Jin yi Cai and Lane A.

Hemachandra. A note on enumerative counting. Information Processing Letters, 38(4):215– 219, May 31, 1991. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yun:1993:SER

[YCH93] In Sig Yun, Kwang Moo Choe, and Taisook Han. Syntactic error repair using repair patterns. Information Processing Letters, 47(4):189–196, September 27, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:1994:PEM

[YDL94] S. B. Yang, S. K. Dhall, and S. Lakshmivarahan. A processor efficient MIS algorithm on random graphs. Information Processing Letters, 49(3):157–163, February 11, 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yen:1991:PTA

[Yen91] Hsu Chun Yen. A polynomial time algorithm to decide pairwise concurrency of transitions for 1-bounded conflict-free Petri nets. Information Processing Letters, 38(2):71–76, April 26, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yen:1995:NFC

[Yen95] Hsu-Chun Yen. A note on fine covers and iterable factors of VAS languages. Information Processing Letters, 56(5): 237–243, December 8, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yen:1996:HSS

[Yen96] I-Ling Yen. A highly safe selfstabilizing mutual exclusion algorithm. Information Processing Letters, 57(6):301–305, March 25, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yan:1994:FCA

[YH94] Jin Tai Yan and Pei Yung Hsiao. A fuzzy clustering algorithm for graph bisection. Information Processing Letters, 52(5):259– 263, December 9, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yellin:1993:FES

[YJ93] Daniel M. Yellin and Charanjit S. Jutla. Finding extremal sets in less than quadratic time.

Information Processing Letters, 48(1):29–34, October 29, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yen:1990:WPD

[YL90] Chain-Chin Yen and R. C. T.

Lee. The weighted perfect domination problem. Information Processing Letters, 35(6):295– 299, September 15, 1990. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:1991:SCQ

[YL91] Jae Dong Yang and Yoon Joon Lee. A sound and complete query evaluation for implicit predicate which is a semantic descriptor of unknown values. Information Processing Letters, 39(5):283–289, September 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:1997:NEC

[YL97a] Ching-Nung Yang and Chi-Sung Laih. A note on error-correcting codes for authentication and subliminal channels. Information Processing Letters, 62(3): 141–143, June 4, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yen:1997:SAT

[YL97b] Sung-Ming Yen and Kuo-Hong Liao. Shared authentication token secure against replay and weak key attacks. Information Processing Letters, 62(2):77–80, May 21, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Young-Lai:1999:ASM

[YL99] Matthew Young-Lai. Adding state merging to the DMC data compression algorithm. Information Processing Letters, 70 (5):223–228, June 21, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/48/21/
                                        18/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/48/21/18/article.
                                        pdf.

Yu:1991:TRG

[YLW91] Bin Yu, Xinggang Lin, and Youshou Wu. The tree representation of the graph used in binary image processing. Information Processing Letters, 37(1): 55–59, January 10, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yan:1998:FNS

[YMC98] Wen-Ming Yan, Wendy Myrvold, and Kuo-Liang Chung. A formula for the number of spanning trees of a multi-star related graph. Information Processing Letters, 68(6):295–298, December 30, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yodaiken:1991:MFC

[Yod91] Victor Yodaiken. Modal functions for concise definition of state machines and products. Information Processing Letters, 40(2):65–72, October 25, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yeh:1998:VLC

[YP98] Chi-Hsiang Yeh and Behrooz Parhami. VLSI layouts of complete graphs and star graphs. Information Processing Letters, 68(1):39–45, October 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yanbe:1996:SCN

[YS96] Akio Yanbe and Kouichi Sakurai. A short certificate of the number of universal optimal strategies for stopping simple stochastic games. Information Processing Letters, 57(1): 17–24, January 15, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Youssef:1995:RBB

[YT95] A. M. Youssef and S. E. Tavares.

Resistance of balanced s-boxes to linear and differential cryptanalysis. Information Processing Letters, 56(5):249–252, December 8, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Youssef:1996:CBN

[YT96] A. M. Youssef and S. E. Tavares.

Comment on: “Bounds on the number of functions satisfying the strict avalanche criterion” [Inform. Process. Lett. 60 (1996), no. 4, 215–219; 1 435 155] by T. W. Cusick and P. Stănică. Information Processing Letters, 60(5):271–275, December 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Cus96, CS96].

Yu:1993:SPA

[YTC93] Ming Shing Yu, Lin Yu Tseng, and Shoe Jane Chang. Sequential and parallel algorithms for the maximum-weight independent set problem on permutation graphs. Information Processing Letters, 46(1):7–11, April 29, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yu:1995:NST

[Yu95] Xiaokang Yu. A new solution for Thue’s problem. Information Processing Letters, 54(4): 187–191, May 26, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yuan:1990:CCD

[Yua90] Shyan-Ming Yuan. The communication complexity for decentralized evaluation of functions. Information Processing Letters, 35(4):177–182, August 7, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yuan:1991:TPS

[Yua91] Shyan Ming Yuan. Topological properties of supercube. Information Processing Letters, 37 (5):241–245, March 14, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yuan:1994:LPS

[Yua94] Li Yan Yuan. Logic program semantics and circumscription of autoepistemic theories. Information Processing Letters, 50(3): 159–164, May 9, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Young:1996:DMB

[YW96] Christopher H. Young and Philip A. Wilsey. A distributed method to bound rollback lengths for fossil collection in time warp simulators. Information Processing Letters, 59 (4):191–196, August 26, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yan:1999:CUU

[YWC99] K. Q. Yan, S. C. Wang, and Y. H. Chin. Consensus under unreliable transmission. Information Processing Letters, 69(5): 243–248, March 12, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/47/21/
21/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/47/21/21/article.
pdf.

Yu:1992:OPA

[YY92] Ming Shing Yu and Cheng-Hsing Yang. An optimal parallel algorithm for the domatic partition problem on an interval graph given its sorted model. Information Processing Letters, 44(1): 15–22, November 9, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yu:1993:TAM

[YY93a] Ming Shing Yu and Cheng-Hsing Yang. An O(n) time algorithm for maximum matching on cographs. Information Processing Letters, 47(2):89–93, August 20, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yu:1993:SOP

[YY93b] Ming Shing Yu and Cheng Hsing Yang. A simple optimal parallel algorithm for the minimum coloring problem on interval graphs. Information Processing Letters, 48(1):47–51, October 29, 1993. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yamakami:1999:NCC

[YY99] T. Yamakami and A. C. Yao. NQP_{C} = co-C_{=}P. Information Processing Letters, 71(2): 63–69, July 30, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/49/18/
                                        19/abstract.html;   http:

//www.elsevier.nl/gej-ng/
29/28/24/49/18/19/article.
pdf.

Zargham:1990:TDF

[ZD90] M. R. Zargham and K. J. Danhof. Toward a definition of fault analysis for Petri nets models. Information Processing Letters, 34(6):299–305, May 28, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zelikovsky:1993:FAA

[Zel93] Alexander Z. Zelikovsky. A faster approximation algorithm for the Steiner tree problem in graphs. Information Processing Letters, 46(2):79–83, May 17, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:1991:OCD

[Zha91] Xubo Zhang. Overlap closures do not suffice for termination of general term rewriting systems. Information Processing Letters, 37(1):9–11, January 10, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:1992:PNP

[Zha92] Louxin Zhang. The pre-NTS property is undecidable for context-free grammars. Information Processing Letters, 44 (4):181–184, December 10, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:1997:SSL

[Zha97] Guochuan Zhang. A simple semi on-line algorithm for P2//C_{max} with a buffer. Information Processing Letters, 61(3):145–148, March 14, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zheng:1995:KAP

[Zhe95] Yuliang Zheng. On key agreement protocols based on tamperproof hardware. Information Processing Letters, 53(1):49–54, January 13, 1995. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhu:1992:CSD

[Zhu92] B. Zhu. Computing the shortest diagonal of a monotone polygon in linear time. Information Processing Letters, 42(6): 303–307, July 24, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zheng:1998:HCE

[ZI98] Yuliang Zheng and Hideki Imai. How to construct efficient signcryption schemes on elliptic curves. Information Processing Letters, 68(5):227–233, December 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zimand:1996:HLK

[Zim96] Marius Zimand. A high-low Kolmogorov complexity law equivalent to the 0-1 law. Information Processing Letters, 57(2):59–64, January 29, 1996. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zimand:1997:LSM

[Zim97] Marius Zimand. Large sets in AC^{0} have many strings with low Kolmogorov complexity. Information Processing Letters, 62 (3):165–170, June 4, 1997. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zimand:1999:RRO

[Zim99] Marius Zimand. Relative to a random oracle, P/poly is not measurable in EXP. Information Processing Letters, 69(2): 83–86, January 29, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/47/18/
22/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/47/18/22/article.
pdf.

Zippel:1997:ZTA

[Zip97] Richard Zippel. Zero testing of algebraic functions. Information Processing Letters, 61(2):63–67, February 28, 1997. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zissimopoulos:1995:PGN

[Zis95] V. Zissimopoulos. On the performance guarantee of neural networks for NP-hard optimization problems. Information Processing Letters, 54(6):317–322, June 23, 1995. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Zivkovic:1991:FAF

[Ziv91] Dejan Zivkovic. A fast algorithm for finding the compact sets. Information Processing Letters, 38 (6):339–342, June 28, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:1994:SMS

[ZJ94] Kaizhong Zhang and Tao Jiang. Some MAX SNP-hard results concerning unordered labeled trees. Information Processing Letters, 49(5):249–254, March 11, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhou:1995:GVR

[ZNN95] Xiao Zhou, Nobuaki Nagai, and Takao Nishizeki. Generalized vertex-rankings of trees. Information Processing Letters, 56 (6):321–328, December 22, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:1994:SRC

[ZÖ94] Xubo Zhang and Z. Meral Özsoyoǧlu. Some results on the containment and minimization of (in)equality queries. Information Processing Letters, 50(5): 259–267, June 10, 1994. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:1992:EDB

[ZSS92] Kaizhong Zhang, Rick Statman, and Dennis Shasha. On the editing distance between unordered labeled trees. Information Processing Letters, 42(3):133–139, May 25, 1992. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zuckerman:1991:TTA

[Zuc91] David Zuckerman. On the time to traverse all edges of a graph. Information Processing Letters, 38(6):335–337, June 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zwick:1991:EKT

[Zwi91] U. Zwick. An extension of Khrapchenko’s theorem. Information Processing Letters, 37 (4):215–217, February 28, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zwick:1996:NAV

[Zwi96] Uri Zwick. On the number of ANDs versus the number of ORs in monotone Boolean circuits. Information Processing Letters, 59(1):29–30, July 8, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).